A Complete Bibliography of the Journal of Computational and Graphical Statistics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

28 March 2025 Version 2.09

Title word cross-reference

(m) [KNN05]. 2 [MBGO11]. 2 × 2 [HVRA96, LH95]. 3 [Dri08, LTW^{+}10]. [0, 1]^{d} [WC94]. _{1} [NST12]. _{2} [CC22a]. χ^{2} [Sta95]. δ^{18}O [RL14]. ℓ_{1} [Kat10, Yua08]. F [Mul98]. G [LD11, vdBBD22, FN16]. H [HPOL14]. H_{0} [Mul98]. K [Ano24a, CC22b, GBS24, LLZ23, MSS08, Gho07, MR10, YK09]. κ [LT94]. L_{0} [FML19, Joh13, KCF21, LW07]. L_{1} [LZ08, LW07, SMB14]. L_{1/2} [KF24]. R^{2} [GŠP^{+}06]. n [HY17, OR99, Paa99]. P [BCH05, GIA22, Hea22, HY17, HB98, NP01, WFCT23]. p > 2 [LMM19]. ψ [LSD05]. Q [LFZ97, RK22]. R [JKD92, MC18, Sta95]. ρ [VT94]. R × C [Fri00, LPM98, VT94]. S [VT94, Rup92, SBY06]. σ [FW13, LFT17]. t [GB02, PLW01, Som98, TEG19, TMMB20, Tsi02]. T_{3} [GB00]. τ [SBWZ08]. U [WDB96, VC21]. X [XMY13]. Y [XMY13].

-An [HR11]. -CDFs [LLZ23]. -Fold [YK09]. -Learning [LSD05]. -Likelihood [HPOL14]. -mean-directions [MR10]. -means [Ano24a, MSS08, CC22b]. -Medians [GBS24]. -Norm [LZ08]. -Penalization [SMB14]. -Penalized [NST12]. -Priors [FN16]. -Regression [SBY06]. -Segmentation [Joh13]. -Stable [FW13, LFT17]. -Statistical [VC21]. -statistics [WDB96]. -Torus [GIA22]. -Value [WFCT23, HB98]. -Values [Hea22]. -Vines [MC18]. -Way [Paa99, Sta95]. -Wishart [LD11, vdBBD22].

1 [Bak92a]. 1-Bit [CZ22]. 1130 [Ano24a]. 15 [Ano12b]. 150-Year-Old [VRH19]. 19 [THM^{+}23, WKT^{+}24].

20 [Goo11].

32 [Ano24a].

40 [Goo11]. 45 [Goo11].

‘50 [HJ17]. 523 [Ano12b]. 526 [OLN11a].

ABC [ER21, ILP19, NFMS14]. ABC-CDE [ILP19]. ABC-SMC [ER21]. Aberrant [MM09]. Abrupt [NST12]. Accelerated [BCH22, CSWZ23, JWL^{+}23, KLH05, SBR22, SNC23, ZL16]. Accelerating [ASND19, HHC03, HV19]. Acceleration [AX23, HV19, KYW19, GS94]. Accept [CR98]. Acceptance [ER21, QTVK18, SGH17, Tan06]. Acceptance-Rejection [Tan06]. Access [SH03]. Accompaniment [Rap01]. Accounting [HDM02, ZHJ16]. Accumulated [BM19]. Accuracy [DNF23, FLMM22]. Accurate [CJ02, HAC^{+}12, HSZ23, LH95, MBY^{+}22, Mil19, TEG19, WFCT23, WK17, ZSW22]. Acoustic [BMO01]. Across [Deg23, GHCU22, SQCK20]. Active [RFV23, WBW21]. Activity [FCL10]. Activity-Event [FCL10]. Acyclic [KB08, ZHZW22]. Adaptation [BCD11, GMTK13]. Adaptive [Ano08a, BMC05, Bil00, BJMD13, BK05, CP22, CKNK06, CTDKL21, CKR19, CGL11, CRC^{+}07, CMJ11, DZ99, Efr05, EHW03, EH09, FG95, FT16, GK97, Gho07, GK10, GT11, Gou98, Gri99, HT17, HCE23, HWF22, IW22, JPvD22, JS13, JBJ^{+}22, KNH23, KSM14, KCK08, LGS13, LAJ13, LSV^{+}09, LW09, LRFK21, LO16, MFMR19, MHDW19, MMV13, MS03, NW10, Oeh98, PTP03, PC06, Pit02, PSRM^{+}23, RR09, SG00, SGH17, SBK23, SW18, SBR10, WB18a, WGDW22, WJL23, WC06, WKC^{+}08, WCR13, YLCR19, Zha97, ZJA16, vDE19, vNvdBvdW22]. Adaptively [WL13a]. AdaptSPEC [BRCC22]. AdaptSPEC-X [BRCC22]. Addendum [Ano12b]. Adding [DS21]. Addition [OSR^{+}18, SWX23]. Additional [SBE17]. Additive [Ano20, BHK15, CLXY23, CR24, CCL21, DMR20, DCV10, EM02, FM12, FNGW20, GKZ10, HHH22, HY19, JPvD22, KO04, KLH99, KSM^{+}18, LZM24, LBCG16, MMRMGMGC08, MHS^{+}14, MH17, OR99, PWS16, PCG^{+}14, SSG15, ST99, SSG23, TV24, TL07, UKZ18, VN18, WYL14, YASR20, ZWZW24]. Additivity [RAC23]. Adjusted [BFM18, CVA08, DG14, DIR02, JLAW22, KTT23, MK92, Tan17, ZCS^{+}19]. Adjustment [BSZ22, RF17a, Sha14, YNTK21]. Adjustments [SS11]. ADMM [RT16, Zhu17]. Advanced [LYG^{+}10]. Advances [EH15]. AdWords [LRR20]. Affine [KO05]. AFT [HJ04]. After [Hub99]. Again [HV17]. Against [PRS11, SŠS07]. Age [JVSS24, SH17]. Age-Specific [JVSS24, SH17]. Agglomerative [Hea11]. Aggregate [OSR^{+}18, Par11]. Aggregated [GY07, HCZMH21, Pro06, WBS21]. Aggregation [HWF22, KSLS22]. Agreement [LT94]. Ahead [Hic17]. Aided [HIT99]. Aids [TN14b]. Air [EO95]. Airline [Wic11c]. Airports [Wic11a]. Akaike [Ano23b, ST99, Sod20]. Algorithm [AB17, ACG14, AT16, Bak92b, BCH09, BL04, BJMD13, CB01, Cap11, CCFM01, Che07a, CC18, CW00, CM08, DASR08, DZL22, DH10, DZFZ18, DMP14, ER04, FSB11, Fea04, FM10, FH15, FN04, GJ97, Hob08, Hoe10, HGK10, HRV12, HL00a, HB16, JK00, Jan06, JS95, JL13, Joh13, Jon98, KHYD23, KB08, Kar21, KPC02, KKK08, KCSA20, KBM23, LW08, LC01, LF16, LD06, LLZ09, LLZ23, LWH23, Maa05, MR10, MPTZ23, MHDW19, MM09, Mey03, MMV13, MR01, Mül99, NG04, OT11, PKH15, Pan99b, PW15, PNLL21, Pie04, PSRM^{+}23, QE96, Res22, RAM19, RL14, SBM22, SBY06, SZ08, SMB14, Smi95, Smy02, SDL19, SC23, TMvD18, VM04, VG01, Wan18, WH98, WK00, XGTB18, YZ13, YH17, YSD13, YZ06, YLW17, ZNYJ14, Zha22, ZL16, Zho05, ZL13, Zhu17, dSCL21, vdBBD22]. algorithm [Bak92a]. Algorithmic [CKR19, WNA20]. Algorithms [AX23, AE92, Béd08, BF02, BK22, BQS94, BDR97, Cai99, CZ16, DJM22, DGS24, FDC^{+}19, GD21, GKZ18, GK06, GV01, GR93, GJR05, HMP96, Hel24, HV19, HLH06, HD23, HHH12, JO02, KMOS92, KA13, KSR19, KSSL21, KYW19, LMST18, LGS09, LL22, LB15, MM10, MWM^{+}23, OSR^{+}18, PLW01, PC06, Pit02, RT16, SS19, SBEG94, Tan15, Tar99, VDC21, WR93, XBW15, XCSP16, Yen20, Yu12, YNTK21, ZJ04, ZC15, ZBS^{+}21, ZL10, ZHZL19, dVTP^{+}17, vlM97, vD00]. Alignment [Ano24a, CC22b, LKR22, RG09]. All-Pairwise [Pie04]. Allocation [BP98, CD19, DGL23, LF16, XD20]. Allocation-Based [LF16]. Allocations [DJA23, PI10]. Allowing [MC18, NST12]. Alternating [CHW08, ZY21, ZYZ10]. Alternative [EM18, Fos01, JU00]. Alternatives [SŠS07, Tuk93a]. Always [PLIK02]. AMlet [ST04]. Among [HM11, Hob01b]. Amplitude [ZKS18]. Analyses [Bri94, GL07, MBS04, PCP11]. Analysis [ACG14, BZ07, BM19, BRS^{+}08, BK03a, BLR22, BZ97, BG00, Buj04, CRGS17, CCS03, CCHS08, CHPR24, CZSL10, CKT21, CR23, CH99, CBM24, CW00, CVQ23, CLCX22, DDM20, Dem23, DH03b, EH17, ES96, Fal99, FG05, Far92, Far99, Far12, FS99, Fie12, FIK19, Fro22, GVBM04, GZQ^{+}22, GEGM03, Gel04a, Gel04b, GD09, GR12b, GM99, GJL^{+}10, GAZ15, HLH15, HAPRD24, HT17, HM15, HV05, Hob08, HM16, HLZH05, Hub00, IG96, IKL08, IW22, JLAW22, JCZG19, Jör09, KXY20, KTZQ20, KNH23, Koo98, KLZ23, LW08, LRB12, LW09, LPS99, LSH16, LRFK21, LZ23, LO16, LP14, LWH23, LH00, MS18a, MHH15, MAJ^{+}98, MHG24, MM09, MW96, MY94, MPH22, MVR02, MBY^{+}22, MDD^{+}12, NP01, NYC15, NW10, NAB16, NFMS14, NBH^{+}15, OC01, Paa99, PM20, PI10, PLH10, Par11, PL12]. Analysis [PRCF24, PTP03, QLCZ15, QWW23, QN00, RA01, ROY20, RDL07, RT98, STM15, SJH^{+}15, Sco99, SS01, SSB^{+}14, SYW^{+}22, SJM^{+}24, Smi01, SC98, SK24b, TS22, Tan15, Tan17, Tre06, TU11, Tu94, TN14b, WH17, WCDS03, WK00, WCSY21, YL22, YBDS02, YSJS23, YHS23, YLM17, YWBA22, YSWS24, Zha97, ZYKC08, ZP14, ZH03, ZCY^{+}11, ZHT06, ZGPM22, dLM04]. Analysts [YL95]. Analytic [Ano06a, Hsu92, KMS23, KGMT06]. Analytical [YT99]. Analytics [Cha24]. Analyzing [HMP96, Mci99, OFR^{+}18, VG01]. Anchors [NJM20]. Ancillarity [HR11, YM11b]. Anderson [HV19]. Angle [Gra13, HK10, XWZ15, ZLWZ16, ZY24]. Angle-Based [ZLWZ16]. Angular [RIC^{+}16]. Animated [SVC^{+}19]. Animation [Far04]. Animations [LTW^{+}10]. Anisotropic [BMR15]. Annealed [KA13]. Annealing [Glu00, Glu04, HW21, IP04, Kar21]. Annotation [MI00]. Anomalous [Ano23b, LCSL14, Sod20, SMN15]. Anomaly [FEF22, OBGC21, THSM^{+}20, THSM21]. ANOVA [GW93, Hoo07, KMS23, QG06, YT21a]. Anthrax [LBM^{+}14]. Antithetic [HJ09]. APL2 [FF94]. Apogee [SUL23]. Appearance [DGL23]. Application [Ano07a, AG18, BK05, CP22, DASR08, DXL12, EH15, Fea04, Fea05, FHVW03, GVBM04, GM15, GJ09, Glu04, Gol09, GX20, GRM09, HMW06, JVSS24, JYYZ18, KTZQ20, Koo98, LACV23, LYHM15, LBM^{+}14, LZ23, Lia07, LMS19, MC05, PAP22, RHC^{+}14, RZ22, RL14, Rya03, SH17, SLR^{+}23, SM13, STIP23, TNKK16, THM^{+}23, Wan20, XP07, YL99, ZSL^{+}22, ZMY^{+}23, ZRL14, ZL16, Zhu17, Fuk22, WDB96]. Applications [ALP^{+}23, BM19, CDM^{+}20, Hof09, HHTB10, Ish99, JKD92, Jör09, KS08, LS93, LPS99, LJ19, MPR^{+}12, MMN^{+}13, Mey03, NP01, NYC15, PvD09, PJS^{+}22, QN00, RRH18, RW10, Wan18, WCKL19, WHH23, WIR01, Wu08, ZCI^{+}17, ZSL23]. Applying [LL15]. Approach [ACS17, ABvS09, ACCQ19, ACIP22, BCD11, BBW20, BEGS17, BK20, Bra21, BDR97, BR02, CHZ23, CS98, CAR^{+}10, CH16, CTW20, CH99, EM18, GD16, GK02, GY07, GKZ10, GH18, GLC20, GB12, HYL^{+}17, HPQ24, Hsu92, KK09, KGM18, LSV^{+}09, LM19, LAB22, LGL16, LDHW23, MVR02, MXM^{+}21, MI00, NM98, Par01, Par05, PH20, PP09, PRHP14, QC24, RA01, SS01, SL99, SBS^{+}23a, SSBW17, SDBD21, SW18, TZ12, Tre06, Tu94, WW12, WCKL19, WZZ^{+}24, WSM^{+}20, WWJ22, WK00, YNH18, YHS23, YOL^{+}22, ZHZ12, ZZW21, ZMY^{+}23, ZJA16, ZDM20, ZF18]. Approaches [KS04, XW18, ZCO15]. Appropriateness [Zha03]. Approximate [Ano06a, BSZ22, Bra99, BC19, DBPW23, FD21, GX01, GF17, HM16, HMZ22, IJ02, ILP19, KGMT06, MMF^{+}19, NFMS14, OW12, Pic14, RNHY12, RHF24, RS01, SG17, TA12, YNTK21]. Approximated [SCG22, YOL^{+}22]. Approximating [Bor11, CB01, ZZWS23]. Approximation [BCE20, BMGRR22, BRS^{+}08, BRS12, BBvdW16, CCX22, CKHG15, CHPR24, CJ02, CHW08, CF05, DZL22, Fea04, GJ09, GA15, HS18, HSZ23, HB16, Jan06, JL13, KSR19, LYXM23, LQ23, MMLLQ^{+}22, MPR^{+}12, MYC^{+}22, Mil19, Mul98, MGF22, NTVK12, ONS18, RYY00, RF17b, SLMN20, SY07, TN14a, ZZW21, ZSW22, ZGTG21, ZLW21, ZBY09]. Approximation-Langevinized [DZL22]. Approximations [BLBCC22, BSX23, Bor11, DNF23, Hes95, HLL07, HWO^{+}17, HK98, KMGJ23, KSM14, NFM16, PB95, RO94, SYNK24, SLBV18, SBS23b, WM19]. Approximator [Hoe10]. Arbitrarily [KLH05, Zho05]. Arbitrary [EM07, HHL19, OLN11a, OLN11b]. Architecture [CCS03, YVMFB03, ZA23]. ArcView^{T M} [SCLK^{+}00]. Area [GGvdV10, XW20]. Area-Proportional [XW20]. Areal [Hug15, VF09]. Areas [JWH16]. Argumentation [Hig01, Hob01a, HJT01, Lev01, Liu01, MG01, WZ01, vDM01b]. ARIES [EFLT94]. ARMA [dLG01]. Arms [PRS11]. Arranging [Mur99]. Array [LVH17]. Arrays [MM14]. Arrows [PRS11]. Art [Hig01, Hob01a, HJT01, Lev01, Liu01, MG01, WZ01, vDM01a, vDM01b]. Artery [BM19]. Article [Ano01b]. Artifacts [BML19]. Artificial [PI10]. ASA [Wic11b]. ASIS [HR11, YM11b]. Aspects [BC93, CRL13, EGO92, JMY^{+}16, LD11, LPS99, QN00, Tuk93a]. Assessing [Laz05, LCTV24, PMB12, RFVJ21, RHM18, TSH16, YKZ18]. Assessment [BG00, BGP03, FBG06, HPZ22, KM03, QM04, SK10, SG17, Yan21, YNTK21, ZJRI10, ZCI^{+}17]. Assignment [MR01]. Assimilation [LPS99]. Assisted [GdTv96, Hui22, Wan20, ZGTG21]. Associate [Ano24b]. Association [CTW20, KNN05, MLP07]. Assumptions [CHLZ22, WW12]. Asymmetric [Hen04, LO16, LL15, RB02]. Asymptotic [BF02, GZQ^{+}22, Hof19, RHF24, Tho93, XSC19]. Asymptotically [VDC21]. Asymptotics [BB03, BO05, SR12]. Asynchronous [SDL19, ZKS22b]. Atmosphere [HMN^{+}03]. Atmosphere-Ocean [HMN^{+}03].

Atmospheric [LPS99]. Attempts [YHS23]. Attracting [TMvD18]. Attribute [KSC^{+}24]. Auctions [SJ05]. Augmentation [BK23, FCWM17, HIS22, HW21, Kel11, PKH15, PRS13, RDL07, VDC21, Wu11, ZKS22b, vDM01a, TYG^{+}20]. Augmentations [XMY13]. Augmented [Zhu17]. author [Tuk93a]. authors [FB94]. Autocorrelated [PP96]. Autocorrelation [CP16]. Autocovariance [LPHK16]. Autocovariances [AV22, McE22]. Autoencoder [FX21]. Autoencoder-Based [FX21]. Autologistic [FP04, MPR^{+}12, ZZ08]. Automated [FHIT20, TGHL14]. Automatic [All13, BJMD13, EN20, HMJ11, HDK^{+}20, PGR19, Rap01, RH94, SAT04, Ska02, ZJA16]. Automating [Sta94]. Autoregressions [Hea23, HTW22]. Autoregressive [DZZ16, HHT21, KC20, MT03, MC05, VM04, WRK11]. Auxiliary [FDBS10, Gua23, MMF^{+}19, PL12, Tüc08, TW96, WRT21, WZZ^{+}24, ZDM20]. Available [BCE20]. Averages [HCB04, Moc98]. Averaging [CLXY23, CGL11, VHSG20, YKW03, YSK^{+}10].

B [FB94, ERPG11]. B-Spline [ERPG11]. Baccalaureate [MHB17]. Backfitting [HY19, KO04, OR99]. Backscoring [Far12]. Backward [KSV23, MÉG^{+}23, PRHP14, WC06]. Baddeley [FR97]. Bagplots [HS10]. Balancing [JS20]. BAMLSS [UKZ18]. Band [BMR15]. Bands [DJ04, HO93, Hut02, LJZD05, LS97, RW99, WASB22]. Bandwidth [FG95, GG17, Her97, KO04, MMRMGMGC08]. Barrier [PL22]. BART [PCGM20, WHH23]. Bartlett [Mul98]. Barycenters [LZ18]. Baseball [STIP23]. Based [Ano20, ABG^{+}21, ACG14, BCD11, Big06, BH02, BSX23, CHW02, CGEMIR18, CS20, CTB19, CHW08, Chr04, CBC93, De 05, DMR20, DCV10, ERPG11, ES19, EH09, FX21, FRW05, FLMMK24, GZQ^{+}22, GR12b, GM20, GG17, GIA22, GH18, GKN23, HSM09, HK10, HHKS11, HHH12, HB16, IW10, JRS11, JTU03, JO02, JRH22, KKLN20, KK09, KKK08, KMGJ23, KSLS22, LG07, LL12, Li05, LF16, LZ18, LDR23, LM23, LJZD05, LD06, LS97, MMF^{+}19, MHG24, MHDW19, MB18, Mey18, MRCM18, MKN23, MGF22, NGH17, OR07, OKL21, PDM23, PI10, PP05, Pie04, Pos01, RAM19, SBR22, SS19, SH03, Sew21, Smi01, SHL12, TSG92, TG13, VM04, WCH20a, Wil96, XBW15, YK09, ZMH07, ZHH08, ZZW21, ZJRI10, ZLWZ16, ZSBS23, ZOR12, ZRL14, ZCO15, ZW10, ZWS23, dLG01, BHvdLD22, CSWZ23, FS18, MXM^{+}21, OH24, PC22]. based [RR22, RHO22, Shi24, SdWDL23, SF23, VHSG20, YSWS24]. Bases [Dob12, Fos01, Kei96]. Basic [Tan15]. Basis [BRS^{+}08, CR23, CVQ23, HLL07, KKB21, KKHB23, MYC^{+}22]. Basketball [YHS23]. Batch [QWW23]. Bayes [BBW20, DH10, FLMM22, FLMMK24, Fri13a, GKN23, Ho06, KDM16, KBM23, LMS19, MPR^{+}12, NFMS14, SL99, SS05, TNKK16, TNK17, WR05, YOL^{+}22]. Bayesian [Ano08a, AB21, ASND19, AVW13, ACCQ19, BMC05, BD23, BS19, BC16, BAMG22, BRB17, BSZ22, Bil00, BRS^{+}08, BSX23, BFM18, BHK15, Bra21, BZ97, BHF02, BK03b, BC19, CDM^{+}20, CPP23, CD12, CBM22, CS99, CCYW16, CD19, CRL13, CGL11, CGR06, CYS09, CRC^{+}07, CM13, DC22, DJM22, DJA23, DV04, De 05, DK18, DGM17, DH03b, DMP14, DCS16, EH05, Eve12, EB02, FK23, FDBS10, FK16, FDC^{+}19, FD21, FPRW09, FR97, GD21, GD16, GK02, GD09, GHC14, Gol09, GD23, GM18, GW93, GQD18, HAC^{+}12, HAPRD24, HJ04, HK22, HZC23, Hil11, HMN^{+}03, HRM02, HLB^{+}21, HDRM05, Ish99, ILP19, JPvD22, JS13, JKD92, JBJ^{+}22, JT21, JP22, Kar21, KFSL17, KNN05, KF24, Kle19, KO05, KSM14, KMQ05, KB20, KC20, KSM22, KY95, LGS09, LGS13, LB04, LG07]. Bayesian [LCW22, Len03, LCRG21, LF16, LRFK21, LZ23, LDR23, LD22, LG20, LLRM24, Loc22, LS05, LHK17, LCD^{+}24, MP16, MWFSG17, MVR14, MKE^{+}21, MMF^{+}19, MHDW19, MMN^{+}13, MW15, MKK18, MV98, MQ17, NMD^{+}20, NW10, NG04, NL04, NFMS14, OKL21, PM20, PI10, Par01, PVH05, Par11, PL12, PJS^{+}22, PMI20, PFBSO22, Pic14, PL22, PCG^{+}14, PGM23, PDLN18, PSF^{+}22, QN00, QM04, RSKZ19, RHF24, RF17a, RW14, RT21, RL14, RT98, SMS^{+}14, Sar22, SG03, Sco99, SR00, SR12, SG23, SG17, SBJ17, SM13, SL10, SX21, SSL17, SK24b, TDM12, TH09, TPA19, TF20a, TRWB16, TKB23, TFS20, TNNK20, Tsi02, Tüc08, UKZ18, VHSG20, WW12, Wan08, WD11, WGDW22, WZZ^{+}24, WBS22, WL24, WRK11, WTW07, XCSP16, XWS24, YL22, YK01, YKW03, YHS23, ZG22, ZBB06, ZCI^{+}17, ZWF^{+}22, ZDM20]. Bayesian [ZICT12, ZLS23]. Be [YM11a]. Beat [VH17]. Beautiful [BY18]. Becker [Tuk93a]. Before [HV17]. Behavior [Dia99, Par11, RT21]. behind [SCLK^{+}00]. Belief [MMA97]. Benchmark [GMP04, HLZH05]. Best [GK06, KPC02, Mül99]. Best-Subset [GK06]. BET [DCS16]. Beta [DW01, EB02, MS18a]. Beta-Binomial [EB02, MS18a]. Better [Bri11, DS21, FK11, Gel11a, Gel11b, LT94, Wai11, Wil11]. Between [BM99, DBH^{+}22, God01, YF22, ZLPT22, LT94, YHH^{+}23]. Beyond [UKZ18, WM19, XC22]. Bi [CDM^{+}20]. Bi-Level [CDM^{+}20]. Biadditive [dR07]. Bias [FG95, Fri13b, GH21, Len09, SZ08]. Biased [CHLZ22]. Biclustered [CHSR19]. Biclustering [MM14, TW14]. Biconvex [CX23]. Big [BCD^{+}24, BW17, DYB^{+}23, FLY21, HLB^{+}21, SBM22, SSBW17]. Billion [LZM24]. Binary [ADY10, BRB17, Che23, CMC^{+}22, FPRW09, LT94, MN97, MM09, MPS92, QN00, QM04, RSDG23, SS06, TL92, TA12, TG13, WKC^{+}08, YK01]. Bingham [Hof09]. Binned [SCAP22]. Binomial [AC95, BHF02, EB02, MS18a]. Biological [FCL10, Mic12]. Biomedical [WZ22]. Bipartite [MCF^{+}20]. Biplot [BL22, CGL13]. Biplots [GGvdV10, Gre13, KSC^{+}24, Fuk22]. Bit [CZ22]. Bivariate [AG05, BL04, GMC02, Gu15, GHCU22, HMS97, KL03, Koo98, LGL16, Maa05, MS93a, MHD20, Mül99, WH98, WKT^{+}24]. BIVAS [CDM^{+}20]. Black [CS98, GKBP15, TVZ^{+}22, Yat23]. Black-Box [CS98, Yat23]. Blind [Kan20]. blink [MKE^{+}21]. Block [BR02, CC13, CAPP23, DV04, DBS06, ESR^{+}14, KTT23, MCB24, PKH15, QTV^{+}21, RW10, SBT00, SLL22a, SK24a, TH09, TRC23, Yen20, ZCS^{+}19, ZGC22]. Block-Poisson [QTV^{+}21]. Block-Split [TRC23]. Block-Wise [SLL22a]. Blockmodels [RAM19]. Blocks [LBM^{+}14]. Blockwise [Efr05]. Board [Ano12a, Ano15a, Ano16b, Ano17a, Ano18b, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano00e, Ano13b, Ano14b]. Boarding [HCK^{+}11]. Boolean [CSY00]. BOOST [KCF21]. Boosted [GH21, WH14]. Boosting [ADY10, BMW17, CKPT04, CMJ11, GH21, HR11, HHKS11, LB06, SBR22, TG10, TG13, Wan13, Wan18, YM11b]. Boosts [HLB^{+}21]. Bootstrap [BP98, CC13, DW02, EN20, Fon19, Hes95, Hut02, KLH99, KS04, LL11, MMRMGMGC08, San23, TK99]. Bootstrapped [BBM24]. Bootstrapping [KCO09, Pan99a]. Both [SQ07]. Bound [GK06, SR00]. Boundaries [FKH07, LACV23, MB18]. Boundary [DS94]. Bounded [WSM^{+}20]. Bounds [Coz99, SJSG22]. Box [CS98, GKBP15, TVZ^{+}22, Yat23]. Boxplot [BC00, RK22]. Boxplots [HWK17, HS10, QG22, RMR^{+}17, SG11]. Bradley [CD12]. Brain [BM19, Dri08, MBY^{+}22, RHC^{+}14, SSL^{+}16, ZSL23]. Branch [GK06]. Branch-and-Bound [GK06]. Branching [GF22]. Breakdown [JP95]. Breakpoints [VH22]. Breaks [OS22, SBM22]. Bridge [Fu98, GLC20, KSV23, MS02]. Brief [Fie12, Fri02, Lev12, Lev13b]. Brownian [BMR15, Ste02]. Broyden [AX23]. Brunelleschi [GVBM04]. Build [KSR19]. Building [BVZ20, DYB^{+}23, HABV06]. Bump [DR10, HH01]. Bumping [TK99]. Bumpy [MMW98]. Burning [LCL22].

C [HLH15]. C-MPPCA [HLH15]. Calculating [BPW93]. Calculation [BCH05, Coz99, GS24, Kab93, McE22, WFCT23]. Calculations [Sta95, TA12]. Calendar [WCH20a]. Calendar-Based [WCH20a]. Calibrated [KNS21]. Calibration [BRS^{+}08, HWLB24, TSH16, WS24]. Can [BLR22, MMO01, Zhe17]. Cancelations [Wic11c]. Canceled [HCK^{+}11]. Cancer [ZCI^{+}17]. Canonical [CKT21]. Capture [BBLG13]. Capture-Recapture [BBLG13]. Cards [LYG^{+}10]. Care [VH15]. Carlo [Ano07a, BGD16, BP98, Bro06, BG00, BC19, CB01, Can99, CGMR04, CS99, CF05, Chr04, CRS06, CMC^{+}22, DG14, DH03b, DBS06, DMP14, FBG06, FKH07, Fea02, Fea05, FHIT20, Fis12, FWR12, GS94, Gel95, Gey95, GVDP99, God01, GF16, GLW07, GR08, Gra07, GMP04, GJR05, HT17, Hel04, HZC23, Hob08, Hol23, HLH06, HW21, HJM22, HK98, Ish99, JN04, JB03, JHD21, JS13, JT21, JO02, Kit96, Kle19, Kle22, LGS13, LSSG15, LYG^{+}10, LC01, Lia07, LJN^{+}17, MP00, MFMR19, MN97, MDS20, ML08, NJM20, OKL21, PP14, PCP11, PH21, RJLW21, SHC00, SNR17, SS06, Spa05, SBF19, Tan06, Tan15, WGDW22, WJG11, XBW15, YSDJ15, YT99, ZC15, ZZ08, ZJA16, dV08]. Carlo-Adjusted [DG14]. Carlo-Based [XBW15]. Carlson [JKD92]. Carr [FB94]. CART [WTW07]. Cartography [WS02b]. CARTscans [NEL04]. Case [AvD20, BZ97, HK17, MK92, RNHY12, SSL^{+}16, WM03, XNN16, ZICT12]. Case-Control [RNHY12]. Case-Specific [XNN16]. Cases [THM^{+}23]. Casting [CSS^{+}09]. Cataloging [FS99]. Categorical [ACC18, Bak92b, FB11, Fri99, LG03, MPS92, Sch24, SS01, TvdV22, TS13, Bak92a]. Category [PGR19]. Causal [CBM22, GMdL23, Hil11, SJSG22, WHH23]. Cave [BCL94]. CDE [ILP19]. CDFs [LLZ23]. cDNA [YBDS02]. Censored [AB17, AE92, BL04, CGL13, CJK22, CPCC24, De 05, DLSS24, GV01, HJ04, KLH05, KS92, LG92, LS05, Maa05, Pan99b, SX23, SS22, Tu94, VL09, Wan23, YNH18, ZJ04, Zho05, ZZL18]. Censored-Data [GV01]. Censored/Truncated [Zho05]. Censoring [DMR20, Ano20]. Census [MK92]. Center [HR11, YM11b]. Centered [AV22]. Central [Ano06a, FLMN23, KGMT06, PSY09]. Certain [MS93a]. Chain [BG00, CB01, Can99, CS98, Fea02, FHIT20, Fis12, FWR12, GVDP99, God01, GF16, GR08, HZC23, Hol23, JS13, JT21, JO02, JV01, LSSG15, MP00, MFMR19, MN97, MDS20, ML08, Nea00, Owe17, SHC00, SK21, SNR17, Tan15, YT99, dV08, Bro06, CF05, CRS06, GS94, Gey95, Gra07, GMP04, GJR05, Hob08, HLH06, JN04, SS06, ZZ08]. Chains [DSD^{+}23, Gué03, HK22, HC98, MB04, MDS20, TDH15]. Challenge [EFG^{+}99]. Challenges [Fie12, FDG^{+}22, Goo11, KP03, Mic12]. Change [CPP23, CR08, DM24a, LKB21, MW04, RL14, XC22, ZZSW23, ZY21]. Change-Point [LKB21, ZY21]. Change-Points [CR08]. Changepoint [HEF17, HT17, SHL19, TEFH20, UTK22, WKM22, YSD13]. Changepoints [YK01]. Changes [FML19, MW04, NST12]. Changing [Gri99]. Characteristics [MBM18, MM20]. Characterization [ABvS09]. Characterizing [MS93a, RMR^{+}17, ZH17]. Chart [SSP^{+}18]. Charts [VRH19]. Cheaper [SD03]. Checking [BWZO20, LCLR23, Par01]. Chernoff [GW01]. Chi [SPV22]. Chi-Square [SPV22]. Chiaromonte [Ano24a]. Child [ZJRI10]. Choice [CGEGMI22, CWZ21, FK23, FH15, Gho07, Her97, LHC17, Par11, PW14, RHO22]. Choices [CCHM18]. Cholesky [DP23, Smi95, Smi01]. Choosing [DBH^{+}22]. Chosen [Ros12]. Circles [DG24]. Circulant [CP18, GF17, HPA14, HKP16]. Circular [DG24, XW20]. Circus [BW17]. Class [ACJ21, BC24, DH10, JHL09, KGM18, KYW19, LAB22, MHS11, SAT04, TvdV22, WP13, WRT21, WZXW24, ZW10]. Class-Distributed [WZXW24]. Classes [SQCK20, SSP^{+}18, WZXW24]. Classical [HY19]. Classification [ADY10, BMO01, BOW02, CS14, DG05, FHVW03, FR97, GTZH16, Gho07, GAZ15, Hen04, Kar21, KL03, KM03, LACV23, LGS09, LCKL05, LZYQ22, MR19, NMD^{+}20, RSDG23, SZ08, SKK22, TMMB20, TL92, Vel08, WR05, Wan13, Wu08, dSCL21]. Classifier [SdWDL23]. Classifiers [FGV22, RR22, WL13a]. Cleveland [Tuk93a]. Climate [RS19]. Clinical [HS08, WC06, ZCI^{+}17]. Closed [SY07]. Closed-Form [SY07]. Cloud [LZ18, MGL18]. Cluster [ACG14, Cai99, CGEGMI22, CW00, Jör09, LCSL14, LFS^{+}17, MWM^{+}23, PMS21, RMH21, SN10, TW05]. Cluster-Weighted [MWM^{+}23]. Clustered [CC13, GM15, LRBS03]. Clusteredness [RHM18]. Clustering [ACC18, ACCK19, Ano24a, BCDL17, BRC^{+}10, BT18, BK03a, CGEMIR18, CH12, CX23, CTDKL21, CL15, CBL20, CC22b, DJM22, DSD^{+}23, DCS16, ES19, FRW05, GR12a, GM20, GLW07, GIA22, HH11, Hea11, Hof19, HLYM15, HWF22, Hur04, HB16, LG07, LL12, Li05, LM23, LKR22, LLZ23, MR10, MM10, MJ18, MSBV22, Mel16, MWM^{+}23, ML08, MQ15, MSS08, MW14, NMD^{+}20, OR07, PQM22, Pos01, PM16, RT21, SS12, SLL22a, Sew21, SBL23, VC21, WZSF18, WNA20, XP07, YL22, YPPM21, YGD22, YHS23, YSWS24, ZSK19, ZGC22, ZZT22, MCB24]. Clustering-by-Segmentation [BCDL17]. clusterMLD [ZZT22]. Clusters [DKH06, DCT08, FSMV02, FRW05, FP20, GBS24, Kou14, VH17]. Clutter [MLP07]. Co [CBL20, MCB24, MDC17, SLL22b]. Co-Clustering [CBL20, MCB24]. Co-Sparse [MDC17]. Co-supervised [SLL22b]. Coarseness [PC22]. Code [CH93, IW22]. Coding [Hes98, JH98]. Coefficient [AGV12, CS20, DXL12, KW21, LG20, WH14, WJL23]. Coefficients [BRMC19, CDFB11, SR23, WKT^{+}24, ZLP23]. Cointegrating [MKN23]. Cokriging [HCB04]. Collaborative [YGD22]. Collaborators [Ano92f, Ano93e, Ano94f, Ano95e, Ano96g, Ano97e, Ano98e, Ano99e, Ano00f, Ano03a, Ano06b, Ano07b, Ano08b, Ano09a, Ano10a, Ano11a, Ano13a, Ano14a, Ano15b, Ano16c, Ano17b, Ano21]. Collapse [SR00]. Collapsed [MJVB18, PvD09, VJ15, YOL^{+}22]. Collections [Deg23]. Collective [FEF22]. Collinearity [LWM21]. Collocation [XWS24]. Color [GM93]. Column [HM16]. Column-Sampling [HM16]. COM [CS20, Tie00]. COM-Poisson [CS20]. Combination [LW07, PGR19]. Combinatorial [MVR02]. Combined [LYLL23]. Combining [BRC^{+}10, DCV10, EH17, PP07]. Comment [Bri11, FK11, Fri14, HR11, Hol93, Kel11, Kos13, McC99, Mur13, Pen17, Tre93, Wai11, Wil93, Wil11]. Comments [Wai93, Tuk93a]. Common [IKL08, LH95, TEFH20]. Communication [RAM19, WLZ^{+}24]. Communication-Efficient [WLZ^{+}24]. Communities [SYF17]. Community [HCCZ23, ZZD23]. Comparative [PCP11]. Comparing [ACIP22, BCL94, TZ12]. Comparison [DJA23, GB02, GB00, GB04, GR93, HH06a, HP94, KS04, PHK14, SS05, SY07, YBDS02, ZJA16]. Comparisons [BC93, HN98, Pie04, STIP23, Tan15, Wai93, Tuk93a]. Compatibility [HC98]. Competing [KK09, KSSL21, WKT^{+}24]. Complete [BK23]. Completion [CHSR19, CKR19, KH23, MWWY24]. Complex [ABG^{+}21, BY18, Buj04, CMC^{+}22, CP18, CYS09, DGS24, EM18, Gel04a, Gel04b, GW15, IW22, ILP19, MWWY24, SH97, SHMD^{+}14]. Complex-Valued [CP18]. Complexity [ES02, FKLW08, YPPM21]. Component [CRGS17, CCFM01, CTB19, CR23, CLCX22, Fal99, Fro22, GW93, GJL^{+}10, GAZ15, HLH15, JLAW22, JCZG19, JU00, JTU03, KXY20, KTZQ20, KNH23, KLZ23, LRB12, LW09, LSH16, LO16, LWH23, MBY^{+}22, PNLL21, SYW^{+}22, SJM^{+}24, WH17, WWL^{+}23, YLCR19, YSWS24, ZSK23, ZP14, ZCY^{+}11, ZHT06, ZGPM22]. Component-Based [CTB19]. Component-Wise [CCFM01, GW93, YLCR19]. Components [BRC^{+}10, Dan97, GCKP08, HM11, HM16, JU00, MHH15, Mel16, Par05, PP09, YKZ18, ZHZL19]. Componentwise [SBR22]. Composite [BCDL17, CHG16, ESR^{+}14, LO16, ZK22]. Composition [FQH16]. Compositional [OPV22]. Compound [QYZ16]. Compressed [HM20]. Compressing [Bra02]. Compressive [CZ22]. Comput. [Bak92a]. Computation [BG92, BSZ22, BC19, CDP04, CDG22, CHZ23, Cli05, DGM17, FLMN23, FF94, GKT18, Gen92, GB02, GD09, GG17, HAC^{+}12, HHC03, HMN^{+}03, ILP19, JRS11, JKD92, JP95, KTS17, KY95, LCW22, LH95, LMM19, LH02, Mar10, MMF^{+}19, MRP22, NFMS14, Pic14, RHF24, RDZ10, SG17, Smi05, Som98, Spa05, SWC^{+}10, TW96, Wan94, WNA20, YMG21, Yua08, vdBBD22]. Computational [Ano06a, Ano07a, Ano08a, Ano12b, Ano24a, Ano24b, CGEGMI22, CH12, CC22a, CRL13, CPV03, GK02, GV01, HKS12, JMY^{+}16, KGMT06, LD11, LPS99, LG92, NFM16, OLN11a, QN00, RT98, SY02, Wat11, Weg95]. Computationally [BRS^{+}08, BRS11, BRS12, BSX23, GH18, HEF17, KLH99, Par22, SS16, ZZL18]. Computations [BAMG22, FCL10, GAS20, HB98, WFS11]. Compute [LSV^{+}09, SD03, SH97]. Computer [BCE20, Cha22, Dri08, GA15, GdTv96, HIT99, IW22, JV01, Kab93, PW14, RFV23, SCG22, WL02, ZLR18]. Computer-Aided [HIT99]. Computer-Assisted [GdTv96]. Computers [Šev04]. Computing [Bak92b, BGD16, FCL10, GDS02, GK06, GI00b, GW01, IJ02, JO02, Lan00, LS05, Mül99, NP01, RTL07, Rup92, SS05, SRE06, VH22, YZ13, YZ06, Bak92a]. Concave [GT11, LLW^{+}22, SF23, Tad17, WASB22]. Concave-Convex [GT11, SF23]. Concavity [DGWR24, ES98]. Concept [LZ18, YSK^{+}10]. Concepts [MS93a]. Conclique [KKLN20]. Conclique-Based [KKLN20]. Concordance [GY22]. Concurvity [GKZ10]. Conditional [BD23, BCH05, Bra99, CB01, Che07b, CDFB11, EC17, FN16, Fri99, GKBP15, GQD18, Hes98, HBG96, IL16, JO02, KSV23, LM19, Mey18, Mil19, MS18b, MW14, RZ22, SG20, WW12, ZMH07, vlM97]. Conditional-Potts [MW14].

Conditionally [DZZ06, SMS^{+}14]. Conditioning [Gey95]. Conditions [TRWB16]. Confidence [CDP04, DJ04, EN20, FSR20, Fon19, GJ09, GW93, GB12, HO93, HMPP19, Hut02, JWH16, KLH99, KHM05, KS04, Kwo98, LT94, LH95, LJZD05, LS97, MS18b, RO94, San23, WASB22]. Confident [Ros93]. Configurations [RG09]. Confirm [Yat23]. Conflicting [WS24]. Confounded [LH15c]. Confounding [CHLZ22]. Conic [GG20]. Conjugate [KN97, LB06]. Connected [ZC15]. Connecting [Kle22, Tie00]. Connectivity [SMN15, XC22, ZSL23]. Conquer [BCD11, CSWZ23, LJN^{+}17, RAM19, XSC19]. Consecutive [PN22]. Consensus [NJM20, RJLW21, XP07]. Consider [Kos13]. Consistency [KMOS92, Tar99]. Consistent [HCG02, Kan20, OR99, SBZ13]. Constant [GN17, GD23, Lee02, WLZB07, ZF18]. Constants [SRE06, dV08]. Constrained [CDP04, CGEMIR18, CT01, DSD^{+}23, GKZ18, GM15, GBW17, Hof00, LJZD05, PA14, TRWB16, ZBS^{+}21, ZL13]. Constraint [HHGG01, Wil96]. Constraint-Based [Wil96]. Constraints [BH01, CGH99, EM00, HP99, HLW24, LL15, LWZ14, OC01, Ros17, SMN15, TLLJ23, WS24, ZNR21, WDB96]. Construct [WZ22]. Constructing [HM02]. Construction [DS94, EN20, FSR20, HMPP19, ZKS22a]. Constructions [VN18]. Contaminated [PM16, TGPM22, Wan23]. Contamination [SK10]. Contexts [TKB23]. Contingency [BL22, EC17, Fis12, Fri00, HM02, LPM98, Sta95, VT94]. Continuous [Big06, BHS00, CMN21, CBL20, Fon19, Lia07, PM20, PRCF24, Zha07]. Continuous-Time [PRCF24]. Contour [BL17, Lia07]. Contrast [Mül99]. Contrastive [LWH23]. Contrasts [HH06a].

Contribution [Gre13]. Control

[BCS96a, BGLM18, GPR99, HV19, LX16, LR04, MDS20, RNHY12]. Controls [CB97, LACV23, MR01]. Conventional [QS09]. Convergence [AE92, BF02, BDR97, BG98, BG00, BGP03, Can99, CKT21, GVDP99, GM20, LLM23, PMB12, RS01, RZ22, SC92, TH09, VS17]. Convex [AE92, BGLM18, CL15, Coz99, GT11, HWS16, HKP16, Jon98, LJ19, Mül99, Pan99b, Som98, SF23, WZSF18, WNA20, YT21b]. ConvexLAR [XWZ15]. Convolution [MPTZ23, ZW10]. Convolution-Based [ZW10]. Convolutional [GMdL23, JRH22]. Convolutions [LCSL14]. Conway [PFBSO22]. Cook [Gel17]. Coordinate [Gri23, LPJ13, Mou09, PW15, SBT00, VGUH23, YH17, ZYKC08]. Coordinates [Far12, RMWH17]. copCAR [Hug15]. Copula [AB21, BC16, Gu15, GKN23, HvHB23, LMS19, MD13, SLMN20, SLM22, TPA19, TGM^{+}14, VN18, XC22]. Copula-Type [TGM^{+}14]. Copulas [AG05, JB22, SYNK24]. Cordillera [RHM18]. Corners [LHC17]. Corrected [CB20]. Correction [Ano92a, Ano94a, Ano96a, Ano96b, Ano00a, Ano01a, Ano06a, Ano07a, Ano08a, Ano16a, Ano18a, Ano19a, Ano19b, Ano19c, Ano19d, Ano20, Ano22, Ano23a, Ano24a, Bak92a, Gel17, Len09, LR15, OLN11a, SZ08]. Corrections [Tho93]. Corrective [SS11]. Correlated [ADY10, HY17, KDM16, LP07, SBZ13, TZ12, ZBB06]. Correlation [CAR^{+}10, CL12, CKT21, Fal99, GDM14, Gra13, HM09, HWQ15, LP14, LD06, MMGG16, SPV22, TDM12, Tro05, ZBB06, ZLP23, ZRL14]. Correlations [YMG21]. Correspondence [dLM04]. CorrPlot [MMGG16]. Cost [Far92, LKR22, NFM16, RHO22, Wan13]. Cost-based [RHO22]. Cost-Sensitive [Wan13]. Costs [BRS11]. Count [BBM24, Wan20, ZZZS17]. Counting [Fis12, ZBY09]. Counts [MOPW20]. Coupled [TFS20]. CoV [CP22]. Covariance [Ano12b, AAN17, BHvdLD22, BSX23, CJ09, CHH10, CW22, Cli05, DP23, DY09, DT13, DRG19, FGN06, GD16, HLH15, HAC^{+}12, HLL07, HSG23, KSM14, LRB12, LCB19, LWZ14, MWDR21, MGF22, NT17, ONS18, PF23, RLZ10, SR12, SY16, SLBV18, Ste13, WL09, WP13, YT21b, vW19]. Covariance-Based [BSX23]. Covariance-Guided [HLH15]. Covariances [GS24, GF17]. Covariate [BCST13, BRCC22, CHZ23, DM24b, Sar22, SHFT18]. Covariate-Dependent [BRCC22]. Covariates [DSD^{+}23, DM24b, Gua23, HPQ24, JS20, MQR11, PQM22, RAM19, SS22, ZDM20]. Covariation [Ram17]. Covariogram [OM19]. Coverage [HGK10, MS18b]. COVID [THM^{+}23, WKT^{+}24]. COVID-19 [THM^{+}23, WKT^{+}24]. Cow [LYHM15]. Cox [CCL21, DBPW23, GD23, JK00, LC24, Pan99b]. Crack [GVBM04]. Create [FF94]. Creating [BY18]. Creation [Wil00]. Credible [CS99, SHL19]. Credit [BO05]. Cremona [Ano24a]. CReSS [SHMD^{+}14]. CriSP [HH01]. Criterion [CS14, CC22a, GBS24, LPPT17, RW10, ST99, SHL12]. Criticism [WFCT23, Wic13]. Cross [BEGS17, BHvdLD22, CSS^{+}09, CLXY23, DPRK23, FP20, GS24, Gv97, Gu92, GX01, KO04, KPC02, KMGJ23, LY16, LHL03, SD03, TSG92, XSC19, YK09, YF14, Zha03, vdWvNR21]. Cross-Entropy [BEGS17]. Cross-Sectional [KMGJ23]. Cross-Validated [BHvdLD22]. Cross-Validating [Gu92, GX01, Zha03]. Cross-Validation [CSS^{+}09, CLXY23, DPRK23, FP20, GS24, Gv97, GX01, KO04, KPC02, LY16, LHL03, SD03, TSG92, XSC19, YF14, vdWvNR21]. Cross-Validations [YK09]. Crossed [BD22, MDW23]. Crowd [CCHM18]. Crowd-Sourced [CCHM18]. CTMCs [BLBCC22]. Cubature [MHDW19]. Cubic [Wan08]. Cucumbers [FW11]. Cumulant [Sta95]. Cumulative [LLZ23]. Cure [ZZL18]. Curriculum [Zhe17]. Curse [LCL22]. Curve [DGW09, FM94, FDBS10, MYC^{+}22, Mül99, SBEG94, TZ12, WH98]. Curved [HH06b]. Curves [Big06, DJ04, GR12a, LX16, PHK14, QG22, SG03, Seg94, STZM09, WGSC24]. Cut [Hof19]. CVX [WZ22]. Cycles [FCL10]. Cyclic [GHCU22].

D [Dri08, LTW^{+}10, MBGO11, MKE^{+}21]. d-blink [MKE^{+}21]. DAG [LZ23]. DAGs [MC18]. Daily [WCH20a]. Damped [HV19]. Daniel [FB94]. Data [ACC18, AE92, Bak92b, BC00, BY18, BS19, Bau19, BCD^{+}24, BML19, BM19, BL04, BK03a, BK23, BB06, BH01, BFE^{+}03, Bri94, BW17, BCS96b, Buj04, BSL^{+}08, CRGS17, CR99, CC13, COvS19, CBS24, CCH16, CH99, CCHM18, CMC^{+}22, CBM24, CVQ23, CB97, CBL20, CC22b, CSK10, DG18, DDM20, DSD^{+}23, De 05, Deg23, Dem23, DYB^{+}23, DLSS24, DGS24, DH03a, DG24, Don17, DH03b, DM24a, DCS16, EH17, ES96, EJPS22, EC14, EFLT94, FK23, FLY21, Far92, Far99, FS99, FCWM17, FB11, FN16, FPRW09, Fri99, FDG^{+}22, GEGM03, GDM14, GD16, Gel04a, Gel04b, GV01, GRG10, Goo99, GY07, GIA22, GW15, Gu92, GX01, Gu15, GG20, GF17, GLMZ15, HQT92, HLP07, HIS22, HJ04, HHC03, HYL^{+}17, HvHB23, Hic17, Hig01, HY17, Hob01a, Hof09, HV17, HWK17, HJ17, HWQ15, HW21]. Data [HCCZ23, HCG02, HLL07, HCZMH21, Hub00, HJT01, Hug15, Huh95, HDK^{+}20, Hur93, Hur00, Hur04, HJNR19, HS10, IG96, ILP19, JBJ^{+}22, JB22, JS20, Jör09, JK21, KDM16, KB99, KS22, KMS23, KT03, KSSL21, Kei96, KH23, KTZQ20, Kle08, KMGJ23, KLH05, KS92, KMQ05, KB20, KKB21, KKHB23, Kus06, LCV20, LCL22, LHL21, Laz05, LRB12, LAJ13, LYHM15, LG03, Lev01, LPS99, LRBS03, LLW^{+}22, LO16, LROI23, LG92, LP14, LPQ^{+}22, LLM23, Liu01, LP07, LS05, LSFI18, LH00, LQ21, MK20, MHH15, Maa05, MR10, MM10, MC17, MM14, MCB24, MCF^{+}20, MMO01, MN97, MM09, MPH22, MHB17, Mci99, MPS92, MWM^{+}23, MPS21, MK92, MG01, MGL18, Mou09, MHS11, MW04, MC05, NM98, OFR^{+}18, OPV22, OS22, PKH15, Pan99b, PRS13, Par11, PJS^{+}22, PP07, PP09, Pen17, PMI20, PC22]. Data [PAP22, QGL09, QTVK18, Rho03, ROY20, RRH18, RDL07, SH18, SBR22, SS12, SJH^{+}15, SK10, SR00, Seg94, SG00, SH03, SS01, SSB^{+}14, SR23, SSBW17, SDBD21, STIP23, SDL19, SC98, Ste92, SK24b, SRL^{+}00, SS22, SCB98, TYG^{+}20, TvdV22, THSM^{+}20, THSM21, TW14, TZ12, TQ16, TKB23, TRS97, Tre06, TU11, Tu94, TG10, TS13, UHHS96, UTK22, VF09, VDC21, WW12, WLZB07, WH17, Wan20, WCH20b, Weg95, WS02a, WM03, WBS21, WCDS03, Wic11b, WZ01, Wu11, WL13b, WW19, WY22, XP07, XCSP16, XW20, XL17, YL22, YBDS02, YMB14, Yat23, YL95, YL99, YLM17, YWBA22, YSWS24, YG22, ZSK19, ZZW21, ZG22, ZH22, Zha97, ZJ04, ZCI^{+}17, ZZZS17, ZSL^{+}22, ZL22, ZW04, Zhe17, Zho05, ZP14, ZZL18, ZKS22b, ZZT22, ZW10, ZICT12, ZWS23, ZCY^{+}11, vDE19, vDM01a, vDM01b, Bak92a, Kel11, Ano24a]. Data-Adaptive [LO16, vDE19]. Data-Dependent [WLZB07]. Data-Driven [NM98]. Data-Generating [Yat23]. Database [Lum19]. Databases [KLSR05]. Datasets [Ano12b, Bra02, BKM99, CY24, CW00, DBH^{+}22, FS99, FRW05, FGN06, GMF16, Goo99, GB04, HM16, HS18, Hub99, HB16, IW10, KP03, KSM14, LCRG21, NBH^{+}15, Par18, Pos01, SS12, SS16, TNKK16, XW18]. Day [YL99]. Dealing [GKZ10]. Deblurring [Kan20]. Decentralized [QC24]. Decision [BK03b, DZFZ18]. Decisions [BLR22]. Decoding [CZ22, PAP22]. Decomposable [PAP22, SYNK24, TG09]. Decomposition [CCX22, HLW22, HJNR19, LXZ22, MPS17, Smi01, YMB14, YG22, ZG22, ZMSZ07, ZZSW23]. Decompositions [ALP^{+}23, GKT18, GB04, HAC^{+}12, HO10, MPS21]. Deconvolution [LS04, OC01, SMS^{+}14, Sar22, YASR20]. DecoupleNets [HPZ22]. Deep [CGLW21, FX21, HL23, KNS21, MCB24, SCG22, SLL22b, TMC23, TNNK20]. Deeply [LROI23]. DeepMoM [HL23]. Default [GD09, Rob19]. Defect [GM93]. Defined [FPRW09, TA12]. Definite [AHNC13, COvS19]. Definition [BF02]. Deformations [SJH^{+}15]. Degree [ABG^{+}21]. Delay [DPS11]. Delayed [ER21, HCK^{+}11, QTVK18, SGH17]. Delayed-Acceptance [SGH17]. Delays [Wic11c]. Deletion [OSR^{+}18]. Deletion/Substitution/Addition [OSR^{+}18]. Delft [GdTv96]. Delineation [DCT08]. Delta [FR97]. Demand [WH14]. Demography [Bri94]. Dendograms [MN97]. Dendrogram [EH15]. Denoising [Gao98, MQ15, SBT00, SM19]. Dense [GM15]. Densely [LQ21, ZC15]. Densities [BMO01, BRS11, BRS12, Bor11, DW01, ES98, HBG96, LS05, SPW95, WL02, vDE19]. Density [ACG14, BJMD13, CGH99, CRL13, GW18, GMTK13, GMC02, GG17, Gu15, GQD18, HO93, HP99, HMW06, Ho06, HB16, IL16, KH22, Kle04, Kle06, KS08, KS92, Koo98, KS04, LW19, Len03, Lia07, LLZ09, MS93b, MV98, MSS08, NTVK12, Ooi02, RW10, SMS^{+}14, Sar22, SN10, Tok07, TGM^{+}14, WASB22, WCKL19, YASR20, ZMY^{+}23]. Density-Based [ACG14, HB16]. Dependence [BD23, CMG21, CM13, HCE23, HM14, HPZ22, MPR^{+}12, MS93a, MC18, OM19, QM04, VHSG20, ZY24, ZH17]. Dependencies [Moc98]. Dependency [LYM23]. Dependent [BRCC22, GRG10, Hob08, Hoo07, Kle22, LG03, LLM23, PQD22, RZ22, SG20, WLZB07]. Depicting [HM11]. Depth [COvS19, FLMN23, NGH17, ND21]. Depth-Based [NGH17]. Depths [EJPS22]. Derivative [BRS12, Ste12, YSWS24]. Derivative-Free [BRS12]. Derivatives [CR08, SQ07]. Deriving [SJSG22]. Dermoscopic [LZYQ22]. Descent [Gri23, Hel24, PW15, QWW23, YH17, Zha07]. Describing [DG14]. Description [TL92]. Design [BCS96a, BB03, DMP14, ES96, GM18, GP11, Gra07, HB92, HLZH05, HS08, IW22, LPQ^{+}22, MPH22, MXM^{+}21, SHL12, SWX23]. Design-based [MXM^{+}21]. Designing [HK17]. Designs [CLHW19, DGM17, HK06, KSR19, LACRM06, MJ18, MHDW19, Rya03, TN14b, WYZ19, WZ22]. Destination [LA12]. Detect [CAR^{+}10, HCZMH21, LN09, Qiu02]. Detected [WJZ09]. Detecting [Ano23b, FML19, FSMV02, MLP07, MM09, Sod20]. Detection [BH02, CPP23, CR99, CC18, DG18, DM22, DM24a, FEF22, GM93, GG20, HL06, HEF17, HCCZ23, JV01, KH21, KH22, LCSL14, LKB21, OBGC21, PP07, PF23, QY97, SBM22, SSL^{+}16, SMN15, SQ07, THSM^{+}20, THSM21, TEFH20, WWG^{+}97, ZB10, ZZSW23, ZZD23]. Detections [MT03]. Detectors [QS09]. Determinant [Hut02]. Determination [HS08]. Determining [HM11]. Deterministic [BCE20, BB21, GR99, HRV12, PW14, PWPS97, RW14, Rou08, WJG11]. Developing [MC17, MDD^{+}12]. Development [IKL08, ZJRI10]. Developmental [RHC^{+}14]. Developments [Tie96]. Deviance [BY04]. Diagnosing [BDR97, Jan06, WW12, WJZ09]. Diagnosis [HIT99, LL11]. Diagnostic [GVDP99, LWM21, PV04, SS11, dSM04].

Diagnostics [BCST13, CK08, FB11, Hoo07, LCTV24, LHC17, NY14, RRSY19, WKM22]. Diagonal [ROY20]. Diagonally [KXY20]. Diagonalwise [CHPR24]. Differences [YHH^{+}23]. Different [GU13a, GY07, PC22, SYL^{+}22]. Differential [CHW12, CGH17, LZ23, Sin02, WGDW22, XWS24]. Differentially [BRMC19]. Differentiation [Ska02, Smi95]. Difficulties [Lee02]. Difficulty [CCHM18]. Diffuse [NM98]. Diffusion [BB21, RIC^{+}16, SY07, YB22]. Diffusions [Gol09, PRS13, WRT21]. Dimension [CL09, DYB^{+}23, HPQ24, Hen04, HV05, HY17, KH21, LMM19, LZCZ22, LWH23, PQM22, SY16, Vel08, WY15, XZYZ24, XY14, YC04, ZYKC08, ZYZ10]. Dimensional [ACJ21, AR22, AG18, BB06, BCS96b, CLLZ17, CB97, DDM20, DM24a, EO95, FM12, FSMV02, GKT18, GF16, HLP07, HM09, HM11, HCE23, Hoo07, HO10, IW10, IL16, ILP19, JVSS24, JT21, KDM16, KB20, Kus06, LCV20, LCL22, LB15, LRFK21, LLW^{+}22, LDR23, LACRM06, LP14, LMW^{+}22, MM14, MMZ22, MBS04, MS18b, NY14, NYC15, NFMS14, OFR^{+}18, PH21, PP07, PW15, Pos95, RRSY19, RSKZ19, ROY20, RSDG23, RFV23, SMB14, SBE17, SYW^{+}22, SJM^{+}24, Shi24, SWZ14, SLMN20, SHFT18, SLH24, SRL^{+}00, SS22, THSM21, TSH16, TRC23, WLWM16, WL23, WCM21, WZXW24, XGTB18, XL17, YMB14, YKW03, YF14, YWL^{+}15, YWBA22, YF22, ZSK23, ZP14, ZDM22, ZCY^{+}11, FB94, HHH22, vdWvNR21]. Dimensionality [FX21, LCL22, SBJ17, YPPM21]. Dimensions [BHvdLD22, HM14, LCBV22, LACV23, Oeh98, PMI20, SLR16, SYL^{+}22]. Diracs [CKPT04]. Direct [CS18, WLZD11, XMY13]. Directed [ABG^{+}21, KB08, ZHZW22]. Direction [CS98, LB06, WLWM16]. Direction-Projection-Permutation [WLWM16]. Directional [ES19, FK23, FGV22, KGM18, RRH18, SJH^{+}15]. Directions [PP07, Tie96, YZ07, MR10]. Directly [BC24]. Dirichlet [CD19, DH03b, GK02, GHW02, IJ02, JN04, LKR22, MM98, MMLLQ^{+}22, MR16, Nea00, WD11, XD20, ZNYJ14]. Disagreement [YF22]. Discontinuous [Koo97, SW23]. Discoverability [BML19]. Discovering [RFV23, WL09]. Discovery [Ano24a, BGLM18, CC22b, HCZMH21, HDK^{+}20, LWM21, LLM23]. Discrete [BCDL17, CWZ21, Gué03, Kos22, KK00, LMS19, MVR14, NS94, RF17b, RO94, Yan21, Yan24, ZJRI10, ZL10]. Discrete-Margined [LMS19]. Discretely [SG23, SY07]. Discriminant [Fos01, HV05, KL03, LW08, LLW^{+}22, QLCZ15, RA01, ROY20, YWBA22, ZH03]. Discriminating [DGM17]. Discrimination [ACJ21, GEH^{+}22, HLD^{+}13, LMST18, LAJ13, Loc22, WZ16, Yat23]. Discussion [BC93, Buj04, Car94, DL95, Eld94, Gel00, GH00, Hig01, Hob01a, HJ17, HJT01, Kie00, Lev01, Liu01, Men00, MG01, Oue94, Pre95, Wu00, WZ01, dLM00, FB94]. Disease [DCT08, GF22, GRM09]. Disparities [DASR08]. Dispersed [AC95]. Dispersion [LACRM06, Rup92]. Display [BC00, BCS96a, Cle93a, Cle93b, EO95, Fri02, Hol93, HO10, Tre93, Wil93]. Displaying [Eic94, GMF16]. Displays [CCH16, Fri99, HH06a]. Dissecting [FCL10]. Dissimilarities [OR07]. Dissimilarity [ACC18, HH11]. Distance [CS14, CCHM18, FHTG05, GEH^{+}22, HLD^{+}13, LMST18, LYXM23, Loc22, SPV22, SY16, SdWDL23, WZ16, WL22, YSWS24, dR07]. Distance-based [YSWS24]. Distance-Preserving [WL22]. Distance-Weighted [HLD^{+}13]. Distances [GR93, HR05]. Distilling [PV23]. Distinct [BR02]. Distortion [Jör09]. Distributable [KYW19]. Distributed [AM07, CZ22, CSWZ23, FLY21, HLW24, KLSR05, LHL21, MKE^{+}21, SX23, SDL19, SX21, TRC23, WZXW24, XSC19, ZKS22b, ZLW21]. Distribution [ABG^{+}21, AE92, AR06, BC24, BHF02, Bre04, CS20, DM24a, EB02, FGH^{+}24, FIK19, GW01, HR05, HK22, Hof09, HZ21, Huh95, KLH05, LLZ23, Maa05, MMO01, MOPW20, Mey18, PP05, PFBSO22, PLW01, RF17b, TMMB20, WWG^{+}97, Yan21, ZKS22a, dSM04]. Distribution-Free [DM24a]. Distributional [HMPP19, KNS21, MS18a, MKK18]. Distributions [CHG16, EGO92, EM00, GDM14, GB00, GD09, GM18, GR08, Gou98, GIA22, GHCU22, HIS23, HC11, Hes98, HPZ21, KR17, KGM18, KW10, LS93, LMW^{+}22, MS93a, MMDL22, MT00, PI10, PH20, RNM18, SAT04, SBF19, SBS23b, TGPM22, WW08, WIR01, YLLZ23, ZSW22, ZL10]. Divergence [RF17b]. Diverse [WCDS03]. Divide [BCD11, CSWZ23, LJN^{+}17, RAM19, XW18, XSC19]. Divide-and-Conquer [LJN^{+}17, RAM19, XSC19]. Divide-and-Conquer-based [CSWZ23]. DNA [Hob08]. Do [KB99]. DOBIN [KH21]. Documents [FHTG05]. Domain [JCZG19, KNH23, McE22, MPS21, XTPO22, ZJRI10]. Domain-Based [ZJRI10]. Dome [GVBM04]. Dominance [ACIP22]. Dominant [GK97, KXY20]. Dose [PPTO22]. Dot [KTT23]. Dotplot [CH93]. Dots [Kle22]. Double [Yan24, YG22, ZBY09]. Double-Matched [YG22]. Double-Saddlepoint [ZBY09]. Doubling [MT00]. Doubly [MMDL22, PH20, PFBSO22, SBF19, YT21a]. Drift [YSK^{+}10]. Driven [NM98, Paa99]. Dropout [ZDM20]. Dual [AT16, OPT00]. Dually [HK10]. Durations [KMGJ23]. Dylan [GI00a]. Dynamic [BRB17, CKHG15, CPV03, CD14, Dob12, EO95, FIK19, FSMV02, HMW06, HABV06, Joh13, LCW22, LC24, LCD^{+}24, MA19, MCB24, PH23, SG05, SJM^{+}24, SNR17, SCB98, VH22, WNA20, WKT^{+}24, XC22, ZLR18, ZZSW23, ZZD23, ZY21]. Dynamical [Dri08, SSBW17, Tre06]. Dynamically [BS19, Kle19, LA12]. Dynamics [LSSG15].

Early [Wat11]. Earth [KB99]. Earthquake [CH99]. Easily [KYW19]. ECM [vlM97]. ECM-Type [vlM97]. Ecosystems [GRVE18]. Edge [BBM24, QS09, Sew21]. Edgeworth [Kab93]. EDI [Yat23]. EDI-Graphic [Yat23]. Editor [Loc11]. Editorial [Ano92f, Ano93e, Ano94f, Ano95e, Ano96g, Ano97e, Ano98e, Ano99e, Ano00f, Ano03a, Ano06b, Ano07b, Ano08b, Ano09a, Ano10a, Ano11a, Ano13a, Ano13b, Ano14a, Ano14b, Ano15b, Ano16c, Ano17b, Ano21, Edd92, Lee14, LTW^{+}10, Lev11, Lev12, Lev13a, Lev13b, Sco03, vD08, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano00e, Ano12a, Ano15a, Ano16b, Ano17a, Ano18b]. Editors [Ano24b]. Education [GdTv96]. Effect [CC13, CTDKL21, GMF16, GM05, HM09, Laz05, LYLL23, LSFI18]. Effective [BLR22, CW00, HABV06, KCF21, ZYKC08]. Effects [BCST13, BT18, BD22, CBM22, CZ02, DH10, ERPG11, EC14, Guh08, HC11, HYL^{+}17, HDK^{+}20, IPH22, JWH16, KW01, KK09, LPJ13, LHC17, LHL03, MDW23, MKK18, NLF^{+}06, PB95, PLW01, RYY00, SJSG22, SY02, SBS^{+}23a, Ska02, SX21, TH09, TG13, VL09, ZBS^{+}21, vD00]. Efficiency [JB03, YM11b, HR11]. Efficient [AB17, AT16, BG92, Béd08, BRS11, BSX23, BK23, CD12, CC18, CVQ23, CDFB11, DC22, DGM17, DBS06, FQH16, FDC^{+}19, FCWM17, FSR20, FWR12, GMdL23, GD09, GH18, HHL19, HMJ11, HEF17, HYL^{+}17, Hes95, HS08, KTS17, KFSL17, KCF21, KLH99, KC20, KSM22, LB15, LYXM23, LGL16, Owe17, Par22, PN22, PNLL21, PLW01, PC06, PSF^{+}22, Rou08, SM19, SBR22, SK21, SLBV18, SHBC^{+}24, SNR17, Smy02, SS16, SK24b, TDM12, WB18a, WFCT23, WLZ^{+}24, YZ13, YZ06, Yua08, ZR21, ZMY^{+}23, ZZL18, ZZT22, ZHZW22, vD00, vdBBD22]. Egocentrically [LWLZ22]. Eigen [JLAW22]. Eigen-Adjusted [JLAW22]. Eigenspaces [HLW24]. Eigenvalue [EM00, GBW17, JAJ19, LWZ14]. Eigenvalues [Fro22]. Eigenvectors [Fro22]. Elastic [Cul13, Hel21, QYZ16, YH17, vNvdBvdW22]. Elastic-Net [YH17]. Elder [FB94]. Electrophoresis [CKPT04]. Element [Mey18]. Elemental [SH00]. Elimination [Ste95]. Ellipsoid [WR93]. Ellipsoids [EM07]. Elliptical [LFZ97]. EM-Like [BCH09, HV19]. EM-Type [vD00]. Emacs [RHS^{+}04]. Embedded [GW15, LDR23]. Embedding [CP18, GF17, HPA14, LPPT17, ZLPT22]. Embeddings [HKP16, San23]. EMICM [AB17]. Emission [JP22]. Empirical [BBW20, CVA08, CL12, Fos01, HO93, Huh95, JS95, Laz05, MS93a, PP05, SL99, SS05, Tan15, Yan21, Zho05, WDB96]. EMS [FSB11]. Emulation [IW22, PH20]. Emulators [Rou08]. Enabling [CCH16]. Encouraging [ZWZW24]. End [MKE^{+}21, PRS11, SCLK^{+}00]. End-to-End [MKE^{+}21]. Endogeneity [CGJ09]. Energy [PSRM^{+}23]. Enforcing [Hea23]. Engine [Paa99]. Enhanced [SSP^{+}18]. Enhancement [SSL^{+}16]. Enhancing [BML19]. Enough [BBKC24]. Ensemble [DZL22, DCS16, LZ18, TSH16, UVM18]. Ensembles [DPRK23, Eld03, RMR^{+}17, XZ12].

Ensembling [ACC18]. Entity [MKE^{+}21]. Entries [CBL20]. Entropy [BEGS17, FGH^{+}24, LL15]. Enumerating [TG09]. Envelope [CZ16, KN97, SLC22]. Environment [FF94, KSC^{+}24, Lan00, ZGL22]. EOV [Ano12a, Ano13b, Ano14b, Ano15a, Ano16b, Ano17a, Ano18b]. Epidemic [FCWM17, ZY21]. Epidemiological [FDG^{+}22]. Equal [Par22]. Equality [LMW^{+}22, TLLJ23]. Equation [KC20, LG03, Par22, SW18]. Equations [ACS17, BMW17, CHW12, CGH17, CL12, CTW20, ER04, LFS^{+}17, Sin02, SS16, WGDW22, XWS24]. Equivalence [MOPW20]. Era [RHV22]. ERGMs [HHH12]. Ergodic [DH10]. Erratum [Ano12b]. Erroneous [DR18]. Error [Che23, HS08, JH98, KLH05, LR05, Par22, RFVJ21, SR23, SM13, SHFT18]. Error-Prone [Che23]. Errorfully [ASPL21, PSTV15]. Errors [BBvdW16, CHZ23, CRC^{+}07, PP96, SMS^{+}14, TEG19, YASR20, ZHJ16]. ESCV [LY16]. Estimate [GH21, PLIK02, Ram17, WGSC24]. Estimated [RW99]. Estimates [Fri00, KH22, KO04, Kle04, Kle06, LPM98, Mar10, SBY06, WB18b, YZ06, Yua08]. Estimating [ACS17, AG18, BMW17, CL12, CTW20, ER04, EGO92, FP20, GLC20, Gué03, HL03, HWQ15, HJR03, HBG96, IP04, Kou14, Lee02, LKW09, LFS^{+}17, LSFI18, MM98, Mai97, Moc98, PMS21, Rya03, Sha14, SW18, SN10, SS16, WCKL19, XTPO22, YP20, YZ07, ZLP23, dSM04]. Estimation [AG05, ABG^{+}21, AE92, ASPL21, AAN17, AG18, BEGS17, BCH09, BOW02, BBLG13, Bre04, BR02, CHW02, CHW12, CBM22, CCHS08, CSY00, CKNK06, CJ09, CHH10, CP16, CW22, CS99, CH16, Che23, CGH99, CR08, CG07, CGJ09, CRL13, CZ16, CDFB11, DJA23, DY09, DT13, DZ99, DH10, DBPW23, DGW09, DIR02, Efr05, ESR^{+}14, Eve12, Fos01, FKLW08, FP04, Fri13a, FN04, GM15, GD16, GW18, GN17, GK10, GMTK13, GMC02, Gou98, GPR99, Gu15, GMP04, GRM09, GG20, HO93, HP99, HHGG01, HMW06, HSM09, HC11, HHH22, HK22, Her97, Ho06, Hsi99, HLL07, IL16, JRS11, JW08, Jon98, JAJ19, Kar21, KCSA20, KBM23, KS08, Koo97, KS92, Koo98, KS04, KKS17, KW10, LW19, Len03, Lia07, LPHK16, LRR20, LY16, Lin99, LTW07, LLZ09, LWZ14, LP07, LMS19, LVH17, MP00]. Estimation [MWFSG17, MVR14, MSBV22, MÉG^{+}23, MBY^{+}22, MMZ22, MBGO11, MRCM18, MV98, MSS08, MGF22, NT17, NFM16, NTVK12, OY23, OPV22, OS22, PA14, PP07, PP09, PLW01, PC22, PWPS97, PSRM^{+}23, Pro06, PF23, RLZ10, SLR21, SBEG94, SR12, SQCK20, SL99, SR23, SCAP22, SNR17, Sin02, Ska02, SS06, SH00, SPW95, SC23, SBR10, SSL17, TCCSB10, Tok07, VM04, WB18a, WLZB07, WP13, WZW19, WDMW22, WH98, WYL14, WY15, XBW15, XWS24, YNH18, Yen20, YSDJ15, ZHZ12, ZMY^{+}23, ZRL14, ZL16, ZY21, ZKS22a, ZL13, ZZL18, ZW07, ZW10, dLG01, dV08, vDE19, vNvdBvdW22, vWB22, SWL23]. Estimations [Kwo98]. Estimator [BHvdLD22, DXL12, KLH99, Len09, LCB19, LCWW17, OR99, QWW23, QTV^{+}21, SBWZ08, SK21, Tad17, ZJ04, vW19]. Estimators [FM94, GG17, HO99, JP95, KSLS22, RDZ10, Rup92, TGM^{+}14, Wan94, WJZ09, YT99, ZZWS23]. Eulerian [HO10]. Evaluating [CH12]. Evaluation [BB21, CS18, DKH06, KSC^{+}24]. Evaluations [Zha22]. Event [CMC^{+}22, Dem23, FCL10, Kle22, MDS20]. Event-Chain [MDS20]. Events [LG20]. Evidence [Fri13a, KSR19, YHH^{+}23]. Evolution [Hob08]. Evolutionary [Glu04, GLW07, Mey03, SS19]. Exact [BCH05, CB01, CR24, CP18, FLMN23, GŠP^{+}06, GD23, HWLB24, HVRA96, HGK10, Hut02, KN06, LH95, LH02, MAJ^{+}98, MPS92, PP14, QTV^{+}21, Res22, RSDG23, SG23, SLR16, Sta95, Ste02, TA12, VT94, VDC21, WRT21]. Exactly [KLZ23]. Examination [MK92]. Examples [Bri94, RR09]. Excesses [SdWDL23]. Exciting [CH16, PH23]. Exhaustive [AZ22, YK09]. Exists [PLIK02]. Expanded [XCSP16]. Expansion [JHD21, LD06, Sin02]. Expansions [BHF02, CBC93, EB02, GJR05, Kab93]. Expectation [Fos01, GKBP15, Hob08, LM19, SDL19, YSD13]. Expectation-Maximization [YSD13]. Expectations [JO02]. Expected [FN16, Rya03]. Expectile [MP16]. Expecting [KB99]. Expensive [BRS^{+}08, BRS11, BRS12, SGH17, WBS22]. Experiences [FDG^{+}22, MC17]. Experimental [CLHW19, JS20, Rya03, WY22]. Experiments [BGL18, GM18, GA15, HLZH05, KSC^{+}24, PW14, SCG22, ZLR18]. Expert [Rap01]. Explanation [MMA97]. Explicit [MWDR21]. Exploit [GS23]. Exploiting [YF22]. Exploration [BJMD13, HH11, HL06, MHS11, WCH20b]. Exploratory [BK03a, Buj04, CRGS17, DDM20, GEGM03, Gel04a, Gel04b, LCKL05, LH00, NAB16, PLH10, Pos95, SMK15, SC98, TU11]. Exploring [BB06, CH93, Far99, Huh95, HOD22, Kei96, LACV23, LX16, MMGG16, OM19, Wic11a]. Expo [Wic11b]. Exponential [BMO01, BH22, BFM18, CL09, Gou98, GPR99, HH06b, STM15, Sue12, TF20a]. Exponential-Family [BH22]. Expression [CAR^{+}10, Jör09, XP07]. Extendable [BY18]. Extended [MMDL22, Sin02, SBS23b]. Extending [CM08, Fri99, KMGJ23, Pan99b, SVC^{+}19, YZ07]. Extensible [Mar96, Saw96].

Extension

[HK10, RNM18, SS12, XWZ15, VT94]. Extensions [ZYKC08]. External [ZH22]. Extract [Bau19]. Extract-Transform-Load [Bau19]. Extracting [FW11, Rho03]. Extraction [SYW^{+}22, TG10, ZH03]. Extrapolated [DPRK23]. Extrapolation [Kan20, WHH23]. Extremal [HSM09, RT21]. Extreme [CMG21]. Extremes [HM02, Mah17, RS19, VHSG20]. Eye [RHV22].

F [Ano24a, FB94]. Face [FHVW03]. Faced [Sha14]. Facilitate [Ska02]. Factor [AR22, CDP04, CD14, DDM20, EH17, Fri13a, GD09, HAPRD24, Hsu92, KFSL17, LCW22, MM09, MDC17, NYC15, ONS18, Paa99, TPA19, TGHL14, TU11]. Factored [Tho93]. Factorization [LSLX22, SNC23, WBQ20, WZZ^{+}24]. Factors [DP23, DH10, GF22, KSR19, SS05, WR05, WYZ19, ZJRI10]. Fail [JP96]. Failure [BCH22, CSWZ23, DLSS24, KLH05, ZJ04]. False [BGLM18, HCZMH21]. Families [Gou98, HH06a]. Family [AG05, BH22, CMC^{+}22, CL09, HH06b]. Far [CSS^{+}09]. Fast [AHS22, BGD16, BL04, CHW08, CGR17, CP18, CLCX22, CLSC22, DBPW23, Dri08, FM94, Fon19, GS24, GK10, GŠP^{+}06, GG17, GX20, HAPRD24, HM15, HCG02, HMZ22, HSZ23, HB16, JT21, Kar21, KN06, KPC02, KCSA20, KCK08, LMST18, LW19, LH95, LMM19, Lum19, LPPT17, MR10, MDS20, Mil19, MMZ22, MBGO11, OLN11a, OLN11b, PDM23, PRHP14, QC24, RNHY12, RT16, RDZ10, SBM22, SBY06, SBWZ08, Sam15, SLR21, SBEG94, SSMN16, Ste02, SBS23b, TKB23, TLLJ23, TW96, VL09, Wan94, WD11, WH98, WNA20, YMG21, ZNYJ14, ZSW22, ZZD23, vNvdBvdW22, vdWvNR21, HCB04]. Fast- [SBWZ08]. Faster [Oeh98, Tok07, WFS11].

Fatigue [Rya03]. Fatigue-Limit [Rya03]. FDR [LX16, SK10]. Feasibility [Weg95]. Feasible [HKP16]. Feature [All13, BMO01, DJA23, Deg23, FMP24, GW04, KCF21, LYM23, RHO22, ROY20, RSDG23, San23, SC08, SYW^{+}22, SBS^{+}23a, TVZ^{+}22, TG10, VH17, XL17, ZH03]. Feature-Inclusion [SC08]. Features [BL22, EH05, FHTG05, GPPVW17, HLW22, LX16, MLP07, MS93b, WLZ^{+}24, WZXW24, ZCS^{+}19]. Feed [KS99]. Feed-Forward [KS99]. Fetching [Bro06]. FFT [HCB04, GG17, WK17]. FFT-Based [GG17]. Fiducial [LHL21, WL23]. Field [HHC03, Mah17, PH21, YHS23, ZF18]. Fields [BMR15, BK20, CK08, De 05, DIR02, Eve12, FPRW09, Fri13a, HPA14, HKP16, HSG23, KKLN20, SS06, Ste12, TA12, YS10]. File [DR18]. Filling [CLHW19]. Filter [DZL22, Kit96, MLJ16, Sin02]. Filtered [MW97]. Filtering [CSK10, FM07, Gol09, GQD18, HZC23, PRHP14, RT16, WJG11]. Filters [Ano07a, BB21, Fea02, Fea05, JK21, KSV23]. Finance [NYC15]. Financial [Ano07a, Fea05, MC05]. Find [BL04, DS21, MHDW19]. Finding [CGEMIR18, GPPVW17, HQT92, LCV20, LCBV22]. Finite [Che07a, GIA22, IJ02, MM10, RNM18, SG23, SRE06, VH22, YKZ18, ZK22]. Finite-Sample [ZK22]. First [MM20, SQ07]. First- [MM20]. First-Order [SQ07]. Fisher [Hof09, Spa05, VT94]. Fisz [FN04]. Fit [AZ22, CZ02, DG14, GPGMFB14, GvDHB08, HM14, KN06, LRO18, NLF^{+}06]. Fitting [BK23, CKHG15, DG24, FG95, FDBS10, FCWM17, GKZ18, HHH12, JL13, KS99, MMLLQ^{+}22, MT00, RHV22, WJL23, XW18, YL99, ZNYJ14, vD00]. Fixed [BT18, CSK10, JWH16, LPJ13, QWW23]. Fixed-Effects [JWH16]. Flamelets [PB22]. Flat [HK10]. Flexible [GMTK13, GKN23, HCB04, Hug15, RT16, SLR21, TGM^{+}14, WW19]. Flight [Wic11a]. Flights [DPS11]. Flipped [Huh95]. Flooding [SLR^{+}23]. Florida [MK92]. Flows [HK06]. Fluvial [ZH17]. Flying [HCK^{+}11]. fMRI [PAP22]. Focused [Fal99]. Focusing [MC17]. Fold [GS24, YK09]. Folded [KR17, LLW^{+}22, MMDL22]. Folding [XY14]. Follow [CDFB11]. Follow-up [CDFB11]. Following [ZL16]. Forecast [AHS22, Dan97]. Forecasting [JVSS24, SH17, THM^{+}23]. Forecasts [HMPP19, TSH16]. Forest [BSZ22, GH21, LSFI18, MMW98, SH06, WSM^{+}20, dSCL21]. Forests [CBM22, CJK22, DZFZ18, FTAW21, GH21, HZ21, JWL^{+}23, LZCZ22, MH17, XNN16, XGTB18]. Form [SY07]. Formal [MH17]. Forms [Ano06a, GGT22, KGMT06, SSG23, ZSW22, Ano06a]. Formula [LS97]. Formulations [ZBS^{+}21]. Forward [FX21, GTZH16, Hui22, KS99, MM09, MDS20, PRHP14]. Forward-Backward [PRHP14]. Four [Hub99]. Fourier [HCB04]. Fourth [YC04]. Fractal [HMP96]. Fractional [BMR15, BSX23, Ste02]. Frailties [JK00, KK09, LRBS03]. Frailty [HL03, HPOL14, TGP03]. Framed [VRH19]. Framework [BML19, CC22a, DH03a, HHKS11, IKL08, KO05, LWM21, MQ15, SKK22, VGUH23, WM19, WGSC24, ZSL^{+}22]. Free [ACS17, BRS12, DDM20, DM24a, DNF23, EGCP24, LD21, LYM23, Lin99, MS03, PDM23, PV23, PSRM^{+}23, Wan08, XL17]. Free-Knot [Lin99, MS03, Wan08]. Frequency [Dem23, HTW22, KTS17, McE22, MC05, QGL09, SH00]. Frequentist [HMPP19]. Friendly [CC22a]. Front [SCLK^{+}00]. Frontiers [Weg95]. Full [GK02, HK06, KK09, Mil19]. Fully [ZDM22]. Function [BCS22, BM19, BRS^{+}08, CBC93, GD23, HB92, HLL07, Huh95, Hut02, JKD92, Kab93, KSM^{+}18, KB20, LPHK16, LO16, LLZ23, LQ19, LQ23, Maa05, MP00, PH20, PP05, PB95, RMWH17, SLC22, WLZB07, WDMW22, XP07, XY14, ZJ04, ZA23, ZHZW22]. Function-on-Function [BCS22, KSM^{+}18, LQ19, LQ23, SLC22, WDMW22]. Function-on-Scalars [KB20]. Functional [Ano24a, ABvS09, BM19, CHZ23, CVQ23, CC22b, CCL21, CLCX22, CLSC22, DG18, EH17, EFG^{+}99, EJPS22, GPGMFB14, GBC^{+}11, GLC20, HLP07, HC98, Hoo07, HMZ22, HS10, JLAW22, JVSS24, JCZG19, KMS23, KSH23, LYHM15, LC24, LSH16, LRR20, LCWW17, LLZ09, LMC22, LQ21, LQ23, Mai97, MSJ^{+}24, MHG24, MHS^{+}14, OY23, PJS^{+}22, PP09, PC22, QGL09, QG22, Ram17, RR23, SKK22, SSG15, SH17, SSB^{+}14, SYW^{+}22, SJM^{+}24, SSBW17, SSG23, SG11, SK24b, TMC23, TG10, WC23, WCR13, WL13b, XC22, YL99, YLM17, YWBA22, YSWS24, ZSK19, ZSK23, ZZF24, ZOR12, ZCO15, ZP14, ZCY^{+}11]. Functionals [GHW02, Vie00]. Functions [CKNK06, CHH10, Coz99, DJM22, Dia99, FS18, GK97, HIS22, Hoo07, Koo97, LHY00, LKW09, LTW07, LJ19, LHK17, MKN23, NGH17, PNLL21, PL22, Rou08, RW99, Sha14, SBR10, Tho93, Yan21, Gel00, GH00, HL00b, Kie00, Men00, Wu00, dLM00]. Fused [BRB17, CJW22, DP23, GN17, GJL^{+}10, Hoe10, JYYZ18, Joh13, LWLY17, MK20, PWS16, SLR21, Yen20, YWL^{+}15]. Fused-Lasso [DP23]. Fused-MCP [JYYZ18]. Fusion [CTDKL21, CGR17, PGR15, PMS21]. Future [HWLB24, Tie96].

G [LHL21]. Gains [Rya03]. Gambler [CJS92]. Game [CR24]. Game-Theoretic [CR24]. GAMlet [ST04]. GAMLSS [Ano23b, Sod20]. GAMLSS-Akaike-Weights-Scoring [Ano23b, Sod20]. Gamma [DW01, DV04, HIS22, Mil19, WIR01]. Gap [OY23]. Garrote [Gao98]. Gauss [GRM09, Ste95]. Gaussian [Ano06a, AB21, ACS17, ACG14, AG18, BMG16, BGL18, BK20, CDG22, CZZ15, CBM24, CP18, CAPP23, DDM20, De 05, DY09, DBPW23, DIR02, EM07, EGCP24, FDC^{+}19, GM15, GAS20, GS23, GN17, GS24, GŠP^{+}06, GP11, GA15, Gu92, GX01, GX20, GF17, HHL19, HIS23, HPA14, HKP16, HD23, JP22, KGMT06, Kit96, KO05, Len03, LD11, LM19, LHK17, MK20, MMZ22, MR16, MC18, NBH^{+}15, ONS18, OW12, PH21, Par22, PC06, PM16, RRG^{+}21, SCG22, SC08, SS19, SQCK20, Ste12, Ste95, SBS23b, SSL17, TEG19, Tok07, VF09, WHH23, Whi01, WC94, XGH11, XGT15, YWBA22, Yua08, YS10, ZLR18, ZSW22, ZGTG21, vdBBD22]. Gaussians [PP14]. GEE [Hui22]. GEE-Assisted [Hui22]. Gelman [Gel17, Kos13]. Gene [CAR^{+}10, CZZ15, Jör09, XP07, ZGL22]. Gene-Environment [ZGL22]. Genealogy [ACCQ19]. General [ACCK19, BK20, BG98, DLSS24, DH03a, HPOL14, Hes98, Hsu92, HN98, HK98, LW09, LQ23, SJSG22, ZSL^{+}22, dVTP^{+}17]. Generalization [Eld03, OY23]. Generalizations [YZ13]. Generalized [AT16, BBW20, BEGS17, BY04, Bil00, BHK15, BD22, CGH17, CD12, CTB19, CZSL10, CL12, CTW20, CLXY23, CR24, CRS06, CZ02, CD14, DMR20, DG05, EM02, EGS^{+}13, FNGW20, Fri00, Fri13b, Fri14, GBW17, Gv97, GX01, GM05, GKZ10, GH18, Guh08, GRM09, HL03, HM09, HIS23, Hel24, Hof01, Hoo07, HLW22, HD23, HWO^{+}17, Ish99, JHL09, JAJ19, KW01, KO04, KM15, KW21, KTT23, LHL21, LMST18, LR05, LROI23, LPM98, LFS^{+}17, LP07, LHL03, Lum19, LVH17, LR14, LXZ22, MPS17, MHS^{+}14, Mou09, NL04, OW12, PW22, PC06, RNM18, RYY00, SD03, SMB14, SK10, SS05, SHFT18, SW18, SN10, TNKK16, TL07, VGUH23, VG01, VN18, Wan13, WL23, WYL14, XSC19, YZ07, YMLY11, Zha03, Zha07, ZNR21, Zha22, ZSL23, ZZL18, Zhu17, dR07, vW19, Ano20]. Generalizing [BZ07]. Generating [BC24, DES93, DZFZ18, MS93a, Yat23]. Generation [AC95, ES98, KK00, Mey18]. Generative [HPZ21, SWL23, WSB24]. Generic [Bra21]. Genetic [HLH06, JS95, Pit02, XCSP16]. Genetically [LYHM15]. Genetics [NP01]. Genomics [DXL12, ZSL^{+}22]. Genotype [KSC^{+}24]. Genotype-Environment [KSC^{+}24]. Geographically [GY07]. Geometric [LKW09, LWH23, PP09]. Geometrically [DH10]. Geometry [HK10, LZ18]. Geometry-Sensitive [LZ18]. Geophysical [Bra02]. Geostatistical [CYS09, GY07, PL12]. Geostatistics [Chr04, GW04]. Germany [FDG^{+}22]. ggplot2 [SVC^{+}19]. Gibbs [CS93, DGL23, FQH16, Fri13a, HIS23, HC98, Hob01b, HJ09, JP96, KKLN20, Lia98, LD22, NM98, PP96, PvD09, RS01, TH09, TRC23, VJ15, WRT21]. Gini [SZ08]. GIS [HHTB10]. Given [Coz99, Fal99, HM02, KK00, WW12]. GLM [TNNK20]. GLMM [TNNK20]. GLMMLasso [SMB14]. GLMMs [RZ22]. Global [BGLM18, Bie05, HCG02, RJLW21, YLLZ23]. Globally [AV22]. Glyphs [TS13]. Go [VGUH23]. Goal [YHS23]. Goals [GU13a]. God [XMY13]. Going [CHSR19]. Good [Rob19]. Goodness [AZ22, DG14, GPGMFB14, HM14, KN06, LRO18]. Goodness-of-Fit [DG14, GPGMFB14, HM14, KN06]. Google [LRR20]. Gossip [YVMFB03]. GPU [BGD16, MR16, SWC^{+}10]. GPU-Powered [MR16]. GPUs [LWLY17]. Graded [YR22]. Gradient [FS18, FGH^{+}24, KKK08, QWW23, SBR22, Zha07]. Gradient-Based [KKK08, FS18]. Gradients [KN97]. Grain [LCBV22]. Grammar [Bau19, VGUH23, Wic10]. Grand [CBCH95]. Granularities [GHCU22]. Graph [ASPL21, Bak92a, BH22, BFM18, CJW22, KS11, PC22, RMR^{+}17, STM15, Shi24, Sue12, TF20a, TG09, Wan20, YK09, YPPM21, ZRL14, ZL16]. Graph-Assisted [Wan20]. Graph-Based [YK09]. Graphic [BC93, Tuk93a, Yat23]. Graphical [ASND19, Ano06a, Ano07a, Ano08a, Ano12b, Ano24a, Ano24b, ACIP22, AAN17, BCL94, CGEGMI22, CLLZ17, CZZ15, CAPP23, DPS11, FK16, Far99, FB11, GM15, GB00, GN17, GLMZ15, HvHB23, HM14, HP94, Hur00, KKB21, KKHB23, LH15a, LD11, LM19, LCB19, LKB21, LH00, MS18a, MMA97, MY94, MBM18, MD13, MMZ22, MR16, OLN11a, SYNK24, SC08, SQCK20, TN14b, WW12, Wic13, WKM22, WFS11, XGH11, XGT15, YP20, Yua08, vdBBD22]. Graphically [Eic94]. Graphics [AZ22, Cle93a, Cle93b, GU13a, GU13b, Hol93, Hur93, IG96, KSC^{+}24, LYG^{+}10, Mur13, SVC^{+}19, Tre93, UHHS96, UVM18, VH17, VGUH23, WCH20a, WL02, Wic10, Wil93, WRCR00]. Graphing [Wai93]. Graphon [SK24a]. Graphs [CAPP23, CPV03, FSMV02, Gel11b, GV01, Gra13, KB08, KTT23, LRO18, LACRM06, ML08, OBGC21, PCP11, PH23, PSTV15, SG20, TAS^{+}17, TA12, Wil96, Wil99, ZLPT22, ZHZW22, Bri11, FK11, Gel11a, Wai11, Wil11]. Greater [Hic17, MHB17]. Grid [CHSR19, MPR^{+}12]. Gridding [BK03b]. Grids [QE96]. Gromov [LYXM23]. Gromov-Wasserstein [LYXM23]. Group [CRGS17, CCYW16, GN17, GLC20, Hel24, LAML19, LH15b, SBZ13, SFHT13, TV24, ZK22, ZWS23, vNvdBvdW22]. Group-Adaptive [vNvdBvdW22]. Group-Fused [GN17]. Group-Lasso [LH15b]. Group-Orthogonal [ZWS23]. Group-Wise [CRGS17]. Grouped [DBH^{+}22, QYZ16, SH17]. Groupings [vlM97]. Groups [CGEMIR18, LR04].

Growth [LYHM15, SG03]. Guarantees [CKT21, LLM23]. GUD [Kou14]. Guide [CGEGMI22]. Guided [HLH15, KN97, OFR^{+}18]. Guiding [YL95]. GWAS [ACCQ19].

Haar [FN04, FT16]. Hamiltonian [BGD16, HO10, Kle19, Kle22, OKL21, PP14, SBF19]. Hammock [Sch24]. Handling [HCE23, LAB22]. Happen [HV17]. Happened [HV17]. HARDI [RIC^{+}16]. Harmonic [Len09]. Hastings [GK10, HMJ11, JRS11, VJ15]. Having [GK97]. Hawkes [CS18, SCAP22]. Haystack [DS21]. Hazard [CHW02, Gra96, Gu15, HHGG01]. Hazards [SS11]. HDLSS [LAJ13]. Headers [WM03]. Heal [GF15]. Healthcare [Goo99]. Heatmaps [BY18]. Heavy [Mah17]. Heavy-Tailed [Mah17]. Heredity [HWS16, TRWB16]. Hermite [BMG16]. Hermitian [COvS19]. Hessian [CH09]. Heterogeneity [BBM24]. Heterogeneous [CBM22, CTDKL21, CZ22, CY24, DCS16, HvHB23, MWWY24, TvdV22, ZH22, ZZD23, ZLS23]. Heteroscedastic [BB03, BGL18, BBvdW16, CRC^{+}07, PCGM20, SMS^{+}14, Smy02, TYG^{+}20]. Heteroscedasticity [DM22]. Heuristic [WR93]. HHH} [MT00]. HHSVM [YZ13]. Hidden [BWZO20, BHS00, Cap11, Dan12, FPRW09, Gué03, JP22, LCBV22, LH02, PM16, RT98, SHBC^{+}24, TFS20, YSDJ15, ZJRI10, ZLS23]. Hierarchical [BS19, BM99, Bra21, BC24, CH12, DH10, GvDHB08, GKT18, GAS20, HL03, HTW22, HMN^{+}03, HS18, JB03, Kle19, LH15b, LH15c, OKL21, Pos01, RG09, SR23, VC21, WCR13, YT21a, ZGL22, ZSL^{+}22, ZZT22, ZWS23]. Hierarchically [SH18]. Hierarchy [VH17]. High [ACJ21, AR22, AG18, BHvdLD22, BCS96b, CHG16, CLLZ17, CB97, DDM20, Dem23, DM24a, FM12, GKT18, GF16, HM09, HM11, HHH22, HCE23, HK22, Hoo07, HO10, IW10, IL16, ILP19, JVSS24, JT21, JV01, JP95, KDM16, KB20, Kus06, LCV20, LCL22, LACV23, LB15, LRFK21, LP14, LMW^{+}22, LWH23, MMZ22, MS18b, MC05, NY14, NYC15, NFMS14, OFR^{+}18, PP07, PW15, PMI20, RRSY19, RSKZ19, RIC^{+}16, RYY00, Rho03, ROY20, RFV23, SMB14, SBE17, Shi24, SLMN20, SHFT18, SLR16, SLH24, SRL^{+}00, SS22, THSM21, TSH16, TRC23, WLWM16, WL23, WZXW24, XGTB18, YMB14, YKW03, YF14, YWL^{+}15, YWBA22, YF22, ZSK23, ZCY^{+}11, vdWvNR21, LDR23]. High-Dimensional [ACJ21, AR22, AG18, BCS96b, CLLZ17, CB97, DDM20, FM12, GKT18, GF16, HM11, HCE23, Hoo07, HO10, IW10, IL16, ILP19, JVSS24, JT21, KDM16, KB20, Kus06, LCV20, LCL22, LB15, LRFK21, LP14, LMW^{+}22, MMZ22, MS18b, NY14, NYC15, NFMS14, OFR^{+}18, PP07, PW15, RRSY19, RSKZ19, ROY20, RFV23, SMB14, SBE17, Shi24, SLMN20, SHFT18, SLH24, SRL^{+}00, THSM21, TSH16, TRC23, WLWM16, WL23, YMB14, YKW03, YF14, YWL^{+}15, YWBA22, YF22, ZCY^{+}11, HHH22, vdWvNR21]. High-Frequency [MC05]. High-Momentum [Rho03]. High-Order [CHG16, HK22, JV01, RYY00]. Higher [BB03, HM14, LXZ22, Mar94, WFCT23]. Higher-Order [LXZ22, Mar94]. Highly [HY17, KKHB23, SK21]. Hilbert [MYC^{+}22, SKK22]. Histogram [BC00, DM09, JB22, Tan17]. Histogram-Valued [JB22]. Histograms [TSH16]. Historical [LC24, Wic13]. History [Fie12, Fri02]. Hit [CS93]. Hit-and-Run [CS93]. Holdout [TVZ^{+}22]. Hole [LCBV22]. Holes [MB18]. Hollowness [LCV20]. Homotopy [OT11]. Honest [TDH15]. Horseshoe [LCB19]. Howard [Tuk93a]. HPD [CS99]. HQ [MT00]. HR [MT00]. Huber [LO16, WLZB07, YH17]. Huge [Weg95]. Human [Far04]. Hunting [DR10, HH01, Ooi02]. Hybrid [ACCK19, CCX22, Ish99, JV01, JP22, LCD^{+}24]. Hydrological [RT21]. Hyperbolic [LLRM24]. Hypergeometric [JKD92]. Hypergraph [TRC23]. Hyperparameter [HBG22, XD20]. Hyperrectangles [DES93]. Hypotheses [SSL^{+}16]. Hypothesis [AvD20, Can99, FKH07, MH17, TAS^{+}17, WLWM16, ZK22].

Icon [EM96]. IDE [VG01]. Idea [Gas03]. Ideas [Gra07]. Identifiability [GM20, MW15, Yat23]. Identification [BK05, PW22, SBZ13, SBK23]. Identify [GF22]. Identifying [CTDKL21, Dia99, HWQ15, MWFSG17, ZC15, ZSL^{+}22]. Identity [Mar10]. iid [WK17]. Ill [KN97, RO94]. Ill-Posed [KN97, RO94]. Illumination [ND21]. Illusion [VH15]. Illustrations [PvD09]. Image [BZ07, FR97, FT16, GJ97, Glu00, Glu04, GHC14, Kan20, KK00, LZYQ22, MQ15, NP01, RRH18, SYW^{+}22, SSL^{+}16, YBDS02]. Images [ABvS09, EM96, KK00, Mai97, QS09, SJM^{+}24, VM04]. Imaginary [FN16]. Imaging [EFG^{+}99, GAZ15, HDK^{+}20, KS22, PDM23, RIC^{+}16]. Imbalance [LAB22]. IMIS [SRE06]. Immersive [WS02a]. Impact [LRR20, WKT^{+}24]. Imperfect [CHLZ22]. Implementation [AB17, BHS00, ES96, FKH07, Tok07, VL09, WC06, XGT15]. Implementations [AT16, FM94, LC01, PMB12]. Implementing [Gas03, Jan06]. Implications [LB15]. Implicit [SYNK24, SLM22]. Import [ZH05]. Importance [BMGRR22, CR24, HB98, IPH22, Ion08, KA13, LYXM23, LZCZ22, MP00, MÉG^{+}23, MMN^{+}13, NP01, Oeh98, OKL21, PV23, SZ08, SK21, SRE06, SC19, TDH15]. Imposing [Ros17]. Improper [HC98, PP96, TH09]. Improve [HDK^{+}20, JRS11]. Improved [Gri23, HHH22, McL99, ST99, XGH11, XBW15, dV08]. Improvements [HO99]. Improving [AAN17, BC19, CKHG15, DNF23, Hof19, HHH12, KMGJ23, LCRG21, LRR20, PKH15]. Imputation [DL23, HJNR19, LBM^{+}14, SS22, ZG22]. Impute [CKR19]. Imputed [DBH^{+}22, LRBS03]. Imputing [FHTG05]. In-Degree [ABG^{+}21]. Including [Paa99]. Inclusion [SC08]. Incomplete [Fri00, HCZMH21, JBJ^{+}22, LPM98, SR23, SSL17, VG01]. Incorporate [Bra21, DMP14]. Incremental [FRW05, MFMR19, SRE06, SC19]. Independence [WW12, ZMH07, ZWZW24]. Independence-Encouraging [ZWZW24]. Independent [BC24, GK10, HMJ11, JRS11, LS93, MBY^{+}22, RRG^{+}21]. Index [Ano92g, Ano93f, Ano94g, Ano95f, Ano96h, Ano97f, Ano98f, Ano00g, Ano08c, Ano09b, Ano10b, Ano11b, CR24, EM96, FXC21, HSM09, Hsi99, LCBV22, PVH05, PPTO22, RDL07, YZ07, ZHZ12, dSM04]. Indexes [CBC93, Sta94]. Indicators [MLP07]. Indices [PP05]. Indirect [XMY13]. Individual [CDFB11, GKBP15, Laz05, LSFI18, Par11, WWL^{+}23]. Individual-Level [CDFB11]. Individualized [DZFZ18, PPTO22, WY22, XTPO22]. Individuals [LYHM15]. Inducing [GDM14, GH20]. Induction [WC06]. Infectious [GF22]. Inference [AZ22, AR22, AVW13, ALP^{+}23, BBM24, BCDL17, BLBCC22, BSX23, BZ97, BHF02, BH01, Bra99, CHZ23, CD12, CHG16, Che07b, CD19, CB20, CJW22, CGR17, CZZ15, CM13, CLSC22, De 05, DG24, DNF23, DGWR24, EC17, EB02, FD21, FP04, FPRW09, GK02, GMdL23, GD23, GF17, HAPRD24, Hil11, Hsu92, HH06b, HMZ22, HSZ23, JK00, JP22, KDM16, KFSL17, KNN05, KC20, LHL21, Laz05, LD11, LRBS03, LL11, MJVB18, MRP22, MN97, MBY^{+}22, MDW23, MPC24, MGL18, MV98, NM98, OW12, PCP11, PFBSO22, Pic14, PV23, PSTV15, RNHY12, RAM19, RW10, SR00, SG23, SHBC^{+}24, SLM22, SX21, Ste92, TEG19, TDM12, TPA19, TF20a, TFS20, Tsi02, UTK22, VC21, WD11, WZLR21, WHH23, Wan23, WBS22, WL23, WM19, YNTK21, FB94, VT94]. Inferences [EH05, NYC15]. Inferential [FLMM22]. Inferring [MCF^{+}20]. Infinite [BLBCC22]. Infinity [LCLR23]. Inflated [MCB24]. Influence [BG92, BZ97, CZSL10, MCF^{+}20, RRSY19, WKM22, ZICT12]. Influential [GRG10, PGM23]. Information [Bak92b, GU13b, GKZ10, HK10, HO10, IJ02, LLM23, LH02, NFM16, Rya03, ST99, Smi05, Spa05, WZZ^{+}24, ZH22, ZZF24, Bak92a]. Information-Theoretic [GKZ10]. Informative [Ano20, DMR20, KLR18]. Informed [Sar22]. Infovis [GU13a, Kos13, Mur13]. Inhomogeneous [DAMC23]. Initialization [CKR19]. Initiating [Tu94]. Inputs [TMC23]. Insight [SCLK^{+}00]. Insights [LRR20, WFS11]. Insomnia [YM11a]. Institutions [MHB17]. Instrumental [CG07]. Integral [HW21, NLF^{+}06, XWS24, Yan24]. Integrals [Zha22]. Integrated [BMGRR22, EJPS22, LSV^{+}09, Len09, MMLLQ^{+}22, SBJ17, ZSBS23]. Integrating [LPQ^{+}22, MHH15]. Integration [GK97, HDK^{+}20, SH97, Tan06, ZH22]. Intelligence [Rho03]. Intelligent [CKR19, SC98]. Intensity [FN04, GD23, MRCM18]. Interacting [BJMD13]. Interaction [HVRA96, IPH22, LQ19, Smi05]. Interactions [Cai99, HWS16, LH15b, SR23, ZGL22, ZSL^{+}22]. Interactive [BCS96b, EM96, HOD22, MMGG16, SH18, SCB98, UHHS96, Wil99, Wil00].

Interactivity [CCH16]. Interest [Fal99]. Interface [Hur93, Goo11]. Interior [CS98]. Interior-Point [CS98]. Internal [Kel11, Wu11]. Internet [CCS03, Rho03, WM03]. Interpolation [Ano12b, BRS11, FGN06, GX20, Zha03]. Interpretable [JWL^{+}23, OH24, WW19, ZA23]. Interpretation [WCM21]. Interpreting [FDG^{+}22]. Interquantile [JWB13]. Intersection [GV01]. Interval [AB17, AE92, BL04, CGL13, CJK22, CPCC24, DLSS24, HJ04, Kwo98, LRB12, LT94, Maa05, Pan99b, WK00, ZL22, ZZL18]. Interval-Censored [AE92, BL04, CGL13, CPCC24, DLSS24, HJ04, Maa05, Pan99b, ZZL18]. Interval-Valued [LRB12]. Intervals [CS99, DES93, EN20, FSR20, Fon19, GJ09, GW93, HWLB24, KLH99, KHM05, KS04, LJZD05, MS18b, RO94]. Interweaving [HR11, SNR17, YM11b]. Intractable [HSZ23, PH20, PFBSO22, SBF19, TNK17, YSDJ15, Zha22, ZF18]. Intransitivity [STIP23]. Intrinsic [COvS19]. Introduction [BKM99, Buj00, ES96, Edi00]. Introductory [HK17]. Intrusion [JV01]. Inuence [RA01]. Invariance [FHVW03]. Invariant [Cai99]. Invariants [PB22]. Inverse [BPW93, DV04, HIS23, LCB19, TDM12, TRC23, Wan20, Wu08, WLM10, ZYZ10, vW19]. Inversion [ACS17, FSR20, KO05]. Inversion-Free [ACS17]. Investigating [ZW04]. Investigation [EFLT94]. Invitation [Kos13]. Involving [GHW02, HIS22]. Irregular [CLHW19, MPR^{+}12, MB18, WL13b]. Irregularly [DKH06, DCT08]. Irreversible [MDS20]. Ising [FK16, KBM23]. Isomap [WL09]. Isotone [FM12]. Isotonic [BQS94, DJ04, FM12, LR14, QE96, WL08]. Issue [Buj00, Edi00]. Itô [Sin02]. Itô-Taylor [Sin02]. Item [PJS^{+}22].

Iterated [BP98, Bor11, CBS24]. Iteration [BPW93, JAJ19]. Iterative [BF02, BG98, CHSR19, CKR19, CM08, Hea11, Jon98, LAML19, Pan99b, PW15, SD03, ST19, SPW95, WZLR21, XGH11, XGT15, ZL16]. Iteratively [HB92, LP07, ZLP23]. IV [FB94].

J [Bak92a]. JADE [VLNO18]. Java [GI00a]. JCGS} [Ano21, Goo11, LTW^{+}10]. JOFC [LPPT17]. John [BC93, FB94]. Joint [CZZ15, Cul13, GM16, HCCZ23, KTS17, KS22, KSC^{+}24, SG05, SQCK20, VS17, WWL^{+}23, ZCI^{+}17]. Journal [Ano06a, Ano07a, Ano08a, Ano12b, Ano24a, Ano24b, OLN11a]. JSS [Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano05a, Ano05b, Ano05c, Ano05d, Ano06c, Ano06d, Ano06e, Ano06f, Ano07c, Ano07d, Ano07e, Ano07f, Ano08d, Ano08e, Ano08f, Ano08g, Ano09c, Ano09d, Ano09e, Ano09f, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano12c, Ano12d, Ano12e, Ano12f, Ano13c, Ano13d, Ano13e, Ano13f, Ano14c, Ano14d, Ano14e, Ano14f, Ano15c, Ano15d, Ano15e, Ano15f, Ano16d]. Jump [BB21, GD21, Gol09, Kan20, KA13, ML08, QY97, SG23, SQ07, ZR21]. Jump-Diffusion [BB21]. Jump-Diffusions [Gol09]. Jump-Preserving [Kan20]. Jumps [Qiu02]. Junction [TG09].

Kalman [DZL22, PRHP14, Sin02]. Kappa [LT94]. Kendall [VT94]. Kernel [CR08, GW18, GM20, GG17, HV05, Her97, JW08, JRH22, KH22, KS08, LW19, MRCM18, MSS08, RDZ10, SKK22, SG20, Wan94, Wu08, XTPO22, XSC19, ZWF^{+}22, ZMY^{+}23, ZY24, ZH05]. Kernel-Based [GM20, JRH22, MRCM18]. Kernels [All13, DS94, Mar94, WCR13]. Kingman [FW13, LFT17]. Knot [Lin99, MS03, Wan08]. Knots [Rup02].

Known [ACCQ19, Bre04]. Kriging [HCB04, MPS21, PW14]. Kronecker [CCX22].

L [CC22a, NST12]. Label [LF16, PI10, RW14]. Label-Switching [LF16]. Labeled [KSH23]. Lagged [MA19]. Lagrangian [LSSG15]. Landau [BJMD13]. Landmark [Big06, LC24]. Landmark-Based [Big06]. Langevin [OY23]. Langevinized [DZL22]. Language [IG96, WRCR00]. Languages [Buj00, Edi00, Hub00]. Laplace [BMGRR22, Bor11, BBvdW16, GJR05, MMLLQ^{+}22, RYY00]. Laplacian [PC06]. Large [Ano12b, BVZ20, BGL18, CW00, DP23, Deg23, DY09, EGCP24, FRW05, FPRW09, FGN06, GMF16, GŠP^{+}06, Gv97, GA15, GF17, HYL^{+}17, HM15, HY17, HWK17, HM16, HLL07, HS18, HB16, IW10, KSSL21, Kei96, KSM14, LMST18, LWLY17, LCRG21, LFS^{+}17, LVH17, MK20, Mou09, NBH^{+}15, Par18, Pos01, RAM19, SS12, SLBV18, SS16, TNKK16, WS24, Wil99, WL13a, WZXW24, XW18, YLW17, YR22, ZLR18, ZGC22, ZW10, ZF18]. Large- [HY17]. Large-Scale [EGCP24, Gv97, HYL^{+}17, LMST18, LWLY17, LVH17, YLW17, ZLR18, ZGC22]. Lasso [ASND19, AT16, AAN17, BRB17, CW22, CJW22, DP23, DS21, Fu98, GKZ18, GN17, Hoe10, HHTB10, Joh13, KKB21, KKHB23, LAML19, LWLY17, LH15b, MPS17, OT11, PWS16, RRSY19, SFHT13, TF20b, TLLJ23, WFS11, YWL^{+}15, ZGL22, ZHZ12, Zhu17, DRG19, FM12, HD23, JTU03, KKK08, OPT00, ZOR12, ZCO15]. LASSO-Like [HD23]. Lasso-Type [ZHZ12]. Lasso-Zero [DS21]. Latent [BCDL17, CD19, CTDKL21, CBL20, Eve12, HWO^{+}17, Hui22, HDK^{+}20, LCW22, LXZ22, Mah17, MVR14, MCB24, OM19, OPV22, PH21, RNHY12, SLL22a, SL10, SBS23b, XD20, XCSP16, YMG21, ZBB06]. Lattice [GŠP^{+}06, MB18, PL12, Smi05, SSL17, WK17, ZW04]. Lattice-Based [MB18]. Lattice-Valued [WK17]. Lattices [FPRW09, GF17, ZF18]. Law [YNTK21]. Layer [HWF22]. Layered [Wic10]. Layers [ZHZW22]. Layouts [Mur99]. Learned [LROI23, San23]. Learners [CM08, LD22]. Learning [CAPP23, GKBP15, GP11, GY22, HHH22, HLW24, HL23, HOD22, IPH22, KSM22, LH15a, LH15b, LP14, LPQ^{+}22, LSD05, LSLX22, MJLK21, OSR^{+}18, Par22, PRHS18, PGR15, QC24, SP23, SH06, SYL^{+}22, TMC23, WC23, WL23, WGSC24, Wu11, WY22, WZXW24, WBW21, XTPO22, YGD22, ZK22, ZHZW22, vWB22, Kel11]. Least [CHW12, DG05, HB92, HK10, HGK10, LP07, Paa99, XWZ15, ZLP23, ZLW21]. Least-Square [ZLW21]. Leave [KH22]. Leave-One-Out [KH22]. Length [MB04]. Lesions [SSL^{+}16]. Less [Zhe17]. Lessons [PSDK22]. Letter [HWK17, Loc11, Pie04]. Letter-Based [Pie04]. Letter-Value [HWK17]. Level [CDM^{+}20, CDFB11, GD23, GX20, Kle04, MBY^{+}22, SLR21, SQCK20, TN14b, YL22]. Level-Set [GD23]. Levels [CDP04]. Leveraging [WZZ^{+}24, WGSC24]. Lexical [GI00b]. Lifetime [BC00]. Like [BCH09, HD23, HV19]. Likelihood [ASND19, AVW13, ASPL21, BCDL17, BB21, Bie05, BLBCC22, CHG16, CJ02, CVA08, CL12, CS18, Chr04, CW00, Coz99, DDM20, DBPW23, DNF23, DIR02, Fea04, FQH16, FD21, FP04, Fri00, GW18, GCKP08, HPOL14, HO93, Hof00, HO99, HSZ23, JW08, JB03, KCSA20, KW10, Laz05, LPM98, LLZ09, Mar10, MMF^{+}19, MBGO11, MR19, NST12, NYC15, Pan99a, Par05, PH21, PP09, PB95, PV23, PDLN18, PSF^{+}22, RNHY12, RYY00, RAM19, SR23, Sin02, Ska02, SDBD21, Smi01, SBJ17, SY07, SSL17, SWF04, Tho93, TNK17, TG13, WZLR21, Wan23, YLLZ23, YZ06, YT99, ZJ04, Zha22, ZZWS23, Zho05, vNvdBvdW22, WDB96]. Likelihood-Based [MMF^{+}19, Smi01, TG13]. Likelihood-Free [DNF23]. Likelihoods [CGEMIR18, ESR^{+}14, HS18, LSV^{+}09, Len09, SGH17, YSDJ15]. Limit [Rya03]. Limited [ILP19]. Limits [LH95]. Linear [AHS22, BBW20, BEGS17, BY04, Bil00, BD22, CTC09, CHZ23, CGH17, CTB19, CZSL10, CLXY23, CPCC24, CRS06, CZ02, EM02, Fon19, FTAW21, Fri00, Fri07, GZQ^{+}22, GPGMFB14, Gra13, GCKP08, GH18, GLC20, Guh08, GRM09, HL03, HHL19, HWLB24, HH06a, Hel24, Hen04, Hes95, HRM02, HM20, Hsu92, HN98, HWO^{+}17, Ish99, JHL09, JPvD22, Kat10, KCF21, KPC02, KL03, KSLS22, LLW^{+}22, LR05, LRR20, LROI23, LG92, LCWW17, LPM98, LJZD05, LBCG16, LH15c, LHC17, LDHW23, Lum19, LVH17, MVR02, MPS92, MXM^{+}21, MDW23, MWDR21, MRCM18, MM20, NFM16, NL04, NFMS14, OW12, OT11, Par22, PLW01, PC06, PC22, RNM18, RYY00, SG05, SD03, SY02, SMB14, SLL22b, SR23, SBS^{+}23a, SNR17, SS05, SDBD21, Smi01, SHFT18, SX21, SLC22, SLH24, TN14a, TNKK16, VG01, VH22, XC22, YZ06, YZ07]. Linear [YOL^{+}22, YF14, Zha03, ZHJ16, ZNR21, Zha22, ZZF24, ZOR12, ZCO15, ZY05, ZWS23]. Linear-Angle [Gra13]. linearization [WDB96]. Lines [CH93, RHV22]. Lineup [CCHM18, LCTV24]. Link [Ano20, DMR20, LWLZ22, MP00, Mar10, PPTO22, RW99, ZWLZ17]. Link-Based [Ano20, DMR20]. Linkage [SN10, TKB23]. Linked [BC93, GR99, LA12, SVC^{+}19, Tuk93a]. Linking [GY07, SCLK^{+}00]. Lisp [Tie00]. Lisp-Stat [Tie00]. LispStat [Tie96]. LLE [GR12b]. Load [Bau19]. Loadings [GJL^{+}10]. Local [Ano24a, BRS12, CCHS08, CZSL10, CC22b, DAMC23, DR10, FSB11, Fos01, FTAW21, Fuk22, GW18, GA15, HPQ24, Her97, KM03, LCRG21, LZCZ22, MMRMGMGC08, MLP07, MQ15, MW04, PF23, Sam15, SG00, TW96, WHH23, ZLR18]. Local-EM [FSB11]. Local-Likelihood [GW18]. Localization [DNF23]. Localized [WLM10, XGT15]. Locally [BCS22, CKNK06, EH09, JPvD22, LCWW17, QGL09, SBR10, TS22, Tan17, WKC^{+}08]. Location [HRV12, KLR18, McL99, PWPS97, Rup92, UKZ18]. Location-Scale [KLR18]. Location/Dispersion [Rup92]. Locations [SHL19]. Log [DBPW23, DGWR24, Fea04, LMW^{+}22, PB95, SH03, WASB22, YT21b]. Log-Concave [WASB22]. Log-Concavity [DGWR24]. Log-Gaussian [DBPW23]. Log-Likelihood [Fea04, PB95]. Log-Rank-Type [LMW^{+}22]. Logarithm [DT13]. Logarithmic [Efr05, GR99]. Logic [RKL03]. Logistic [Bra99, CDFB11, ET99a, ET99b, LRO18, Len03, LAB22, LCLR23, McC99, PGR19, Tok07, Tüc08, WBQ20, WBS21, WZXW24, ZH05]. Logistic-Gaussian [Len03]. Loglinear [Bra99, TL99]. Logspline [KS92, KS04]. Long [MW96]. Long-Memory [MW96]. Longitudinal [CCH16, CDFB11, CLSC22, EC14, Far99, GDM14, GD16, HWQ15, HLL07, JS20, KS22, KH23, KTZQ20, MCF^{+}20, PJS^{+}22, PP09, PM16, Seg94, SJM^{+}24, SK24b, TQ16, WL13b, Zha97, ZCI^{+}17, ZDM20, ZZT22]. Look [GDS02]. Looking [Fal99]. Looks [GU13a]. Loss [BHvdLD22, DJM22, FLMMK24, FR97, Wan18, YH17]. Loss-Based [FLMMK24, BHvdLD22]. Losses [Kat10]. Lost [GF15]. LOTUS [CL04]. Low [CHPR24, HLP07, HS18, LM23, Oeh98, PRHP14, PAP22, QC24]. Low-Dimensional [HLP07]. Low-Rank [CHPR24, LM23, PRHP14, QC24, PAP22]. LowCon [MXM^{+}21]. Lower [CHW08, WCM21]. Lower-Dimensional [WCM21].

M [Ano24a, SWL23, FKLW08]. M-estimation [SWL23, FKLW08]. Machine [HOD22, IPH22, LSD05, WGSC24, ZK22, ZH05]. Machines [LL22, LY11, SWZ14, TZ06, WZW19, WL13b, Yat23, ZLWZ16]. Macro [UTK22]. Magnetic [EFG^{+}99]. Majorization [Wan18, YWL^{+}15]. Majorization-Minimization [YWL^{+}15]. Making [EH05]. MALA [RZ22]. Mammographic [HIT99]. Manage [GGT22]. MANET [UHHS96]. Manifold [JHD21, PC22, ZGTG21]. Manifolds [SSG23]. Manual [CB97, LACV23]. Manuscript [Lev12, Lev13b]. Many [CTC09, LGS09, SYF17]. Map [OKL21]. Mapper [DGS24]. Mapper-Type [DGS24]. Mapping [RHC^{+}14]. Maps [BL17, Mei11, MM97, SLR^{+}23]. Margin [Wan13, WL13a]. Marginal [BLBCC22, CVQ23, DNF23, Fri99, GMTK13, Lia07, LFT17, WCKL19, YC04, YT99, vNvdBvdW22]. Marginalized [CYS09, Sue12]. Marginally [KNS21]. Marginals [HM02]. Margined [LMS19]. Marked [DAMC23, LA12]. Markers [TZ12]. Marketing [KTZQ20]. Markov [Bro06, BG00, BWZO20, BHS00, CB01, Can99, Cap11, CS98, CF05, CRS06, CK08, Dan12, DSD^{+}23, DY09, Dob12, DIR02, Eve12, Fea02, FHIT20, Fis12, FWR12, FB11, FPRW09, GS94, Gey95, GVDP99, God01, GF16, GR08, Gra07, Gué03, GMP04, GJR05, HHC03, HK22, HZC23, HC98, Hob08, Hol23, HLH06, JN04, JS13, JT21, JO02, JV01, JP22, KKLN20, LSSG15, LH02, MP00, MB04, MFMR19, MN97, MDS20, ML08, Nea00, Owe17, PM16, RS19, RT98, SHC00, SK21, Sco99, SG23, SHBC^{+}24, SNR17, SS06, TDH15, Tan15, TA12, TFS20, YSDJ15, YT99, YS10, ZJRI10, ZR21, ZZ08, ZF18, ZLS23, dV08]. Masking [JP96].

Masses [HIT99]. Massive [Bra02, BKM99, DH03a, FS99, Goo99, HLB^{+}21, Hub99, JK21, KP03, KKHB23, LPS99, SDL19, SWC^{+}10, YSJS23, ZKS22b]. Massively [LYG^{+}10, RHC^{+}14, SWC^{+}10]. Matched [BVZ20, CHLZ22, YG22]. Matching [ASPL21, CHLZ22, CCHM18, DR18, DASR08, Deg23, GR93, HK06, LR04, MR01, Ros12, Ros17, YR22]. Mathematical [MI00]. Matrices [ACC18, ALP^{+}23, CJ09, COvS19, CHW08, Cli05, DP23, GDM14, GAS20, GG17, HLL07, LCB19, MCB24, PMS21, SG23, SLBV18, TDM12, TU11, ZBB06]. Matrix [AAN17, Bak92b, BK22, BHvdLD22, CHSR19, CKR19, CP18, DDM20, DS00, DY09, DT13, EGCP24, Fal99, HIS23, HPA14, HKP16, Hof09, HHT21, KH23, KKS17, LD21, LD06, LWZ14, LH02, MWWY24, MR19, NFM16, OS22, PDM23, QC24, SNC23, Smi05, Spa05, TMMB20, TGPM22, WP13, WZZ^{+}24, WL22, YT21b, YHS23, YG22, ZSL23, ZRL14, vW19, Bak92a]. Matrix-Free [DDM20, EGCP24, PDM23]. Matrix-Log [YT21b]. Matrix-Logarithm [DT13]. Matrix-Valued [LD21, MR19]. Matrix-Variate [TGPM22]. Matrix-Variate- [TMMB20]. Max [CHG16, CMG21, HSZ23]. Max-Stable [CHG16, CMG21, HSZ23]. Maximization [Hob08, LM19, LLZ09, SDL19, YSD13]. Maximizations [vlM97]. Maximizing [CW00]. Maximum [ASPL21, Bie05, Chr04, DBPW23, DIR02, Fri00, Hof00, HO99, JW08, JB03, KCSA20, KW10, LPM98, Mar10, MBGO11, PH21, PP09, RYY00, SR23, Sin02, Ska02, SSL17, SWF04, Wan23, YZ06, ZJ04, Zha22]. Maxwell [PFBSO22]. MCMC [HR11, AV22, BCD11, BAMG22, BK22, BK05, BDR97, BGP03, CMN21, CJ09, DH10, GPR99, HHC03, JL13, KA13, Kos22, LK02, LS05, MJVB18, PMB12, QTVK18, QTV^{+}21, RR09, SK21, Sha14, SGH17, SK24b, VS17, WB18a, YM11b]. MCP [JYYZ18]. MCVIS [LWM21]. Mean [Bre04, CKNK06, Far99, HH06a, HB16, Len09, LZ18, MWDR21, PSY09, XY14, MR10]. Mean-Covariance [MWDR21]. Mean-Shift [HB16]. Means [Dia99, HL23, MSS08, Ano24a, CC22b]. Measure [BHS04, GW11, LL15, RRH18, SZ08, YHH^{+}23]. Measurement [CHZ23, CD14, LR05, SMS^{+}14, SR23, SHFT18, YASR20, ZNR21]. Measurements [CZ22, ZJRI10]. Measures [Coz99, FHTG05, HIT99, KM03, NY14, YSG24, ZY24, ZICT12]. Measuring [CCHM18, LT94, PRHS18]. Mechanism [Mur99]. Median [DJ04, HL23, LMM19, YSG24]. Median-of-Means [HL23]. Medians [GBS24]. Medical [NP01]. Medium [Bau19]. Meets [DGWR24]. Mellin [GGT22]. Membership [LKR22, MSJ^{+}24]. Memory [MW96]. Merge [JN04]. Merging [Mel16]. Meta [YGD22]. Metaheuristic [MHDW19, SWX23]. Method [AX23, Bak92b, BCD^{+}24, BMR15, BK03b, CDM^{+}20, CKHG15, CSWZ23, CDFB11, DDM20, DM24a, Far99, FWR12, GJ09, Gel95, GS23, GM20, HLP07, Hes98, HP94, JH98, JP95, Kab93, KMOS92, KN06, MR19, PW22, PFBSO22, PWPS97, RMR^{+}17, SJSG22, SBK23, SN10, TEFH20, Vel08, XWS24, XL17, YMB14, YMLY11, ZLP23, ZZT22, Bak92a, LHL21]. Methodology [FHTG05]. Methods [BG98, CR08, CL15, CRS06, CMC^{+}22, Cle93a, Cle93b, CPV03, CMJ11, DH03b, DBS06, DBH^{+}22, FNGW20, FM07, FX21, GVBM04, GZQ^{+}22, GB02, God01, GM18, GLW07, GR93, Hol93, HJ09, HM16, HKS12, JRH22, KS99, Lee07, LYG^{+}10, LG92, LSFI18, MP00, Nea00, NW10, NTVK12, OLN11a, OLN11b, PDM23, PVH05, PTP03, PV04, RSKZ19, SD03, SBT00, SS05, Tre93, WL02, Wil93, YBDS02]. Methylation [GX20]. Metric [ZY24]. Metrics [CCHM18]. Metropolis [Béd08, CS93, GK10, HMJ11, JRS11, NW10, TMvD18, VJ15, YLCR19]. MFAI [WZZ^{+}24]. Micro [UTK22]. Micro-Macro [UTK22]. Microarray [QS09, YBDS02]. Microbiome [ZZW21, Fuk22]. Millions [CH93]. Mind [PRS11]. Mini [QWW23]. Mini-Batch [QWW23]. Minibatch [GZQ^{+}22]. Minimax [MJ18, Ros17]. Minimization [BGLM18, YWL^{+}15]. Minimizing [LJ19, PNLL21]. Minimum [CS14, DPS11, SP23, WR93]. Mining [DH03a, FS99, Goo99, WS02a]. Minorant [AE92, Jon98, Pan99b]. MIP [KCF21]. MIP-BOOST [KCF21]. Misclassified [Che23]. Mises [Hof09]. Missing [CBL20, FM10, LYLL23, LROI23, LKB21, SY02, SR00, Tu94, UHHS96, ZG22, ZICT12]. Missing-Data [Tu94]. Missingness [MWWY24]. Misspecified [MXM^{+}21]. Mixed [BD22, CHW02, CTB19, CZSL10, CLLZ17, CRS06, CZ02, CD14, ERPG11, EC14, GCKP08, GM05, GH18, Guh08, GRM09, HTW22, HvHB23, HMZ22, HJNR19, JHL09, KW01, LH15a, LKR22, LHC17, LHL03, LCD^{+}24, MWWY24, MSJ^{+}24, MDW23, NLF^{+}06, OW12, PB95, PLW01, PC06, Pro06, SY02, SSG15, SMB14, SBS^{+}23a, SS05, SX21, TN14a, TNKK16, VL09, Zha03, Zha22, ZBS^{+}21, ZWS23, vD00]. Mixed-Effect [GM05]. Mixed-Effects [EC14, LHC17, NLF^{+}06, PB95, PLW01, SY02, SX21, vD00]. Mixed-Frequency [HTW22]. Mixed-Measurement [CD14]. Mixed-Membership [LKR22]. Mixed-Variable [LCD^{+}24]. Mixing [LLZ09, PRS11, PMB12]. Mixture [ACG14, BCD11, BRC^{+}10, BAMG22, CD19, CW00, CM13, DSD^{+}23, EGO92, FW13, Fea04, FM07, FWR12, GK02, GB04, GPR99, GKN23, HLH15, HK22, JN04, Kar21, KCSA20, Li05, LF16, LFT17, MM98, MFMR19, MM10, Mel16, MW15, Nea00, NW10, OC01, PTP03, RNM18, RW14, RF17b, SŠS07, SLL22a, SLL22b, SPW95, SRE06, SC19, TEG19, Tan17, TQ16, Tüc08, WWG^{+}97, WD11, WM19, YKZ18, ZNYJ14, ZKS22a, ZZL18]. Mixture-Based [BCD11]. Mixturegram [YKZ18]. Mixtures [BCH09, CCFM01, DGL23, DV04, DH03b, GK10, GR08, GIA22, GW11, Hof00, HJR03, IJ02, JHL09, KLR18, MWFSG17, MR16, PI10, STM15, SG03, SS19, SWC^{+}10, TN14a, TGPM22, TGM^{+}14, WR05, WIR01, WRK11]. MM [AX23, HL00a, KHYD23, LW08, ZL10, ZHZL19]. Mode [HMW06, Kle08, MW97, MS93b, MMW98, Ooi02]. Model [AR22, BCE20, BCH22, BCDL17, BY04, BCH05, BK23, BFM18, BHK15, Bra21, BGP03, BWZO20, CHW02, CGEGMI22, CKHG15, CGEMIR18, CS20, CH09, CD19, CLXY23, CSWZ23, CHSR19, Chr04, Cle93a, Cle93b, CGL11, CM13, CCL21, DXL12, DYB^{+}23, DLSS24, DS21, DG24, DMP14, DCV10, ES02, ERPG11, FK23, FK16, FH15, FRW05, FP04, Fri00, GD21, GPGMFB14, GR99, God01, GIA22, GM93, Gua23, Guh08, HJ04, HMN^{+}03, HML18, HPZ22, HHKS11, Hol93, HDM02, Hsu92, HN98, HHTB10, Hug15, JN04, JK00, JPvD22, JBJ^{+}22, Jör09, JV01, JP22, KNN05, KK09, KBM23, KLH05, Kus06, LG07, LYHM15, Len03, LG03, Li05, LD21, LDR23, LM23, LCLR23, LYM23, LYLL23, Lia07, LPM98, LG20, LKR22, LSLX22, LHC17, LQ19, LCD^{+}24, LXZ22, MK20, Mah17, MCB24, MPR^{+}12, MMN^{+}13, MW15, MXM^{+}21, MC05, MQR11]. Model [MSS08, MW14, MQ17, NBH^{+}15, OC01, OR07, OH24, Paa99, PDM23, Pan99a, Pan99b, Par01, PPTO22, PWS16, PB95, Pos01, PM16, QYZ16, QGL09, RNHY12, RHF24, RHO22, RS19, Rya03, SS11, Sew21, SG17, ST99, SMK15, SK24a, SNB^{+}01, SM13, SC23, SLC22, TPA19, TRWB16, TK99, Tre93, Tsi02, VHSG20, WBQ20, WDMW22, Wan23, WJL23, Wil93, WCM21, XC22, XL17, YPPM21, YKW03, YSK^{+}10, ZHH08, ZZW21, ZHJ16, ZCI^{+}17, ZLR18, ZCS^{+}19, ZSBS23, ZZF24, ZZ08, ZJA16, ZZL18, ZY05, dVTP^{+}17]. Model-Based [CHW02, CGEMIR18, Chr04, FRW05, LG07, LM23, OR07, PDM23, Pos01, Sew21, ZHH08, ZZW21, ZSBS23]. Model-Free [LD21, LYM23, XL17]. Modeling [AM07, BD23, BRCC22, BGL18, BH22, CHZ23, CD19, CMG21, CG07, DR18, DZZ16, DK18, EC14, EGCP24, GM16, GRM09, HWS16, HYL^{+}17, HCE23, Hil11, Hof01, HZ21, KS22, KT03, KH23, KSM14, KMQ05, KKB21, KKHB23, LCSL14, LGL16, MM10, NW10, NAB16, OR99, PQD22, PH23, PW14, PC22, QN00, RR23, STIP23, TQ16, TGM^{+}14, TL07, WCH20b, WKT^{+}24, XCSP16, YT21a, ZCI^{+}17, ZKS18]. Models [AB21, ACS17, Ano20, AGV12, ACG14, AHS22, BCST13, BBW20, BS19, BEGS17, BB21, BAMG22, BM99, BF02, BT18, Bil00, BH22, BRS^{+}08, BLR22, BFM18, BHK15, BC24, Bra99, BWZO20, Buj04, BHS00, BD22, CMN21, CDG22, Cap11, CD12, CKNK06, Cha22, CTB19, CZSL10, CLXY23, CLLZ17, CWZ21, CG07, CGJ09, CR24, CRS06, CMC^{+}22, CZZ15, CZ02, CAPP23, CGR06, CYS09, CD14, CLSC22, DZZ06, Dan12, DG14, DMR20, DZL22, DH10, Dri08, ESR^{+}14, FNGW20, FW13, Fea04, FM07, FXC21, FM10, Fie12, FCWM17, FWR12, FMM^{+}19, FB11, FLMM22, Fri07, GF22, GM15, GZQ^{+}22, GK06, GK02, Gel04a, Gel04b, GvDHB08, GS23, GN17, GP11, GCKP08, GM05, GKZ10, GH18, GLC20, GRM09, GLMZ15, HL03, HPOL14, HIS22, HAC^{+}12, HHC03, HC11, HHH22, HK22, HM15, Hel24, HvHB23, HCB04]. Models [HM14, HHT21, HD23, HO99, HY19, HWO^{+}17, Hui22, HH06b, HKS12, HMZ22, HOD22, HK98, HSZ23, IPH22, Ish99, JB03, JHL09, JPvD22, JWB13, JL13, JT21, Kab93, Kar21, KW01, KFSL17, KO04, KL03, KW21, Kit96, Kle19, KTT23, KS99, KC20, LGS13, LRO18, LCW22, LH15a, LD11, LCRG21, LF16, LM19, LCLR23, LZM24, LR05, LROI23, LCWW17, LMS19, LFT17, LKB21, LBCG16, LH15c, LHC17, LHL03, LDHW23, Lum19, LVH17, LH02, MS18a, MM98, MJVB18, MA19, MM14, MVR14, MSJ^{+}24, MMF^{+}19, MÉG^{+}23, MMLLQ^{+}22, MMRMGMGC08, MHDW19, MT03, MHS^{+}14, MDW23, MBM18, MMZ22, MR16, Nea00, NLF^{+}06, NFM16, NW10, NL04, OW97, OY23, OW12, OKL21, PVH05, PL12, PH21, PJS^{+}22, Par22, Pic14, PTP03, PLW01, PC06, PGM23, Pro06, QG06, RNM18, RYY00, RW14, RDL07, RG09]. Models [RFV23, RT98, SBM22, SYNK24, STM15, SG05, SY02, SSG15, SMB14, SC08, SŠS07, SLL22a, SLL22b, SQCK20, SR23, SBS^{+}23a, SHBC^{+}24, SSBW17, SNR17, SS05, Ska02, Smi05, Smi01, SL10, SHFT18, SX21, SRE06, SSG23, SBS23b, Sue12, TDM12, TH09, TN14a, TF20a, TVZ^{+}22, TGP03, TL99, TV24, TFS20, TNKK16, Tüc08, TG13, TS13, UKZ18, VL09, VM04, VG01, VH22, VN18, VF09, VDC21, WK09, WD11, WH14, WHH23, WBS22, WM19, WBS21, WKM22, WYL14, WYZ19, WRK11, WW19, XGH11, XGT15, XW18, YP20, Yan21, Yan24, Yat23, YSD13, YSDJ15, YZ06, YZ07, YKZ18, YF14, Yua08, ZHZ12, Zha03, ZBB06, ZJRI10, ZNYJ14, ZZZS17, ZNR21, Zha22, ZGC22, ZWZW24, ZBS^{+}21, ZKS22a, ZHZL19, ZW07, ZW10, ZICT12, ZWS23, ZLS23, dLG01, dR07, dVTP^{+}17, vD00, vWB22, vdBBD22, TYG^{+}20]. Modern [FNGW20, RSKZ19, RHV22, ST19]. Modes [HLW22]. Modifications [Glu00]. Modified [JTU03, XGTB18, YF14]. Modular [HK17]. Modularity [MPC24]. Modulated [Sco99]. Modulo [FHVW03]. Moment [HLW24]. Moments [Gel95, KR17, MMDL22, Tho93, YC04, YMLY11]. Momentum [GZQ^{+}22, Rho03, SBR22]. Momentum-Based [SBR22]. Monitoring [BG98, SG05, VS17]. Monotone [FX21, LT98, LBM^{+}14, Wan08]. Monotonic [BLBCC22, TL07]. Monotonically [Yu12]. Monotonicity [BJG98, CGH99, HHGG01, HV19]. Monte [Ano07a, BG00, FHIT20, HZC23, JS13, MFMR19, OKL21, PP14, SHC00, SBF19, BGD16, BP98, Bro06, BC19, CB01, Can99, CGMR04, CS99, CF05, Chr04, CRS06, CMC^{+}22, DG14, DH03b, DBS06, DMP14, FBG06, FKH07, Fea02, Fea05, Fis12, FWR12, GS94, Gel95, Gey95, GVDP99, God01, GF16, GLW07, GR08, Gra07, GMP04, GJR05, HT17, Hel04, Hob08, Hol23, HLH06, HW21, HJM22, HK98, Ish99, JN04, JB03, JHD21, JT21, JO02, Kit96, Kle19, Kle22, LGS13, LSSG15, LYG^{+}10, LC01, Lia07, LJN^{+}17, MP00, MN97, MDS20, ML08, NJM20, PCP11, PH21, RJLW21, SNR17, SS06, Spa05, Tan06, Tan15, WGDW22, WJG11, XBW15, YSDJ15, YT99, ZC15, ZZ08, ZJA16, dV08]. Morse [GRB^{+}13]. Mortality [JVSS24, SH17]. Mosaic [Fri99, Fri02]. Motif [Ano24a, CC22b, CC18]. Movies [LK02, LTW^{+}10]. Moving [HCB04]. MPPCA [HLH15]. MR [PDM23]. MSiCOR [DSD^{+}23]. Much [Bri11, FK11, Gel11a, Gel11b, Kos13, Wai11, Wil11]. Multi [ACJ21, BVZ20, HAPRD24, HWF22, JK21, KSC^{+}24, LZ23, MS18a, MHS11, PNLL21, SJM^{+}24, Shi24, SWL23, THM^{+}23, WL23, WYZ19, YG22, vdWvNR21]. Multi-Attribute [KSC^{+}24]. Multi-Class [ACJ21, MHS11]. Multi-Dimensional [SJM^{+}24]. Multi-Layer [HWF22]. Multi-penalty [vdWvNR21]. Multi-Period [THM^{+}23]. Multi-Purpose [SWL23]. Multi-Resolution [JK21]. Multi-Response [WYZ19]. Multi-Sample [Shi24]. Multi-Scale [MS18a]. Multi-Study [HAPRD24]. Multi-Task [LZ23, WL23]. Multi-Valued [BVZ20]. Multi-View [YG22]. Multicategory [LW08, LSD05, LY11, ZLWZ16]. Multiclass [HLD^{+}13, TZ06, Wan13, WZW19]. Multidimensional [AvD20, BD23, BSL^{+}08, CMG21, CVQ23, HLB^{+}21, Huh95, Hur04, LLRM24, MY94, Par18, RB02, SBL23, ZSK19, Fuk22]. Multifold [CLXY23]. Multilayer [Li05]. Multilevel [AM07, BEGS17, CLCX22, HJNR19, PC06, ZCY^{+}11]. Multilinear [Paa99]. Multimodal [YLLZ23]. Multimodality [TMvD18]. Multinomial [KN06, Kwo98, PGR19, Res22, RDL07, WBQ20, WZXW24, YKW03]. Multinomials [DS94]. Multiobjective [DCT08, SMK15]. Multiple [AvD20, BRCC22, BEM13, CR99, DSD^{+}23, DL23, GS24, GM16, HWLB24, HT17, HH06a, HM15, HP94, HN98, HLW22, JKD92, KTS17, KCB19, LCBV22, LWLY17, LBM^{+}14, LJZD05, MA19, MJLK21, PHK14, PMI20, PMS21, QN00, RAC23, RG09, SS12, SQCK20, SSP^{+}18, SHL19, TDH15, TZ12, Wai93, WZ22, XTPO22, YK01, YLCR19, ZHJ16, ZY21]. Multiple-Fold [GS24]. Multiple-Runs [ZHJ16]. Multiple-Try [YLCR19]. Multiple-Use [HWLB24]. Multiplicative [PCGM20, SM13]. Multiply [DBH^{+}22]. Multiply-Imputed [DBH^{+}22]. Multiresolution [FMM^{+}19, HLH06, NBH^{+}15, ZY05]. Multiscale [ABvS09, EH05, LX16, PLH10, RW10]. Multiset [HML18, HSC18, KM15]. Multistage [Fis12]. Multitapers [BPW93]. Multitype [LG20]. Multivariate [BCH09, CHPR24, CW22, CTB19, Che23, CCH16, DG18, DZZ06, DK18, DRG19, EM07, FXC21, FEF22, Fri07, GDS02, Gas03, GKT18, Gen92, GB02, GMTK13, GG17, GIA22, GR93, HPZ21, Hof09, HSG23, HB98, JHL09, JP95, KR17, KFSL17, Kle04, Kle06, Kle08, KMQ05, KKHB23, Kus06, LW19, LN09, LL10, LMC22, LS05, MVR14, MKK18, MWDR21, MMDL22, Mou09, PP14, PM20, PRCF24, PP07, PFBSO22, PLW01, PC22, PV04, PWPS97, PM16, QG22, RYY00, RLZ10, Rou08, Rup92, SM19, Sar22, SY02, SBE17, SK10, Som98, SH97, SW18, SLH24, TDM12, Tho93, TGM^{+}14, Tre06, Unw19, VHSG20, Wan94, WWG^{+}97, Wan23, WL23, WJZ09, YT21a, Zha97, ZJRI10, ZZZS17, ZL10, ZZT22]. Multivariate- [Som98]. Multiway [GEH^{+}22]. Musical [Rap01]. Mutagrams [MW04]. Mutually [GR08, PH23].

National [JVSS24]. Natural [Wil00]. Nearest [FDC^{+}19, Gho07, SdWDL23]. Need [VH15]. Needles [DS21]. Needs [MC17]. Negative [BHF02, Gao98]. Neighbor [FDC^{+}19, Gho07, Hob08, SdWDL23]. Neighbor-Dependent [Hob08]. Neighborhood [FWR12]. Nested [BMGRR22, GD21, GLC20, LMC22, MMLLQ^{+}22, RYY00, YL22, YLM17]. Net [Cul13, Hel21, QYZ16, TNNK20, YH17, vNvdBvdW22, CCS03]. Nets [RR22]. Network [AR22, ALP^{+}23, BH22, CKHG15, CC18, CZZ15, EM18, Fie12, FMM^{+}19, GM16, HK06, HCCZ23, HKS12, JL13, KS99, Lee02, MA19, MOPW20, Mci99, MBY^{+}22, MBM18, OH24, OPV22, PJS^{+}22, RNHY12, RHO22, SBK23, WW12, ZZF24]. Network-based [OH24]. Networks [ABG^{+}21, ASPL21, BC16, BRB17, CGLW21, Eve12, GM15, GMdL23, GY22, HABV06, HPZ21, HWF22, HH06b, Ish99, JRH22, KSM22, LWLZ22, LXZ22, MMA97, MCF^{+}20, Mic12, MPC24, MJLK21, MRCM18, MM20, PH23, QC24, RR23, STM15, SLL22b, WL24, ZZSW23, ZZD23, ZA23, ZWLZ17, ZH17]. Neural [CGLW21, GMdL23, HPZ21, Ish99, JRH22, KS99, Lee02, LZYQ22, MJLK21, OH24, RR23, RR22, SLL22b, SBK23, WL24, ZA23]. Neurophysiology [Bri94]. Neutral [Lee07]. Newton [AX23, HYL^{+}17, YH17]. Next [SH03]. NicheWorks [Wil99]. NIMBLE [dVTP^{+}17]. no [Bak92a]. Nobler [PRS11]. Node [KL03, ZCS^{+}19]. Nodes [KM03]. Noise [GS23, HMP96, YB22]. Noisy [DG24, DIR02, GG20, HQT92, HCCZ23, SBF19]. Nomination [ZLPT22]. Non [BGLM18, CBM24, DM24a, Gao98, Gu92, GX01, Kit96, MC18, PNLL21]. Non-Convex [BGLM18]. Non-Gaussian [CBM24, Gu92, GX01, Kit96, MC18]. Non-Negative [Gao98]. Non-Smooth [PNLL21]. Non-Sparse [DM24a]. Noncentral [Ano06a, KGMT06, Tsi02]. Noncollapsing [CLHW19]. Nonconvex [LB15, PW15, Wan18, YLW17]. Noncrossing [KCB19, RF17a]. Nonlinear [BB21, BB03, CHW12, CH09, CAR^{+}10, CGLW21, ERPG11, FH15, FX21, Kit96, KC20, LQ23, MHDW19, NLF^{+}06, NW10, PB95, RR23, Sin02, Ska02, WK09, WC23, WYZ19, ZYKC08]. Nonlinearity [FSMV02]. Nonlinearly [Pro06]. Nonnegative [PA14]. Nonnormal [GM18, NLF^{+}06, OM19]. Nonnormality [LN09]. Nonparametric [AVW13, BBM24, BD23, BCH09, BH01, BGP03, CAR^{+}10, CH16, CGH99, CH99, DXL12, DK18, DZ99, DM22, DM24b, FM94, Far04, FH15, GK02, Gra96, GM05, HHGG01, HC11, Hel21, HV05, Hil11, Hof00, HB16, IL16, Jon98, KHYD23, Kar21, KW01, KT03, KLH99, KMQ05, KS11, LZM24, LD22, LLZ09, LLZ23, LHL03, MS93b, MKN23, NMD^{+}20, OBGC21, OLN11a, OLN11b, PHK14, Qiu02, QN00, QM04, RHC^{+}14, RMWH17, SBT00, SG03, SK10, SŠS07, SBEG94, Shi24, WDMW22, WLZ^{+}24, WH98, WKC^{+}08, YT21a, YKW03, ZJ04, ZWZW24, ZH03]. Nonparametrically [RW99]. Nonproportional [OH24]. Nonregular [TN14b]. Nonreversible [GD21, Kos22]. Nonseparable [GX20]. Nonsmooth [FS18]. Nonstandard [Spa05]. Nonstationary [AHNC13, BRCC22, CHH10, CMG21, DP23, DGW09, Gri99, HMP96, IP04, KNH23, KSM14, KKB21, KKHB23, LRFK21, MGF22, THSM^{+}20, YWBA22, YS10, ZW10]. Norm [LZ08, SLH24]. Normal [CGEMIR18, DW01, DV04, EGO92, FIK19, GGT22, GDS02, Gas03, GKT18, Gen92, GIA22, IJ02, KR17, KW10, LS93, LN09, LL10, LH15c, MMDL22, SYNK24, SG05, SG03, Som98, SH97, TGPM22, VL09, Wan23]. Normal-Inverse [DV04]. Normal/Independent [LS93]. Normalized [GW11, Hof19]. Normalizing [Lee02, SRE06, ZF18, dV08]. Normals [GK10]. Note [Fri00, MR01]. Notes [Wic13]. Novel [LQ21, OSR^{+}18, ZF18]. NPMLE [AB17, Maa05]. NSHP [VM04]. Nuclear [SLH24]. Nugget [PW14]. Nuggets [BCD^{+}24]. Nuisance [DM24b]. Null [SSL^{+}16]. Number [CGEMIR18, CD19, FP20, GBS24, Kou14, LCLR23, Rup02, WS24, WZXW24, YKZ18, ZLS23]. Numbers [Lia98]. Numerical [ALP^{+}23, EGO92, Gen92, JS95, Kle22, Sch24, Som98, Zha22]. Numerically [KN06]. Nyström [HM16, SS12].

O3 [Unw19]. OAK [Wic11a]. Oberon [Mar96, Saw96]. Object [FF94]. Object-Oriented [FF94]. Objective [BHK15, Gel00, GH00, HL00b, Kie00, LHY00, Men00, Wu00, dLM00]. Objectives [WZ22]. Objects [SSB^{+}14]. Oblique [JWL^{+}23]. Observation [KB99]. Observational [BVZ20, CHLZ22, KSR19, LSFI18, Ros12, Ros17, YR22]. Observations [FM10, GRG10, LGS09, LZM24, PGM23, PM16, SSP^{+}18, TU11, VLNO18]. Observed [ASPL21, Bak92b, EJPS22, LQ21, LH02, NFM16, PSTV15, SG23, SY07, YHH^{+}23, YB22, ZWLZ17, Bak92a]. Obtaining [Fri00, KS04, LPM98]. Occurrences [CH99]. Occurring [EM18]. Ocean [ES19, HMN^{+}03]. Odds [Hof01, LH95, OH24, PLIK02, ZZL18]. Off [CHSR19]. Offset [FIK19, KW10]. Offset-Normal [KW10]. Old [Gas03, VRH19]. Omegahat [Lan00]. One [AvD20, GH21, KH22, LFS^{+}17, Lum19, Ste12, Tad17, TPA19]. One-Factor [TPA19]. One-Step [GH21, LFS^{+}17, Lum19, Tad17]. Online [Cap11, KDM16, LP23, MÉG^{+}23, SJ05, WCSY21, YSD13, YB22]. Onto [HLP07]. Open [Fie12, Sha14, SdWDL23]. Open-Faced [Sha14]. Open-Set [SdWDL23]. Operations [Bau19]. Operators [KO05]. Opportunities [Kos13]. Opposing [PRS11]. OPTICS [RHM18]. Optimal [CDP04, CLHW19, CR08, DASR08, DGM17, GM18, HK06, KSH23, KYW19, LPQ^{+}22, LR04, MHDW19, MR01, Ros12, Ros17, SHL12, WYZ19, WZ22, Zha03, ZNR21, ZZWS23, ZMY^{+}23]. Optimality [XSC19]. Optimally [QTV^{+}21, Ros12, Tan17]. Optimization [ACJ21, BRS^{+}08, CC08, CJS92, DSD^{+}23, DCT08, GP11, HKP16, HBG22, KKK08, KS99, LHY00, LPJ13, LL15, LCD^{+}24, OW97, PN22, PL22, SBR22, ST19, SHBC^{+}24, SC19, WS24, XBW15, ZGTG21, Gel00, GH00, HL00b, Kie00, Men00, Wu00, dLM00]. Optimization-Assisted [ZGTG21]. Optimized [RMH21]. Optimizer [JS95]. Optimizing [PPTO22]. Optional [JMY^{+}16]. Oracle [KLH99, ZH22]. Orca [SRL^{+}00]. Order [BB03, CHG16, CTW20, DGL23, EHW03, FG95, Fea04, HK22, JV01, LXZ22, Mar94, MM20, QM04, RYY00, SWX23, SQ07]. Order-of-Addition [SWX23]. Ordered [BL22, BQS94, DGL23, LBM^{+}14, LKW09, MPS92, MC05, QE96]. Ordering [CR99, LLM23]. Orderings [vlM97]. Ordinal [CBL20, DK18, GLMZ15, KT03, KMQ05, OH24, TG13]. Ordinary [CHW12, CGH17, Gra96, XWS24]. Ordination [ZZW21]. Organizing [MM97].

Oriented [FF94, Kei96]. Origin [LA12]. Origin-Destination [LA12]. Orthant [AG18]. Orthogonal [JAJ19, ZWS23]. Orthogonally [CT01, PAP22]. Orthonormal [CBC93]. Other [BML19, RW10]. Oue [FB94]. Our [YM11a]. Outcome [RNM18, WSM^{+}20, XTPO22]. Outcome-Weighted [XTPO22]. Outcomes [LT94, VG01, XGTB18, Yan21, Yan24]. Outer [EM96]. Outlier [CB20, DG18, HL06, JP96, KH21, KH22, MT03, PP07, PF23, WWG^{+}97, ZB10]. Outliers [AR06, BH02, CR99, MM97, RA01, Unw19, WJZ09]. Outlying [NGH17]. Outlyingness [RRH18]. Output [BCE20, FBG06, Hel04, LK02, dV08]. Outrageous [PRS11]. Over-Dispersed [AC95]. Overlap [Mel16]. Overlapping [ZSBS23]. Overparameterized [OY23]. Overrelaxed [Yu12].

P [Ano08a, AGV12, BMC05, LB04, SL10]. P-splines [Ano08a, AGV12, LB04, SL10, BMC05]. Package [BY18]. Packet [CCS03, GVBM04, WM03]. Page [Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95a, Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e, Ano96f, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00b, Ano00c, Ano00d, Ano00e, Mur99]. Painless [Bau19]. Pair [BC16, LR04, VN18]. Pair-Copula [BC16, VN18]. Pairs [EGS^{+}13, Fri14]. Pairwise [HO10, Mel16, Pie04, STIP23]. Pandemic [BLR22, PSDK22]. Panel [LG03]. Paradox [Eld03]. PARAFAC [Gua23]. Parallel [Bro06, Hol23, JRS11, LYG^{+}10, LWLY17, MJVB18, MMV13, Mou09, MLJ16, Paa99, PCG^{+}14, RHC^{+}14, RTL07, Šev04, SWC^{+}10, TNKK16, VGUH23, YWL^{+}15, YLW17]. Parallelizable [KYW19, RAM19]. Parallelization [CMN21, HLB^{+}21, TEFH20, YSJS23]. Parameter [BRS11, CGEGMI22, Eve12, FS18, HM15, KBM23, LD06, LCD^{+}24, MBM18, NFM16, SCAP22, Sin02, TS13, XCSP16, XWS24, Yat23, YSDJ15, ZR21]. Parameterization [MWDR21]. Parameterizations [BS19, GvDHB08]. Parameterized [TS22]. Parameters [CHW12, CTW20, DZL22, EGO92, LCLR23, Mil19, ZW07]. Parametric [CSY00, CS14, DG14, KS04]. Parent [MMO01]. Parsimonious [CH16, FQH16, XZYZ24]. Part [KK09]. Partial [CLXY23, DS00, DG05, Fri99, GKZ10, Hea22, JPvD22, LLM23, WR05, ZZWS23]. Partially [BQS94, EJPS22, LR05, LBCG16, LDHW23, MJVB18, PvD09, SDBD21, VJ15, YZ07, ZHJ16, ZWLZ17]. Particle [Ano07a, CMN21, Fea02, Fea05, GP11, KSV23, KP03, MLJ16, NFM16]. Particular [Fal99]. Partition [GZQ^{+}22, GD16, Mei11, MQR11, MGF22, PN22, Sta94, WCKL19]. Partition-Based [MGF22]. Partitional [HH11]. Partitioning [HDRM05, HHZ06, IW10, IW22, KSH23, OSR^{+}18, ZHH08]. Partitions [PQD22, PN22]. Path [AT16, EFLT94, Hoe10, HW21, LSV^{+}09, RMR^{+}17, SUL23, VH22, ZL16, ZL13]. Path-Following [ZL16]. Path-Integral [HW21]. Paths [Ho06, Tad17]. Pathwise [Gri23, LPJ13]. Pattern [HC07]. Patterns [DG14, DM24b, EM18, ES19, LA12, MM09, MRCM18, SMN15, YL99]. PC [KB08]. PC-Algorithm [KB08]. PCA [JWH16, SBJ17]. PDMP [SF23]. PDMP-based [SF23]. Peak [GK97]. Pearson [Sta95]. Peeking [GKBP15]. Penalization [LYLL23, SMB14]. Penalized [Ano08a, BMC05, BHK15, CHW12, CRC^{+}07, DT13, FLY21, FKLW08, Fu98, GBW17, GTZH16, GBS24, GBC^{+}11, HPOL14, HM20, JAJ19, KCO09, KCK08, LPHK16, Lin99, LLZ09, LVH17, MPTZ23, MBGO11, MR19, NST12, Par05, PW15, RMWH17, Rup02, SBJ17, SSMN16, TCCSB10, TGP03, YT21a, YSK^{+}10, YH17, YF14, YLW17, vWB22]. Penalties [CGR17, GH20, Gri23, HEF17, LB15, LW07, MPS17, SP23, YT21b, vNvdBvdW22]. Penalty [Efr05, FML19, FS18, HD23, KK09, LLW^{+}22, SLR21, SBR10, WSB24, Yen20, vdWvNR21]. Penguins [VGUH23]. People [WCH20a]. Perfect [GHW02]. Perform [LYG^{+}10]. Performance [BBKC24, CS93, MM10, MWM^{+}23]. Performing [YK09]. Period [THM^{+}23]. Periodic [UTK22]. Permutation [KMS23, WLWM16]. Persistence [BM19, MGL18, PB22]. Personalized [PDM23, WGSC24]. Perspective [ST19, dR07]. Perturbation [CZSL10]. Perturbative [PRHP14]. Phase [ZKS18]. Phase-Amplitude [ZKS18]. Phases [FCL10]. Phylodynamics [CP22]. Phylogenetic [CH12, Kos22, MN97, WB18b]. Physical [LPS99]. Physics [KP03]. PICTs [JH98]. Piecewise [GN17, GD23, Kat10, KO05, MVR02, OT11, WJG11]. Piling [LAJ13]. Pillai [Mul98]. Pivotal [LF16]. Pixel [Kei96]. Pixel-Oriented [Kei96]. Plackett [Gas03]. Plaid [MM14]. Planar [SSG23]. Plane [AHNC13]. Planets [EM96]. Pleiotropy [XCSP16]. Pleistocene [RL14]. Pliable [TF20b]. Plio [RL14]. Plio-Pleistocene [RL14]. Plot [BY04, BBLG13, EGS10, EGS^{+}13, Fri13b, GB00, GF15, Hur93, LCTV24, LACRM06, LN09, Rob19, Unw19, dSM04, Fri14]. Plot-Data [Hur93]. Plots [BCL94, BEM13, Fri07, GKBP15, Gra13, GW15, HH11, HWK17, HMS97, HS10, KSC^{+}24, LFZ97, MHD20, Mou09, MHS11, Mur99, MI00, PV04, RR22, SS11, Sch24, VGUH23]. Plotting [GMF16, Tre95]. Plug [BF02]. Plug-In [BF02]. Plus [BHS00]. Point [BCST13, BAMG22, BM19, Cai99, CPP23, CS98, CH16, DAMC23, DG14, DM24a, DM24b, EM18, FEF22, HC07, JP95, LCSL14, LKB21, LA12, MLP07, MGL18, MRCM18, MM20, PH23, RL14, SG17, WR05, XC22, ZZSW23, ZY21]. Points [CR08, Laz05, LRR20]. Pointwise [Efr05]. Poisson [Fri00, CMN21, CS20, DC22, FW13, FN04, GM20, GRM09, HL03, LPM98, LFT17, Mar10, OC01, PFBSO22, QYZ16, QTV^{+}21, Sco99, YK01]. Polar [JHD21]. Policy [BLR22]. Polishing [Lum19]. Polya [HMJ11, JHL09, JMY^{+}16, NBNA21]. Polychoric [ZLP23]. Polyclass [KS99]. Polygons [MHD20]. Polynomial [BL22, BHF02, EB02, FG95, Sam15, SG00, TW96]. Polyserial [ZLP23]. Pooling [GR99]. Popularity [KTT23]. Population [CGMR04, HJM22]. Populations [GRVE18, KTS17, ZH22]. Portfolios [BO05]. Posed [KN97, RO94]. Position [RZ22]. Position-Dependent [RZ22]. Positive [AHNC13, COvS19, GW18, LS04, MS93a, ZRL14]. Positivity [OC01]. Possibilities [Lan00]. Possible [BL04]. Possibly [DR18]. Post [CR98]. Post-Processing [CR98]. Posterior [BRS11, BRS12, BC24, CGR06, CYS09, Coz99, DC22, GD09, GW11, Guh08, Hel04, Lee02, SBS23b, Tho93, Vie00, WCKL19, WCM21, ZB10, vdBBD22]. Posteriors [BK22, HC98]. Postsmoothing [QS09]. Potatoes [GM93]. Potts [MSS08, MW14, MQ17]. Power [AVW13, BMO01, FN16, Gri23, PCP11]. Power-Conditional-Expected [FN16]. Powered [MR16]. Powerful [CJW22]. Practical [BGL18, BH22, GF16, KHM05]. Pre [Bro06, JS20]. Pre-Experimental [JS20]. Pre-Fetching [Bro06]. Precision [AAN17, Fri13b, KKS17, Mai97, PMS21, SLBV18, WB18a]. Predetermined [RS01]. Predicted [DPS11]. Predicting [BCE20, SH03]. Prediction [BMW17, De 05, DCS16, ESR^{+}14, FH15, FLMMK24, HDRM05, HCG02, LC24, LWLZ22, MA19, PQM22, SJM^{+}24, Ste92, TW05, THM^{+}23, WH14, WM19, XP07, ZWLZ17, ZW10]. Predictive [CDG22, Cha22, Cha24, CYS09, FLMM22, HZ21, LS05, ZB10]. Predictor [KPC02, LJZD05]. Predictors [CTC09, Che23, CL09, HY17, LD21, MR19, SBZ13, SH06, WCR13, XGTB18, ZSK23]. Preferential [CP22]. Presence [BBvdW16, CGJ09, Dem23, DM24b, LKB21, OS22, SMS^{+}14]. Preserving [BCD^{+}24, Kan20, WL22]. Prevalence [FCWM17]. Preview [Lev12, Lev13b]. Primary [Ano01b]. Principal [CRGS17, CR23, CBM24, CLCX22, Dan97, Fal99, Far12, Fro22, GJL^{+}10, GAZ15, HLH15, HLW24, HM16, JLAW22, JCZG19, JU00, JTU03, KXY20, KTZQ20, KNH23, KLZ23, LRB12, LW09, LSH16, LO16, LWH23, MHH15, Par05, PP09, RMWH17, SYW^{+}22, SJM^{+}24, WH17, YLM17, YSWS24, ZSK23, ZP14, ZCY^{+}11, ZHT06, ZGPM22]. Principles [MPH22, Tuk93a]. Prior [Coz99, GDM14, GD09, GM18, Hea23, IJ02, LD11, LLM09, LHK17, RRG^{+}21, ZZF24]. Priors [CDG22, CPP23, DV04, DH03b, FN16, GD16, GR99, GH23, HHL19, JT21, LD22, MBY^{+}22, NM98, PP96, TH09, Tok07, Vie00, WP13]. Private [BRMC19]. Probabilistic [CC22b, Gua23, GG20, HLH15, Rap01, Ano24a]. Probabilities [AG18, CB01, CDG22, GDS02, Gas03, GKT18, Gen92, GB02, Som98, SH97, WK17]. Probability [Bor11, Coz99, FKH07, GHCU22, Hel04, LFZ97, NLF^{+}06, WZW19, Yan24, YSG24]. Probit [CDG22, DLSS24, LSLX22, MC05, TDM12]. Problem [CC08, CJS92, JAJ19, LF16, LAB22, LH15c, PI10, TAS^{+}17, TLLJ23, Zhu17]. Problems [BCDL17, Bri94, BK03b, CDP04, CCHS08, FS18, GBW17, GV01, Gv97, HM09, JP96, Kat10, KN97, LHL21, LGS09, MS18b, OC01, OT11, Paa99, PRS11, RW10, RO94, SWX23, TRC23, Vel08, Yen20]. Procedure [JN04, Qiu02, SMK15, XGH11, XGT15, YK01, ZG22]. Procedures [AG05, LG07]. Process [ACS17, BCST13, BGL18, BM19, CDG22, CSY00, CS18, DH03b, GF22, GK02, GAS20, GS23, GS24, GP11, GA15, GX20, GHW02, HD23, IJ02, JN04, JP22, Len03, LCSL14, LHK17, MK20, MM98, MR16, Nea00, NBH^{+}15, OW97, PH23, RRG^{+}21, Sco99, Tok07, WD11, WHH23, YK01, YWBA22, ZNYJ14, ZLR18]. Processes [BAMG22, BMG16, Cai99, CHG16, CH16, CMG21, CP18, DAMC23, DZZ06, DBPW23, EGCP24, FDC^{+}19, GD23, HCG02, Lee07, LCSL14, MP16, MLP07, MW96, MM20, NST12, SCG22, SG23, SG17, SCAP22, SY07, SSL17, WR05, Whi01, WJG11, WC94, YB22, ZR21, ZSBS23]. Processing [CR98, FT16, JYYZ18, Lev12, Lev13b, YWL^{+}15]. Processor [Rho03]. Product [CCX22, CVQ23, KTT23, MQR11, RRG^{+}21, WH14]. Products [Hea22]. Professional [YHS23]. Profiles [CJ02]. Program [CH93, LBM^{+}14, MC17, Paa99]. Programmers [Cha00]. Programming [Joh13, KCSA20, LLW^{+}22, LKW09, SWC^{+}10, VH22, WYZ19, YASR20, ZY21, ZLPT22, dVTP^{+}17]. Projected [NBNA21]. Projecting [HLP07]. Projection [BK03a, CBC93, CBCH95, GH18, HPQ24, IW10, LACV23, LCKL05, MJ18, PP05, Pos95, QLCZ15, RH94, SS19, WS02a, WLWM16, dSCL21]. Projection-Based [GH18, IW10]. Projections [CB97, DGWR24, OFR^{+}18, SLR16, FB94]. Projective [GM99]. Prone [Che23]. Prony [KMOS92]. Proof [LZ18]. Proof-of-Concept [LZ18]. Propagation [MBS04]. Proper [NM98]. Properties [BF02, CVA08, DS94, ES98, GM20, HSG23, JS95, MKN23, RHF24, RS01, Ste13, VDC21, Wan23]. Property [PN22]. Proportional [SS11, ST19, XGH11, XGT15, XW20]. Proportions [KCSA20, Kwo98]. Proposal [vdBBD22]. Proposals [HMJ11]. Prosection [FB94]. Protocol [LCTV24]. Providing [MI00]. Proximal [HZC23, TZ06]. Proximity [RMH21]. Proxy [RL14]. Pruned [ZY21]. Pruning [FWR12]. Pseudo [BLBCC22, DIR02, HO99, Len09, SDBD21]. Pseudo-Bias [Len09]. Pseudo-Likelihood [DIR02, HO99, SDBD21]. Pseudo-Marginal [BLBCC22]. Pseudolikelihoods [BFM18]. Publications [Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano05a, Ano05b, Ano05c, Ano05d, Ano06c, Ano06d, Ano06e, Ano06f, Ano07c, Ano07d, Ano07e, Ano07f, Ano08d, Ano08e, Ano08f, Ano08g, Ano09c, Ano09d, Ano09e, Ano09f, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano12c, Ano12d, Ano12e, Ano12f, Ano13c, Ano13d, Ano13e, Ano13f, Ano14c, Ano14d, Ano14e, Ano14f, Ano15c, Ano15d, Ano15e, Ano15f, Ano16d]. published [Ano12b, OLN11a]. Publishing [LTW^{+}10]. Purpose [SWL23]. Pursuit [BK03a, CBC93, CBCH95, LCBV22, LCKL05, PP05, Pos95, RH94, SS19, dSCL21]. Pyramid [RDBF19].

QGPCP [Mou09]. Quadratic [Ano06a, GGT22, KGMT06, KCSA20, Vel08, Wan18, YASR20, ZSW22, ZLPT22, ZHZW22]. Quadrature [CZ02, GM18, PC06]. Quantification [LHL21, RSKZ19, XBW15]. Quantifying [BL17, CHLZ22, RHM18]. Quantile [BY04, BEM13, Che07a, CY24, CPCC24, Dia99, DGW09, DS96, FLY21, FGV22, FSR20, HL00a, Hut02, KHYD23, KCB19, LZ08, MPTZ23, OLN11a, OLN11b, OT11, PLH10, PW15, PW22, QC24, RF17a, RDBF19, Ros17, SK10, STZM09, SX23, WLZ^{+}24, WSB24, YNH18, YH17, YLW17]. Quantiles [Bre04, CGR06, KHM05, MJLK21, MT00, PLH10]. Quantities [TW96]. Quantization [Bra02, BFE^{+}03]. Quantized [Mou09]. Quantizing [SLR^{+}23]. Quantum [Hol23, HW21]. Quartet [BBKC24]. Quasi [AX23, Ano07a, BCH05, BC19, CMC^{+}22, Fea05, HYL^{+}17, HPZ21, HJM22, Lia98, RR22, Yan21]. Quasi-Empirical [Yan21]. Quasi-Monte [BC19, CMC^{+}22, HJM22]. Quasi-Monte-Carlo [Ano07a, Fea05]. Quasi-Newton [AX23, HYL^{+}17]. Quasi-Random [HPZ21]. Quasi-Symmetry [BCH05]. Quasicomplete [Ish99]. Question [HR11, YM11b]. Questionnaire [LPQ^{+}22]. Quick [Kwo98].

R [BY18, BML19, IG96, RTL07]. Radial [BRS^{+}08, CZ02, ZDM22]. Rainbow [HS10]. RAMlet [ST04]. Random [Ano07a, ALP^{+}23, BSZ22, BH22, BK20, BFM18, CC13, CS98, CBL20, CDFB11, CK08, De 05, Dem23, DES93, DH10, DZFZ18, DIR02, DGWR24, ES98, Eve12, Fea05, FN16, FPRW09, Fri13a, GZQ^{+}22, GGT22, GW18, GH21, GW11, Guh08, HHC03, HC11, HPA14, HPZ21, HSG23, JWL^{+}23, KKLN20, KTT23, LRO18, Lia98, LPJ13, LZCZ22, LSFI18, Mah17, MRP22, MPS21, MDW23, MH17, MKK18, ML08, NLF^{+}06, NJM20, PQD22, PCP11, PH21, PP07, RYY00, Rya03, STM15, SH06, SR23, SLR^{+}23, Ska02, SS06, SLR16, Ste12, Sue12, TH09, TF20a, TAS^{+}17, TA12, TG13, WLZ^{+}24, WSM^{+}20, WK17, XNN16, XGTB18, YK09, YS10, ZF18]. Random-Direction [CS98]. Randomization [CHLZ22, TVZ^{+}22]. Randomized [DPRK23, DS96, Kle22, ZGC22]. Randomness [MDS20]. Range [HEF17, HGK10]. Rank [BMG16, CHPR24, CHW08, CSK10, HS18, Kus06, LM23, LMW^{+}22, LMC22, MPS92, MS93a, MS18b, PRHP14, QC24, TSH16, WWJ22, ZRL14, PAP22]. Rank-Based [ZRL14]. Rapid [WBS22]. RAPTT [SLR16]. Rare [SLR^{+}23]. Rashomon [BBKC24]. Rate [BGLM18, Gra96, HHGG01, Jör09, SG23, ZZL18]. Rate-Distortion [Jör09]. Raters [LT94]. Rates [GRM09, Hob08, HCZMH21, Kle22, SH17]. Ratio [ACJ21, BBLG13, FKH07, GCKP08, LH95, PLIK02, Zho05]. Rational [BK20, BSX23]. Ratios [Ano06a, Hof01, KGMT06, SdWDL23]. Raw [LPPT17]. RC [KNN05]. Real [Dem23, HMW06, LBW14, MDD^{+}12, WY22]. Real-Time [Dem23, HMW06, LBW14, MDD^{+}12]. Real-World [WY22]. Realized [SM13]. Really [Bri11, FK11, Gel11a, Gel11b, Wai11, Wil11]. Recapture [BBLG13]. Recentered [MMN^{+}13]. Reciprocal [HIS22]. Reclassification [Hea11]. Recognition [BMO01, FHVW03, NGH17]. Recognize [MMO01]. Recombine [XW18]. Recommender [WBQ20]. Reconciliation [AHS22]. Reconstructing [MK92]. Reconstruction [HCCZ23]. Record [RL14, TKB23]. Records [MHG24]. Rectangular [QE96]. Recurrent [Dem23, LG20]. Recursive [CJK22, HHZ06, OSR^{+}18, TA12, VM04, ZHH08]. Recycling [CR98]. Redistricting [FHIT20]. Reduce [GH21]. Reduced [GR12a, Kus06, LMC22, MPR^{+}12, PH21, SHBC^{+}24, WWJ22]. Reduced-Dimensional [PH21]. Reduced-Rank [Kus06, LMC22, WWJ22]. Reducing [BCD^{+}24]. Reduction [CL09, DYB^{+}23, FG95, FX21, Hen04, HV05, HY17, KH21, Lia98, LZCZ22, Maa05, SY16, SBJ17, Vel08, WY15, XZYZ24, YC04, ZYKC08, ZYZ10]. Redundant [GvDHB08]. Reemployment [KK09]. Reference [RW99]. Regeneration [BK05, TH09]. Regimes [XTPO22]. Region [GB12, SHMD^{+}14]. Regional [BCD11]. Regions [BL04, CDP04, CLHW19, FLMN23, LMM19, MB18, San23, SHL19, Som98]. Registration [Big06, EH17, LHK17, ZSK19]. Regressing [XMY13]. Regression [AG96, AB21, Ano08a, AR06, BMC05, BCH22, BG92, BRMC19, BB03, BCS22, BRB17, Bil00, BQS94, BJG98, BBvdW16, CTC09, CHZ23, CKNK06, CW22, Che07a, CHW08, CY24, CC22a, CPCC24, CT01, CDFB11, DC22, DR18, DK18, DLSS24, DCV10, ET99b, EC14, FG05, FLY21, FM12, Far04, FSR20, Fon19, GZQ^{+}22, GK06, GRB^{+}13, GBC^{+}11, GHC14, Gra96, HHL19, HWLB24, HL06, HCE23, HB92, HPQ24, Hel21, Her97, HK10, HRM02, HM20, Hug15, Hui22, HL00a, IL16, JWB13, JT21, KHYD23, KLSR05, KCO09, KF24, KCF21, KLH99, KSM^{+}18, KNS21, KHM05, Koo97, KS11, KB20, LGS13, LS93, LRO18, LB15, LC24, LZ08, LX16, LAB22, LP23, LPHK16, LRR20, LG92, LCWW17, LJZD05, LMC22, Loc18, LQ19, LQ21, LQ23, LR14, LBW14, MP16, MA19, MPTZ23, MSBV22, MAJ^{+}98, Mar10, MMLLQ^{+}22, MKK18]. Regression [MDC17, MWDR21, MKN23, MQR11, MQ17, NY14, NTVK12, OW97, OLN11a, OLN11b, OH24, OT11, Par01, PLH10, PHK14, PW15, PMI20, PAP22, PCG^{+}14, PCGM20, PGM23, PGR19, QLCZ15, QE96, QC24, QY97, Qiu02, RRSY19, RSKZ19, RHC^{+}14, RMWH17, RAC23, RF17a, RDBF19, RLZ10, RKL03, Rup92, SS11, SBY06, SBWZ08, Sam15, Seg94, SG00, SBZ13, SBK23, SMK15, SX23, SDBD21, Smy02, SW18, SWF04, SLC22, SLH24, SQ07, TRS97, TCCSB10, TW96, TL07, TG10, Wan08, WL08, WK09, WH14, Wan20, WDMW22, WWL^{+}23, Wan23, WLZ^{+}24, WSB24, WL23, WWJ22, WBS21, WS02b, WYZ19, WKC^{+}08, WCR13, Wu08, WLM10, WZXW24, XWZ15, XSC19, XY14, YNH18, Yan21, YT21a, Yan24, YKW03, YSK^{+}10, YH17, YC04, YOL^{+}22, YF14, YWL^{+}15, YLW17, ZZZS17, ZSL23, ZSK23, ZOR12, ZCO15, ZZ08, ZH05, ZYZ10, vWB22, vdWvNR21, ET99a]. Regression [McC99]. Regressions [CL09, Fu98, HDM02, HJR03, KCB19]. Regularization [All13, CTB19, CMJ11, FS18, Hel24, Kat10, KN97, LH15b, LMC22, LQ21, MBM18, MBGO11, PMS21, SLH24, Tad17, VH22, WWJ22, WNA20]. Regularized [CH09, DP23, GN17, Hel21, Kus06, LLW^{+}22, WH17, Yua08, ZZSW23]. Regularizers [HTW22]. Regularizing [RIC^{+}16]. Regulation [CZZ15]. Reinforced [LY11, ZLWZ16]. Reject [CR98]. Rejection [GT11, Tan06]. Rejoinder [BF94, Cle93b, ET99a, Gel04b, Gel11a, HL00b, Tuk93b, You95, YM11a, vDM01b, FB94, Tuk93a]. Relabeling [LAB22, RW14]. Related [HK06, KMOS92, KSR19, LYHM15, LD11, OT11]. Relational [Hof09, LSLX22, MCF^{+}20]. Relations [DGS24]. Relationship [BM99, God01]. Relationships [HM11, Hob01b]. Relative [FGH^{+}24]. Relaxation [SBT00, SBS^{+}23a]. Reliability [KM03, KY95]. REML [LPJ13, Smy02]. Remote [BFE^{+}03]. Removal [Ano13g, CB20]. Renewable [CY24]. Renewal [CS18]. Reparameterization [LD06]. Reparameterizations [KLR18]. Reparameterized [CYS09]. Repeated [TRS97]. Repelling [TMvD18]. Repelling-Attracting [TMvD18]. Replicated [DP23]. Replicates [BC24]. Reported [XGTB18]. Representation [FLY21, HABV06, Kel11, Pie04, QG06, SBR22, SSP^{+}18, Wu11]. Representations [HP94, Wil96]. Representative [BVZ20, Seg94, ZMY^{+}23]. Representing [MC18]. Reproducibility [BML19]. Reproducible [Bau19, GL07, HBG22]. Reproducing [SKK22, VRH19]. Repulsive [BAMG22, SG17]. Rerandomization [JS20]. Resampling [CMN21, HSM09, MLJ16, Tan15]. Rescaled [Kle19]. Rescaling [CR98]. Research [GL07, HMN^{+}03, LBM^{+}14]. Residual [BCST13, CHLZ22, Dia99, LCTV24, LH15c, MHD20, RR22, XMY13, Yan21, ZMH07]. Residual-Based [ZMH07]. Residuals [BY04, CWZ21, DS96, GS24, Yan24]. Resolution [HMS97, HCG02, JK21, MKE^{+}21, RIC^{+}16, WL24]. Resolution-Consistent [HCG02]. Resonance [EFG^{+}99]. Resources [BP98]. Response [CDP04, GPGMFB14, MM09, MWDR21, OW97, PJS^{+}22, PGR19, PSDK22, SHL12, TS13, VL09, WYZ19, ZSL23]. Responses [Che23, FXC21, MKK18, OM19, PM20, WBQ20, ZSK23]. Restarts [HV19]. Restoration [CSY00, GJ97, Glu00, Glu04]. Restore [VM04]. Restricted [CJ09, GCKP08, RF17b]. Results [PRHS18]. RETAS [SC23]. Retraction [Ano23b]. Return [SLR21]. Returns [NAB16]. Revealing [Gra13]. Reversible [KA13, ML08]. Reversing [LCL22]. Review [Lev11]. Revisited [CT01, GX01, ST19]. Revisiting [JRH22, VRH19]. Reweighted [HB92, LP07, ZLP23]. Reweighting [ZB10]. Rhythms [FCL10]. Rich [HCCZ23]. Richard [Tuk93a]. Ridge [Fri13b, HQT92, PGR15, XSC19, vW19, vdWvNR21]. Ridgelines [LL12]. Ridging [SG00]. Riemannian [MPS21]. Right [Lee07]. Ring [BW17]. Risk [BO05, BOW02, GF22, KK09, MAJ^{+}98, WKT^{+}24, ZJRI10]. Risks [KSSL21, PP96]. Road [MMW98]. Robust [BHS04, CGEGMI22, CHPR24, CC08, CP16, CW22, CHW08, CS14, CC22a, CRS06, FD21, HL06, HR05, HB92, HL23, HRV12, JK00, KN06, LS93, PP07, PLW01, PV04, PWPS97, PF23, QC24, RAC23, SS01, SH00, TEG19, VM04, WLZB07, WJZ09, WYL14, WL13b, XL17, dLG01]. Robustification [FG95, KB08]. Robustifying [JHL09]. Robustness [Hsi99, LS97, LL11]. ROC [TZ12]. RODEO [Sam15]. Role [OC01, STM15]. Root [OR99]. Root- [OR99]. Roots [AAN17, Gri99]. ROPE [KKS17]. Rotated [JU00]. Rotation [Par05, RB02]. Rotational [SJH^{+}15]. Roughening [SL99]. Roughness [SBR10]. Routine [DSD^{+}23, WCDS03]. Rubin [Gel17]. Ruin [CJS92]. Rule [GF16, LLW^{+}22, SD03]. Rules [LB15, PPTO22, WY22]. Run [CS93]. Runs [ZHJ16].

S [Tuk93a, BHS00, CCS03, HB92, Ho06, NS94, TCCSB10, MMGG16]. s-CorrPlot [MMGG16]. S-Estimation [TCCSB10]. S-Net [CCS03]. S-Paths [Ho06]. S-Plus [BHS00]. Saddlepoint [ZBY09]. Sage [LCL22]. Same [XMY13]. Sample [AAN17, BBM24, BO05, CGJ09, DYB^{+}23, GB00, KTS17, Kwo98, LYLL23, Mey18, SS12, Shi24, SLR16, TAS^{+}17, YSJS23, ZB10, ZK22, ZWF^{+}22]. Sample-Size [YSJS23]. Sample-Wise [LYLL23]. Sampled [LWLZ22, MRP22]. Sampler [DGL23, FGH^{+}24, HIS23, HMJ11, Hob01b, HSC18, KKLN20, KM15, Lia98, LFT17, Owe17, RS01, SUL23, SWL23, TRC23, WRT21, YLLZ23]. Samplers [CS93, HJ09, PvD09, VS17, VJ15]. Samples [ACIP22, BVZ20, CR98, GMP04, HM15, Mül99, Pan99a]. Sampling [ABG^{+}21, ACG14, Béd08, BMGRR22, BR02, CP22, CS98, CC18, CGL11, CYS09, DC22, DW01, DGL23, DBS06, EC17, FW13, FQH16, GM20, GT11, HHL19, HC98, HPZ21, HM16, HB98, Ion08, JP96, KA13, KSV23, Kos22, KSM22, LGS09, Lee07, LSV^{+}09, LL10, Lum19, MWWY24, MÉG^{+}23, MMN^{+}13, MS02, MDS20, NP01, NM98, Nea00, NL04, Oeh98, OKL21, PP96, Par01, PV23, PSRM^{+}23, SC92, SK21, SG20, SNB^{+}01, SRE06, SC19, SF23, TH09, TDH15, Tan17, TGHL14, Tüc08, WLZD11, YLCR19, ZBB06, ZR21, ZNR21]. Sampling-Based [OKL21]. Sandwich [Sha14]. Sarmanov [PFBSO22]. SARS [CP22]. SARS-CoV-2 [CP22]. Satellite [HCG02]. Satellites [KB99]. Satisfying [MS93a]. Scagnostics [WW08]. Scalability [EK02]. Scalable [BBW20, CDM^{+}20, CDG22, CHZ23, Deg23, DGWR24, FNGW20, GAS20, GS23, JP22, KSSL21, LYM23, NMD^{+}20, PMI20, RSKZ19, SBM22, SMN15, SBS23b, TFS20, WZZ^{+}24, XD20, ZGL22]. Scalar [GPGMFB14, GHC14, RMWH17]. Scalar-on-Function [RMWH17]. Scalar-on-Image [GHC14]. Scalars [KB20]. Scale [EGCP24, GMC02, Gv97, HAC^{+}12, HYL^{+}17, KLR18, LMST18, LWLY17, LVH17, MS18a, OM19, PB22, PHK14, RAM19, STZM09, UKZ18, YLW17, ZLR18, ZGC22]. Scale-Space [PHK14]. Scale-Usage [HAC^{+}12]. Scaled [ZGPM22]. Scales [PC22]. Scaling [BSL^{+}08, CMG21, HSM09, HLB^{+}21, LLRM24, MY94, Par18, RB02, RMH21, ST19, SBL23, XGH11, XGT15, Fuk22]. Scan [DKH06, NP01, PN22, PTP03]. Scanning [DM22, SSMN16]. Scatter [HRV12]. Scatterplot [DS00, MMGG16]. Scatterplots [ET99b, EH05, Tre95, ET99a, McC99]. Scenarios [DW02]. Schedules [WCH20a]. Schemes [BK05, NL04, SNB^{+}01]. Science [BW17, Don17, FS99, Hic17, HV17, MC17, MHB17, Pen17, Zhe17, HJ17]. Scientific [Bri94]. Scope [GI00b]. Score [FBG06, MGF22, NFM16]. Scoring [Ano23b, Sod20]. Screening [FMP24, GS23, LDR23, LYM23, LP23, LP14, RSDG23, XL17, ZCO15]. SDE [Pic14]. Sea [PRS11]. Search [DJM22, Kar21, LCD^{+}24, MM09, MR16, Ros93, SC08, TK99, WR93]. Seasonal [McE22]. Second [CTW20, Fea04, MM20, SQ07]. Second-Order [CTW20, Fea04, MM20, SQ07]. Section [BKM99, ES96, LCBV22]. Sectional [KMGJ23]. Sections [GG20, FB94]. Secure [KLSR05]. Seemingly [HDM02]. Segment [CSY00]. Segmentation [BCDL17, Joh13, NST12, QS09, YK01]. Seidel [GRM09]. Seismology [Bri94]. Select [ES02, Yat23]. Selected [VH22]. Selecting [GBS24, GKZ10, Rup02, ZMY^{+}23]. Selection [AB21, All13, AGV12, BBW20, BHvdLD22, BFM18, BHK15, BGP03, BMW17, CTC09, CDM^{+}20, CHH10, Cha22, CZSL10, CL12, CCYW16, CGLW21, Che23, CS14, CHSR19, CGJ09, CGL11, DYB^{+}23, DS21, DRG19, DBH^{+}22, EHW03, FK16, FS18, FN16, GD21, GTZH16, GHC14, GB12, HPOL14, HM09, HM11, HK22, HM15, Hel24, HY17, HML18, HRM02, HPZ22, HHKS11, HD23, HHTB10, Hur00, JPvD22, JS13, JBJ^{+}22, Jör09, KO04, KCF21, LGS09, LGS13, LL12, LD21, LZ23, LDR23, LPJ13, LD22, LW07, LDHW23, LQ19, LCD^{+}24, MYC^{+}22, MBM18, MD13, NY14, NG04, NL04, OPV22, Pan99a, PM20, RHO22, ROY20, RL14, SM19, SG20, SLL22a, SBZ13, SBS^{+}23a, ST99, SMK15, SHFT18, Su15, TVZ^{+}22, TRWB16, TSG92, TRS97, TV24, Tüc08, WDMW22, WWJ22, Wil00, XD20, XZ12, YPPM21, YKW03, YOL^{+}22, YWBA22, ZHZ12]. Selective [CJW22, SS22]. Selectivity [LG03]. Selector [MMRMGMGC08]. Selectors [YF22]. Self [CH16, CH93, HSC18, MM97, Tar99, XGTB18]. Self-Consistency [Tar99]. Self-Exciting [CH16]. Self-Multiset [HSC18]. Self-Organizing [MM97]. Self-Reported [XGTB18]. Self-Similarity [CH93]. Semi [BCH09, BK23, BFE^{+}03, CS20, CM08, DZ99, Gué03, KT03, SBJ17, TFS20]. Semi- [BCH09, KT03]. Semi-Adaptive [DZ99]. Semi-Complete [BK23]. Semi-Integrated [SBJ17]. Semi-Markov [Gué03, TFS20]. Semi-Streaming [BFE^{+}03]. Semi-Supervised [CM08]. Semi-Varying [CS20]. Semidefinite [WYZ19, ZRL14]. SEMIFAR [BF02]. Semiparametric [AG05, Bil00, BBvdW16, CKNK06, CHH10, CG07, CGJ09, Dan12, DLSS24, ERPG11, Guh08, HJ04, KHYD23, Len03, LG92, LBW14, MSBV22, MQ17, OR99, PM20, PTP03, SMS^{+}14, Sar22, ST99, SM13, SL10, TZ12, TAS^{+}17, TV24, WK09, WGDW22, ZRL14, ZL16, ZDM20]. Semismooth [YH17]. Semisupervised [Cul13]. Sensing [BFE^{+}03]. Sensitive [LZ18, Wan13]. Sensitivity [BLR22, CCHS08, MY94, VG01]. Separability [LL12, Tre95]. Separable [MMZ22]. Separation [Ish99, TZ12, ZKS18]. Sequence [DSD^{+}23, Hob08, MW04]. Sequences [Gué03, PQD22, QN00, QM04, UTK22]. Sequencing [Wan20]. Sequential [BK03b, DBS06, DMP14, FKH07, GF16, GP11, HT17, HS08, KCSA20, LJN^{+}17, MDC17, WGDW22, WDB96, WBW21, YSDJ15, ZNYJ14, ZJA16, vWB22]. Sequentially [WC06]. Seriation [EH15]. Series [Ano23b, BCL94, BM99, BRCC22, BH02, CH09, CCH16, CD14, DP23, Dan97, DGW09, Fea05, FCL10, GVBM04, Gri99, HMP96, JVSS24, KNH23, LRFK21, LMS19, MHG24, MVR02, NYC15, OBGC21, PVH05, PSY09, QGL09, SBE17, SH17, Sod20, ZHJ16, Ano07a]. Set [AG18, GD23, Kle04, SdWDL23, VH22]. Sets [Coz99, HMPP19, SK10, SH00, UHHS96, Weg95]. Setting [CM08, IL16]. Settings [HAPRD24, LDR23, Spa05, ZKS22b]. Setup [PKH15]. Several [BC93, HVRA96, LH95, SG05, WYZ19, YT99, Tuk93a]. SFO [Wic11a]. Shadings [ZMH07]. Shadow [LLM09]. SHAH [FT16]. Shall [KB99]. Shape [BB06, FIK19, GM99, GM93, Kle06, KW10, Mil19, NGH17, PA14, PWPS97, UKZ18, FT16]. SHape-Adaptive [FT16]. Shape-Constrained [PA14]. Shaped [DKH06, DCT08]. Shapes [SSG23, TS22]. Shared [NJM20]. Sharpening [BH01, CBS24]. Shift [Cai99, HB16]. Shingo [FB94]. Shocks [SG05]. Shorth [EGS10]. Shotgun [MR16, SC19]. Shrinkage [CBM22, DXL12, DV04, Efr05, Gao98, GH23, JWB13, JT21, LT98, LHL03, PJS^{+}22, RRG^{+}21, TV24, WP13, ZL16]. Shrinking [YT21b]. Shuffled [SDBD21]. Sided [HWLB24]. Sign [DJ04]. Signal [Hoe10, JYYZ18, TG10]. Signals [HCZMH21, SW23]. Signature [FCL10]. Significance [BRMC19, DM24b, GW04, GMC02, HLYM15, Jör09, SBL23]. Signs [VH15]. Silhouettes [RR22]. Similar [KK00]. Similarity [CH93, MBS04]. Simple [Bak92b, CJ02, CSWZ23, Res22, RTL07, SSP^{+}18, Bak92a]. Simplex [DSD^{+}23]. Simplified [JU00, LACV23]. Simulated [Glu00, Glu04, IP04, JB03, Sin02, SY07]. Simulating [Hes98, KKLN20, LD06, MM10, MWM^{+}23, Whi01]. Simulation [BMR15, BGL18, Bre04, BK05, Bro06, CP18, EM07, EM00, FHIT20, Gel95, GR99, GŠP^{+}06, GW11, GHW02, Guh08, Hof09, HHH12, JHD21, KGM18, LYG^{+}10, Len09, LJZD05, MA19, MHD20, RFVJ21, RZ22, SHL12, Ste02, Ste12, Vie00, WC94, dLG01]. Simulation-Based [HHH12, LJZD05, SHL12, dLG01]. Simulation/Prediction [MA19]. Simulations [AR06, BG98, FBG06, Hes95, ILP19, LZ18, Šev04]. Simultaneous [DRG19, GD16, HWLB24, Hel04, Hsu92, Jör09, Kwo98, LG03, LJZD05, MSBV22, RFVJ21, SHL19, YPPM21, ZSK19].

Simultaneously [DCV10]. SinaPlot

[SSP^{+}18]. Sine [VH15]. Single

[FXC21, HAPRD24, PVH05, PPTO22, SSP^{+}18, SN10, YZ07, ZHZ12, ZHJ16]. Single- [ZHJ16]. Single-Index [PVH05, PPTO22, YZ07]. Singular [CHPR24, GPPVW17, GR08, HJNR19, YMB14, ZG22, ZMSZ07]. Singularities [WL24]. Size [Kwo98, YSJS23]. Sizer [PLH10, HL06]. Sizes [GMF16, LFS^{+}17]. Sketch [WL22]. Skew [MMDL22, SYNK24]. Skew-Normal [MMDL22]. Skewed [Mah17, MWM^{+}23, MKK18]. Skewness [BHS04]. Slab [CPP23, DRG19]. Slack [FLY21]. Slice [HOD22, LCV20, LACV23, LL10, TGHL14, FW13]. Sliced [LYM23, Wu08, WLM10, ZYKC08]. Sliced-Wasserstein [LYM23]. Slings [PRS11]. Slope [FML19, JBJ^{+}22]. Smale [GRB^{+}13]. Small [BO05, FRW05, HY17, Pan99a, WK17]. Small- [HY17]. smashGP [EGCP24]. SMC [ER21]. Smooth [EM02, GHC14, Gua23, HY19, Kab93, KK09, PNLL21, SK24a, TL07, THM^{+}23, ZW07]. Smooth-Backfitting [HY19]. Smoothed [AZ22, KLH05, MPTZ23, PF23]. Smoother [CGH17, Kit96, MB18, SHMD^{+}14]. Smoothers [Gra96, Hel21]. Smoothing [BM99, Che07a, DS94, DS03, EH05, EHW03, FHTG05, GW93, HPA14, HM15, IJ02, LX16, LTW07, MYC^{+}22, NST12, OLN11a, OLN11b, PRHP14, QG06, RH94, SAT04, SL99, SW23, WL08, WK09, XW18, YB22, ZSBS23]. Smoothness [BK20]. Social [Eve12, HKS12, JL13]. Soft [HLYM15, SSL^{+}16]. Software [BB03, Cha00, CGR06, Gra07, HK17, KY95, Mar96, Saw96]. Solution [PI10, SWZ14, TLLJ23]. Solutions [SWX23]. Solving [Kat10, LF16, Paa99, Yen20]. Some [CKR19, CK08, LFZ97, SS11, Wai93, WM03, Wic13, ZL10]. Sourced [CCHM18].

Sources [GY07, PMI20]. Space [BK23, CLHW19, CH99, DZZ06, DZL22, EM18, FLMM22, GMC02, HLP07, HK98, Kit96, KC20, LXZ22, MMF^{+}19, MÉG^{+}23, MHS11, NFM16, NW10, PHK14, QG06, RNHY12, SKK22, Smi01, ZW07]. Space-Filling [CLHW19]. Space-Time [CH99, DZZ06]. Spaces [BK22, BLBCC22, HK10, LMW^{+}22, PB22, TRWB16, ZY24]. Sparse [BCS22, CCYW16, CR23, CZZ15, DZZ16, DXL12, DR10, DS94, DM24a, FXC21, Fro22, GTZH16, GJL^{+}10, GEH^{+}22, Hel21, JAJ19, KLZ23, LL22, LAML19, LW09, LS04, LSH16, LP23, LCWW17, LWZ14, LBCG16, MJVB18, MDC17, MC18, NT17, PP09, QLCZ15, QG22, RLZ10, SJM^{+}24, SLBV18, SSBW17, SFHT13, TDM12, TW14, WL09, WZ16, WZSF18, WL13b, YMB14, ZHT06]. Sparse-Group [LAML19, SFHT13]. Sparsely [MRP22]. Sparsification [LYXM23]. Sparsity [GDM14, GH20, KYW19, KKB21, LWLY17, TV24]. Sparsity-Inducing [GH20]. Spatial [Ano12b, BCST13, BS19, BM19, BC24, Cai99, CHZ23, CR99, CHH10, CRS06, Cli05, CM13, DG14, DKH06, DM24b, ESR^{+}14, ES02, EGCP24, FM10, FGN06, GHC14, GH18, HHC03, HC07, HCB04, HO99, HCG02, HHTB10, HS18, Hui22, HSZ23, IP04, KSM14, KKB21, KKHB23, LCRG21, Lia07, LGL16, LA12, MK20, Mah17, MP16, MLP07, MBY^{+}22, Moc98, MRCM18, NBH^{+}15, OW97, OM19, OS22, PTP03, RS19, RT21, RRG^{+}21, SLR21, SHMD^{+}14, SG17, Smi05, Ste92, SS16, WR05, WH17, YS10, Zha03, Zha22, ZZ08, ZY05, ZW10, ZH17]. Spatial-Temporal [ZZ08]. Spatially [Ano08a, BMC05, BS19, CRC^{+}07, GRM09, GAZ15, HCZMH21, KW21, PF23, RIC^{+}16, SLR21]. Spatio [CSK10, HHT21, HSG23, JK21, MM20]. Spatio-Temporal [CSK10, HHT21, HSG23, JK21, MM20].

Spatiotemporal [OS22, VF09]. SPDE [BK20, LGL16]. SPDEs [BSX23]. Speaks [RHS^{+}04]. Spearman [VT94]. Special [BKM99, Buj00, ES96, Edi00]. Species [ACG14]. Specific [Hes95, JVSS24, PP07, SH17, TvdV22, XNN16, BD22]. Specification [RHF24]. Specified [DZZ06, Vie00]. Spectra [ES19, KTS17, Whi01]. Spectra-Based [ES19]. Spectral [ALP^{+}23, BPW93, BRCC22, CRL13, DZ99, FM10, Hof19, HWF22, LRFK21, YPPM21, ZGC22, ZLPT22, vDE19]. Spectrally [Hel21]. Spectroscopic [SS12]. Spectrum [CHPR24]. Speech [BMO01]. Speeding [QTVK18]. SPEW [GRVE18]. Sphere [GM20, MR10]. Spheres [YLM17]. Spherical [CZ02, LFZ97, ZKS18]. Spherical-Radial [CZ02]. Spike [CPP23, DRG19]. Spike-and-Slab [CPP23, DRG19]. Spikes [LS04]. Spline [DS03, ERPG11, EHW03, GW93, HM15, KCO09, Koo97, LX16, LTW07, QG06, RH94, Wan08, WL08, WK09, WJL23, XW18]. Splines [Ano08a, AGV12, BMC05, Bil00, BHK15, CRC^{+}07, JPvD22, KCK08, LB04, Lin99, LGL16, MYC^{+}22, MS03, PA14, Pit02, Rup02, SL10, SW18, SW23, TCCSB10, Zha97, ZSBS23]. Split [JN04, LACRM06, LPQ^{+}22, TRC23]. Split-Merge [JN04]. Split-Plot [LACRM06]. Splitting [CL15, TRS97, Yen20]. Spotted [QS09]. Spread [McL99]. Spread-Location [McL99]. SpreadPlots [YVMFB03]. Spreadsheets [GPR99]. Square [PSRM^{+}23, SPV22, ZLW21]. Squares [CHW12, DG05, HB92, HGK10, LP07, Paa99, ZLP23]. SR [RIC^{+}16]. SR-HARDI [RIC^{+}16]. Stability [Hob01b, LY16, PRHS18, WWJ22]. Stabilization [RAC23]. Stable [CHG16, CMG21, CDFB11, FW13, HM15, HSZ23, LW19, LFT17, MA19, Mar10, WH98, WY15].

Stacked [DBH^{+}22]. Stage [ACCQ19, JB03]. Standardized [Hea22]. Stat [Tie00]. State [BK23, DZL22, FLMM22, HK98, Kit96, Kle22, KC20, MMF^{+}19, MÉG^{+}23, NFM16, NW10, QG06, Sco99, SG20, Smi01, ZW07]. State-Dependent [Kle22, SG20]. State-Space [HK98, NFM16, NW10, Smi01]. Statement [Ano13g, Ano23b]. Statements [Hel04]. States [SLL22a, ZLS23]. Stationarity [Hea23]. Stationary [CMG21, CP18, HPA14, HKP16, MW96, QGL09, SD03, Whi01, WC94, ZKS22a]. Statist [Bak92a]. Statistic [BM19, FBG06, Kou14, NP01]. Statistical [ABvS09, Cha00, CKT21, CKR19, Cle93a, Cle93b, DG24, Dri08, EFLT94, Fie12, FF94, FCL10, GU13a, GI00b, GL07, GKBP15, HIT99, Hob08, Hol93, HO99, HLYM15, HKS12, Hur93, IKL08, KP03, Lan00, LPS99, Lia07, Mic12, Mur13, PCP11, PRHS18, PGR15, PSTV15, QWW23, RTL07, Saw96, Šev04, SBL23, SP23, SR23, Ste13, SWC^{+}10, TS22, Tre93, VC21, VH17, Wil93, Wil96, YL95, Zhe17, ZICT12, dVTP^{+}17]. Statistically [KBM23, Owe17, SS16]. Statistician [GF15]. Statistics [Ano06a, Ano07a, Ano08a, Ano24a, Ano24b, CC08, DAMC23, DKH06, ES02, Fea02, GdTv96, HK17, Hub00, JO02, Lev12, Lev13b, LHC17, Mey03, MPC24, MKN23, OLN11a, PN22, RHS^{+}04, Sta95, Ste95, Wat11, WDB96, Ano12b]. Steck [Hut02]. Steinian [NT17]. Step [CDFB11, GH21, GR12b, LFS^{+}17, Lum19, MA19, Tad17]. Stepwise [FX21, XZ12]. STERGM [KMGJ23]. Stiefel [JHD21]. STIMA [DCV10]. Stimulus [PAP22]. Stochastic [AR22, ACIP22, BCE20, CSY00, CTW20, CJS92, DZL22, FCWM17, GJ09, Gol09, GM93, GX20, HB16, Jan06, JL13, KFSL17, KS99, MR16, MGF22, NTVK12, RAM19, SC08, SHBC^{+}24, Sin02, SK24a, WBS22, XZ12, YMLY11, ZCS^{+}19, ZGC22].

Stochastically [LKW09]. Stock [NAB16]. Stopping [GF16, SD03]. Story [BBKC24]. Straight [RHV22]. Strategies [CCS03, DBS06, JS20, LS05, RW14, SY02, SNR17, YK09, YL95]. Strategy [HR11, TL92, YM11b]. Stratification [PMB12]. Stratified [AR22, Cha24, MPS92]. Straw [DS21]. Stream [CCS03, Rho03, ZH17]. Streaming [BFE^{+}03, CY24, MDD^{+}12, THSM^{+}20, TKB23, WM03]. Streamlined [MDW23]. Streams [CCS03, DH03a]. Street [LHC17]. Strength [TW05, YHH^{+}23]. Stress [LPPT17]. Strong [HWS16, JP96, LB15]. Structural [Che07b, DG24, LLM23, MVR14, OS22, Par22, SBM22]. Structure [BCD^{+}24, BQS94, CL12, Far99, Gra13, HWQ15, IP04, KSM22, LCBV22, LH15a, MVR14, MH17, OM19, ONS18, OFR^{+}18, SBE17, SBK23, TRC23, WCH20b]. Structured [BT18, CC22a, CAPP23, DXL12, EC14, GH23, HHC03, KYW19, LWLY17, LZCZ22, MPC24, SYNK24, SH18, SHC00, SNC23, TV24, YL22, ZY05]. Structures [BD23, DY09, EM18, EM02, GR93, HL06, Kos22, LH15c, VHSG20, WL09, dVTP^{+}17]. Student [TEG19, Tsi02]. Student- [Tsi02]. Studies [BVZ20, CHLZ22, KSR19, Ros12, Ros17, SWC^{+}10, WZ22, YR22, ZSL23, ZDM20]. Study [HAPRD24, HK17, MM10, MK92, SSL^{+}16, WM03, Yat23, YT99]. Studying [Cle93a, Cle93b, Hol93, MWM^{+}23, SPW95, Tre93, Wil93]. Style [LZYQ22]. Sub [Gou98, JVSS24]. Sub-Families [Gou98]. Sub-National [JVSS24]. Subchains [Hob01b]. Subdata [Cha22]. Subgraph [MOPW20]. Subgraphs [ZC15]. Subgroup [PW22]. Subject [BH01, CCHM18, MBY^{+}22, ZZF24]. Subject-Level [MBY^{+}22]. Subordination [BMG16]. Subpixel [GJ97]. Subpopulations [YHH^{+}23]. Subregion [GK97]. Subregion-Adaptive [GK97]. Subsample [ZMY^{+}23]. Subsamples [MKN23]. Subsampling [Bra21, Cha24, Dem23, GVDP99, JP95, MXM^{+}21, QTVK18, QTV^{+}21, ZZWS23, ZWZW24, ZWS23]. Subset [BG92, DM22, FEF22, GK06, Ros12, SSMN16, Wil00]. Subsets [NJM20]. Subspace [GR12a, HHH22, PSY09]. Subspaces [RFV23, SYL^{+}22, WBW21]. Substitution [Hob08, OSR^{+}18, SC92]. Subtle [Su15]. Successive [SC92]. Suffer [PRS11]. Sufficiency [HR11, YM11b]. Sufficient [Fea02, HY17, JO02, SY16, XY14]. suggested [Tuk93a]. Sum [LW19, LJ19]. Summarization [WCM21]. Summary [BM19, DAMC23, GMF16]. Sums [Hea22, WK17]. Sunbeams [FW11]. Superheat [BY18]. Supervised [Cha24, CTDKL21, CM08, LCKL05, LSH16, PRHS18, ZSK23, dSCL21, SLL22b]. Support [BL04, BLR22, GTZH16, LL22, LSD05, LY11, SWZ14, SC98, TZ06, WZW19, WCH20b, WL13b, ZLWZ16]. Surface [CDP04, Fea04, PPTO22, SWZ14]. Surface-Link [PPTO22]. Surfaces [DS03, MRP22, OW97, QY97, Qiu02, Ste02, SQ07]. Surrogate [CWZ21, Gel00, GH00, HL00b, Kie00, LHY00, Men00, Wu00, dLM00]. Surrogates [Dri08]. Surveillance [FDG^{+}22]. Survey [BD22, MWWY24, WS24]. Survival [Ano20, BCH22, DASR08, DMR20, DH03b, Gu15, HYL^{+}17, JWL^{+}23, KS22, Koo98, LRBS03, LKW09, LH00, SJM^{+}24, TGP03, WGSC24, WW19, WCSY21, XGTB18, YSJS23, ZCI^{+}17]. Swendsen [BZ07, NG04]. Switching [LF16, PI10, RW14]. Symbolic [LRB12, SJSG22]. Symbols [Rob19, Tre95]. Symmetric [Cai99, Ho06]. Symmetry [BCH05, GM15, LFZ97]. Symmetry-Constrained [GM15]. Synthesis [HPA14, HKP16, MW96]. Synthetic [ASND19, FD21, GRVE18, PDM23, PDLN18, PSF^{+}22]. System [CCS03, Rap01, Tre06, ZLW21, SCB98]. Systems [Buj00, CVQ23, ES96, Edi00, GŠP^{+}06, Mar96, MDD^{+}12, WBQ20].

Table [Paa99]. Table-Driven [Paa99]. Tables [BL22, Bri11, Che07b, EC17, Fis12, Fri00, FK11, Gel11a, Gel11b, HVRA96, HM02, LH95, LPM98, Sta95, Wai11, Wil11, ZBY09, VT94]. Tail [Dia99, Hes95, Hsi99, WK17, dSM04]. Tail-Specific [Hes95]. Tailed [Mah17]. Take [PRS11]. Tale [Wic11a]. Tapered [BH22, DASR08, SR12]. Tapering [AHNC13, Ano12b, FGN06]. Tapers [Ste13]. TARGET [FG05]. Targets [SGH17]. Task [LZ23, WL23]. Taylor [Sin02]. Teach [Zhe17]. Teaching [Zhe17]. Technique [ACCK19, BCL94, JU00, JTU03, SPW95, TG10]. Techniques [Kei96, RB02, WM03]. Technology [WS02a, GdTv96]. Telephone [Mci99]. Tempering [MMV13]. Template [MBY^{+}22]. Temporal [CSK10, GHCU22, HHT21, HSG23, JK21, MM20, PQD22, THSM^{+}20, WCH20b, ZZ08]. Tendencies [TvdV22]. Tends [LCLR23]. Tensor [CKT21, HLW22, LM23, LSLX22, Loc18, LXZ22, MPS17, MMZ22, PAP22, VLNO18, XZYZ24, ZZSW23]. Tensor-on-Tensor [Loc18]. Tensor-Valued [VLNO18]. Tent [BW17]. Terminating [Tu94]. Terrace [MGL18]. Terry [CD12]. Test [DG14, DW02, DGWR24, FKH07, FSR20, GPGMFB14, HM14, HSG23, KN06, LFZ97, MKN23, SPV22, SLR16, TVZ^{+}22, TSG92, WWG^{+}97, VT94]. Test-Based [TSG92]. Testing [AvD20, BMG16, Bie05, BJG98, CHLZ22, DM24b, GCKP08, GH20, HC07, KMS23, LP07, MOPW20, RHF24, SŠS07, TAS^{+}17, VH17, ZK22, ZWF^{+}22]. Tests [AR06, BBM24, BRMC19, Can99, DJ04, FKH07, Gri99, HVRA96, HB98, LCTV24, LMW^{+}22, MPS92, MH17, Res22, Shi24, WFCT23, WLWM16]. Text [CH93, Eic94]. Thank [XMY13]. Their [LS93, LB15, GS24]. Them [PRS11]. Theoretic [CR24, GKZ10]. Theoretical [GR12b, JS95]. Theory [AR06, CKR19, Jör09, Shi24]. There [SYF17]. Thinning [Owe17]. Thomson [BPW93]. Thousand [LCTV24]. Three [BB06, BW17, EO95, FSMV02, LACRM06, MM14, ZDM22]. Three-Dimensional [BB06, EO95, FSMV02, LACRM06, MM14, ZDM22]. ThrEEBoost [BMW17]. Threshold [Fon19, SM19]. Thresholded [BMW17]. Thresholding [CKR19, EH09, HLYM15, ZL16]. Thyself [GF15]. Tidy [WCH20b]. Tie [KMGJ23]. Tight [SJSG22]. Tilted [LP14]. Time [Ano07a, Ano23b, BCH22, BCL94, BM99, BRCC22, BH02, BHS00, CMN21, CHW12, CH09, CSWZ23, CCH16, CH99, CD14, DP23, DZZ06, Dan97, Dem23, DLSS24, DGW09, EM18, Fea05, FCL10, GVBM04, Gri99, HMP96, HMW06, HYL^{+}17, HCK^{+}11, JVSS24, KK09, KNH23, KLH05, LRFK21, LLM23, LG20, LMS19, LBW14, MA19, MHG24, MVR02, MDD^{+}12, NYC15, OBGC21, PSY09, PRCF24, QGL09, SBE17, SH17, Sod20, VH22, YP20, YL99, YLM17, ZHJ16, ZW07, vDE19]. Time-Dependent [LLM23]. Time-Dynamic [HMW06]. Time-Frequency [QGL09]. Time-of-Day [YL99]. Time-Varying [CHW12, HYL^{+}17, LG20, YP20, ZW07, vDE19]. Time-Warped [YLM17]. Times [AvD20, BK05, Lev11, PSRM^{+}23, Tu94]. ’tis [PRS11]. Told [BBKC24]. Tolerance [ALP^{+}23, HWLB24]. Tool [DPS11, HH11, HH01, MD13, MS93b, OM19, WZLR21, Yat23, YKZ18]. Toolkit [SRL^{+}00]. Tools [CGEGMI22, CH12, GEGM03, Pos95, WGSC24]. Topic [MJVB18]. Topics [CD19]. Topological [BM19, MGL18, PB22, ZHZW22]. Torgegram [ZH17]. Torus [GIA22, ZGPM22]. Total [BDR97, SM19, YT21a, YNTK21]. Tour [CBCH95, LCV20]. Tours [HO10, LACV23]. Trace [ACJ21, Fri13b, Mul98]. Tracking [HMW06, YSK^{+}10]. trackr [BML19]. Tradeoffs [GU13b]. Traffic [EO95, Wic11a]. Trained [Cul13]. Training [BMO01]. Trajectories [ZKS18]. Transdimensional [LGS09]. Transfer [Gel00, GH00, HL00b, Kie00, LHY00, LZYQ22, Men00, Wu00, WY22, dLM00]. Transform [Bau19, Big06, GGT22, HCB04, NS94, Yan24]. Transformation [DT13, GW18, HRM02, HZ21, HDK^{+}20, KS08, MHH15, NLF^{+}06, SLMN20, YT21b, YZ06]. Transformations [PSF^{+}22, RA01, RAC23]. Transformed [ES98, MHS11]. transforming [TYG^{+}20]. Transition [HK22, ZKS22a]. Transmission [GF22]. Transport [OKL21, ZMY^{+}23]. Transposable [TW14]. Treatment [CBM22, DZFZ18, LSFI18, WY22, XTPO22]. Treatments [BVZ20]. Tree [BT18, CS20, CGR17, DCV10, EC14, EH09, FG05, HMJ11, JMY^{+}16, KCB19, KSLS22, MW97, MS93b, NBNA21, PW22, PGM23, RR22, SSB^{+}14, SN10, VHSG20, WS02b, WB18b, YL22, ZY05]. Tree-Based [CS20, DCV10, EH09, KSLS22, RR22, VHSG20]. Tree-Structured [BT18, EC14, YL22, ZY05]. Treed [AG96, STY06]. Trees [BM19, BOW02, CH12, DZFZ18, DCS16, IW22, JHL09, KL03, Kle04, Kle06, Kle08, KM03, LT98, Mey18, MW04, Ooi02, PCG^{+}14, PCGM20, SZ08, Seg94, SWF04, TG09, YL99]. Trellis [BCS96a]. Trend [HZC23, RT16, VH17]. TREX [BGLM18]. Trial [HS08]. Trials [WC06, ZCI^{+}17]. Triangular [GY22]. Trimmed [HGK10, LMM19, ZBS^{+}21]. Trimming [GEGM03]. Troubles [PRS11]. Truncated [DW01, FKH07, GLC20, Ion08, KR17, LJ19, MMDL22, PP14, Ste92, Zho05]. Truncation [JL13]. Truthful [SSP^{+}18]. Try [YLCR19]. Tube [LS97]. Tukey [BC93, FLMN23, HP94, LMM19, Wai93]. Tumors [Dri08]. Tuned [QTV^{+}21]. Tuning [GB12, WLZB07]. Turing [DW02]. Turning [BMR15]. Turning-Band [BMR15]. TV [MBGO11]. Tweedie [QYZ16]. Two [ACCQ19, ACIP22, BBM24, BCDL17, BM99, BG00, CDFB11, GB00, HWLB24, JB03, LT94, LR04, Mül99, PKH15, Pos95, Sco99, SQCK20, SYW^{+}22, SWZ14, SLR16, TAS^{+}17, TN14b, Wic11a, YL22, ZK22, ZWF^{+}22, ZP14]. Two-Block [PKH15]. Two-Dimensional [Pos95, SYW^{+}22, SWZ14, ZP14]. Two-Group [ZK22]. Two-Level [SQCK20, TN14b, YL22]. Two-Sample [BBM24, GB00, SLR16, TAS^{+}17, ZWF^{+}22]. Two-Sided [HWLB24]. Two-Stage [ACCQ19, JB03]. Two-State [Sco99]. Two-Step [CDFB11]. Two-Way [BCDL17, BG00]. Type [DGS24, LPQ^{+}22, LMW^{+}22, TGM^{+}14, ZHZ12, vD00, vlM97]. Types [SG05, WCDS03].

U.S. [EO95]. Ultra [HMZ22, LDR23, SK24b, WFCT23]. Ultra-Efficient [SK24b, WFCT23]. Ultra-Fast [HMZ22]. Ultra-high [LDR23]. Ultrahigh [LLW^{+}22, RSDG23, XL17]. Ultrahigh-Dimensional [LLW^{+}22, RSDG23, XL17]. Unbiased [HHKS11, HHZ06, KPC02]. Unbounded [Whi01]. Uncertainties [MBS04]. Uncertainty [BRS^{+}08, BL17, DMP14, God01, HDM02, LHL21, RMR^{+}17, RSKZ19, WB18b, XBW15, YF22]. Unconstrained [GG17]. Uncorrelated [KLZ23]. Under-Dispersed [AC95]. Understand [WKT^{+}24]. Understanding [Mar94, SWC^{+}10]. Unified [KGM18, KCB19, LDHW23, MPTZ23, RA01, WZLR21]. Uniformity [HC07]. Unimodal [Ho06]. Unimodality [CGH99].

Unit [BD22, Gri99]. Unit-specific [BD22]. Univariate [CLSC22, KC20]. Universal [DGWR24, LQ23]. University [GdTv96]. Unknown [DZL22, RDL07, ZLS23]. Unleashed [BB03]. Unnormalized [FGH^{+}24]. Unrelated [HDM02]. Unspecified [HWQ15]. Unsupervised [Kel11, LKR22, SH06, Wu11]. Unwin [Kos13]. Upcoming [KB99]. Updated [BS19]. Updating [LW19, WH98, WCSY21]. Uprooting [Su15]. USA [HCK^{+}11]. Usage [HAC^{+}12, Rho03]. Use [BBLG13, HWLB24, LO16, MBS04, NLF^{+}06, SP23, ZH22, ZW04]. Useful [BM19]. User [CC22a, SH03]. User-Friendly [CC22a]. Users [Cha00]. Using [Ano07a, AGV12, ACG14, AHS22, Bak92b, BGD16, Béd08, BB21, BAMG22, BRS^{+}08, Bra02, CTDKL21, CF05, CZ02, CVQ23, CGR06, DR18, DSD^{+}23, DT13, DYB^{+}23, DG05, DH03b, DH10, DJ04, ES02, ES98, FG05, FLY21, Far04, FKH07, Fea05, FHIT20, FSR20, Fos01, FN16, FF94, Fri00, FP20, Gao98, Gel95, Gel00, GvDHB08, GV01, GAS20, Gho07, GVDP99, Glu00, Gou98, GB04, Gra96, GH00, HPOL14, HM09, HHH22, HB92, HPA14, HCB04, Hof19, HZ21, HLYM15, HL00b, HMZ22, HB16, IP04, JRS11, JHL09, JL13, JS20, JP22, KH21, Kan20, Kar21, KSR19, KK09, KH23, Kie00, KCSA20, KSC^{+}24, KSM14, KW10, LHL21, LHY00, Len03, LC24, LRBS03, LYXM23, LCSL14, Lia98, LPM98, LGL16, LHC17, LZCZ22, LSFI18, LPPT17, MJ18, MWFSG17, MN97, MOPW20, Men00, MYC^{+}22, MBM18, MR01]. Using [MHS11, MGF22, NP01, NJM20, Ooi02, PP96, PA14, PLH10, PLW01, QS09, RNHY12, RHF24, RR23, RDL07, RG09, STM15, SG03, SBR22, SMB14, SS12, SK10, SR00, Sha14, SG23, SJM^{+}24, ST99, Smi01, SS06, SH00, SBJ17, SHL12, SH97, SLR16, SQ07, TSH16, TGM^{+}14, Tüc08, WLZB07, WZ22, Wu00, XBW15, YASR20, Yan21, YT21b, YSDJ15, YWL^{+}15, YNTK21, ZB10, ZBY09, dLM00, vD00, Bak92a, WL23]. Utility [LYG^{+}10].

Vaccine [LBM^{+}14]. Valid [CB20, GMdL23, KBM23, NM98]. Validated [BHvdLD22]. Validating [Gu92, GX01, Zha03]. Validation [CSS^{+}09, CLXY23, CGR06, DPRK23, FP20, GS24, Gv97, GX01, KO04, KPC02, LY16, LHL03, MMN^{+}13, SD03, TSG92, TW05, XSC19, YF14, vdWvNR21]. Validations [YK09]. Validity [WW12]. Value [BCH05, BR02, HWK17, Hof19, HJNR19, WFCT23, YMB14, ZG22, ZMSZ07, HB98]. Valued [BVZ20, CP18, JB22, LRB12, LD21, MR19, VLNO18, WK17]. Values [HWLB24, Hea22, HGK10, LKB21, NP01, SY02, UHHS96]. VAR [SBM22]. Variability [WSB24]. Variable [AB21, AGV12, BBW20, BCDL17, BRS11, BMW17, CTC09, CDM^{+}20, CGLW21, Che23, CS14, CG07, CR24, CGL11, DRG19, DBH^{+}22, ES98, Fal99, FLY21, FK16, FN16, FCL10, GHC14, GB12, HPOL14, HM09, HM11, HPQ24, Hel24, HY17, HRM02, HD23, HMS97, HWO^{+}17, Hui22, IPH22, JS13, JCZG19, LGS09, LGS13, LL12, LG03, LD21, LZ23, LDR23, LP23, LD22, LW07, LZCZ22, LDHW23, LCD^{+}24, MB04, MR01, NY14, NG04, NL04, OPV22, PQM22, PM20, RL14, SZ08, SLL22a, SBZ13, SL10, Su15, TRWB16, TSG92, TRS97, Tüc08, WRT21, XZ12, XCSP16, YKW03, YOL^{+}22, YWBA22, YF22, ZHZ12]. Variable-Dimension [HPQ24]. Variable-Domain [JCZG19]. Variables [CBM24, CBL20, DR18, FDBS10, GGT22, GW18, GKZ10, Hoo07, LGS09, LJZD05, RT21, Sch24, TU11, WSM^{+}20, WK17, ZBB06, ZSW22]. Variance [BG00, CKNK06, Fos01, GVBM04, GH21, GCKP08, HM15, LACRM06, Lia98, LTW07, OY23, RAC23, SHBC^{+}24, SHL12, STY06, YNTK21, ZHZL19, ZHZW22]. Variance-Reduced [SHBC^{+}24]. Variances [Par22]. Variate [TMMB20, TGPM22]. Variates [AC95]. Variation [BB06, BDR97, GMF16, MS18a, SM19]. Variational [AR22, BBvdW16, FLMM22, FLMMK24, GKN23, HAPRD24, HWO^{+}17, HMZ22, KDM16, KBM23, LMS19, MPR^{+}12, MDW23, NTVK12, ONS18, OW12, OPV22, SYNK24, SLMN20, SLM22, TN14a, TF20a, TNKK16, TNK17, WL24, WM19, YOL^{+}22, ZZW21, ZGTG21]. Variations [DM09, YT21a]. Variogram [Zha03, ZW04]. Variography [ZH17]. Various [HO99]. VARMA [McE22]. Varying [AGV12, BS19, CHW12, CS20, DXL12, GRM09, HYL^{+}17, JL13, KW21, LG20, SBR10, WH14, WJL23, WKT^{+}24, YP20, ZW07, vDE19]. Varying-Coefficient [AGV12, DXL12, WH14]. Vecchia [SCG22, HSZ23]. Vecchia-Approximated [SCG22]. Vector [Bra02, DZZ16, GTZH16, Hea23, HTW22, LL22, LSD05, LY11, SWZ14, TZ06, WZW19, WL13b, ZLWZ16, ZH05]. Vectors [AG18, BMO01]. Verification [Len03]. Versatile [Hel24]. Versus [Fu98, GBW17, Sin02]. Vertex [LW08, ML08, ZLPT22]. Very [HM09, Kei96, MK20, Par18, Wil99]. Via [AZ22, IW22, JPvD22, LMC22, MM97, OPV22, PL22, RRG^{+}21, SLR21, SLL22a, Su15, TRS97, WGDW22, AX23, All13, ACC18, Big06, BLBCC22, BH22, BFM18, BHF02, BG00, BMW17, BC19, CGEMIR18, CL12, CD19, CGLW21, CBL20, DV04, Dem23, DNF23, EGCP24, EB02, FBG06, FDBS10, Fis12, GHC14, HL03, HK06, HHC03, Hel21, HZC23, Ho06, HPZ21, Hof00, HDRM05, HW21, HWF22, HL00a, HO10, Hut02, JHD21, Jör09, KHYD23, Kou14, KKB21, LWLY17, LYM23, LPHK16, LKW09, LH15b, LW07, LLZ09, LLZ23, LQ23, MLP07, MPS21, ML08, MQ17, OY23, PCP11, PMB12, PN22, Pic14, PTP03, PRHP14, PCGM20, RSKZ19, RYY00, RMWH17, RAC23, RF17b, Seg94, SL99, SY16, SJM^{+}24, SRE06, TEG19, TSG92, VG01, WR05, WJL23, WLZ^{+}24, WL24, WNA20, WK17, WYZ19, Yen20, YSK^{+}10, YC04, YHS23, ZK22]. via [ZZWS23, ZZSW23, ZLPT22, ZHZW22, ZYZ10]. Video [RRH18]. View [BM19, YG22]. Viewpoint [JRH22]. Views [Fri99, GM16, Hur00, FB94]. Vine [XC22]. Vines [MC18]. Virginia [PSDK22]. Visual [BCS96a, EK02, EFLT94, GMF16, GB04, HH11, Hof01, Mar94, PAP22, Tre95, WS02a, YL95]. Visualization [BCS96b, BSL^{+}08, DG18, DGW09, EH15, FNGW20, HMPP19, HSG23, HDK^{+}20, HOD22, Kei96, Kle04, Kle06, LCL22, LK02, LRB12, LWM21, MHG24, McL99, MS93b, Mou09, MHS11, MW04, MM97, Ooi02, SH18, SLR^{+}23, SNB^{+}01, SPW95, SRL^{+}00, SCB98, TS13, WNA20, Wil99, XW20, YKZ18, YF22, ZMSZ07, ZL22, ZA23, ZDM22]. Visualizations [BML19, Hur04, LTW^{+}10]. Visualize [DPS11]. Visualizing [BY18, BM99, BEM13, ES19, Fri13b, GRG10, GKBP15, GW15, GHCU22, HBG96, IPH22, LA12, RFVJ21, Sch24, SJ05, TvdV22, TL99, Tro05, WCH20a, WL02, Wic11c, ZMH07]. Visually [YHH^{+}23]. Visuals [Goo11]. Volatility [Gol09, KFSL17, NST12, SM13]. Volume [Ano92g, Ano93f, Ano94g, Ano95f, Ano96h, Ano97f, Ano98f, Ano00g, Ano08c, Ano09b, Ano10b, Ano11b, Lev11, Lev12, Lev13b, WR93]. Voyage [Saw96].

W [BC93]. Wainer [Tuk93a]. Wang [BZ07, BJMD13, NG04]. Warp [MS02]. Warped [YLM17]. Wasserstein [LZ18, LYXM23, LYM23, YSG24]. Wavelet [Big06, BH02, DV04, Efr05, FMP24, GVBM04, Gao98, KNH23, LHL03, MAJ^{+}98, MW96, NS94, PVH05, SBT00, ZOR12, ZCO15]. Wavelet-Based [BH02, ZOR12, ZCO15]. Wavelets [FT16, MV98, SAT04]. Waves [ES19]. Way [BCDL17, BG00, Paa99, SD03, Sta95]. Weak [LD22]. Weakly [KLR18]. Web [SH03, SVC^{+}19]. Weber [dLM04]. Weighted [All13, CGEGMI22, CGR17, DAMC23, GAZ15, GEH^{+}22, HLD^{+}13, LMST18, Loc22, MWM^{+}23, SWZ14, SWL23, SLH24, Tan17, WZ16, WCKL19, WL13a, XTPO22, ZCO15, vdBBD22]. Weighting [GR12b, OS22]. Weights [Ano23b, BD22, SG03, Sod20, WS24]. West [PSDK22]. Whether [PRS11]. While [BCD^{+}24]. Whitening [PSF^{+}22]. Wide [KGM18, SAT04]. Widths [GVBM04]. Wiener [ZSBS23]. Will [HV17, JP96]. William [Tuk93a]. Window [SCB98]. Windows [HPA14, LA12]. Winds [ES19, HMN^{+}03]. Wise [CRGS17, CCFM01, GW93, LYLL23, SLL22a, YLCR19]. Wishart [EM00, LD11, vdBBD22]. Within [Ano07a, Fea05, VJ15, vlM97]. without [HLW24, Zha22]. Working [CL12]. Workshop [Hub99]. World [GRVE18, WY22]. Worth [LCTV24]. Wrapped [GIA22]. Writing [dVTP^{+}17].

X [BRCC22, SCB98]. XGBoost

[BCH22, DL23]. XGobi [SCB98, SCLK^{+}00].

Year [VRH19]. Years

[Don17, HJ17, Hub99, Pen17].

Zag [Kos22]. Zelterman [Sta95]. Zero [DS21, GCKP08, MP16, MCB24]. Zero-Inflated [MCB24]. Zeros [Che07b, MVR14]. Zig [Kos22]. Zig-Zag [Kos22].

References

Avagyan:2017:IGL

[AAN17] Vahe Avagyan, Andrés M. Alonso, and Francisco J. Nogales. Improving the graphical lasso estimation for the precision matrix through roots of the sample covariance matrix. Journal of Computational and Graphical Statistics, 26(4):865–872, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anderson-Bergman:2017:EIE

[AB17] Clifford Anderson-Bergman.

An efficient implementation of the EMICM algorithm for the interval censored NPMLE. Journal of Computational and Graphical Statistics, 26 (2):463–467, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Alexopoulos:2021:BVS

[AB21] Angelos Alexopoulos and Leonardo Bottolo. Bayesian variable selection for Gaussian copula regression models. Journal of Computational and Graphical Statistics, 30 (3):578–593, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1840997.

Antunes:2021:SBE

[ABG^{+}21] Nelson Antunes, Shankar Bhamidi, Tianjian Guo, Vladas Pipiras, and Bang Wang. Sampling based estimation of in-degree distribution for directed complex networks. Journal of Computational and Graphical Statistics, 30 (4):863–876, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1873143.

Antoniadis:2009:MAS

[ABvS09] Anestis Antoniadis, Jérémie Bigot, and Rainer von Sachs. A multiscale approach for statistical characterization of functional images. Journal of Computational and Graphical Statistics, 18(1):216–237, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ahn:1995:GDU

[AC95] Hongshik Ahn and James J. Chen. Generation of over-dispersed and underdispersed binomial variates. Journal of Computational and Graphical Statistics, 4(1):55–64, March 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474665.

Amiri:2018:CCD

[ACC18] Saeid Amiri, Bertrand S.

Clarke, and Jennifer L. Clarke. Clustering categorical data via ensembling dissimilarity matrices. Journal of Computational and Graphical Statistics, 27(1):195–208, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Amiri:2019:GHC

[ACCK19] Saeid Amiri, Bertrand S.

Clarke, Jennifer L. Clarke, and Hoyt Koepke. A general hybrid clustering technique. Journal of Computational and Graphical Statistics, 28 (3):540–551, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Armero:2019:TSB

[ACCQ19] Carmen Armero, Stefano Cabras, María Eugenia Castellanos, and Alicia Quirós. Two-stage Bayesian approach for GWAS with known genealogy. Journal of Computational and Graphical Statistics, 28(1):197–204, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Argiento:2014:DBA

[ACG14] Raffaele Argiento, Andrea Cremaschi, and Alessandra Guglielmi. A “density-based” algorithm for cluster analysis using species sampling Gaussian mixture models. Journal of Computational and Graphical Statistics, 23(4):1126– 1142, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Arza:2022:CTS

[ACIP22] Etor Arza, Josu Ceberio, Ekhiñe Irurozki, and Aritz Pérez. Comparing two samples through stochastic dominance: a graphical approach. Journal of Computational and Graphical Statistics, 32 (2):551–566, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2084405.

Ahn:2021:TRO

[ACJ21] Jeongyoun Ahn, Hee Cheol Chung, and Yongho Jeon. Trace ratio optimization for high-dimensional multi-class discrimination. Journal of Computational and Graphical Statistics, 30(1):192– 203, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2020.1807352.

Anitescu:2017:IFE

[ACS17] Mihai Anitescu, Jie Chen, and Michael L. Stein. An inversion-free estimating equations approach for Gaussian process models. Journal of Computational and Graphical Statistics, 26(1):98–107, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1164056.

Adewale:2010:BCB

[ADY10] Adeniyi J. Adewale, Irina Dinu, and Yutaka Yasui. Boosting for correlated binary classification. Journal of Computational and Graphical Statistics, 19 (1):140–153, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Aragon:1992:CCM

[AE92] Jorge Aragón and David Eberly. On convergence of convex minorant algorithms for distribution estimation with intervalcensored data. Journal of Computational and Graphical Statistics, 1(2): 129–140, June 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477009.

Alexander:1996:TR

[AG96] William P. Alexander and Scott D. Grimshaw. Treed regression. Journal of Computational and Graphical Statistics, 5(2): 156–175, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474702.

Amblard:2005:EPS

[AG05] Cécile Amblard and Stéphane Girard. Estimation procedures for a semiparametric family of bivariate copulas. Journal of Computational and Graphical Statistics, 14(2):363–377, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Azzimonti:2018:EOP

[AG18] Dario Azzimonti and David Ginsbourger. Estimating orthant probabilities of high-dimensional Gaussian vectors with an application to set estimation. Journal of Computational and Graphical Statistics, 27(2):255–267, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Antoniadis:2012:VSV

[AGV12] Anestis Antoniadis, Irène Gijbels, and Anneleen Verhasselt. Variable selection in varying-coefficient models using P-splines. Journal of Computational and Graphical Statistics, 21(3):638–661, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anderes:2013:NPD

[AHNC13] Ethan Anderes, Raphaël Huser, Douglas Nychka, and Marc Coram. Nonstationary positive definite tapering on the plane. Journal of Computational and Graphical Statistics, 22(4):848–865, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ashouri:2022:FFR

[AHS22] Mahsa Ashouri, Rob J.

Hyndman, and Galit Shmueli. Fast forecast reconciliation using linear models. Journal of Computational and Graphical Statistics, 31(1):263– 282, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1939038.

Allen:2013:AFS

[All13] Genevera I. Allen. Automatic feature selection via weighted kernels and regularization. Journal of Computational and Graphical Statistics, 22 (2):284–299, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Athreya:2023:NTS

[ALP^{+}23] Avanti Athreya, Zachary Lubberts, Carey E. Priebe, Youngser Park, Minh Tang, Vince Lyzinski, Michael Kane, and Bryan W. Lewis. Numerical tolerance for spectral decompositions of random matrices and applications to network inference. Journal of Computational and Graphical Statistics, 32 (1):145–156, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2082972.

Afshartous:2007:DMM

[AM07] David Afshartous and George Michailidis. Distributed multilevel modeling. Journal of Computational and Graphical Statistics, 16(4):901–924, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:1992:C

[Ano92a] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 1(2):180, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477013.

Anonymous:1992:EBPa

[Ano92b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 1 (1):??, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474571.

Anonymous:1992:EBPb

[Ano92c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 1 (2):??, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477007.

Anonymous:1992:EBPc

[Ano92d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 1 (3):??, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474579.

Anonymous:1992:EBPd

[Ano92e] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 1 (4):??, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474585.

Anonymous:1992:EC

[Ano92f] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 1(4): 385–386, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474592.

Anonymous:1992:IV

[Ano92g] Anonymous. Index to volume 1 (1992). Journal of Computational and Graphical Statistics, 1(4): 387–388, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474593.

Anonymous:1993:EBPa

[Ano93a] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 2 (1):??, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474594.

Anonymous:1993:EBPb

[Ano93b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 2 (2):??, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474603.

Anonymous:1993:EBPc

[Ano93c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 2 (3):??, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474609.

Anonymous:1993:EBPd

[Ano93d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 2 (4):??, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474615.

Anonymous:1993:EC

[Ano93e] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 2(4): 421–422, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474624.

Anonymous:1993:IV

[Ano93f] Anonymous. Index to volume 2 (1993). Journal of Computational and Graphical Statistics, 2(4): 423–424, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474625.

Anonymous:1994:C

[Ano94a] Anonymous. Correction. Journal of Computational and Graphical Statistics, 3(2):234, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474640.

Anonymous:1994:EBPa

[Ano94b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 3 (1):??, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474626.

Anonymous:1994:EBPb

[Ano94c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 3 (2):??, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474633.

Anonymous:1994:EBPc

[Ano94d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 3 (3):??, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1080/10618600.
                                        1994.10474641.

Anonymous:1994:EBPd

[Ano94e] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 3 (4):??, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474648.

Anonymous:1994:EC

[Ano94f] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 3(4): 473–474, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474659.

Anonymous:1994:IV

[Ano94g] Anonymous. Index to volume 3 (1994). Journal of Computational and Graphical Statistics, 3(4): 475–476, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474660.

Anonymous:1995:EBPa

[Ano95a] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 4 (1):??, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474661.

Anonymous:1995:EBPb

[Ano95b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 4 (2):??, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474667.

Anonymous:1995:EBPc

[Ano95c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 4 (3):??, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474673.

Anonymous:1995:EBPd

[Ano95d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 4 (4):??, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474679.

Anonymous:1995:EC

[Ano95e] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 4(4): 344–345, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474689.

Anonymous:1995:IV

[Ano95f] Anonymous. Index for volume 4(1995). Journal of Computational and Graphical Statistics, 4(4): 346–347, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474690.

Anonymous:1996:Ca

[Ano96a] Anonymous. Correction. Journal of Computational and Graphical Statistics, 5(1):112, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/abs/10.1080/10618600.
1996.10474698.

Anonymous:1996:Cb

[Ano96b] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 5(4):415, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474721.

Anonymous:1996:EBPa

[Ano96c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 5 (1):??, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474691.

Anonymous:1996:EBPb

[Ano96d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 5 (2):??, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474699.

Anonymous:1996:EBPc

[Ano96e] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 5 (3):??, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474705.

Anonymous:1996:EBPd

[Ano96f] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 5 (4):??, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474714.

Anonymous:1996:EC

[Ano96g] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 5(4): 416–417, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474722.

Anonymous:1996:IV

[Ano96h] Anonymous. Index to volume 5 (1996). Journal of Computational and Graphical Statistics, 5(4): 418–419, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474723.

Anonymous:1997:EBPa

[Ano97a] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 6 (1):??, 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474724.

Anonymous:1997:EBPb

[Ano97b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 6 (2):??, 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474733.

Anonymous:1997:EBPc

[Ano97c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 6 (3):??, 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474740.

Anonymous:1997:EBPd

[Ano97d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 6 (4):??, 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474747.

Anonymous:1997:EC

[Ano97e] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 6(4): 481–482, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474755.

Anonymous:1997:IV

[Ano97f] Anonymous. Index to volume 6 (1997). Journal of Computational and Graphical Statistics, 6(4): 483–484, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474756.

Anonymous:1998:EBPa

[Ano98a] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 7 (1):??, 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1998.10474757.

Anonymous:1998:EBPb

[Ano98b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 7 (2):??, 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1998.10474766.

Anonymous:1998:EBPc

[Ano98c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 7 (3):??, 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1998.10474774.

Anonymous:1998:EBPd

[Ano98d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 7 (4):??, 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1998.10474785.

Anonymous:1998:EC

[Ano98e] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 7(4): 559–560, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1998.10474795.

Anonymous:1998:IV

[Ano98f] Anonymous. Index to volume 7 (1998). Journal of Computational and Graphical Statistics, 7(4): 561–563, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1998.10474796.

Anonymous:1999:EBPa

[Ano99a] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 8 (1):??, 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1999.10474797.

Anonymous:1999:EBPb

[Ano99b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 8 (2):??, 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474807.

Anonymous:1999:EBPc

[Ano99c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 8 (3):??, 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474819.

Anonymous:1999:EBPd

[Ano99d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 8 (4):??, 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474841.

Anonymous:1999:EC

[Ano99e] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 8(4): 913–914, 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474856.

Anonymous:2000:C

[Ano00a] Anonymous. Correction. Journal of Computational and Graphical Statistics, 9(1):215, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474875.

Anonymous:2000:EBPa

[Ano00b] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 9 (1):??, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474857.

Anonymous:2000:EBPb

[Ano00c] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 9 (2):??, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/abs/10.1080/10618600.
2000.10474876.

Anonymous:2000:EBPc

[Ano00d] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 9 (3):??, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474888.

Anonymous:2000:EBPd

[Ano00e] Anonymous. Editorial Board page. Journal of Computational and Graphical Statistics, 9 (4):??, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474902.

Anonymous:2000:EC

[Ano00f] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 9(4): 801–802, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474914.

Anonymous:2000:IV

[Ano00g] Anonymous. Index to volume 9 (2000). Journal of Computational and Graphical Statistics, 9(4): 803–806, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474915.

Anonymous:2001:C

[Ano01a]  Anonymous.  Correc-

tion. Journal of Computational and Graphical Statistics, 10(3):615, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2001:PA

[Ano01b] Anonymous. Primary article. Journal of Computational and Graphical Statistics, 10(3): 615, 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317115144.

Anonymous:2001:RPJa

[Ano01c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 10 (2):401, 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628121.

Anonymous:2001:RPJb

[Ano01d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 10 (3):616, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317115027.

Anonymous:2002:RPJa

[Ano02a] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 11 (1):259–260, March 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s14/p259;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375721.

Anonymous:2002:RPJb

[Ano02b] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 11 (2):483–484, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=

                                        asa&reqidx=/cw/asa/10618600/
                                        v11n2/s14/p483;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002760180635.

Anonymous:2002:RPJc

[Ano02c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 11(3):733–734, September 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s13/p733;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002320604191.

Anonymous:2002:RPJd

[Ano02d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 11(4):996–997, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n4/s15/p996;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002321018902.

Anonymous:2003:EC

[Ano03a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 12(4):971–974, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032689.

Anonymous:2003:RPJa

[Ano03b] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 12 (1):243–248, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186003321335099.

Anonymous:2003:RPJb

[Ano03c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 12(2):472–474, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s12/p472;   http:
//www.tandfonline.com/
doi/abs/10.1198/10618600360681860.

Anonymous:2003:RPJc

[Ano03d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 12(4):975–976, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186003322691958.

Anonymous:2004:RPJa

[Ano04a] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 13 (2):534–536, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2004:RPJb

[Ano04b] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 13 (3):751–754, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2004:RPJc

[Ano04c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 13 (4):1018–1019, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2005:RPJa

[Ano05a] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 14 (1):251–253, March 2005.

CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2005:RPJb

[Ano05b] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 14 (2):505–510, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2005:RPJc

[Ano05c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 14 (3):746–749, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2005:RPJd

[Ano05d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 14 (4):947–951, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2006:CAC

[Ano06a] Anonymous. Correction:

Analytic, Computational, and Approximate Forms for Ratios of Noncentral and Central Gaussian Quadratic Forms, Journal of Computational and Graphical Statistics 15, 443–459 (2006). Journal of Computational and Graphical Statistics, 15 (4):956, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [KGMT06].

Anonymous:2006:EC

[Ano06b] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 15 (4):961–965, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2006:RPJa

[Ano06c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 15 (1):262–264, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2006:RPJb

[Ano06d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 15 (2):477–481, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2006:RPJc

[Ano06e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 15 (3):753–756, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2006:RPJd

[Ano06f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 15 (4):957–960, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2007:CUR

[Ano07a] Anonymous. Correction:

Using Random Quasi-Monte-Carlo Within Particle Filters, With Application to Financial Time Series, Journal of Computational and Graphical Statistics 15, 751–769 (2005). Journal of Computational and Graphical Statistics, 16(1):255, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Fea05].

Anonymous:2007:EC

[Ano07b] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 16 (4):999–1002, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2007:RPJa

[Ano07c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 16 (1):256–263, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2007:RPJb

[Ano07d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 16 (2):503–506, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2007:RPJc

[Ano07e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 16 (3):752–759, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2007:RPJd

[Ano07f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 16 (4):991–998, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:CSA

[Ano08a]  Anonymous.  Correc-

tion: Spatially Adaptive Bayesian Penalized Regression Splines (Psplines), Journal of Computational and Graphical Statistics 14, 378–394 (2005). Journal of Computational and Graphical Statistics, 17(2):515, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [BMC05].

Anonymous:2008:EC

[Ano08b] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 17 (4):1016–1019, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:IV

[Ano08c] Anonymous. Index to volume 17 (2008). Journal of Computational and Graphical Statistics, 17 (4):1020–1023, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:RPJa

[Ano08d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 17 (1):263–268, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:RPJb

[Ano08e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 17 (2):516–525, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:RPJc

[Ano08f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 17 (3):770–772, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2008:RPJd

[Ano08g] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 17 (4):1010–1015, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:EC

[Ano09a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 18 (4):1037–1040, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:IV

[Ano09b] Anonymous. Index to volume 18 (2009). Journal of Computational and Graphical Statistics, 18 (4):1041–1044, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:RPJa

[Ano09c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 18 (1):238–242, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:RPJb

[Ano09d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 18 (2):527–532, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:RPJc

[Ano09e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 18 (3):792–796, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2009:RPJd

[Ano09f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 18 (4):1035–1036, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:EC

[Ano10a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 19 (4):1044–1046, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:IV

[Ano10b] Anonymous. Index to volume 19 (2010). Journal of Computational and Graphical Statistics, 19 (4):1047–1051, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:RPJa

[Ano10c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 19 (1):240–241, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:RPJb

[Ano10d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 19 (2):493–495, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:RPJc

[Ano10e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 19 (3):766–767, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2010:RPJd

[Ano10f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 19 (4):1041–1043, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:EC

[Ano11a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 20 (4):1031–1033, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:IV

[Ano11b] Anonymous. Index to volume 20. Journal of Computational and Graphical Statistics, 20(4):1034– 1038, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:RPJa

[Ano11c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 20 (1):279–280, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:RPJb

[Ano11d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 20 (2):527–529, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:RPJc

[Ano11e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 20 (3):789–792, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2011:RPJd

[Ano11f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 20 (4):1029–1030, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2012:EBE

[Ano12a] Anonymous. Editorial Board EOV. Journal of Computational and Graphical Statistics, 21 (4):ebi, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2012:EAC

[Ano12b]  Anonymous.  Erratum

and addendum to: “Covariance Tapering for Interpolation of Large Spatial Datasets” published in the Journal of Computational and Graphical Statistics, 15, 502–523. Journal of Computational and Graphical Statistics, 21(3):823–824, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [FGN06].

Anonymous:2012:RPJa

[Ano12c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 21 (1):272–273, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2012:RPJb

[Ano12d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 21 (2):558–560, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2012:RPJc

[Ano12e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 21 (3):821–822, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2012:RPJd

[Ano12f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 21 (4):979–980, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:EC

[Ano13a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 22 (4):989–992, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:EEB

[Ano13b] Anonymous. EOV Editorial Board. Journal of Computational and Graphical Statistics, 22 (4):ebi, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:RPJa

[Ano13c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 22 (1):229–230, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:RPJb

[Ano13d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 22 (2):513–514, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:RPJc

[Ano13e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 22 (3):774–775, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:RPJd

[Ano13f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 22 (4):987–988, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2013:SR

[Ano13g] Anonymous. Statement of removal. Journal of Computational and Graphical Statistics, 22(4):994, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:EC

[Ano14a] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 23 (4):1182–1185, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:EEB

[Ano14b] Anonymous. EOV Editorial Board. Journal of Computational and Graphical Statistics, 23 (4):ebi, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:RPJa

[Ano14c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 23 (1):292–293, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:RPJb

[Ano14d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 23 (2):586–588, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:RPJc

[Ano14e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 23 (3):893–894, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2014:RPJd

[Ano14f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 23 (4):1179–1181, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2015:EBE

[Ano15a] Anonymous. Editorial Board EOV. Journal of Computational and Graphical Statistics, 24 (4):ebi, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2015.1116909.

Anonymous:2015:EC

[Ano15b] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 24 (4):1234, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2015.1120102.

Anonymous:2015:RPJa

[Ano15c] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 24 (1):297–300, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2015:RPJb

[Ano15d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 24 (2):600–601, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2015.1055220.

Anonymous:2015:RPJc

[Ano15e] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 24 (3):905–908, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2015:RPJd

[Ano15f] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 24 (4):1231–1233, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/abs/10.1080/10618600.
2015.1116340.

Anonymous:2016:C

[Ano16a] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 25(3):980, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1189257.

Anonymous:2016:EBE

[Ano16b] Anonymous. Editorial Board EOV. Journal of Computational and Graphical Statistics, 25 (4):ebi, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2016:EC

[Ano16c] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 25 (4):1317–1320, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2016:RPJ

[Ano16d] Anonymous. Recent publications in JSS. Journal of Computational and Graphical Statistics, 25 (1):321–323, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2016.1153373;   http:
                                        //www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2016.1153373.

Anonymous:2017:EBE

[Ano17a] Anonymous. Editorial Board EOV. Journal of Computational and Graphical Statistics, 26 (4):ebi, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2017:EC

[Ano17b] Anonymous. Editorial collaborators. Journal of Computational and Graphical Statistics, 26 (4):941–943, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2018:C

[Ano18a] Anonymous. Correction. Journal of Computational and Graphical Statistics, 27(4):947, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2018:EBE

[Ano18b] Anonymous. Editorial Board EOV. Journal of Computational and Graphical Statistics, 27 (4):ebi, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2019:Ca

[Ano19a] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 28(2):i, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2019:Cb

[Ano19b] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 28(4):i, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2019:Cc

[Ano19c] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 28(4):ii, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2019:Cd

[Ano19d] Anonymous. Correction.

Journal of Computational and Graphical Statistics, 28(4):1017, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Anonymous:2020:CGL

[Ano20] Anonymous. Correction:

Generalized Link-Based Additive Survival Models with Informative Censoring. Journal of Computational and Graphical Statistics, 29(3):??, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1760564.  See

[DMR20].

Anonymous:2021:JEC

[Ano21] Anonymous. JCGS editorial collaborators. Journal of Computational and Graphical Statistics, 30 (1):??, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1875769.

Anonymous:2022:C

[Ano22] Anonymous. Correction. Journal of Computational and Graphical Statistics, 31(4):1439, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2129213.

Anonymous:2023:C

[Ano23a] Anonymous. Correction. Journal of Computational and Graphical Statistics, 32(1):??, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2165501.

Anonymous:2023:SRD

[Ano23b] Anonymous. Statement of retraction: Detecting anomalous time series by GAMLSS-Akaike-weights-scoring. Journal of Computational and Graphical Statistics, 33 (1):329, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2285648.

Anonymous:2024:CPM

[Ano24a]  Anonymous.  Correc-

tion: Probabilistic Kmeans with Local Alignment for Clustering and Motif Discovery in Functional Data, Cremona, M. A., and Chiaromonte, F., Journal of Computational and Graphical Statistics, 32(3) 1119–1130. Journal of Computational and Graphical Statistics, 33 (3):1129, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2356159.  See

[CC22b].

Anonymous:2024:JCG

[Ano24b] Anonymous. The journal of computational and graphical statistics 2023 associate Editors. Journal of Computational and Graphical Statistics, 33 (1):??, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2319476.

Atkinson:2006:DTS

[AR06] Anthony C. Atkinson and Marco Riani. Distribution theory and simulations for tests of outliers in regression. Journal of Computational and Graphical Statistics, 15(2):460–476, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Aliverti:2022:SSV

[AR22] Emanuele Aliverti and Massimiliano Russo. Stratified stochastic variational inference for highdimensional network factor model. Journal of Computational and Graphical Statistics, 31 (2):502–511, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1984929.

An:2019:ABS

[ASND19] Ziwen An, Leah F. South, David J. Nott, and Christopher C. Drovandi. Accelerating Bayesian synthetic likelihood with the graphical lasso. Journal of Computational and Graphical Statistics, 28 (2):471–475, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Arroyo:2021:MLE

[ASPL21] Jesús Arroyo, Daniel L.

Sussman, Carey E. Priebe, and Vince Lyzinski. Maximum likelihood estimation and graph matching in errorfully observed networks. Journal of Computational and Graphical Statistics, 30(4):1111– 1123, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1872582.

Arnold:2016:EIG

[AT16] Taylor B. Arnold and Ryan J. Tibshirani. Efficient implementations of the generalized lasso dual path algorithm. Journal of Computational and Graphical Statistics, 25 (1):1–27, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2015.1008638;   http:
                                        //www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1008638.

Agarwal:2022:GCA

[AV22] Medha Agarwal and Dootika Vats. Globally centered autocovariances in MCMC. Journal of Computational and Graphical Statistics, 31 (3):629–638, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2037433.

Algeri:2020:TOH

[AvD20] Sara Algeri and David A. van Dyk. Testing one hypothesis multiple times: The multidimensional case. Journal of Computational and Graphical Statistics, 29(2):358– 371, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1677474.

Antoniano-Villalobos:2013:BNI

[AVW13] Isadora Antoniano-Villalobos and Stephen G. Walker. Bayesian nonparametric inference for the power likelihood. Journal of Computational and Graphical Statistics, 22 (4):801–813, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Agarwal:2023:QNA

[AX23] Medha Agarwal and Jason Xu. Quasi-Newton acceleration of EM and MM algorithms via Broyden’s method. Journal of Computational and Graphical Statistics, 33 (2):393–406, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257261.

Algeri:2022:EGF

[AZ22] Sara Algeri and Xiangyu Zhang. Exhaustive goodness of fit via smoothed inference and graphics. Journal of Computational and Graphical Statistics, 31(2):378–389, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1999826.

Baker:1992:CSM

[Bak92a] S. G. Baker. Correction: “A simple method for computing the observed information matrix when using the EM algorithm with categorical data” [J. Comput. Graph. Statist. 1 (1992), no. 1, 63–76]. Journal of Computational and Graphical Statistics, 1(2):180, June 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Bak92b].

Baker:1992:SMC

[Bak92b] Stuart G. Baker. A simple method for computing the observed information matrix when using the EM algorithm with categorical data. Journal of Computational and Graphical Statistics, 1(1): 63–76, March 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474576. See cor-

rection [Bak92a].

Beraha:2022:MCB

[BAMG22] Mario Beraha, Raffaele Argiento, Jesper Møller, and Alessandra Guglielmi. MCMC computations for Bayesian mixture models using repulsive point processes. Journal of Computational and Graphical Statistics, 31(2):422– 435, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000424.

Baumer:2019:GRP

[Bau19] Benjamin S. Baumer. A grammar for reproducible and painless extracttransform-load operations on medium data. Journal of Computational and Graphical Statistics, 28 (2):256–264, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bellio:2003:HOA

[BB03] Ruggero Bellio and Alessandra R. Brazzale. Higher order asymptotics unleashed: Software design for nonlinear heteroscedastic regression. Journal of Computational and Graphical Statistics, 12(3):682–697, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s10/p682;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032003.

Bowman:2006:EVT

[BB06] Adrian W. Bowman and Mitchum T. Bock. Exploring variation in threedimensional shape data.

Journal of Computational and Graphical Statistics, 15(3):524–541, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Begin:2021:LEJ

[BB21] Jean-François Bégin and Mathieu Boudreault. Likelihood evaluation of jumpdiffusion models using deterministic nonlinear filters. Journal of Computational and Graphical Statistics, 30(2):452– 466, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1840995.

Biecek:2024:PES

[BBKC24] Przemysław Biecek, Hubert Baniecki, Mateusz Krzyziński, and Dianne Cook. Performance is not enough: The story told by a Rashomon quartet. Journal of Computational and Graphical Statistics, 33(3):1118–1121, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2344616.

Bohning:2013:URP

[BBLG13] Dankmar Böhning, M. Fazil Baksh, Rattana Lerdsuwansri, and James Gallagher. Use of the ratio plot in capture-recapture estimation. Journal of Computational and Graphical Statistics, 22 (1):135–155, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Banerjee:2024:BEC

[BBM24] Trambak Banerjee, Bhaswar B.

Bhattacharya, and Gourab Mukherjee. Bootstrapped edge count tests for nonparametric two-sample inference under heterogeneity. Journal of Computational and Graphical Statistics, 34(1):306– 317, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2374583.

Bugbee:2016:LVA

[BBvdW16] Bruce D. Bugbee, F. Jay Breidt, and Mark J. van der Woerd. Laplace variational approximation for semiparametric regression in the presence of heteroscedastic errors. Journal of Computational and Graphical Statistics, 25(1):225–245, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1080/10618600.
                                        2014.983642;   http:/
                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2014.983642.

Bar:2020:SEB

[BBW20] Haim Y. Bar, James G. Booth, and Martin T. Wells. A scalable empirical Bayes approach to variable selection in generalized linear models. Journal of Computational and Graphical Statistics, 29(3):535–546, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1706542.

Becker:1993:DSG

[BC93] Richard A. Becker and William S. Cleveland. Discussion of “Graphic Comparisons of Several Linked Aspects” by John W. Tukey. Journal of Computational and Graphical Statistics, 2(1): 41–48, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474597.  See

[Tuk93a].

Barnett:2000:HBD

[BC00] Ofra Barnett and Ayala Cohen. The histogram and boxplot for the display of lifetime data. Journal of Computational and Graphical Statistics, 9(4):759–778, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Barnett.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474912.

Bauer:2016:PCB

[BC16] Alexander Bauer and Claudia Czado. Paircopula Bayesian networks. Journal of Computational and Graphical Statistics, 25(4):1248– 1271, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Buchholz:2019:IAB

[BC19] Alexander Buchholz and Nicolas Chopin. Improving approximate Bayesian computation via quasi-Monte Carlo. Journal of Computational and Graphical Statistics, 28 (1):205–219, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bradley:2024:GIR

[BC24] Jonathan R. Bradley and Madelyn Clinch. Generating independent replicates directly from the posterior distribution for a class of spatial hierarchical models. Journal of Computational and Graphical Statistics, 34 (1):123–139, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2365728.

Bai:2011:DCM

[BCD11] Yan Bai, Radu V. Craiu, and Antonio F. Di Narzo. Divide and conquer: a mixture-based approach to regional adaptation for MCMC. Journal of Computational and Graphical Statistics, 20(1):63– 79, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Beavers:2024:DNM

[BCD^{+}24] Traymon E. Beavers, Ge Cheng, Yajie Duan, Javier Cabrera, Mariusz Lubomirski, Dhammika Amaratunga, and Jeffrey E. Teigler. Data nuggets: a method for reducing big data while preserving data structure. Journal of Computational and Graphical Statistics, 34(1):330–342, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2024.2341896.

Bartolucci:2017:CLI

[BCDL17] Francesco Bartolucci, Francesca Chiaromonte, Prabhani Kuruppumullage Don, and Bruce G. Lindsay. Composite likelihood inference in a discrete latent variable model for two-way “clustering-by-segmentation” problems. Journal of Computational and Graphical Statistics, 26(2):388–402, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Baker:2020:POS

[BCE20] Evan Baker, Peter Challenor, and Matt Eames. Predicting the output from a stochastic computer model when a deterministic approximation is available. Journal of Computational and Graphical Statistics, 29 (4):786–797, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1750416.

Booth:2005:ECV

[BCH05] James G. Booth, Marinela Capanu, and Ludwig Heigenhauser. Exact conditional P value calculation for the quasisymmetry model. Journal of Computational and Graphical Statistics, 14 (3):716–725, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Benaglia:2009:ELA

[BCH09] Tatiana Benaglia, Didier Chauveau, and David R. Hunter. An EM-like algorithm for semi- and nonparametric estimation in multivariate mixtures. Journal of Computational and Graphical Statistics, 18(2):505–526, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Barnwal:2022:SRA

[BCH22] Avinash Barnwal, Hyunsu Cho, and Toby Hocking. Survival regression with accelerated failure time model in XGBoost. Journal of Computational and Graphical Statistics, 31 (4):1292–1302, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2067548.

Becker:1994:CPG

[BCL94] Richard A. Becker, Linda A. Clark, and Diane Lambert. Cave plots: a graphical technique for comparing time series. Journal of Computational and Graphical Statistics, 3(3): 277–283, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474645.

Becker:1996:VDC

[BCS96a] Richard A. Becker, William S.

Cleveland, and Ming-Jen Shyu. The visual design and control of trellis display. Journal of Computational and Graphical Statistics, 5(2):123– 155, June 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474701.

Buja:1996:IHD

[BCS96b] Andreas Buja, Dianne Cook, and Deborah F. Swayne Research Scientist. Interactive highdimensional data visualization. Journal of Computational and Graphical Statistics, 5(1):78– 99, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474696.

Bernardi:2022:LSF

[BCS22] Mauro Bernardi, Antonio Canale, and Marco Stefanucci. Locally sparse function-on-function regression. Journal of Computational and Graphical Statistics, 32(3):985– 999, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2130926.

Baddeley:2013:RDC

[BCST13] Adrian Baddeley, Ya-Mei Chang, Yong Song, and Rolf Turner. Residual diagnostics for covariate effects in spatial point process models. Journal of Computational and Graphical Statistics, 22 (4):886–905, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Burgard:2022:GLM

[BD22] Jan Pablo Burgard and Patricia Dörr. Generalized linear mixed models with crossed effects and unit-specific survey weights. Journal of Computational and Graphical Statistics, 31(4):1037– 1050, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2001342.

Barone:2023:BNM

[BD23] Rosario Barone and Luciana Dalla Valle. Bayesian nonparametric modeling of conditional multidimensional dependence structures. Journal of Computational and Graphical Statistics, 32 (4):1361–1370, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2173604.

Brooks:1997:ADT

[BDR97] S. P. Brooks, P. Dellaportas, and G. O. Roberts. An approach to diagnosing total variation convergence of MCMC algorithms. Journal of Computational and Graphical Statistics, 6(3):251–265, September 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/brooks.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474741.

Bedard:2008:ESU

[Béd08] Mylène Bédard. Efficient sampling using Metropolis algorithms. Journal of Computational and Graphical Statistics, 17 (2):312–332, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bee:2017:CEA

[BEGS17] Marco Bee, Giuseppe Espa, Diego Giuliani, and Flavio Santi. A crossentropy approach to the estimation of generalized linear multilevel models. Journal of Computational and Graphical Statistics, 26(3):695–708, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Boon:2013:VMQ

[BEM13] Marko A. A. Boon, John H. J. Einmahl, and Ian W. McKeague. Visualizing multiple quantile plots. Journal of Computational and Graphical Statistics, 22(1):69–78, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Buja:1994:R

[BF94] Andreas Buja and George W. Furnas. Rejoinder. Journal of Computational and Graphical Statistics, 3(4): 377–385, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/abs/10.1080/10618600.
1994.10474653.

Beran:2002:IPA

[BF02] Jan Beran and Yuanhua Feng. Iterative plug-in algorithms for SEMIFAR models — definition, convergence, and asymptotic properties. Journal of Computational and Graphical Statistics, 11(3):690–713, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s11/p690;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002420.

Braverman:2003:SSQ

[BFE^{+}03] Amy Braverman, Eric Fetzer, Annmarie Eldering, Silvia Nittel, and Kelvin Leung. Semistreaming quantization for remote sensing data. Journal of Computational and Graphical Statistics, 12(4):759–780, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032535.

Bouranis:2018:BMS

[BFM18] Lampros Bouranis, Nial Friel, and Florian Maire.

Bayesian model selection for exponential random graph models via adjusted pseudolikelihoods. Journal of Computational and Graphical Statistics, 27 (3):516–528, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Barrett:1992:ECS

[BG92] Bruce E. Barrett and J. Brian Gray. Efficient computation of subset influence in regression. Journal of Computational and Graphical Statistics, 1(4):271–286, December 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474586.

Brooks:1998:GMM

[BG98] Stephen P. Brooks and Andrew Gelman. General methods for monitoring convergence of iterative simulations. Journal of Computational and Graphical Statistics, 7(4):434–455, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/brooks.
html;   http://www.tandfonline.

com/doi/abs/10.1080/
10618600.1998.10474787.

Brooks:2000:MCM

[BG00] S. P. Brooks and P. Giudici. Markov Chain Monte Carlo convergence assessment via two-way analysis of variance. Journal of Computational and Graphical Statistics, 9(2): 266–285, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474880.

Beam:2016:FHM

[BGD16] Andrew L. Beam, Sujit K.

Ghosh, and Jon Doyle. Fast Hamiltonian Monte Carlo using GPU computing. Journal of Computational and Graphical Statistics, 25(2):536–548, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Binois:2018:PHG

[BGL18] Mickaël Binois, Robert B.

Gramacy, and Mike Ludkovski. Practical heteroscedastic Gaussian process modeling for large simulation experiments. Journal of Computational and Graphical Statistics, 27(4):808–821, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bien:2018:NCG

[BGLM18] Jacob Bien, Irina Gaynanova, Johannes Lederer, and Christian L. Müller. Non-convex global minimization and false discovery rate control for the TREX. Journal of Computational and Graphical Statistics, 27 (1):23–33, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Brooks:2003:NCA

[BGP03] S. P. Brooks, P. Giudici, and A. Philippe. Nonparametric convergence assessment for MCMC model selection. Journal of Computational and Graphical Statistics, 12(1):1–22, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

                                        URL http://giorgio.
                                        ingentaselect.com/cgi-
                                        bin/linker?ini=asa{\&
                                        }reqidx=/cw/asa/10618600/
                                        v12n1/s1/p1;   http:/
                                        /www.tandfonline.com/
                                        doi/abs/10.1198/1061860031347.

Braun:2001:DSN

[BH01] W. John Braun and Peter Hall. Data sharpening for nonparametric inference subject to constraints. Journal of Computational and Graphical Statistics, 10(4):786–806, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243458.

Bilen:2002:WBD

[BH02] Canan Bilen and S. Huzurbazar. Wavelet-based detection of outliers in time series. Journal of Computational and Graphical Statistics, 11(2):311– 327, June 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s4/p311;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180536.

Blackburn:2022:PNM

[BH22] Bart Blackburn and Mark S.

Handcock. Practical network modeling via tapered exponential-family random graph models. Journal of Computational and Graphical Statistics, 32(2):388–401, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2116444.

Bradlow:2002:BIN

[BHF02] Eric T. Bradlow, Bruce G. S. Hardie, and Peter S.

Fader. Bayesian inference for the negative binomial distribution via polynomial expansions. Journal of Computational and Graphical Statistics, 11 (1):189–201, March 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s9/p189;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375677.

Bove:2015:OBM

[BHK15] Daniel Sabanés Bové, Leonhard Held, and Göran Kauermann. Objective Bayesian model selection in generalized additive models with penalized splines. Journal of Computational and Graphical Statistics, 24 (2):394–415, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.912136.

Bureau:2000:PIH

[BHS00] Alexandre Bureau, James P. Hughes, and Stephen C. Shiboski. An S-Plus implementation of hidden Markov models in continuous time. Journal of Computational and Graphical Statistics, 9(4):621–632, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Bureau.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474903.

Brys:2004:RMS

[BHS04] G. Brys, M. Hubert, and A. Struyf. A robust measure of skewness. Journal of Computational and Graphical Statistics, 13 (4):996–1017, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Boileau:2022:CVL

[BHvdLD22] Philippe Boileau, Nima S.

Hejazi, Mark J. van der Laan, and Sandrine Dudoit. Cross-validated loss-based covariance matrix estimator selection in high dimensions. Journal of Computational and Graphical Statistics, 32 (2):601–612, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2110883.

Biernacki:2005:TGM

[Bie05]  Christophe  Biernacki.

Testing for a global maximum of the likelihood.

Journal of Computational and Graphical Statistics, 14(3):657–674, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bigot:2006:LBR

[Big06] Jérémie Bigot. Landmarkbased registration of curves via the continuous wavelet transform. Journal of Computational and Graphical Statistics, 15 (3):542–564, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Biller:2000:ABR

[Bil00] Clemens Biller. Adaptive Bayesian regression splines in semiparametric generalized linear models. Journal of Computational and Graphical Statistics, 9(1):122–140, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Biller.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474869.

Bowman:1998:TMR

[BJG98] A. W. Bowman, M. C. Jones, and I. Gijbels. Testing monotonicity of regression. Journal of Computational and Graphical Statistics, 7(4): 489–500, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts98/bowman.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474790.

Bornn:2013:AIW

[BJMD13] Luke Bornn, Pierre E. Jacob, Pierre Del Moral, and Arnaud Doucet. An adaptive interacting Wang–Landau algorithm for automatic density exploration. Journal of Computational and Graphical Statistics, 22 (3):749–773, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bolton:2003:PPC

[BK03a] R. J. Bolton and W. J.

Krzanowski. Projection pursuit clustering for exploratory data analysis. Journal of Computational and Graphical Statistics, 12(1):121–142, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s6/p121;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031374.

Brockwell:2003:GMB

[BK03b] Anthony E. Brockwell and Joseph B. Kadane. A gridding method for Bayesian sequential decision problems. Journal of Computational and Graphical Statistics, 12 (3):566–584, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

                                        URL http://giorgio.
                                        ingentaselect.com/cgi-
                                        bin/linker?ini=asa{\&
                                        }reqidx=/cw/asa/10618600/
                                        v12n3/s5/p566;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/1061860032274.

Brockwell:2005:IRT

[BK05] Anthony E. Brockwell and Joseph B. Kadane. Identification of regeneration times in MCMC simulation, with application to adaptive schemes. Journal of Computational and Graphical Statistics, 14(2):436–458, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bolin:2020:RSA

[BK20] David Bolin and Kristin Kirchner. The rational SPDE approach for Gaussian random fields with general smoothness. Journal of Computational and Graphical Statistics, 29 (2):274–285, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1665537.

Beskos:2022:MAP

[BK22] Alexandros Beskos and Kengo Kamatani. MCMC algorithms for posteriors on matrix spaces. Journal of Computational and Graphical Statistics, 31 (3):721–738, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2058953.

Borowska:2023:SCD

[BK23] Agnieszka Borowska and Ruth King. Semicomplete data augmentation for efficient state space model fitting. Journal of Computational and Graphical Statistics, 32 (1):19–35, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2077350.

Buja:1999:ISS

[BKM99] Andreas Buja and Sallie Keller-McNulty. Introduction to the special section on massive datasets. Journal of Computational and Graphical Statistics, 8(3):544, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1999.10474831.

Bogaerts:2004:NFA

[BL04] Kris Bogaerts and Emmanuel Lesaffre. A new, fast algorithm to find the regions of possible support for bivariate intervalcensored data. Journal of Computational and Graphical Statistics, 13 (2):330–340, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bolin:2017:QUC

[BL17] David Bolin and Finn Lindgren. Quantifying the uncertainty of contour maps. Journal of Computational and Graphical Statistics, 26(3):513–524, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Beh:2022:FPB

[BL22] Eric J. Beh and Rosaria Lombardo. Features of the polynomial biplot for ordered contingency tables. Journal of Computational and Graphical Statistics, 31(2):403– 412, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1990773.

Biron-Lattes:2022:PMI

[BLBCC22] Miguel Biron-Lattes, Alexandre Bouchard-Côté, and Trevor Campbell. Pseudomarginal inference for CTMCs on infinite spaces via monotonic likelihood approximations. Journal of Computational and Graphical Statistics, 32 (2):513–527, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2118750.

Borgonovo:2022:SAP

[BLR22]  Emanuele  Borgonovo,

Xuefei Lu, and Giovanni Rabitti. Sensitivity analysis of pandemic models can support effective policy decisions. Journal of Computational and Graphical Statistics, 32 (3):767–768, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2126483.

Beran:1999:VRB

[BM99] Jan Beran and Guerino Mazzola. Visualizing the relationship between two time series by hierarchical smoothing models. Journal of Computational and Graphical Statistics, 8(2):213– 238, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/hierarchical.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/beran.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474811.

Biscio:2019:APF

[BM19] Christophe A. N. Biscio and Jesper Møller. The accumulated persistence function, a new useful functional summary statistic for topological data analysis, with a view to brain artery trees and spatial point process applications. Journal of Computational and Graphical Statistics, 28 (3):671–681, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Baladandayuthapani:2005:SAB

[BMC05] Veerabhadran Baladandayuthapani, Bani K. Mallick, and Raymond J. Carroll. Spatially adaptive Bayesian penalized regression splines (Psplines). Journal of Computational and Graphical Statistics, 14(2):378– 394, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See corrections [Ano08a].

Beran:2016:THR

[BMG16] Jan Beran, Sven Möhrle, and Sucharita Ghosh. Testing for Hermite rank in Gaussian subordination processes. Journal of Computational and Graphical Statistics, 25 (3):917–934, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1056345.

Berild:2022:ISI

[BMGRR22] Martin Outzen Berild, Sara Martino, Virgilio Gómez-Rubio, and Håvard Rue. Importance sampling with the integrated nested Laplace approximation. Journal of Computational and Graphical Statistics, 31(4):1225– 1237, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2067551.

Becker:2019:TFE

[BML19] Gabriel Becker, Sara E. Moore, and Michael Lawrence. trackr: A framework for enhancing discoverability and reproducibility of data visualizations and other artifacts in R. Journal of Computational and Graphical Statistics, 28 (3):644–658, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Basu:2001:PED

[BMO01] Sankar Basu, Charles A. Micchelli, and Peder Olsen. Power exponential densities for the training and classification of acoustic feature vectors in speech recognition. Journal of Computational and Graphical Statistics, 10 (1):158–184, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/10618600152418818.

Bierme:2015:TBM

[BMR15] Hermine Biermé, Lionel Moisan, and Frédéric Richard. A turning-band method for the simulation of anisotropic fractional Brownian fields. Journal of Computational and Graphical Statistics, 24 (3):885–904, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Brown:2017:TTB

[BMW17] Ben Brown, Christopher J. Miller, and Julian Wolfson. ThrEEBoost: Thresholded boosting for variable selection and prediction via estimating equations. Journal of Computational and Graphical Statistics, 26 (3):579–588, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Beran:2005:SSA

[BO05] Jan Beran and Dirk Ocker. Small sample asymptotics for credit risk portfolios. Journal of Computational and Graphical Statistics, 14 (2):339–351, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bornkamp:2011:APD

[Bor11] Björn Bornkamp. Approximating probability densities by iterated Laplace approximations. Journal of Computational and Graphical Statistics, 20(3):656–669, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bloch:2002:REC

[BOW02] Daniel A. Bloch, Richard A. Olshen, and Michael G. Walker. Risk estimation for classification trees. Journal of Computational and Graphical Statistics, 11(2):263–288, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s1/p263;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180509.

Booth:1998:AMC

[BP98] James Booth and Brett Presnell. Allocation of Monte Carlo resources for the iterated bootstrap. Journal of Computational and Graphical Statistics, 7(1):92–112, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.research.att.com/
~andreas/jcgs/abstracts/
1998/march/booth.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474763.

Bell:1993:CTS

[BPW93] B. Bell, D. B. Percival, and A. T. Walden. Calculating Thomson’s spectral multitapers by inverse iteration. Journal of Computational and Graphical Statistics, 2(1):119– 130, March 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474602.

Block:1994:SAP

[BQS94] H. Block, S. Qian, and A. Sampson. Structure algorithms for partially ordered isotonic regression. Journal of Computational and Graphical Statistics, 3(3):285–300, September 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474646.

Brutlag:2002:BSA

[BR02] Jake D. Brutlag and Thomas S. Richardson. A block sampling approach to distinct value estimation. Journal of Computational and Graphical Statistics, 11(2):389– 404, June 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s8/p389;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180572.

Brazzale:1999:ACI

[Bra99] Alessandra R. Brazzale. Approximate conditional inference in logistic and loglinear models. Journal of Computational and Graphical Statistics, 8(3): 653–661, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1999.10474839.

Braverman:2002:CMG

[Bra02] Amy Braverman. Compressing massive geophysical datasets using vector quantization. Journal of Computational and Graphical Statistics, 11 (1):44–62, March 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n1/s3/p44;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002317375613.

Bradley:2021:AIS

[Bra21] Jonathan R. Bradley. An approach to incorporate subsampling into a generic Bayesian hierarchical model. Journal of Computational and Graphical Statistics, 30 (4):889–905, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923518.

Betancourt:2017:BFL

[BRB17] Brenda Betancourt, Abel Rodríguez, and Naomi Boyd. Bayesian fused lasso regression for dynamic binary networks. Journal of Computational and Graphical Statistics, 26(4):840–850, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Baudry:2010:CMC

[BRC^{+}10] Jean-Patrick Baudry, Adrian E.

Raftery, Gilles Celeux, Kenneth Lo, and Raphaël Gottardo. Combining mixture components for clustering. Journal of Computational and Graphical Statistics, 19 (2):332–353, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bertolacci:2022:AXC

[BRCC22] Michael Bertolacci, Ori Rosen, Edward Cripps, and Sally Cripps. AdaptSPEC-X: Covariate-dependent spectral modeling of multiple nonstationary time series. Journal of Computational and Graphical Statistics, 31(2):436– 454, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2000870.

Breidt:2004:SEQ

[Bre04] F. Jay Breidt. Simulation estimation of quantiles from a distribution with known mean. Journal of Computational and Graphical Statistics, 13 (2):487–498, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Brillinger:1994:ESP

[Bri94] David R. Brillinger. Examples of scientific problems and data analyses in demography, neurophysiology, and seismology. Journal of Computational and Graphical Statistics, 3(1):1–22, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474627.

Briggs:2011:CWT

[Bri11] William M. Briggs. Comment: Why Tables Are Really Much Better Than Graphs. Journal of Computational and Graphical Statistics, 20(1):16– 17, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Gel11b].

Barrientos:2019:DPS

[BRMC19] Andrés F. Barrientos, Jerome P. Reiter, Ashwin Machanavajjhala, and Yan Chen. Differentially private significance tests for regression coefficients. Journal of Computational and Graphical Statistics, 28(2):440–453, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Brockwell:2006:PMC

[Bro06] A. E. Brockwell. Parallel Markov chain Monte Carlo simulation by prefetching. Journal of Computational and Graphical Statistics, 15(1):246–261, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bliznyuk:2008:BCU

[BRS^{+}08] Nikolay Bliznyuk, David Ruppert, Christine Shoemaker, Rommel Regis, Stefan Wild, and Pradeep Mugunthan. Bayesian calibration and uncertainty analysis for computationally expensive models using optimization and radial basis function approximation. Journal of Computational and Graphical Statistics, 17 (2):270–294, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bliznyuk:2011:EIC

[BRS11] Nikolay Bliznyuk, David Ruppert, and Christine A. Shoemaker. Efficient interpolation of computationally expensive posterior densities with variable parameter costs. Journal of Computational and Graphical Statistics, 20(3):636–655, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bliznyuk:2012:LDF

[BRS12] Nikolay Bliznyuk, David Ruppert, and Christine A. Shoemaker. Local derivative-free approximation of computationally expensive posterior densities. Journal of Computational and Graphical Statistics, 21 (2):476–495, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bass:2019:DUS

[BS19] Mark R. Bass and Sujit K. Sahu. Dynamically updated spatially varying parameterizations of hierarchical Bayesian models for spatial data. Journal of Computational and Graphical Statistics, 28 (1):105–116, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Buja:2008:DVM

[BSL^{+}08] Andreas Buja, Deborah F. Swayne, Michael L. Littman, Nathaniel Dean, Heike Hofmann, and Lisha Chen. Data visualization with multidimensional scaling. Journal of Computational and Graphical Statistics, 17 (2):444–472, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bolin:2023:CBR

[BSX23] David Bolin, Alexandre B. Simas, and Zhen Xiong. Covariance-based rational approximations of fractional SPDEs for computationally efficient Bayesian inference. Journal of Computational and Graphical Statistics, 33 (1):64–74, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2231051.

Bi:2022:RFA

[BSZ22] Jiefeng Bi, Weining Shen, and Weixuan Zhu. Random forest adjustment for approximate Bayesian computation. Journal of Computational and Graphical Statistics, 31 (1):64–73, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1981341.

Berger:2018:TSC

[BT18] Moritz Berger and Gerhard Tutz. Tree-structured clustering in fixed effects models. Journal of Computational and Graphical Statistics, 27(2):380–392, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Buja:2000:ISI

[Buj00]  Scandreas Buja.  In-

troduction: Special issue on systems and languages. Journal of Computational and Graphical Statistics, 9(3):403, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Buja.
                                        htm.

Buja:2004:DED

[Buj04] Andreas Buja. Discussion: Exploratory data analysis for complex models. Journal of Computational and Graphical Statistics, 13(4):780–784, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bennett:2020:BRM

[BVZ20] Magdalena Bennett, Juan Pablo Vielma, and José R. Zubizarreta. Building representative matched samples with multi-valued treatments in large observational studies. Journal of Computational and Graphical Statistics, 29 (4):744–757, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1753532.

Bryan:2017:DST

[BW17] Jennifer Bryan and Hadley Wickham. Data science: A three ring circus or a big tent? Journal of Computational and Graphical Statistics, 26 (4):784–785, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Buckby:2020:MCH

[BWZO20] Jodie Buckby, Ting Wang, Jiancang Zhuang, and Kazushige Obara. Model checking for hidden Markov models. Journal of Computational and Graphical Statistics, 29(4):859– 874, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1743295.

Ben:2004:QQP

[BY04] Marta García Ben and Víctor J. Yohai. Quantile– quantile plot for deviance residuals in the generalized linear model. Journal of Computational and Graphical Statistics, 13 (1):36–47, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Barter:2018:SRP

[BY18] Rebecca L. Barter and

                                        Bin Yu. Superheat: An

R package for creating beautiful and extendable heatmaps for visualizing complex data. Journal of Computational and Graphical Statistics, 27 (4):910–922, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Bradlow:1997:CIA

[BZ97] Eric T. Bradlow and Alan M. Zaslavsky. Case influence analysis in Bayesian inference. Journal of Computational and Graphical Statistics, 6(3):314– 331, September 1997. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts/bradlow.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474745.

Barbu:2007:GSW

[BZ07] Adrian Barbu and Song-Chun Zhu. Generalizing Swendsen–Wang for image analysis. Journal of Computational and Graphical Statistics, 16 (4):877–900, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cai:1999:CAS

[Cai99] Haiyan Cai. Cluster algorithms for spatial point processes with symmetric and shift invariant interactions. Journal of Computational and Graphical Statistics, 8(2):353– 372, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/../../../publications/
jcgs/pdf99/cai.pdf;
http://www.amstat.org/
publications/jcgs/abstracts99/
algorithms.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474818.

Canty:1999:HTC

[Can99] Angelo J. Canty. Hypothesis tests of convergence in Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 8(1):93–108, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/upcoming.html;
                                        http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1999.10474803.

Cappe:2011:OEA

[Cap11] Olivier Cappé. Online EM algorithm for hidden Markov models. Journal of Computational and Graphical Statistics, 20 (3):728–749, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Colombi:2023:LBS

[CAPP23] Alessandro Colombi, Raffaele Argiento, Lucia Paci, and Alessia Pini. Learning block structured graphs in Gaussian graphical models. Journal of Computational and Graphical Statistics, 33 (1):152–165, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2023.2210184.

Carr:1994:D

[Car94] Daniel B. Carr. Discussion. Journal of Computational and Graphical Statistics, 3(4):369– 376, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474652.

Chen:2010:NAD

[CAR^{+}10] Y. Ann Chen, Jonas S.

Almeida, Adam J. Richards, Peter Müller, Raymond J. Carroll, and Baerbel Rohrer. A nonparametric approach to detect nonlinear correlation in gene expression. Journal of Computational and Graphical Statistics, 19 (3):552–568, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cook:1997:MCH

[CB97] Dianne Cook and Andreas Buja. Manual controls for high-dimensional data projections. Journal of Computational and Graphical Statistics, 6(4):464–480, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts/cook.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1997.10474754.

Caffo:2001:MCM

[CB01] Brian S. Caffo and James G. Booth. A Markov chain Monte Carlo algorithm for approximating exact conditional probabilities. Journal of Computational and Graphical Statistics, 10(4):730–745, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317243421.

Chen:2020:VIC

[CB20] Shuxiao Chen and Jacob Bien. Valid inference corrected for outlier removal. Journal of Computational and Graphical Statistics, 29(2):323–334, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1660180.

Cook:1993:PPI

[CBC93] Dianne Cook, Andreas Buja, and Javier Cabrera. Projection pursuit indexes based on orthonormal function expansions. Journal of Computational and Graphical Statistics, 2(3):225–250, September 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474610.

Cook:1995:GTP

[CBCH95] Dianne Cook, Andreas Buja, Javier Cabrera, and Catherine Hurley. Grand tour and projection pursuit. Journal of Computational and Graphical Statistics, 4(3):155– 172, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474674.

Corneli:2020:CCO

[CBL20] Marco Corneli, Charles Bouveyron, and Pierre Latouche. Co-clustering of ordinal data via latent continuous random variables and not missing at random entries. Journal of Computational and Graphical Statistics, 29 (4):771–785, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1739533.

Caron:2022:SBC

[CBM22] Alberto Caron, Gianluca Baio, and Ioanna Manolopoulou. Shrinkage Bayesian causal forests for heterogeneous treatment effects estimation. Journal of Computational and Graphical Statistics, 31 (4):1202–1214, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2067549.

Clark-Boucher:2024:PVA

[CBM24] Dylan Clark-Boucher and Jeffrey W. Miller. Principal variables analysis for non-Gaussian data. Journal of Computational and Graphical Statistics, 34 (1):374–383, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2367098.

Chen:2024:IDS

[CBS24] Hanxiao Chen, W. John Braun, and Xiaoping Shi. Iterated data sharpening. Journal of Computational and Graphical Statistics, 34(1):162–174, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2024.2362219.

Chakraborty:2008:OPR

[CC08] Biman Chakraborty and Probal Chaudhuri. On an optimization problem in robust statistics. Journal of Computational and Graphical Statistics, 17 (3):683–702, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chambers:2013:REB

[CC13] Raymond Chambers and Hukum Chandra. A random effect block bootstrap for clustered data. Journal of Computational and Graphical Statistics, 22(2):452–470, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2018:ESA

[CC18] Yinghan Chen and Yuguo Chen. An efficient sampling algorithm for network Motif detection. Journal of Computational and Graphical Statistics, 27(3):503–515, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chi:2022:UFC

[CC22a] Jocelyn T. Chi and Eric C. Chi. A userfriendly computational framework for robust structured regression with the l_{2} criterion. Journal of Computational and Graphical Statistics, 31 (4):1051–1062, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2035232.

Cremona:2022:PML

[CC22b] Marzia A. Cremona and Francesca Chiaromonte. Probabilistic K-means with local alignment for clustering and motif discovery in functional data. Journal of Computational and Graphical Statistics, 32(3):1119–1130, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2156522. See cor-

rection [Ano24a].

Celeux:2001:CWE

[CCFM01] Gilles Celeux, Stéphane Chrétien, Florence Forbes, and Abdallah Mkhadri. A component-wise EM algorithm for mixtures. Journal of Computational and Graphical Statistics, 10(4):697–712, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243403.

Cheng:2016:EID

[CCH16] Xiaoyue Cheng, Dianne Cook, and Heike Hofmann. Enabling interactivity on displays of multivariate time series and longitudinal data. Journal of Computational and Graphical Statistics, 25 (4):1057–1076, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chowdhury:2018:MLD

[CCHM18] Niladri Roy Chowdhury, Dianne Cook, Heike Hofmann, and Mahbubul Majumder. Measuring lineup difficulty by matching distance metrics with subject choices in crowdsourced data. Journal of Computational and Graphical Statistics, 27 (1):132–145, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Castillo:2008:LSA

[CCHS08] Enrique Castillo, Carmen Castillo, Ali S. Hadi, and Jose Maria Sarabia. Local sensitivity analysis in estimation problems. Journal of Computational and Graphical Statistics, 17 (3):703–725, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cui:2021:AFC

[CCL21] Erjia Cui, Ciprian M. Crainiceanu, and Andrew Leroux. Additive functional Cox model. Journal of Computational and Graphical Statistics, 30 (3):780–793, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1853550.

Cao:2003:NSI

[CCS03] Jin Cao, William S. Cleveland, and Don X. Sun. The S-Net system for Internet packet streams: Strategies for stream analysis and system architecture. Journal of Computational and Graphical Statistics, 12(4):865–892, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/1061860032607.

Cai:2022:HKP

[CCX22] Chencheng Cai, Rong Chen, and Han Xiao. Hybrid Kronecker product decomposition and approximation. Journal of Computational and Graphical Statistics, 32 (3):838–852, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2134873.

Chen:2016:BSG

[CCYW16] Ray-Bing Chen, Chi-Hsiang Chu, Shinsheng Yuan, and Ying Nian Wu. Bayesian sparse group selection. Journal of Computational and Graphical Statistics, 25 (3):665–683, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1041636.

Caron:2012:EBI

[CD12] François Caron and Arnaud Doucet. Efficient Bayesian inference for generalized Bradley– Terry models. Journal of Computational and Graphical Statistics, 21 (1):174–196, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cui:2014:GDF

[CD14] Kai Cui and David B.

Dunson. Generalized dynamic factor models for mixed-measurement time series. Journal of Computational and Graphical Statistics, 23(1):169–191, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2019:INT

[CD19] Zhe Chen and Hani Doss. Inference for the number of topics in the latent Dirichlet allocation model via Bayesian mixture modeling. Journal of Computational and Graphical Statistics, 28 (3):567–585, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Craiu:2011:CLR

[CDFB11] Radu V. Craiu, Thierry Duchesne, Daniel Fortin, and Sophie Baillargeon. Conditional logistic regression with longitudinal follow-up and individuallevel random coefficients: a stable and efficient twostep estimation method. Journal of Computational and Graphical Statistics, 20(3):767–784, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cao:2022:SCP

[CDG22] Jian Cao, Daniele Durante, and Marc G. Genton. Scalable computation of predictive probabilities in probit models with Gaussian process priors. Journal of Computational and Graphical Statistics, 31(3):709– 720, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2036614.

Cai:2020:BSB

[CDM^{+}20] Mingxuan Cai, Mingwei Dai, Jingsi Ming, Heng Peng, Jin Liu, and Can Yang. BIVAS: A scalable Bayesian method for bi-level variable selection with applications. Journal of Computational and Graphical Statistics, 29 (1):40–52, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1624365.

Cahya:2004:CCR

[CDP04] Suntara Cahya, Enrique Del Castillo, and John J. Peterson. Computation of confidence regions for optimal factor levels in constrained response surface problems. Journal of Computational and Graphical Statistics, 13 (2):499–518, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Christen:2005:MCM

[CF05] J. Andrés Christen and Colin Fox. Markov chain Monte Carlo using an approximation. Journal of Computational and Graphical Statistics, 14 (4):795–810, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chib:2007:SME

[CG07] Siddhartha Chib and Edward Greenberg. Semiparametric modeling and estimation of instrumental variable models. Journal of Computational and Graphical Statistics, 16 (1):86–114, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cappozzo:2022:GCT

[CGEGMI22] Andrea Cappozzo, Luis Angel García-Escudero, Francesca Greselin, and Agustín Mayo-Iscar. Graphical and computational tools to guide parameter choice for the cluster weighted robust model. Journal of Computational and Graphical Statistics, 32 (3):1195–1214, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2154218.

Cerioli:2018:FNN

[CGEMIR18] Andrea Cerioli, Luis Angel García-Escudero, Agustín Mayo-Iscar, and Marco Riani. Finding the number of normal groups in model-based clustering via constrained likelihoods. Journal of Computational and Graphical Statistics, 27(2):404–416, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cheng:1999:NDE

[CGH99] Ming-Yen Cheng, Theo Gasser, and Peter Hall. Nonparametric density estimation under unimodality and monotonicity constraints. Journal of Computational and Graphical Statistics, 8(1): 1–21, March 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/upcoming.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474798.

Carey:2017:GSL

[CGH17] Michelle Carey, Eugene G. Gath, and Kevin Hayes. A generalized smoother for linear ordinary differential equations. Journal of Computational and Graphical Statistics, 26(3):671–681, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chib:2009:ESM

[CGJ09] Siddhartha Chib, Edward Greenberg, and Ivan Jeliazkov. Estimation of semiparametric models in the presence of endogeneity and sample selection. Journal of Computational and Graphical Statistics, 18(2):321–348, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Clyde:2011:BAS

[CGL11] Merlise A. Clyde, Joyee Ghosh, and Michael L. Littman. Bayesian adaptive sampling for variable selection and model averaging. Journal of Computational and Graphical Statistics, 20(1):80–101, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cecere:2013:ICB

[CGL13] Silvia Cecere, Patrick J. F. Groenen, and Emmanuel Lesaffre. The interval-censored biplot. Journal of Computational and Graphical Statistics, 22(1):123–134, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2021:NVS

[CGLW21] Yao Chen, Qingyi Gao, Faming Liang, and Xiao Wang. Nonlinear variable selection via deep neural networks. Journal of Computational and Graphical Statistics, 30 (2):484–492, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1814305.

Cappe:2004:PMC

[CGMR04] O. Cappé, A. Guillin, J. M. Marin, and C. P. Robert. Population Monte Carlo. Journal of Computational and Graphical Statistics, 13 (4):907–929, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cook:2006:VSB

[CGR06] Samantha R. Cook, Andrew Gelman, and Donald B. Rubin. Validation of software for Bayesian models using posterior quantiles. Journal of Computational and Graphical Statistics, 15 (3):675–692, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See correction [Gel17].

Chiquet:2017:FTI

[CGR17] Julien Chiquet, Pierre Gutierrez, and Guillem Rigaill. Fast tree inference with weighted fusion penalties. Journal of Computational and Graphical Statistics, 26 (1):205–216, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1096789.

Church:1993:DPE

[CH93] Kenneth Ward Church and Jonathan Isaac Helfman. Dotplot: a program for exploring selfsimilarity in millions of lines of text and code. Journal of Computational and Graphical Statistics, 2(2):153–174, 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474605.

Choi:1999:NAA

[CH99] Edwin Choi and Peter Hall. Nonparametric approach to analysis of space-time data on earthquake occurrences. Journal of Computational and Graphical Statistics, 8(4):733–748, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/choi.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474846.

Chen:2009:HRN

[CH09] Jie Chen and Xiaoming Huo. A Hessian regularized nonlinear time series model. Journal of Computational and Graphical Statistics, 18(3):694–716, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chakerian:2012:CTE

[CH12] John Chakerian and Susan Holmes. Computational tools for evaluating phylogenetic and hierarchical clustering trees. Journal of Computational and Graphical Statistics, 21(3):581–599, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2016:NES

[CH16] Feng Chen and Peter Hall. Nonparametric estimation for self-exciting point processes — a parsimonious approach. Journal of Computational and Graphical Statistics, 25 (1):209–224, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.1001491;   http:
//www.tandfonline.com/
doi/full/10.1080/10618600.
2014.1001491.

Chambers:2000:UPS

[Cha00] John M. Chambers. Users, programmers, and statistical software. Journal of Computational and Graphical Statistics, 9(3):404–422, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Chambers.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474890.

Chang:2022:PSS

[Cha22] Ming-Chung Chang. Predictive subdata selection for computer models. Journal of Computational and Graphical Statistics, 32(2):613–630, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2097247.

Chang:2024:SSS

[Cha24] Ming-Chung Chang. Supervised stratified subsampling for predictive analytics. Journal of Computational and Graphical Statistics, 33(3): 1017–1036, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2304075.

Chen:2007:FSA

[Che07a] Colin Chen. A finite smoothing algorithm for quantile regression. Journal of Computational and Graphical Statistics, 16 (1):136–164, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2007:CIT

[Che07b] Yuguo Chen. Conditional inference on tables with structural zeros. Journal of Computational and Graphical Statistics, 16 (2):445–467, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2023:VSE

[Che23] Li-Pang Chen. Variable selection and estimation for misclassified binary responses and multivariate error-prone predictors. Journal of Computational and Graphical Statistics, 33(2):407– 420, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2218428.

Castruccio:2016:HOC

[CHG16] Stefano Castruccio, Raphaël Huser, and Marc G. Genton. High-order composite likelihood inference for max-stable distributions and processes. Journal of Computational and Graphical Statistics, 25 (4):1212–1229, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chang:2010:SES

[CHH10] Ya-Mei Chang, Nan-Jung Hsu, and Hsin-Cheng Huang. Semiparametric estimation and selection for nonstationary spatial covariance functions. Journal of Computational and Graphical Statistics, 19(1):117–139, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2022:TBR

[CHLZ22] Kan Chen, Siyu Heng, Qi Long, and Bo Zhang. Testing biased randomization assumptions and quantifying imperfect matching and residual confounding in matched observational studies. Journal of Computational and Graphical Statistics, 32 (2):528–538, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2116447.

Centofanti:2024:MSS

[CHPR24] Fabio Centofanti, Mia Hubert, Biagio Palumbo, and Peter J. Rousseeuw. Multivariate singular spectrum analysis by robust diagonalwise low-rank approximation. Journal of Computational and Graphical Statistics, 34 (1):360–373, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2362222.

Christensen:2004:MCM

[Chr04] Ole F. Christensen. Monte Carlo maximum likelihood in model-based geostatistics. Journal of Computational and Graphical Statistics, 13 (3):702–718, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chi:2019:GGI

[CHSR19] Eric C. Chi, Liuyi Hu, Arvind K. Saibaba, and Arvind U. K. Rao. Going off the grid: Iterative model selection for biclustered matrix completion. Journal of Computational and Graphical Statistics, 28(1):36–47, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cai:2002:MMB

[CHW02] T. Cai, R. J. Hyndman, and M. P. Wand. Mixed model-based hazard estimation. Journal of Computational and Graphical Statistics, 11(4):784–798, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s3/p784;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002862.

Chen:2008:LRA

[CHW08] Colin Chen, Xuming He, and Ying Wei. Lower rank approximation of matrices based on fast and robust alternating regression. Journal of Computational and Graphical Statistics, 17(1):186–200, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cao:2012:PNL

[CHW12] Jiguo Cao, Jianhua Z.

Huang, and Hulin Wu. Penalized nonlinear least squares estimation of time-varying parameters in ordinary differential equations. Journal of Computational and Graphical Statistics, 21 (1):42–56, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cao:2023:SLR

[CHZ23] Jiahao Cao, Shiyuan He, and Bohai Zhang. Spatial linear regression with covariate measurement errors: Inference and scalable computation in a functional modeling approach. Journal of Computational and Graphical Statistics, 32(4):1588– 1599, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2175686.

Chen:2002:SAA

[CJ02] Jian-Shen Chen and Robert I. Jennrich. Simple accurate approximation of likelihood profiles. Journal of Computational and Graphical Statistics, 11(3):714–732, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s12/p714;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002493.

Chan:2009:MER

[CJ09] Joshua Chi-Chun Chan and Ivan Jeliazkov. MCMC estimation of restricted covariance matrices. Journal of Computational and Graphical Statistics, 18 (2):457–480, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cho:2022:ICR

[CJK22] Hunyong Cho, Nicholas P. Jewell, and Michael R. Kosorok. Interval censored recursive forests. Journal of Computational and Graphical Statistics, 31(2):390–402, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1987253.

Consultant:1992:SOG

[CJS92] I. O. Bohachevsky Consultant, M. E. Johnson, and M. L. Stein. Stochastic optimization and the Gambler’s Ruin problem. Journal of Computational and Graphical Statistics, 1(4):367–384, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474591.

Chen:2022:MPS

[CJW22] Yiqun Chen, Sean Jewell, and Daniela Witten. More powerful selective inference for the graph fused lasso. Journal of Computational and Graphical Statistics, 32 (2):577–587, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2097246.

Cressie:2008:SDM

[CK08] Noel Cressie and Prasenjit Kapat. Some diagnostics for Markov random fields. Journal of Computational and Graphical Statistics, 17(3):726–749, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Carnegie:2015:AMI

[CKHG15] Nicole Bohme Carnegie, Pavel N. Krivitsky, David R. Hunter, and Steven M. Goodreau. An approximation method for improving dynamic network model fitting. Journal of Computational and Graphical Statistics, 24 (2):502–519, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2014.903087.

Chan:2006:LAS

[CKNK06] David Chan, Robert Kohn, David Nott, and Chris Kirby. Locally adaptive semiparametric estimation of the mean and variance functions in regression models. Journal of Computational and Graphical Statistics, 15 (4):915–936, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cardot:2004:BDE

[CKPT04] Hervé Cardot, Ja-Yong Koo, Heon Jin Park, and Alain Trubuil. Boosting diracs for electrophoresis. Journal of Computational and Graphical Statistics, 13(3):659–673, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cho:2019:IIA

[CKR19] Juhee Cho, Donggyu Kim, and Karl Rohe. Intelligent initialization and adaptive thresholding for iterative matrix completion: Some statistical and algorithmic theory for Adaptive–Impute. Journal of Computational and Graphical Statistics, 28 (2):323–333, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2021:TCC

[CKT21] You-Lin Chen, Mladen Kolar, and Ruey S. Tsay. Tensor canonical correlation analysis with convergence and statistical guarantees. Journal of Computational and Graphical Statistics, 30(3):728– 744, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1856118.

Chan:2004:L

[CL04] Kin-Yee Chan and Wei-Yin Loh. LOTUS. Journal of Computational and Graphical Statistics, 13 (4):826–852, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cook:2009:DRR

[CL09] R. Dennis Cook and Lexin Li. Dimension reduction in regressions with exponential family predictors.

Journal of Computational and Graphical Statistics, 18(3):774–791, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2012:SWC

[CL12] Jien Chen and Nicole A. Lazar. Selection of working correlation structure in generalized estimating equations via empirical likelihood. Journal of Computational and Graphical Statistics, 21 (1):18–41, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chi:2015:SMC

[CL15] Eric C. Chi and Kenneth Lange. Splitting methods for convex clustering. Journal of Computational and Graphical Statistics, 24(4):994– 1013, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2014.948181.

Cui:2022:FMF

[CLCX22] Erjia Cui, Ruonan Li, Ciprian M. Crainiceanu, and Luo Xiao. Fast multilevel functional principal component analysis. Journal of Computational and Graphical Statistics, 32(2):366–377, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2115500.

Cleveland:1993:MSD

[Cle93a] William S. Cleveland. A model for studying display methods of statistical graphics. Journal of Computational and Graphical Statistics, 2(4): 323–343, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474616.  See

comments [Tre93, Hol93, Wil93] and rejoinder [Cle93b].

Cleveland:1993:RMS

[Cle93b] William S. Cleveland. Rejoinder: “A Model for Studying Display Methods of Statistical Graphics”. Journal of Computational and Graphical Statistics, 2(4):361– 364, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474620.  See

[Cle93a].

Chen:2019:ONS

[CLHW19] Ray-Bing Chen, Chi-Hao Li, Ying Hung, and Weichung Wang. Optimal noncollapsing space-filling designs for irregular experimental regions. Journal of Computational and Graphical Statistics, 28 (1):74–91, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Clifford:2005:CSC

[Cli05] David Clifford. Computation of spatial covariance matrices. Journal of Computational and Graphical Statistics, 14 (1):155–167, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cheng:2017:HDM

[CLLZ17] Jie Cheng, Tianxi Li, Elizaveta Levina, and Ji Zhu. High-dimensional mixed graphical models. Journal of Computational and Graphical Statistics, 26(2):367–378, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cui:2022:FUI

[CLSC22] Erjia Cui, Andrew Leroux, Ekaterina Smirnova, and Ciprian M. Crainiceanu. Fast univariate inference for longitudinal functional models. Journal of Computational and Graphical Statistics, 31(1):219– 230, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1950006.

Chen:2023:MCV

[CLXY23] Ze Chen, Jun Liao, Wangli Xu, and Yuhong Yang. Multifold cross-validation model averaging for generalized additive partial linear models. Journal of Computational and Graphical Statistics, 32 (4):1649–1659, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2174127.

Culp:2008:IAE

[CM08] Mark Culp and George Michailidis. An iterative algorithm for extending learners to a semisupervised setting. Journal of Computational and Graphical Statistics, 17 (3):545–571, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cucala:2013:BIM

[CM13] Lionel Cucala and Jean-Michel Marin. Bayesian inference on a mixture model with spatial dependence. Journal of Computational and Graphical Statistics, 22(3):584–597, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Christoffersen:2022:QMC

[CMC^{+}22] Benjamin Christoffersen, Behrang Mahjani, Mark Clements, Hedvig Kjellström, and Keith Humphreys. Quasi-Monte Carlo methods for binary event models with complex family data. Journal of Computational and Graphical Statistics, 32(4):1393– 1401, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2151454.

Chevalier:2021:MNE

[CMG21] Clément Chevalier, Olivia Martius, and David Ginsbourger. Modeling nonstationary extreme dependence with stationary max-stable processes and multidimensional scaling. Journal of Computational and Graphical Statistics, 30(3):745–755, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2020.1844213.

Culp:2011:ARM

[CMJ11] Mark Culp, George Michailidis, and Kjell Johnson. On adaptive regularization methods in boosting. Journal of Computational and Graphical Statistics, 20(4):937–955, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cakala:2021:PMP

[CMN21] Tomasz Cakała, Błazej Miasojedow, and Wojciech Niemiro. Particle MCMC with Poisson resampling: Parallelization and continuous time models. Journal of Computational and Graphical Statistics, 30 (3):671–684, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1840998.

Chau:2019:IDD

[COvS19] Joris Chau, Hernando Ombao, and Rainer von Sachs. Intrinsic data depth for Hermitian positive definite matrices. Journal of Computational and Graphical Statistics, 28(2):427–439, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cozman:1999:CPB

[Coz99] Fabio Gagliardi Cozman. Calculation of posterior bounds given convex sets of prior probability measures and likelihood functions. Journal of Computational and Graphical Statistics, 8(4):824–838, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/cozman.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474851.

Chang:2016:RAE

[CP16] Christopher C. Chang and Dimitris N. Politis. Robust autocorrelation estimation. Journal of Computational and Graphical Statistics, 25 (1):144–166, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.969431;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.969431.

Coeurjolly:2018:FES

[CP18] Jean-Francois Coeurjolly and Emilio Porcu. Fast and exact simulation of complex-valued stationary Gaussian processes through embedding circulant matrix. Journal of Computational and Graphical Statistics, 27 (2):278–290, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cappello:2022:APS

[CP22] Lorenzo Cappello and Julia A. Palacios. Adaptive preferential sampling in phylodynamics with an application to SARS-CoV-2. Journal of Computational and Graphical Statistics, 31(2):541– 552, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1987256.

Choi:2024:ICL

[CPCC24] Taehwa Choi, Seohyeon Park, Hunyong Cho, and Sangbum Choi. Intervalcensored linear quantile regression. Journal of Computational and Graphical Statistics, 34 (1):187–198, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2365740.

Cappello:2023:BCP

[CPP23] Lorenzo Cappello, Oscar Hernan Madrid Padilla, and Julia A. Palacios. Bayesian change point detection with spike-and-slab priors. Journal of Computational and Graphical Statistics, 32 (4):1488–1500, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2182312.

Cortes:2003:CMD

[CPV03] Corinna Cortes, Daryl Pregibon, and Chris Volinsky. Computational methods for dynamic graphs. Journal of Computational and Graphical Statistics, 12(4):950–970, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/1061860032742.

Casella:1998:PPA

[CR98] George Casella and Christian P. Robert. Postprocessing accept–reject samples: Recycling and rescaling. Journal of Computational and Graphical Statistics, 7(2):139– 157, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/casella.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474767.

Cerioli:1999:OSD

[CR99] Andrea Cerioli and Marco Riani. The ordering of spatial data and the detection of multiple outliers. Journal of Computational and Graphical Statistics, 8(2):239– 258, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/ordering.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/ceriol.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474812.

Cheng:2008:KMO

[CR08] Ming-Yen Cheng and Marc Raimondo. Kernel methods for optimal change-points estimation in derivatives. Journal of Computational and Graphical Statistics, 17 (1):56–75, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2023:NBS

[CR23] Fan Chen and Karl Rohe.

A new basis for sparse principal component analysis. Journal of Computational and Graphical Statistics, 33(2):421– 434, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2256502.

Chokami:2024:EGT

[CR24] Amir Khorrami Chokami and Giovanni Rabitti. An exact game-theoretic variable importance index for generalized additive models. Journal of Computational and Graphical Statistics, 33(4):1276– 1285, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2327577.

Crainiceanu:2007:SAB

[CRC^{+}07] Ciprian M. Crainiceanu, David Ruppert, Raymond J. Carroll, Adarsh Joshi, and Billy Goodner. Spatially adaptive Bayesian penalized splines with heteroscedastic errors. Journal of Computational and Graphical Statistics, 16 (2):265–288, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Camacho:2017:GWP

[CRGS17] José Camacho, Rafael A.

Rodríguez-Gómez, and Edoardo Saccenti. Groupwise principal component analysis for exploratory data analysis. Journal of Computational and Graphical Statistics, 26 (3):501–512, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chopin:2013:CAB

[CRL13] N. Chopin, J. Rousseau, and B. Liseo. Computational aspects of Bayesian spectral density estimation. Journal of Computational and Graphical Statistics, 22(3):533–557, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Christensen:2006:RMC

[CRS06] Ole F. Christensen, Gareth O.

Roberts, and Martin Sköld. Robust Markov chain Monte Carlo methods for spatial generalized linear mixed models. Journal of Computational and Graphical Statistics, 15(1):1–17, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:1993:PGH

[CS93] Ming-Hui Chen and Bruce Schmeiser. Performance of the Gibbs, hit-and-run, and Metropolis samplers. Journal of Computational and Graphical Statistics, 2(3):251–272, September 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474611.

Chen:1998:TBB

[CS98] Ming-Hui Chen and Bruce Schmeiser. Toward blackbox sampling: a randomdirection interior-point Markov chain approach. Journal of Computational and Graphical Statistics, 7(1):1–22, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL

http://www.research.
att.com/~andreas/jcgs/
abstracts/1998/march/
chen.html; http://www.
tandfonline.com/doi/
abs/10.1080/10618600.
1998.10474758.

Chen:1999:MCE

[CS99] Ming-Hui Chen and Qi-Man Shao. Monte Carlo estimation of Bayesian credible and HPD intervals. Journal of Computational and Graphical Statistics, 8(1):69–92, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/upcoming.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474802.

Chi:2014:RPC

[CS14] Eric C. Chi and David W.

Scott. Robust parametric classification and variable selection by a minimum distance criterion. Journal of Computational and Graphical Statistics, 23(1):111–128, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2018:DLE

[CS18] Feng Chen and Tom Stindl. Direct likelihood evaluation for the renewal Hawkes process. Journal of Computational and Graphical Statistics, 27 (1):119–131, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chatla:2020:TBS

[CS20] Suneel Babu Chatla and Galit Shmueli. A treebased semi-varying coefficient model for the COM-Poisson distribution. Journal of Computational and Graphical Statistics, 29(4):827– 846, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1753530.

Cressie:2010:FRF

[CSK10] Noel Cressie, Tao Shi, and Emily L. Kang. Fixed rank filtering for spatiotemporal data. Journal of Computational and Graphical Statistics, 19 (3):724–745, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Carmack:2009:FCC

[CSS^{+}09] Patrick S. Carmack, William R. Schucany, Jeffrey S. Spence, Richard F. Gunst, Qihua Lin, and Robert W. Haley. Far casting cross-validation. Journal of Computational and Graphical Statistics, 18(4):879–893, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2023:SDC

[CSWZ23] Lanjue Chen, Jin Su, Alan T. K. Wan, and Yong Zhou. A simple divide-and-conquer-based distributed method for the accelerated failure time model. Journal of Computational and Graphical Statistics, 33 (2):681–698, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2252028.

Chadoeuf:2000:PEB

[CSY00] J. Chadoeuf, R. Senoussi, and J. F. Yao. Parametric estimation of a Boolean segment process with stochastic restoration estimation. Journal of Computational and Graphical Statistics, 9(2): 390–402, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474887.

Chu:2001:OCR

[CT01] Moody T. Chu and Nickolay T. Trendafilov. The orthogonally constrained regression revisited. Journal of Computational and Graphical Statistics, 10(4):746–771, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243430.

Chauvet:2019:CBR

[CTB19] Jocelyn Chauvet, Catherine Trottier, and Xavier Bry. Component-based regularization of multivariate generalized linear mixed models. Journal of Computational and Graphical Statistics, 28 (4):909–920, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Cai:2009:VSL

[CTC09] Airong Cai, Ruey S. Tsay, and Rong Chen. Variable selection in linear regression with many predictors. Journal of Computational and Graphical Statistics, 18(3):573–591, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2021:IHE

[CTDKL21] Jingxiang Chen, Quoc Tran-Dinh, Michael R. Kosorok, and Yufeng Liu. Identifying heterogeneous effect using latent supervised clustering with adaptive fusion. Journal of Computational and Graphical Statistics, 30 (1):43–54, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1763808.

Chen:2020:SSO

[CTW20] Tom Chen, Eric J. Tchetgen Tchetgen, and Rui Wang. A stochastic second-order generalized estimating equations approach for estimating association parameters. Journal of Computational and Graphical Statistics, 29(3):547–561, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1710156.

Culp:2013:SJT

[Cul13]  Mark Culp.  On the

semisupervised joint trained elastic net. Journal of Computational and Graphical Statistics, 22 (2):300–318, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chen:2008:AEL

[CVA08] Jiahua Chen, Asokan Mulayath Variyath, and Bovas Abraham. Adjusted empirical likelihood and its properties. Journal of Computational and Graphical Statistics, 17 (2):426–443, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Consagra:2023:EMF

[CVQ23] William Consagra, Arun Venkataraman, and Xing Qiu. Efficient multidimensional functional data analysis using marginal product basis systems. Journal of Computational and Graphical Statistics, 33(2):567–577, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257256.

Coleman:2000:CAL

[CW00] Daniel A. Coleman and David L. Woodruff. Cluster analysis for large datasets: an effective algorithm for maximizing the mixture likelihood. Journal of Computational and Graphical Statistics, 9(4):672–688, 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474907.

Chang:2022:RML

[CW22] Le Chang and A. H. Welsh. Robust multivariate lasso regression with covariance estimation. Journal of Computational and Graphical Statistics, 32(3):961– 973, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2118752.

Cheng:2021:SRD

[CWZ21] Chao Cheng, Rui Wang, and Heping Zhang. Surrogate residuals for discrete choice models. Journal of Computational and Graphical Statistics, 30 (1):67–77, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1775618.

Chakraborty:2023:BC

[CX23] Saptarshi Chakraborty and Jason Xu. Biconvex clustering. Journal of Computational and Graphical Statistics, 32 (4):1524–1536, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2197474.

Chen:2024:RQR

[CY24] Xuerong Chen and Senlin Yuan. Renewable quantile regression with heterogeneous streaming datasets. Journal of Computational and Graphical Statistics, 33(4):1185–1201, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2309343.

Cowles:2009:RMP

[CYS09] Mary Kathryn Cowles, Jun Yan, and Brian Smith. Reparameterized and marginalized posterior and predictive sampling for complex Bayesian geostatistical models. Journal of Computational and Graphical Statistics, 18(2):262–282, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Clarkson:2002:USR

[CZ02] Douglas B. Clarkson and Yihui Zhan. Using spherical-radial quadrature to fit generalized linear mixed effects models. Journal of Computational and Graphical Statistics, 11(3):639–659, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s8/p639;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002439.

Cook:2016:AEE

[CZ16] R. Dennis Cook and Xin Zhang. Algorithms for envelope estimation. Journal of Computational and Graphical Statistics, 25 (1):284–300, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2015.1029577;   http:
//www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1029577.

Chen:2022:DDH

[CZ22] Canyi Chen and Liping Zhu. Distributed decoding from heterogeneous 1bit compressive measurements. Journal of Computational and Graphical Statistics, 32(3):884– 894, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2118751.

Chen:2010:PSL

[CZSL10] Fei Chen, Hong-Tu Zhu, Xin-Yuan Song, and Sik-Yum Lee. Perturbation selection and local influence analysis for generalized linear mixed models. Journal of Computational and Graphical Statistics, 19(4):826–842, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Chun:2015:GRN

[CZZ15] Hyonho Chun, Xianghua Zhang, and Hongyu Zhao. Gene regulation network inference with joint sparse Gaussian graphical models. Journal of Computational and Graphical Statistics, 24(4):954– 974, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.956876.

DAngelo:2023:LIW

[DAMC23] Nicoletta D’Angelo, Giada Adelfio, Jorge Mateu, and Ottmar Cronie. Local inhomogeneous weighted summary statistics for marked point processes. Journal of Computational and Graphical Statistics, 33(2):588–602, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2206441.

Danilov:1997:PCT

[Dan97] D. L. Danilov. Principal components in time series forecast. Journal of Computational and Graphical Statistics, 6(1):112–121, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
danilov.html;   http:
//www.tandfonline.com/

doi/abs/10.1080/10618600.
1997.10474730.

Dannemann:2012:SHM

[Dan12] Jörn Dannemann. Semiparametric hidden Markov models. Journal of Computational and Graphical Statistics, 21(3):677–692, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Daniel:2008:AOT

[DASR08] Shoshana R. Daniel, Katrina Armstrong, Jeffrey H. Silber, and Paul R. Rosenbaum. An algorithm for optimal tapered matching, with application to disparities in survival. Journal of Computational and Graphical Statistics, 17(4):914–924, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Du:2022:VSM

[DBH^{+}22] Jiacong Du, Jonathan Boss, Peisong Han, Lauren J. Beesley, Michael Kleinsasser, Stephen A. Goutman, Stuart Batterman, Eva L. Feldman, and Bhramar Mukherjee. Variable selection with multiply-imputed datasets: Choosing between stacked and grouped methods. Journal of Computational and Graphical Statistics, 31(4):

1063–1075, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2035739.

Dovers:2023:FAM

[DBPW23] Elliot Dovers, Wesley Brooks, Gordana C. Popovic, and David I. Warton. Fast, approximate maximum likelihood estimation of log-Gaussian Cox processes. Journal of Computational and Graphical Statistics, 32(4):1660–1670, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2182311.

Doucet:2006:EBS

[DBS06] Arnaud Doucet, Mark Briers, and Stéphane Sénécal. Efficient block sampling strategies for sequential Monte Carlo methods. Journal of Computational and Graphical Statistics, 15(3):693– 711, September 2006. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

DAngelo:2022:EPS

[DC22] Laura D’Angelo and Antonio Canale. Efficient posterior sampling for Bayesian Poisson regression. Journal of Computational and Graphical Statistics, 32(3):917– 926, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2123337.

Duan:2016:BET

[DCS16] Leo L. Duan, John P.

Clancy, and Rhonda D. Szczesniak. Bayesian ensemble trees (BET) for clustering and prediction in heterogeneous data. Journal of Computational and Graphical Statistics, 25(3):748–761, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1089774.

Duczmal:2008:DIS

[DCT08] Luiz Duczmal, André L. F. Cançado, and Ricardo H. C. Takahashi. Delineation of irregularly shaped disease clusters through multiobjective optimization. Journal of Computational and Graphical Statistics, 17 (1):243–262, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dusseldorp:2010:CAT

[DCV10] Elise Dusseldorp, Claudio Conversano, and Bart Jan Van Os. Combining an additive and tree-based regression model simultaneously: STIMA. Journal of Computational and Graphical Statistics, 19 (3):514–530, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dai:2020:MFL

[DDM20] Fan Dai, Somak Dutta, and Ranjan Maitra. A matrix-free likelihood method for exploratory factor analysis of high-dimensional Gaussian data. Journal of Computational and Graphical Statistics, 29 (3):675–680, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1704296.

DeOliveira:2005:BIP

[De 05] Victor De Oliveira. Bayesian inference and prediction of Gaussian random fields based on censored data. Journal of Computational and Graphical Statistics, 14(1):95–115, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Degras:2023:SFM

[Deg23] David Degras. Scalable feature matching across large data collections. Journal of Computational and Graphical Statistics, 32(1):196–212, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2074429.

Dempsey:2023:REA

[Dem23] Walter Dempsey. Recurrent event analysis in the presence of real-time high frequency data via random subsampling. Journal of Computational and Graphical Statistics, 33 (2):525–537, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2276114.

Devroye:1993:GRI

[DES93] Luc Devroye, Peter Epstein, and Jörg-Rüdiger Sack. On generating random intervals and hyperrectangles. Journal of Computational and Graphical Statistics, 2(3): 291–307, September 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1080/10618600.
                                        1993.10474613.

Ding:2005:CUG

[DG05] Beiying Ding and Robert Gentleman. Classification using generalized partial least squares. Journal of Computational and Graphical Statistics, 14 (2):280–298, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dao:2014:MCA

[DG14] Ngoc Anh Dao and Marc G. Genton. A Monte Carlo-adjusted goodness-of-fit test for parametric models describing spatial point patterns. Journal of Computational and Graphical Statistics, 23(2):497–517, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dai:2018:MFD

[DG18] Wenlin Dai and Marc G. Genton. Multivariate functional data visualization and outlier detection. Journal of Computational and Graphical Statistics, 27(4):923–934, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Donner:2024:SIC

[DG24] A. Donner and A. Goldenshluger. Statistical inference in circular structural model and fitting circles to noisy data. Journal of Computational and Graphical Statistics, 33 (4):1350–1369, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2315953.

DeBlasi:2023:GSM

[DGL23] Pierpaolo De Blasi and María F. Gil-Leyva. Gibbs sampling for mixtures in order of appearance: The ordered allocation sampler. Journal of Computational and Graphical Statistics, 32(4):1416– 1424, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2177298.

Dette:2017:ECB

[DGM17] Holger Dette, Roman Guchenko, and Viatcheslav B. Melas. Efficient computation of Bayesian optimal discriminating designs. Journal of Computational and Graphical Statistics, 26(2):424–433, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dlotko:2024:MTA

[DGS24] Paweł Dłotko, Davide Gurnari, and Radmila Sazdanovic. Mapper-type algorithms for complex data and relations. Journal of Computational and Graphical Statistics, 33 (4):1383–1396, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2343321.

Draghicescu:2009:QCE

[DGW09] Dana Draghicescu, Serge Guillas, and Wei Biao Wu. Quantile curve estimation and visualization for nonstationary time series. Journal of Computational and Graphical Statistics, 18(1):1–20, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dunn:2024:UIM

[DGWR24] Robin Dunn, Aditya Gangrade, Larry Wasserman, and Aaditya Ramdas. Universal inference meets random projections: a scalable test for log-concavity. Journal of Computational and Graphical Statistics, 34 (1):267–279, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2347338.

Domingos:2003:GFM

[DH03a] Pedro Domingos and Geoff Hulten. A general framework for mining massive data streams. Journal of Computational and Graphical Statistics, 12(4):945–949, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032544.

Doss:2003:MCM

[DH03b] Hani Doss and Fred W.

Huffer. Monte Carlo methods for Bayesian analysis of survival data using mixtures of Dirichlet process priors. Journal of Computational and Graphical Statistics, 12(2):282–307, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s3/p282;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031680.

Doss:2010:EBF

[DH10] Hani Doss and James P.

Hobert. Estimation of Bayes factors in a class of hierarchical random effects models using a geometrically ergodic MCMC algorithm. Journal of Computational and Graphical Statistics, 19(2):295–312, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Diaz:1999:ITB

[Dia99] Francisco J. Diaz. Identifying tail behavior by means of residual quantile functions. Journal of Computational and Graphical Statistics, 8(3): 493–509, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474828.

Dryden:2002:AML

[DIR02] Ian Dryden, Luigi Ippoliti, and Luca Romagnoli. Adjusted maximum likelihood and pseudolikelihood estimation for noisy Gaussian Markov random fields. Journal of Computational and Graphical Statistics, 11 (2):370–388, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/

v11n2/s7/p370;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180563.

Dumbgen:2004:CBI

[DJ04] Lutz Dümbgen and Robert B.

Johns. Confidence bands for isotonic median curves using sign tests. Journal of Computational and Graphical Statistics, 13 (2):519–533, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dahl:2023:CBE

[DJA23] David B. Dahl, Devin J.

Johnson, and R. Jacob Andros. Comparison and Bayesian estimation of feature allocations. Journal of Computational and Graphical Statistics, 32 (4):1371–1382, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2204136.

Dahl:2022:SAL

[DJM22] David B. Dahl, Devin J.

Johnson, and Peter Müller. Search algorithms and loss functions for Bayesian clustering. Journal of Computational and Graphical Statistics, 31 (4):1189–1201, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2069779.

DeYoreo:2018:BNM

[DK18] Maria DeYoreo and Athanasios Kottas. Bayesian nonparametric modeling for multivariate ordinal regression. Journal of Computational and Graphical Statistics, 27 (1):71–84, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Duczmal:2006:ESS

[DKH06] Luiz Duczmal, Martin Kulldorff, and Lan Huang. Evaluation of spatial scan statistics for irregularly shaped clusters. Journal of Computational and Graphical Statistics, 15(2):428–442, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dumouchel:1995:D

[DL95] William Dumouchel and Thomas Lane. Discussion. Journal of Computational and Graphical Statistics, 4(4):251–256, 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1995.10474681.

Deng:2023:MIT

[DL23] Yongshi Deng and Thomas Lumley. Multiple imputation through XGBoost. Journal of Computational and Graphical Statistics, 33(2):352–363, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2252501.

deLuna:2001:RSB

[dLG01] Xavier de Luna and Marc G. Genton. Robust simulation-based estimation of ARMA models. Journal of Computational and Graphical Statistics, 10(2):370–387, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628347.

deLeeuw:2000:DOT

[dLM00] Jan de Leeuw and George Michailidis. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 26–31, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.

                                        2000.10474860.  See

[LHY00].

deLeeuw:2004:WCA

[dLM04] Jan de Leeuw and George Michailidis. Weber correspondence analysis. Journal of Computational and Graphical Statistics, 13 (4):946–953, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Deng:2024:SPR

[DLSS24] Yi Deng, Shuwei Li, Liuquan Sun, and Xinyuan Song. Semiparametric probit regression model with general intervalcensored failure time data. Journal of Computational and Graphical Statistics, 33(4):1413– 1423, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2330523.

Denby:2009:VH

[DM09] Lorraine Denby and Colin Mallows. Variations on the histogram. Journal of Computational and Graphical Statistics, 18 (1):21–31, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Doss:2022:NSS

[DM22] Charles R. Doss and Edward Mcfowland III. Nonparametric subset scanning for detection of heteroscedasticity. Journal of Computational and Graphical Statistics, 31 (3):813–823, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2026779.

Drikvandi:2024:DFM

[DM24a] Reza Drikvandi and Reza Modarres. A distributionfree method for change point detection in nonsparse high dimensional data. Journal of Computational and Graphical Statistics, 34(1):290– 305, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2365733.

Dvorak:2024:NTC

[DM24b] Jirí Dvorák and Tomás Mrkvicka. Nonparametric testing of the covariate significance for spatial point patterns under the presence of nuisance covariates. Journal of Computational and Graphical Statistics, 33 (4):1434–1445, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2357626.

Drovandi:2014:SMC

[DMP14] Christopher C. Drovandi, James M. McGree, and Anthony N. Pettitt. A sequential Monte Carlo algorithm to incorporate model uncertainty in Bayesian sequential design. Journal of Computational and Graphical Statistics, 23(1):3–24, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dettoni:2020:GLB

[DMR20] Robinson Dettoni, Giampiero Marra, and Rosalba Radice. Generalized link-based additive survival models with informative censoring. Journal of Computational and Graphical Statistics, 29 (3):503–512, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1724544. See cor-

rection [Ano20].

Drovandi:2023:IAM

[DNF23] Christopher Drovandi, David J. Nott, and David T. Frazier. Improving the accuracy of marginal approximations in likelihood-free inference via localization. Journal of Computational and Graphical Statistics, 33(1):101–111, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2223574.

Dobra:2012:DMB

[Dob12] Adrian Dobra. Dynamic Markov bases. Journal of Computational and Graphical Statistics, 21 (2):496–517, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Donoho:2017:YDS

[Don17] David Donoho. 50 years of data science. Journal of Computational and Graphical Statistics, 26 (4):745–766, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comments [Pen17, HJ17].

Dallakyan:2023:FLR

[DP23] Aramayis Dallakyan and Mohsen Pourahmadi. Fusedlasso regularized Cholesky factors of large nonstationary covariance matrices of replicated time series. Journal of Computational and Graphical Statistics, 32(1):157– 170, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2090367.

Du:2023:ECV

[DPRK23] Jin-Hong Du, Pratik Patil, Kathryn Roeder, and Arun Kumar Kuchibhotla. Extrapolated crossvalidation for randomized ensembles. Journal of Computational and Graphical Statistics, 33 (3):1061–1072, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2288194.

Dey:2011:GTV

[DPS11] Tanujit Dey, David J. Phillips, and Patrick Steele. A graphical tool to visualize predicted minimum delay flights. Journal of Computational and Graphical Statistics, 20 (2):294–297, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

deRooij:2007:DPG

[dR07] Mark de Rooij. The distance perspective of generalized biadditive models. Journal of Computational and Graphical Statistics, 16(1):210–227, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dazard:2010:LSB

[DR10] Jean-Eudes Dazard and J. Sunil Rao. Local sparse bump hunting. Journal of Computational and Graphical Statistics, 19 (4):900–929, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dalzell:2018:RMF

[DR18] Nicole M. Dalzell and Jerome P. Reiter. Regression modeling and file matching using possibly erroneous matching variables. Journal of Computational and Graphical Statistics, 27(4):728–738, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Deshpande:2019:SVC

[DRG19] Sameer K. Deshpande, Veronika Rocková, and Edward I. George. Simultaneous variable and covariance selection with the multivariate spikeand-slab LASSO. Journal of Computational and Graphical Statistics, 28 (4):921–931, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Drignei:2008:FSS

[Dri08]  Dorin Drignei.  Fast

statistical surrogates for dynamical 3D computer models of brain tumors. Journal of Computational and Graphical Statistics, 17(4):844–859, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dong:1994:CPB

[DS94] Jianping Dong and Jeffrey S. Simonoff. The construction and properties of boundary kernels for smoothing sparse multinomials. Journal of Computational and Graphical Statistics, 3(1): 57–66, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474630.

Dunn:1996:RQR

[DS96] Peter K. Dunn and Gordon K. Smyth. Randomized quantile residuals. Journal of Computational and Graphical Statistics, 5(3):236–244, 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474708.

Davison:2000:PSM

[DS00] A. C. Davison and S. Sardy. The partial scatterplot matrix. Journal of Computational and Graphical Statistics, 9(4):750–758, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Davison.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474911.

Duchamp:2003:SSS

[DS03] Tom Duchamp and Werner Stuetzle. Spline smoothing on surfaces. Journal of Computational and Graphical Statistics, 12(2):354–381, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s6/p354;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031743.

Descloux:2021:MSL

[DS21] Pascaline Descloux and Sylvain Sardy. Model selection with lasso-zero:

Adding straw to the haystack to better find needles. Journal of Computational and Graphical Statistics, 30(3):530– 543, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1869026.

daSilva:2021:PPF

[dSCL21] Natalia da Silva, Dianne Cook, and Eun-Kyung Lee. A projection pursuit forest algorithm for supervised classification. Journal of Computational and Graphical Statistics, 30 (4):1168–1180, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1870480.

Das:2023:CSD

[DSD^{+}23] Priyam Das, Deborshee Sen, Debsurya De, Jue Hou, Zahra S. H. Abad, Nicole Kim, Zongqi Xia, and Tianxi Cai. Clustering sequence data with mixture Markov chains with covariates using multiple simplex constrained optimization routine (MSiCOR). Journal of Computational and Graphical Statistics, 33 (2):379–392, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257258.

deSousa:2004:DPE

[dSM04] Bruno de Sousa and George Michailidis. A diagnostic plot for estimating the tail index of a distribution. Journal of Computational and Graphical Statistics, 13 (4):974–995, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Deng:2013:PCM

[DT13] Xinwei Deng and Kam-Wah Tsui. Penalized covariance matrix estimation using a matrixlogarithm transformation. Journal of Computational and Graphical Statistics, 22(2):494–512, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

DeCanditiis:2004:WBB

[DV04] Daniela De Canditiis and Brani Vidakovic. Wavelet Bayesian block shrinkage via mixtures of normalinverse gamma priors. Journal of Computational and Graphical Statistics, 13(2):383–398, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

deValpine:2008:IEN

[dV08] Perry de Valpine. Improved estimation of normalizing constants from Markov chain Monte Carlo output. Journal of Computational and Graphical Statistics, 17 (2):333–351, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

deValpine:2017:PMW

[dVTP^{+}17] Perry de Valpine, Daniel Turek, Christopher J. Paciorek, Clifford Anderson-Bergman, Duncan Temple Lang, and Rastislav Bodik. Programming with models: Writing statistical algorithms for general model structures with NIMBLE. Journal of Computational and Graphical Statistics, 26 (2):403–413, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Damien:2001:STN

[DW01] Paul Damien and Stephen G. Walker. Sampling truncated normal, beta, and gamma densities. Journal of Computational and Graphical Statistics, 10 (2):206–215, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152627906.

Demirel:2002:TTB

[DW02] Omer F. Demirel and Thomas R. Willemain. A Turing test of bootstrap scenarios. Journal of Computational and Graphical Statistics, 11(4):896–909, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s10/p896;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002691.

Daye:2012:SSS

[DXL12] Z. John Daye, Jichun Xie, and Hongzhe Li. A sparse structured shrinkage estimator for nonparametric varying-coefficient model with an application in genomics. Journal of Computational and Graphical Statistics, 21(1):110–133, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Deng:2009:LGC

[DY09] Xinwei Deng and Ming Yuan. Large Gaussian covariance matrix estimation with Markov structures. Journal of Computational and Graphical Statistics, 18(3):640–657, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Deng:2023:BDM

[DYB^{+}23] Lih-Yuan Deng, Ching-Chi Yang, Dale Bowman, Dennis K. J. Lin, and Henry Horng-Shing Lu. Big data model building using dimension reduction and sample selection. Journal of Computational and Graphical Statistics, 33(2):435–447, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2260052.

DiRienzo:1999:SAN

[DZ99] A. Gregory DiRienzo and Igor G. Zurbenko. Semi-adaptive nonparametric spectral estimation. Journal of Computational and Graphical Statistics, 8(1):41–59, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/upcoming.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474800.

Doubleday:2018:AGI

[DZFZ18] Kevin Doubleday, Hua Zhou, Haoda Fu, and Jin Zhou. An algorithm for generating individualized treatment decision trees and random forests. Journal of Computational and Graphical Statistics, 27 (4):849–860, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Dong:2022:SAL

[DZL22] Tianning Dong, Peiyi Zhang, and Faming Liang. A stochastic approximationlangevinized ensemble Kalman filter algorithm for state space models with unknown parameters. Journal of Computational and Graphical Statistics, 32(2):448– 469, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2107531.

Daniels:2006:CSS

[DZZ06] Michael J. Daniels, Zhigang Zhou, and Hui Zou. Conditionally specified space-time models for multivariate processes. Journal of Computational and Graphical Statistics, 15(1):157–177, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Davis:2016:SVA

[DZZ16] Richard A. Davis, Pengfei Zang, and Tian Zheng. Sparse vector autoregressive modeling. Journal of Computational and Graphical Statistics, 25 (4):1077–1096, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Everson:2002:BIB

[EB02] Philip J. Everson and Eric T. Bradlow. Bayesian inference for the betabinomial distribution via polynomial expansions. Journal of Computational and Graphical Statistics, 11(1):202–207, March 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n1/s10/p202;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002317375686.

Eo:2014:TSM

[EC14] Soo-Heang Eo and HyungJun Cho. Tree-structured mixed-effects regression modeling for longitudinal data. Journal of Computational and Graphical Statistics, 23(3):740–760, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eisinger:2017:SCI

[EC17] Robert D. Eisinger and Yuguo Chen. Sampling for conditional inference on contingency tables. Journal of Computational and Graphical Statistics, 26(1):79–87, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1153478.

Eddy:1992:E

[Edd92] William F. Eddy. Editorial. Journal of Computational and Graphical Statistics, 1(1):1–2, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474572.

Editor:2000:ISI

[Edi00] Andreas Buja Editor. Introduction: Special issue on systems and languages. Journal of Computational and Graphical Statistics, 9(3):403, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/abs/10.1080/10618600.
2000.10474889.

Eddy:1999:CFM

[EFG^{+}99] William F. Eddy, Mark Fitzgerald, Christopher Genovese, Nicole Lazar, Audris Mockus, and Joel Welling. The challenge of functional magnetic resonance imaging. Journal of Computational and Graphical Statistics, 8(3): 545–558, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474832.

Esposito:1994:AVP

[EFLT94] Richard Esposito, James K. Fox, Dongyow Lin, and Kevin Tidemann. ARIES: a visual path in the investigation of statistical data. Journal of Computational and Graphical Statistics, 3(2):113– 125, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474634.

Efromovich:2005:LPA

[Efr05] Sam Efromovich. On logarithmic penalty in adaptive pointwise estimation and blockwise wavelet shrinkage. Journal of Computational and Graphical Statistics, 14 (1):20–40, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Erlandson:2024:SLS

[EGCP24] Lucas Erlandson, Ana María Estrada Gómez, Edmond Chow, and Kamran Paynabar. smashGP: Large-scale spatial modeling via matrix-free Gaussian processes. Journal of Computational and Graphical Statistics, 34 (1):15–33, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2353653.

Evans:1992:NAE

[EGO92] Michael Evans, Irwin Guttman, and Ingram Olkin. Numerical aspects in estimating the parameters of a mixture of normal distributions. Journal of Computational and Graphical Statistics, 1(4): 351–365, December 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474590.

Einmahl:2010:SP

[EGS10] John H. J. Einmahl, Maria Gantner, and Günther Sawitzki. The shorth plot. Journal of Computational and Graphical Statistics, 19 (1):62–73, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Emerson:2013:GPP

[EGS^{+}13] John W. Emerson, Walton A. Green, Barret Schloerke, Jason Crowley, Dianne Cook, Heike Hofmann, and Hadley Wickham. The generalized pairs plot. Journal of Computational and Graphical Statistics, 22 (1):79–91, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comment [Fri14].

Erasto:2005:BMS

[EH05] Panu Erästö and Lasse Holmström. Bayesian multiscale smoothing for making inferences about features in scatterplots. Journal of Computational and Graphical Statistics, 14(3):569–589, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Evers:2009:LAT

[EH09] L. Evers and T. J. Heaton. Locally adaptive treebased thresholding. Journal of Computational and Graphical Statistics, 18 (4):961–977, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Earle:2015:ADS

[EH15] Denise Earle and Catherine B. Hurley. Advances in dendrogram seriation for application to visualization. Journal of Computational and Graphical Statistics, 24(1):1–25, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Earls:2017:CFD

[EH17] Cecilia Earls and Giles Hooker. Combining functional data registration and factor analysis. Journal of Computational and Graphical Statistics, 26 (2):296–305, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eubank:2003:AOS

[EHW03] R. L. Eubank, Chunfeng Huang, and Suojin Wang. Adaptive order selection for spline smoothing. Journal of Computational and Graphical Statistics, 12(2):382–397, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.

ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s7/p382;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031707.

Eick:1994:GDT

[Eic94] Stephen G. Eick. Graphically displaying text. Journal of Computational and Graphical Statistics, 3(2):127–142, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474635.

Elias:2022:IDP

[EJPS22] Antonio Elías, Raúl Jiménez, Anna M. Paganoni, and Laura M. Sangalli. Integrated depths for partially observed functional data. Journal of Computational and Graphical Statistics, 32(2):341– 352, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2070171.

Eick:2002:VS

[EK02] Stephen G. Eick and Alan F. Karr. Visual scalability. Journal of Computational and Graphical Statistics, 11(1):22– 43, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s2/p22;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375604.

Elder:1994:D

[Eld94] John F. Elder IV. Discussion. Journal of Computational and Graphical Statistics, 3(4):355– 362, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474650.

Elder:2003:GPE

[Eld03] John F. Elder IV. The generalization paradox of ensembles. Journal of Computational and Graphical Statistics, 12(4):853–864, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032733.

Eddy:1996:III

[EM96] William F. Eddy and Audris Mockus. An interactive icon index: Images of the outer planets. Journal of Computational and Graphical Statistics, 5(1):100–111, 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474697.

Everson:2000:SWD

[EM00] Philip J. Everson and Carl N. Morris. Simulation from Wishart distributions with eigenvalue constraints. Journal of Computational and Graphical Statistics, 9(2): 380–389, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474886.

Eilers:2002:GLA

[EM02] Paul H. C. Eilers and Brian D. Marx. Generalized linear additive smooth structures. Journal of Computational and Graphical Statistics, 11(4):758–783, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s2/p758;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002844.

Ellis:2007:MGS

[EM07] Nick Ellis and Ranjan Maitra. Multivariate Gaussian simulation outside arbitrary ellipsoids. Journal of Computational and Graphical Statistics, 16(3):692–708, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eckardt:2018:PPO

[EM18] Matthias Eckardt and Jorge Mateu. Point patterns occurring on complex structures in space and space–time: An alternative network approach. Journal of Computational and Graphical Statistics, 27(2):312–322, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Efron:2020:ACB

[EN20] Bradley Efron and Balasubramanian Narasimhan. The automatic construction of bootstrap confidence intervals. Journal of Computational and Graphical Statistics, 29 (3):608–619, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1714633.

Eddy:1995:DTD

[EO95] William F. Eddy and Shingo Oue. Dynamic three-dimensional display of U.S. air traffic. Journal of Computational and Graphical Statistics, 4(4): 261–280, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1995.10474684.

Elashoff:2004:EAE

[ER04] Michael Elashoff and Louise Ryan. An EM algorithm for estimating equations. Journal of Computational and Graphical Statistics, 13 (1):48–65, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Everitt:2021:DAA

[ER21] Richard G. Everitt and Paulina A. Rowińska. Delayed acceptance ABC-SMC. Journal of Computational and Graphical Statistics, 30(1):55– 66, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1775617.

Elmi:2011:BSB

[ERPG11] Angelo Elmi, Sarah J.

Ratcliffe, Samuel Parry, and Wensheng Guo. A B-spline based semiparametric nonlinear mixed effects model. Journal of Computational and Graphical Statistics, 20 (2):492–509, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eddy:1996:ISS

[ES96] William F. Eddy and Günther Sawitzki. Introduction to the special section on design and implementation of data analysis systems. Journal of Computational and Graphical Statistics, 5 (1):57, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474694.

Evans:1998:RVG

[ES98] M. Evans and T. Swartz.

Random variable generation using concavity properties of transformed densities. Journal of Computational and Graphical Statistics, 7(4):514–528, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/evans.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474792.

Ellner:2002:USS

[ES02] Stephen P. Ellner and Yodit Seifu. Using spatial statistics to select model complexity. Journal of Computational and Graphical Statistics, 11 (2):348–369, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n2/s6/p348;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002760180554.

Euan:2019:DSB

[ES19] Carolina Euán and Ying Sun. Directional spectrabased clustering for visualizing patterns of ocean waves and winds. Journal of Computational and Graphical Statistics, 28 (3):659–670, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eidsvik:2014:EPS

[ESR^{+}14] Jo Eidsvik, Benjamin A. Shaby, Brian J. Reich, Matthew Wheeler, and Jarad Niemi. Estimation and prediction in spatial models with block composite likelihoods. Journal of Computational and Graphical Statistics, 23 (2):295–315, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Eno:1999:RSL

[ET99a] Daniel R. Eno and George R. Terrell. Rejoinder: Scatterplots for Logistic Regression. Journal of Computational and Graphical Statistics, 8(3): 428–430, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474824.  See

[ET99b].

Eno:1999:SLR

[ET99b] Daniel R. Eno and George R. Terrell. Scatterplots for logistic regression. Journal of Computational and Graphical Statistics, 8(3):413–425, September 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474822. See com-

ment [McC99] and rejoinder [ET99a].

Everitt:2012:BPE

[Eve12] Richard G. Everitt. Bayesian parameter estimation for latent Markov random fields and social networks. Journal of Computational and Graphical Statistics, 21(4):940–960, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Falissard:1999:FPC

[Fal99] Bruno Falissard. Focused principal component analysis: Looking at a correlation matrix with a particular interest in a given variable. Journal of Computational and Graphical Statistics, 8(4):906–912, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts99/falis.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1999.10474855.

Faraway:1992:CDA

[Far92] Julian J. Faraway. On the cost of data analysis. Journal of Computational and Graphical Statistics, 1(3):213–229, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1992.10474582.

Faraway:1999:GME

[Far99] Julian J. Faraway. A graphical method of exploring the mean structure in longitudinal data analysis. Journal of Computational and Graphical Statistics, 8(1):60–68, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/upcoming.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474801.

Faraway:2004:HAU

[Far04] Julian Faraway. Human animation using nonparametric regression. Journal of Computational and Graphical Statistics, 13 (3):537–553, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Faraway:2012:BPC

[Far12] Julian J. Faraway. Backscoring in principal coordinates analysis. Journal of Computational and Graphical Statistics, 21 (2):394–412, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Furnas:1994:PVD

[FB94] George W. Furnas and Andreas Buja. Prosection views: dimensional inference through sections and projections. With a discussion by John F. Elder IV, Shingo Oue and Daniel B. Carr and a rejoinder by the authors. Journal of Computational and Graphical Statistics, 3(4):323–353, December 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474649.

Foster:2011:GDM

[FB11] Scott D. Foster and Mark V. Bravington. Graphical diagnostics for Markov models for categorical data. Journal of Computational and Graphical Statistics, 20 (2):355–374, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fan:2006:OAM

[FBG06] Y. Fan, S. P. Brooks, and A. Gelman. Output assessment for Monte Carlo simulations via the score statistic. Journal of Computational and Graphical Statistics, 15(1):178–206, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fushing:2010:SCB

[FCL10] Hsieh Fushing, Shuchun Chen, and How-Jing Lee. Statistical computations on biological rhythms I: Dissecting variable cycles and computing signature phases in activityevent time series. Journal of Computational and Graphical Statistics, 19 (1):221–239, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fintzi:2017:EDA

[FCWM17] Jonathan Fintzi, Xiang Cui, Jon Wakefield, and Vladimir N. Minin. Efficient data augmentation for fitting stochastic epidemic models to prevalence data. Journal of Computational and Graphical Statistics, 26 (4):918–929, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Frazier:2021:RAB

[FD21] David T. Frazier and Christopher Drovandi. Robust approximate Bayesian inference with synthetic likelihood. Journal of Computational and Graphical Statistics, 30 (4):958–976, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1875839.

Fan:2010:BCF

[FDBS10] Y. Fan, J.-L. Dortet-Bernadet, and S. A. Sisson. On Bayesian curve fitting via auxiliary variables. Journal of Computational and Graphical Statistics, 19(3):626–644, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Finley:2019:EAB

[FDC^{+}19] Andrew O. Finley, Abhirup Datta, Bruce D. Cook, Douglas C. Morton, Hans E. Andersen, and Sudipto Banerjee. Efficient algorithms for Bayesian nearest neighbor Gaussian processes. Journal of Computational and Graphical Statistics, 28(2):401–414, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fritz:2022:CIE

[FDG^{+}22] Cornelius Fritz, Giacomo De Nicola, Felix Günther, David Rügamer, Martje Rave, Marc Schneble, Andreas Bender, Maximilian Weigert, Ralph Brinks, Annika Hoyer, Ursula Berger, Helmut Küchenhoff, and Göran Kauermann. Challenges in interpreting epidemiological surveillance data

— experiences from Germany. Journal of Computational and Graphical Statistics, 32(3):765– 766, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2126482.

Fearnhead:2002:MCM

[Fea02] Paul Fearnhead. Markov chain Monte Carlo, sufficient statistics, and particle filters. Journal of Computational and Graphical Statistics, 11(4):848–862, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s7/p848;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002835.

Fearnhead:2004:SOA

[Fea04]  Paul Fearnhead.  A

second-order approximation to the log-likelihood surface for mixture models, with application to the EM algorithm. Journal of Computational and Graphical Statistics, 13 (3):739–750, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fearnhead:2005:URQ

[Fea05]  Paul Fearnhead.  Us-

ing random quasi-Monte-Carlo within particle filters, with application to financial time series. Journal of Computational and Graphical Statistics, 14(4):751–769, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See corrections [Ano07a].

Fisch:2022:SMC

[FEF22] Alexander T. M. Fisch, Idris A. Eckley, and Paul Fearnhead. Subset multivariate collective and point anomaly detection. Journal of Computational and Graphical Statistics, 31(2):574–585, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1987257.

Friendly:1994:UAC

[FF94] Michael Friendly and John Fox. Using APL2 to create an object-oriented environment for statistical computation. Journal of Computational and Graphical Statistics, 3(4):

387–407, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474654.

Fan:1995:AOP

[FG95] Jianqing Fan and Irène Gijbels. Adaptive order polynomial fitting: Bandwidth robustification and bias reduction. Journal of Computational and Graphical Statistics, 4(3): 213–227, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474678.

Fan:2005:RTA

[FG05] Guangzhe Fan and J. Brian Gray. Regression tree analysis using TARGET. Journal of Computational and Graphical Statistics, 14(1):206–218, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Feng:2024:REG

[FGH^{+}24] Xingdong Feng, Yuan Gao, Jian Huang, Yuling Jiao, and Xu Liu. Relative entropy gradient sampler for unnormalized distribution. Journal of Computational and Graphical Statistics, 34 (1):211–221, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2340523.

Furrer:2006:CTI

[FGN06] Reinhard Furrer, Marc G. Genton, and Douglas Nychka. Covariance tapering for interpolation of large spatial datasets. Journal of Computational and Graphical Statistics, 15(3):502–523, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See erratum and addendum [Ano12b].

Farcomeni:2022:DQC

[FGV22] Alessio Farcomeni, Marco Geraci, and Cinzia Viroli. Directional quantile classifiers. Journal of Computational and Graphical Statistics, 31 (3):907–916, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2021209.

Ferraty:2015:ANN

[FH15] Frédéric Ferraty and Peter Hall. An algorithm for nonlinear, nonparametric model choice and prediction. Journal of Computational and Graphical Statistics, 24(3):695–714, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fifield:2020:ARS

[FHIT20] Benjamin Fifield, Michael Higgins, Kosuke Imai, and Alexander Tarr. Automated redistricting simulation using Markov Chain Monte Carlo. Journal of Computational and Graphical Statistics, 29 (4):715–728, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1739532.

Feuerverger:2005:DMS

[FHTG05] Andrey Feuerverger, Peter Hall, Gelila Tilahun, and Michael Gervers. Distance measures and smoothing methodology for imputing features of documents. Journal of Computational and Graphical Statistics, 14 (2):255–262, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fraser:2003:CMI

[FHVW03] Andrew M. Fraser, Nicolas W. Hengartner, Kevin R. Vixie, and Brendt E. Wohlberg. Classification modulo invariance, with application to face recognition. Journal of Computational and Graphical Statistics, 12(4):829–852, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/1061860032634.

Fienberg:2012:BHS

[Fie12] Stephen E. Fienberg. A brief history of statistical models for network analysis and open challenges. Journal of Computational and Graphical Statistics, 21(4):825–839, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fontanella:2019:ONS

[FIK19] Lara Fontanella, Luigi Ippoliti, and Alfred Kume. The offset normal shape distribution for dynamic shape analysis. Journal of Computational and Graphical Statistics, 28 (2):374–385, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fishman:2012:CCT

[Fis12] George S. Fishman. Counting contingency tables via multistage Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 21 (3):713–738, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Friendly:2011:CWT

[FK11] Michael Friendly and Ernest Kwan. Comment: Why Tables Are Really Much Better Than Graphs. Journal of Computational and Graphical Statistics, 20(1):18– 27, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Gel11b].

Fang:2016:BIG

[FK16] Zaili Fang and Inyoung Kim. Bayesian Ising graphical model for variable selection. Journal of Computational and Graphical Statistics, 25 (2):589–605, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fallaize:2023:BMC

[FK23] Christopher J. Fallaize and Theodore Kypraios. Bayesian model choice for directional data. Journal of Computational and Graphical Statistics, 33 (1):25–34, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2023.2206076.

Fay:2007:UTS

[FKH07] Michael P. Fay, Hyune-Ju Kim, and Mark Hachey. On using truncated sequential probability ratio test boundaries for Monte Carlo implementation of hypothesis tests. Journal of Computational and Graphical Statistics, 16(4):946–967, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Friedrich:2008:CPM

[FKLW08] F. Friedrich, A. Kempe, V. Liebscher, and G. Winkler. Complexity penalized M-estimation. Journal of Computational and Graphical Statistics, 17 (1):201–224, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Frazier:2022:VBS

[FLMM22] David T. Frazier, Rubén Loaiza-Maya, and Gael M. Martin. Variational Bayes in state space models: Inferential and predictive accuracy. Journal of Computational and Graphical Statistics, 32 (3):793–804, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2134875.

Frazier:2024:LBV

[FLMMK24] David T. Frazier, Rubén Loaiza-Maya, Gael M. Martin, and Bonsoo Koo. Loss-based variational Bayes prediction. Journal of Computational and Graphical Statistics, 34 (1):84–95, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2341899.

Fojtik:2023:ECT

[FLMN23] Vít Fojtík, Petra Laketa, Pavlo Mozharovskyi, and Stanislav Nagy. On exact computation of Tukey depth central regions. Journal of Computational and Graphical Statistics, 33(2):699–713, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257781.

Fan:2021:PQR

[FLY21] Ye Fan, Nan Lin, and Xianjun Yin. Penalized quantile regression for distributed big data using the slack variable representation. Journal of Computational and Graphical Statistics, 30 (3):557–565, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1840996.

Fan:1994:FIN

[FM94] Jianqing Fan and James S. Marron. Fast implementations of nonparametric curve estimators. Journal of Computational and Graphical Statistics, 3(1): 35–56, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474629.

Fearnhead:2007:FMM

[FM07] Paul Fearnhead and Loukia Meligkotsidou. Filtering methods for mixture models. Journal of Computational and Graphical Statistics, 16 (3):586–607, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fernandez-Macho:2010:MOS

[FM10] Javier Fernández-Macho. Missing observations in spatial models: a spectral EM algorithm. Journal of Computational and Graphical Statistics, 19 (3):684–701, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fang:2012:LIH

[FM12] Zhou Fang and Nicolai Meinshausen. LASSO isotone for high-dimensional additive isotonic regression. Journal of Computational and Graphical Statistics, 21(1):72–91, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fearnhead:2019:DCS

[FML19] Paul Fearnhead, Robert Maidstone, and Adam Letchford. Detecting changes in slope with an L_{0} penalty. Journal of Computational and Graphical Statistics, 28 (2):265–275, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fosdick:2019:MNM

[FMM^{+}19] Bailey K. Fosdick, Tyler H.

McCormick, Thomas Brendan Murphy, Tin Lok James Ng, and Ted Westling. Multiresolution network models. Journal of Computational and Graphical Statistics, 28(1):185–196, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fonseca:2024:WFS

[FMP24] Rodney Fonseca, Pedro Morettin, and Aluísio Pinheiro. Wavelet feature screening. Journal of Computational and Graphical Statistics, 33 (4):1310–1319, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2342984.

Fryzlewicz:2004:HFA

[FN04] Piotr Fryzlewicz and Guy P. Nason. A Haar– Fisz algorithm for Poisson intensity estimation. Journal of Computational and Graphical Statistics, 13(3):621–638, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fouskakis:2016:PCE

[FN16] Dimitris Fouskakis and Ioannis Ntzoufras. Powerconditional-expected priors: Using g-priors with random imaginary data for variable selection. Journal of Computational and Graphical Statistics, 25(3):647–664, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1036996.

Fasiolo:2020:SVM

[FNGW20] Matteo Fasiolo, Raphaël Nedellec, Yannig Goude, and Simon N. Wood. Scalable visualization methods for modern generalized additive models. Journal of Computational and Graphical Statistics, 29(1):78–86, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1629942.

Fong:2019:FBC

[Fon19] Youyi Fong. Fast bootstrap confidence intervals for continuous threshold linear regression. Journal of Computational and Graphical Statistics, 28 (2):466–470, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fossgaard:2001:ALD

[Fos01]  Eirik Fossgaard.  Al-

ternative local discriminant bases using empirical expectation and variance estimation. Journal of Computational and Graphical Statistics, 10 (2):329–349, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628202.

Friel:2004:LEI

[FP04] N. Friel and A. N. Pettitt. Likelihood estimation and inference for the autologistic model. Journal of Computational and Graphical Statistics, 13 (1):232–246, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fu:2020:ENC

[FP20] Wei Fu and Patrick O. Perry. Estimating the number of clusters using cross-validation. Journal of Computational and Graphical Statistics, 29 (1):162–173, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1647846.

Friel:2009:BIH

[FPRW09] N. Friel, A. N. Pettitt, R. Reeves, and E. Wit. Bayesian inference in hidden Markov random fields for binary data defined on large lattices. Journal of Computational and Graphical Statistics, 18 (2):243–261, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ferrari:2016:PEL

[FQH16] Davide Ferrari, Guoqi Qian, and Tane Hunter. Parsimonious and efficient likelihood composition by Gibbs sampling. Journal of Computational and Graphical Statistics, 25(3):935–953, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1058799.

Frigessi:1997:BIC

[FR97] Arnoldo Frigessi and Håvard Rue. Bayesian image classification with Baddeley’s delta loss. Journal of Computational and Graphical Statistics, 6(1):55–73, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
frigessi.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474727.

Friendly:1999:EMD

[Fri99] Michael Friendly. Extending mosaic displays: Marginal, conditional, and partial views of categorical data. Journal of Computational and Graphical Statistics, 8(3): 373–395, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1999.10474820.

Friendly:2000:NSO

[Fri00] Michael Friendly. Note on “Obtaining the Maximum Likelihood Estimates in Incomplete R × C Contingency Tables Using a Poisson Generalized Linear Model”. Journal of Computational and Graphical Statistics, 9(1):158–166, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Friendly.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474871.

Friendly:2002:BHM

[Fri02] Michael Friendly. A brief history of the Mosaic display. Journal of Computational and Graphical Statistics, 11(1):89– 107, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n1/s5/p89;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002317375631.

Friendly:2007:PML

[Fri07] Michael Friendly. HE plots for multivariate linear models. Journal of Computational and Graphical Statistics, 16 (2):421–444, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Friel:2013:EBF

[Fri13a] Nial Friel. Evidence and Bayes factor estimation for Gibbs random fields. Journal of Computational and Graphical Statistics, 22(3):518–532, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Friendly:2013:GRT

[Fri13b] Michael Friendly. The generalized ridge trace plot: Visualizing bias and precision. Journal of Computational and Graphical Statistics, 22 (1):50–68, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Friendly:2014:CGP

[Fri14] Michael Friendly. Comment on The Generalized Pairs Plot. Journal of Computational and Graphical Statistics, 23 (1):290–291, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [EGS^{+}13].

Frost:2022:EES

[Fro22] H. Robert Frost. Eigenvectors from eigenvalues sparse principal component analysis. Journal of Computational and Graphical Statistics, 31 (2):486–501, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1987254.

Fraley:2005:IMB

[FRW05] Chris Fraley, Adrian Raftery, and Ron Wehrens. Incremental model-based clustering for large datasets with small clusters. Journal of Computational and Graphical Statistics, 14 (3):529–546, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fayyad:1999:CMM

[FS99] Usama M. Fayyad and Padhraic Smyth. Cataloging and mining massive datasets for science data analysis. Journal of Computational and Graphical Statistics, 8(3):589–610, September 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474835.

Feng:2018:GBR

[FS18] Jean Feng and Noah Simon. Gradient-based regularization parameter selection for problems with nonsmooth penalty functions. Journal of Computational and Graphical Statistics, 27(2):426–435, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fan:2011:LEE

[FSB11] Chun-Po Steve Fan, Jamie Stafford, and Patrick E. Brown. Local-EM and the EMS algorithm. Journal of Computational and Graphical Statistics, 20 (3):750–766, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fox:2002:DCN

[FSMV02] John Fox, Robert Stine, Georges Monette, and Neeru Vohra. Detecting clusters and nonlinearity in three-dimensional dynamic graphs. Journal of Computational and Graphical Statistics, 11(4):875–895, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s9/p875;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002718.

Fisher:2020:ECT

[FSR20] Eyal Fisher, Regev Schweiger, and Saharon Rosset. Efficient construction of test inversion confidence intervals using quantile regression. Journal of Computational and Graphical Statistics, 29(1):140– 148, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1647215.

Fryzlewicz:2016:SSA

[FT16] Piotr Fryzlewicz and Catherine Timmermans. SHAH: SHape-adaptive Haar wavelets for image processing. Journal of Computational and Graphical Statistics, 25 (3):879–898, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1048345.

Friedberg:2021:LLF

[FTAW21] Rina Friedberg, Julie Tibshirani, Susan Athey, and Stefan Wager. Local linear forests. Journal of Computational and Graphical Statistics, 30 (2):503–517, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1831930.

Fu:1998:PRB

[Fu98] Wenjiang J. Fu. Penalized regressions: The bridge versus the lasso. Journal of Computational and Graphical Statistics, 7(3):397–416, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/fu.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474784.

Fukuyama:2022:LBM

[Fuk22] Julia Fukuyama. Local biplots for multidimensional scaling, with application to the microbiome. Journal of Computational and Graphical Statistics, 31 (3):779–789, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2024438.

Feinberg:2011:ESC

[FW11] Richard A. Feinberg and Howard Wainer. Extracting sunbeams from cucumbers. Journal of Computational and Graphical Statistics, 20(4):793–810, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Favaro:2013:SSS

[FW13] Stefano Favaro and Stephen G. Walker. Slice sampling σstable Poisson–Kingman mixture models. Journal of Computational and Graphical Statistics, 22 (4):830–847, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fong:2012:EMC

[FWR12] Youyi Fong, Jon Wakefield, and Kenneth Rice. An efficient Markov chain Monte Carlo method for mixture models by neighborhood pruning. Journal of Computational and Graphical Statistics, 21 (1):197–216, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Fong:2021:FSD

[FX21] Youyi Fong and Jun Xu. Forward stepwise deep autoencoder-based monotone nonlinear dimensionality reduction methods. Journal of Computational and Graphical Statistics, 30(3):519–529, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1856119.

Feng:2021:SSI

[FXC21] Yuan Feng, Luo Xiao, and Eric C. Chi. Sparse single index models for multivariate responses. Journal of Computational and Graphical Statistics, 30 (1):115–124, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1779080.

Gramacy:2015:LGP

[GA15] Robert B. Gramacy and Daniel W. Apley. Local Gaussian process approximation for large computer experiments. Journal of Computational and Graphical Statistics, 24 (2):561–578, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.914442.

Gao:1998:WSD

[Gao98] Hong-Ye Gao. Wavelet shrinkage denoising using the non-negative garrote. Journal of Computational and Graphical Statistics, 7(4):469–488, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/gao.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474789.

Gassmann:2003:MNP

[Gas03] H. I. Gassmann. Multivariate normal probabilities: Implementing an old idea of Plackett’s. Journal of Computational and Graphical Statistics, 12(3):731–752, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

                                        URL http://giorgio.
                                        ingentaselect.com/cgi-
                                        bin/linker?ini=asa{\&
                                        }reqidx=/cw/asa/10618600/
                                        v12n3/s13/p731;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/1061860032283.

Geoga:2020:SGP

[GAS20] Christopher J. Geoga, Mihai Anitescu, and Michael L. Stein. Scalable Gaussian process computations using hierarchical matrices. Journal of Computational and Graphical Statistics, 29 (2):227–237, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1652616.

Guo:2015:SWP

[GAZ15] Ruixin Guo, Mihye Ahn, and Hongtu Zhu. Spatially weighted principal component analysis for imaging classification. Journal of Computational and Graphical Statistics, 24(1):274–296, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ghosh:2000:TSP

[GB00] Sucharita Ghosh and Jan Beran. Two-sample T_{3} plot: a graphical comparison of two distributions. Journal of Computational and Graphical Statistics, 9(1):167–179, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Ghosh.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474872.

Genz:2002:CMC

[GB02] Alan Genz and Frank Bretz. Comparison of methods for the computation of multivariate t probabilities. Journal of Computational and Graphical Statistics, 11(4):950–971, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n4/s13/p950;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002394.

Gous:2004:VCD

[GB04] Alan Gous and Andreas Buja. Visual comparison of datasets using mixture decompositions. Journal of Computational and Graphical Statistics, 13(1):1–19, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gunes:2012:CRA

[GB12] Funda Gunes and Howard D. Bondell. A confidence region approach to tuning for variable selection. Journal of Computational and Graphical Statistics, 21(2):295–314, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Goldsmith:2011:PFR

[GBC^{+}11] Jeff Goldsmith, Jennifer Bobb, Ciprian M. Crainiceanu, Brian Caffo, and Daniel Reich. Penalized functional regression. Journal of Computational and Graphical Statistics, 20(4):830–851, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Godichon-Baggioni:2024:PCS

[GBS24] Antoine Godichon-Baggioni and Sobihan Surendran. A penalized criterion for selecting the number of clusters for K-medians. Journal of Computational and Graphical Statistics, 33(4):1298–1309, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2325458.

Gaynanova:2017:PVC

[GBW17] Irina Gaynanova, James G.

Booth, and Martin T. Wells. Penalized versus constrained generalized eigenvalue problems. Journal of Computational and Graphical Statistics, 26(2):379–387, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Greven:2008:RLR

[GCKP08] Sonja Greven, Ciprian M.

Crainiceanu, Helmut Küchenhoff, and Annette Peters. Restricted likelihood ratio testing for zero variance components in linear mixed models. Journal of Computational and Graphical Statistics, 17 (4):870–891, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ghosh:2009:DPD

[GD09] Joyee Ghosh and David B. Dunson. Default prior distributions and efficient posterior computation in Bayesian factor analysis. Journal of Computational and Graphical Statistics, 18(2):306–320, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gaskins:2016:CPP

[GD16] J. T. Gaskins and M. J. Daniels. Covariance partition priors: A Bayesian approach to simultaneous covariance estimation for longitudinal data. Journal of Computational and Graphical Statistics, 25 (1):167–186, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2015.1028549;   http:
//www.tandfonline.com/

doi/full/10.1080/10618600.
2015.1028549.

Gagnon:2021:NJA

[GD21] Philippe Gagnon and Arnaud Doucet. Nonreversible jump algorithms for Bayesian nested model selection. Journal of Computational and Graphical Statistics, 30(2):312– 323, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1826955.

Goncalves:2023:EBI

[GD23] Flávio B. Gonçalves and Bárbara C. C. Dias. Exact Bayesian inference for level-set Cox processes with piecewise constant intensity function. Journal of Computational and Graphical Statistics, 32 (1):1–18, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2092117.

Gaskins:2014:SIP

[GDM14] J. T. Gaskins, M. J.

Daniels, and B. H. Marcus. Sparsity inducing prior distributions for correlation matrices of longitudinal data. Journal of Computational and Graphical Statistics, 23 (4):966–984, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gassmann:2002:CMN

[GDS02] H. I. Gassmann, I. Deák, and T. Szántai. Computing multivariate normal probabilities: a new look. Journal of Computational and Graphical Statistics, 11(4):920–949, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s12/p920;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002385.

Groeneboom:1996:CAS

[GdTv96] Piet Groeneboom, Peter de Jong, Dimitri B. Tischenko, and Bert C. van Zomeren. Computerassisted statistics education at Delft University of Technology. Journal of Computational and Graphical Statistics, 5(4): 386–399, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474719.

Garcia-Escudero:2003:TTE

[GEGM03] Luis Angel García-Escudero, Alfonso Gordaliza, and Carlos Matrán. Trimming tools in exploratory data analysis. Journal of Computational and Graphical Statistics, 12(2):434–449, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s10/p434;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031806.

Guo:2022:MSD

[GEH^{+}22] Bin Guo, Lynn E. Eberly, Pierre-Gilles Henry, Christophe Lenglet, and Eric F. Lock. Multiway sparse distance weighted discrimination. Journal of Computational and Graphical Statistics, 32(2):730–743, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2099404.

Gelman:1995:MMU

[Gel95] Andrew Gelman. Method of moments using Monte Carlo simulation. Journal of Computational and Graphical Statistics, 4(1): 36–54, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1995.10474664.

Gelman:2000:DOT

[Gel00] Andrew Gelman. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 49–51, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474864.  See

[LHY00].

Gelman:2004:EDA

[Gel04a] Andrew Gelman. Exploratory data analysis for complex models. Journal of Computational and Graphical Statistics, 13 (4):755–779, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gelman:2004:RED

[Gel04b] Andrew Gelman. Rejoinder: Exploratory data analysis for complex models. Journal of Computational and Graphical Statistics, 13(4):785–787, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gelman:2011:RWT

[Gel11a] Andrew Gelman. Rejoinder: Why Tables Are Really Much Better Than Graphs. Journal of Computational and Graphical Statistics, 20(1):36– 40, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Gel11b].

Gelman:2011:WTR

[Gel11b] Andrew Gelman. Why tables are really much better than graphs. Journal of Computational and Graphical Statistics, 20 (1):3–7, March 2011. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comments [Wai11, Bri11, FK11, Wil11] and rejoinder [Gel11a].

Gelman:2017:CCG

[Gel17] Andrew Gelman. Correction to Cook, Gelman, and Rubin (2006). Journal of Computational and Graphical Statistics, 26 (4):940, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [CGR06].

Genz:1992:NCM

[Gen92] Alan Genz. Numerical computation of multivariate normal probabilities. Journal of Computational and Graphical Statistics, 1(2):141–149, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477010.

Geyer:1995:CMC

[Gey95] Charles J. Geyer. Conditioning in Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 4(2): 148–154, June 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474672.

Gordon:2015:SHT

[GF15] Ian Gordon and Sue Finch. Statistician heal thyself: Have we lost the plot? Journal of Computational and Graphical Statistics, 24(4):1210– 1229, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.989324.

Gong:2016:PSS

[GF16] Lei Gong and James M. Flegal. A practical sequential stopping rule for high-dimensional Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 25 (3):684–700, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1044092.

Guinness:2017:CEA

[GF17] Joseph Guinness and Montserrat Fuentes. Circulant embedding of approximate covariances for inference from Gaussian data on large lattices. Journal of Computational and Graphical Statistics, 26(1):88–97, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1164534.

Gallagher:2022:BPM

[GF22] Shannon K. Gallagher and Dean Follmann. Branching process models to identify risk factors for infectious disease transmission. Journal of Computational and Graphical Statistics, 31 (2):529–540, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000871.

Gramacki:2017:FBF

[GG17] Artur Gramacki and Jaroslaw Gramacki. FFTbased fast computation of multivariate kernel density estimators with unconstrained bandwidth matrices. Journal of Computational and Graphical Statistics, 26 (2):459–462, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Guha:2020:PDE

[GG20] Subharup Guha and Sujit K. Ghosh. Probabilistic detection and estimation of conic sections from noisy data. Journal of Computational and Graphical Statistics, 29 (3):513–522, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1737084.

Gardini:2022:MTM

[GGT22] Aldo Gardini, Fedele Greco, and Carlo Trivisano. The Mellin transform to manage quadratic forms in normal random variables. Journal of Computational and Graphical Statistics, 31 (4):1416–1425, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2034639.

Gower:2010:AB

[GGvdV10] J. C. Gower, P. J. F.

Groenen, and M. van de Velden. Area biplots. Journal of Computational and Graphical Statistics, 19(1):46–61, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Groenen:2000:DOT

[GH00] Patrick J. F. Groenen and Willem J. Heiser. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 44–48, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474863.  See

[LHY00].

Guan:2018:CEP

[GH18] Yawen Guan and Murali Haran. A computationally efficient projectionbased approach for spatial generalized linear mixed models. Journal of Computational and Graphical Statistics, 27(4):701–714, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Griffin:2020:TSI

[GH20] Maryclare Griffin and Peter D. Hoff. Testing sparsity-inducing penalties. Journal of Computational and Graphical Statistics, 29(1):128– 139, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1637749.

Ghosal:2021:BRF

[GH21] Indrayudh Ghosal and Giles Hooker. Boosting random forests to reduce bias; one-step boosted forest and its variance estimate. Journal of Computational and Graphical Statistics, 30(2):493– 502, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1820345.

Griffin:2023:SSP

[GH23] Maryclare Griffin and Peter D. Hoff. Structured shrinkage priors. Journal of Computational and Graphical Statistics, 33 (1):1–14, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2233577.

Goldsmith:2014:SSI

[GHC14] Jeff Goldsmith, Lei Huang, and Ciprian M. Crainiceanu. Smooth scalar-on-image regression via spatial Bayesian variable selection. Journal of Computational and Graphical Statistics, 23(1):46–64, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gupta:2022:VPD

[GHCU22] Sayani Gupta, Rob J.

Hyndman, Dianne Cook, and Antony Unwin. Visualizing probability distributions across bivariate cyclic temporal granularities. Journal of Computational and Graphical Statistics, 31(1):14– 25, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1938588.

Ghosh:2007:NNC

[Gho07] Anil K. Ghosh. On nearest neighbor classification using adaptive choice of k. Journal of Computational and Graphical Statistics, 16(2):482–502, June 2007.

CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Guglielmi:2002:PSI

[GHW02] Alessandra Guglielmi, Chris C. Holmes, and Stephen G. Walker. Perfect simulation involving functionals of a Dirichlet process. Journal of Computational and Graphical Statistics, 11(2):306– 310, June 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n2/s3/p306;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002760180527.

Gentleman:2000:JD

[GI00a] Robert Gentleman and Ross Ihaka. Java and Dylan. Journal of Computational and Graphical Statistics, 9(3):452–458, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Gentleman.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474892.

Gentleman:2000:LSS

[GI00b] Robert Gentleman and Ross Ihaka. Lexical scope and statistical computing. Journal of Computational and Graphical Statistics, 9(3):491–508, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Ihaka.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474895.

Greco:2022:FMM

[GIA22] Luca Greco, Pier Luigi Novi Inverardi, and Claudio Agostinelli. Finite mixtures of multivariate wrapped normal distributions for model based clustering of p-torus data. Journal of Computational and Graphical Statistics, 32(3):1215–1228, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2128808.

Gavin:1997:SIR

[GJ97] John Gavin and Christopher Jennison. A subpixel image restoration algorithm. Journal of Computational and Graphical Statistics, 6(2):182– 201, June 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/

                                        jcgs/abstracts/june/
                                        gavin.html;   http:/
                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474736.

Garthwaite:2009:SAM

[GJ09] Paul H. Garthwaite and M. C. Jones. A stochastic approximation method and its application to confidence intervals. Journal of Computational and Graphical Statistics, 18 (1):184–200, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Guo:2010:PCA

[GJL^{+}10] Jian Guo, Gareth James, Elizaveta Levina, George Michailidis, and Ji Zhu. Principal component analysis with sparse fused loadings. Journal of Computational and Graphical Statistics, 19(4):930–946, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Guihenneuc-Jouyaux:2005:LEM

[GJR05] Chantal Guihenneuc-Jouyaux and Judith Rousseau. Laplace expansions in Markov chain Monte Carlo algorithms. Journal of Computational and Graphical Statistics, 14 (1):75–94, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Genz:1997:SAI

[GK97] Alan Genz and Robert E.

Kass. Subregion-adaptive integration of functions having a dominant peak. Journal of Computational and Graphical Statistics, 6(1):92–111, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
genz.html;   http://
www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474729.

Gelfand:2002:CAF

[GK02] Alan E. Gelfand and Athanasios Kottas. A computational approach for full nonparametric Bayesian inference under Dirichlet process mixture models. Journal of Computational and Graphical Statistics, 11(2):289– 305, June 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s2/p289;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180518.

Gatu:2006:BBA

[GK06] Cristian Gatu and Erricos John Kontoghiorghes. Branch-and-bound algorithms for computing the best-subset regression models. Journal of Computational and Graphical Statistics, 15(1):139–156, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Giordani:2010:AIM

[GK10] Paolo Giordani and Robert Kohn. Adaptive independent Metropolis–Hastings by fast estimation of mixtures of normals. Journal of Computational and Graphical Statistics, 19 (2):243–259, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Goldstein:2015:PIB

[GKBP15] Alex Goldstein, Adam Kapelner, Justin Bleich, and Emil Pitkin. Peeking inside the black box: Visualizing statistical learning with plots of individual conditional expectation. Journal of Computational and Graphical Statistics, 24(1):44–65, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gunawan:2023:FVB

[GKN23] David Gunawan, Robert Kohn, and David Nott. Flexible variational Bayes based on a copula of a mixture. Journal of Computational and Graphical Statistics, 33(2):665– 680, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2262080.

Genton:2018:HDC

[GKT18] Marc G. Genton, David E.

Keyes, and George Turkiyyah. Hierarchical decompositions for the computation of high-dimensional multivariate normal probabilities. Journal of Computational and Graphical Statistics, 27(2):268–277, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:2010:PGA

[GKZ10] Hong Gu, Toby Kenney, and Mu Zhu. Partial generalized additive models: An informationtheoretic approach for dealing with concurvity and selecting variables. Journal of Computational and Graphical Statistics, 19(3):531–551, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gaines:2018:AFC

[GKZ18] Brian R. Gaines, Juhyun Kim, and Hua Zhou. Algorithms for fitting the constrained lasso. Journal of Computational and Graphical Statistics, 27 (4):861–871, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gentleman:2007:SAR

[GL07] Robert Gentleman and Duncan Temple Lang. Statistical analyses and reproducible research. Journal of Computational and Graphical Statistics, 16(1):1–23, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Guan:2020:ETF

[GLC20] Tianyu Guan, Zhenhua Lin, and Jiguo Cao. Estimating truncated functional linear models with a nested group bridge approach. Journal of Computational and Graphical Statistics, 29(3):620– 628, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1713797.

Guo:2015:GMO

[GLMZ15] Jian Guo, Elizaveta Levina, George Michailidis, and Ji Zhu. Graphical models for ordinal data. Journal of Computational and Graphical Statistics, 24(1):183–204, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gluhovsky:2000:IRU

[Glu00] Ilya Gluhovsky. Image restoration using modifications of simulated annealing. Journal of Computational and Graphical Statistics, 9(4):653– 671, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474906.

Gluhovsky:2004:ESA

[Glu04] Ilya Gluhovsky. Evolutionary simulated annealing with application to image restoration. Journal of Computational and Graphical Statistics, 13 (4):871–885, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Goswami:2007:EMC

[GLW07] Gopi Goswami, Jun S.

Liu, and Wing H. Wong. Evolutionary Monte Carlo methods for clustering. Journal of Computational and Graphical Statistics, 16(4):855–876, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Grenander:1993:SSC

[GM93] Ulf Grenander and Kevin M. Manbeck. A stochastic shape and color model for defect detection in potatoes. Journal of Computational and Graphical Statistics, 2(2):131– 151, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474604.

Goodall:1999:PSA

[GM99] Colin R. Goodall and Kanti V. Mardia. Projective shape analysis. Journal of Computational and Graphical Statistics, 8(2):143–168, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/projective.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/mardia.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474808.

Gu:2005:GNM

[GM05] Chong Gu and Ping Ma.

Generalized nonparametric mixed-effect models. Journal of Computational and Graphical Statistics, 14(2):485–504, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gao:2015:ESC

[GM15] Xin Gao and Hélène Massam. Estimation of symmetry-constrained Gaussian graphical models: Application to clustered dense networks. Journal of Computational and Graphical Statistics, 24(4):909–929, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.937811.

Gollini:2016:JMM

[GM16] Isabella Gollini and Thomas Brendan Murphy. Joint modeling of multiple network views. Journal of Computational and Graphical Statistics, 25(1):246– 265, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.978006;   http:/
/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2014.978006.

Goos:2018:QMB

[GM18] Peter Goos and Kalliopi Mylona. Quadrature methods for Bayesian optimal design of experiments with nonnormal prior distributions. Journal of Computational and Graphical Statistics, 27 (1):179–194, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Golzy:2020:PKB

[GM20] Mojgan Golzy and Marianthi Markatou. Poisson kernel-based clustering on the sphere: Convergence properties, identifiability, and a method of sampling. Journal of Computational and Graphical Statistics, 29(4):758– 770, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1740713.

Godtliebsen:2002:SSS

[GMC02] F. Godtliebsen, J. S. Marron, and Probal Chaudhuri. Significance in scale space for bivariate density estimation. Journal of Computational and Graphical Statistics, 11 (1):1–21, March 2002.

CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s1/p1;   http:/
/www.tandfonline.com/
doi/abs/10.1198/106186002317375596.

Ghasempour:2023:CNN

[GMdL23] Mohammad Ghasempour, Niloofar Moosavi, and Xavier de Luna. Convolutional neural networks for valid and efficient causal inference. Journal of Computational and Graphical Statistics, 33 (2):714–723, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257247.

Gloor:2016:DVL

[GMF16] Gregory B. Gloor, Jean M.

Macklaim, and Andrew D. Fernandes. Displaying variation in large datasets: Plotting a visual summary of effect sizes. Journal of Computational and Graphical Statistics, 25(3):971– 979, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1131161.

Guha:2004:BEM

[GMP04] Subharup Guha, Steven N. MacEachern, and Mario Peruggia. Benchmark estimation for Markov chain Monte Carlo samples. Journal of Computational and Graphical Statistics, 13(3):683–701, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Giordani:2013:FMD

[GMTK13] Paolo Giordani, Xiuyan Mun, Minh-Ngoc Tran, and Robert Kohn. Flexible multivariate density estimation with marginal adaptation. Journal of Computational and Graphical Statistics, 22 (4):814–829, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gibberd:2017:REP

[GN17] Alexander J. Gibberd and James D. B. Nelson. Regularized estimation of piecewise constant Gaussian graphical models: The group-fused graphical lasso. Journal of Computational and Graphical Statistics, 26 (3):623–634, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Godsill:2001:RBM

[God01] Simon J. Godsill. On the relationship between Markov chain Monte Carlo methods for model uncertainty. Journal of Computational and Graphical Statistics, 10 (2):230–248, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152627924.

Golightly:2009:BFJ

[Gol09] Andrew Golightly. Bayesian filtering for jump-diffusions with application to stochastic volatility. Journal of Computational and Graphical Statistics, 18 (2):384–400, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Goodall:1999:DMM

[Goo99] Colin R. Goodall. Data mining of massive datasets in healthcare. Journal of Computational and Graphical Statistics, 8(3): 620–634, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474837.

Goodman:2011:JVI

[Goo11] Arnold Goodman. JCGS@20, Visuals@40, Interface@45 & !!Challenges!! Journal of Computational and Graphical Statistics, 20 (4):818–829, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gous:1998:AED

[Gou98] Alan Gous. Adaptive estimation of distributions using exponential subfamilies. Journal of Computational and Graphical Statistics, 7(3):388–396, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/gous.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474783.

Gramacy:2011:PLG

[GP11] Robert B. Gramacy and Nicholas G. Polson. Particle learning of Gaussian process models for sequential design and optimization. Journal of Computational and Graphical Statistics, 20 (1):102–118, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Garcia-Portugues:2014:GFT

[GPGMFB14] Eduardo García-Portugués, Wenceslao González-Manteiga, and Manuel Febrero-Bande. A goodness-of-fit test for the functional linear model with scalar response. Journal of Computational and Graphical Statistics, 23(3):761–778, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Genovese:2017:FSF

[GPPVW17] Christopher Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. Finding singular features. Journal of Computational and Graphical Statistics, 26(3):598–609, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gruet:1999:MCS

[GPR99] Marie-Anne Gruet, Anne Philippe, and Christian P. Robert. MCMC control spreadsheets for exponential mixture estimation. Journal of Computational and Graphical Statistics, 8(2):298– 317, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/mcmc_
control.html;   http:
//www.amstat.org/publications/
jcgs/pdf99/gruet.pdf;
 http://www.tandfonline.

                                        com/doi/abs/10.1080/
                                        10618600.1999.10474815.

Guhaniyogi:2018:BCD

[GQD18] Rajarshi Guhaniyogi, Shaan Qamar, and David B. Dunson. Bayesian conditional density filtering. Journal of Computational and Graphical Statistics, 27(3):657–672, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:1993:CMM

[GR93] Xing Sam Gu and Paul R. Rosenbaum. Comparison of multivariate matching methods: Structures, distances, and algorithms. Journal of Computational and Graphical Statistics, 2(4):405–420, 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474623.

Givens:1999:LPP

[GR99] Geof H. Givens and Paul J. Roback. Logarithmic pooling of priors linked by a deterministic simulation model. Journal of Computational and Graphical Statistics, 8(3): 452–478, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/abs/10.1080/10618600.
1999.10474826.

Gottardo:2008:MCM

[GR08] Raphael Gottardo and Adrian E. Raftery. Markov chain Monte Carlo with mixtures of mutually singular distributions. Journal of Computational and Graphical Statistics, 17 (4):949–975, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gattone:2012:CCR

[GR12a] Stefano Antonio Gattone and Roberto Rocci. Clustering curves on a reduced subspace. Journal of Computational and Graphical Statistics, 21 (2):361–379, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Goldberg:2012:TAL

[GR12b] Yair Goldberg and Ya’acov Ritov. Theoretical analysis of LLE based on its weighting step. Journal of Computational and Graphical Statistics, 21 (2):380–393, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gray:1996:HRR

[Gra96] Robert J. Gray. Hazard rate regression using ordinary nonparametric regression smoothers. Journal of Computational and Graphical Statistics, 5(2): 190–207, June 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474704.

Graves:2007:DIM

[Gra07] Todd L. Graves. Design ideas for Markov chain Monte Carlo software. Journal of Computational and Graphical Statistics, 16(1):24–43, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Graffelman:2013:LAC

[Gra13] Jan Graffelman. Linearangle correlation plots: New graphs for revealing correlation structure. Journal of Computational and Graphical Statistics, 22(1):92–106, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gerber:2013:MSR

[GRB^{+}13] Samuel Gerber, Oliver Rübel, Peer-Timo Bremer, Valerio Pascucci, and Ross T. Whitaker. Morse–Smale regression. Journal of Computational and Graphical Statistics, 22(1):193–214, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Greenacre:2013:CB

[Gre13] Michael Greenacre. Contribution biplots. Journal of Computational and Graphical Statistics, 22 (1):107–122, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Genton:2010:VIO

[GRG10] Marc G. Genton and Anne Ruiz-Gazen. Visualizing influential observations in dependent data. Journal of Computational and Graphical Statistics, 19(4):808–825, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Grillenzoni:1999:ATC

[Gri99] Carlo Grillenzoni. Adaptive tests for changing unit roots in nonstationary time series. Journal of Computational and Graphical Statistics, 8(4):763–778, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/grillen.
html;   http://www.tandfonline.

com/doi/abs/10.1080/
10618600.1999.10474848.

Griffin:2023:IPC

[Gri23] Maryclare Griffin. Improved pathwise coordinate descent for power penalties. Journal of Computational and Graphical Statistics, 33(1):310– 315, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2256807.

Guha:2009:GSE

[GRM09] Subharup Guha, Louise Ryan, and Michele Morara. Gauss–Seidel estimation of generalized linear mixed models with application to Poisson modeling of spatially varying disease rates. Journal of Computational and Graphical Statistics, 18(4):818–837, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gallagher:2018:SSP

[GRVE18] Shannon Gallagher, Lee F. Richardson, Samuel L. Ventura, and William F. Eddy. SPEW: Synthetic populations and ecosystems of the world. Journal of Computational and Graphical Statistics, 27 (4):773–784, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gelfand:1994:MCM

[GS94] Alan E. Gelfand and Sujit K. Sahu. On Markov chain Monte Carlo acceleration. Journal of Computational and Graphical Statistics, 3(3):261–276, September 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474644.

Geoga:2023:SME

[GS23] Christopher J. Geoga and Michael L. Stein. A scalable method to exploit screening in Gaussian process models with noise. Journal of Computational and Graphical Statistics, 33(2):603–613, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2218027.

Ginsbourger:2024:FCG

[GS24] David Ginsbourger and Cédric Schärer. Fast calculation of Gaussian process multiple-fold crossvalidation residuals and their covariances. Journal of Computational and Graphical Statistics, 34 (1):1–14, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2353633.

Gneiting:2006:FES

[GŠP^{+}06] Tilmann Gneiting, Hana Ševčíková, Donald B. Percival, Martin Schlather, and Yindeng Jiang. Fast and exact simulation of large Gaussian lattice systems in R^{2}. Journal of Computational and Graphical Statistics, 15 (3):483–501, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gorur:2011:CCA

[GT11] Dilan Görür and Yee Whye Teh. Concave-convex adaptive rejection sampling. Journal of Computational and Graphical Statistics, 20(3):670–691, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ghosal:2016:SPF

[GTZH16] Subhashis Ghosal, Bradley Turnbull, Hao Helen Zhang, and Wook Yeon Hwang. Sparse penalized forward selection for support vector classification.

Journal of Computational and Graphical Statistics, 25(2):493–514, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:1992:CVN

[Gu92]  Chong Gu.  Cross-

validating non-Gaussian data. Journal of Computational and Graphical Statistics, 1(2):169– 179, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477012.

Gelman:2013:ISG

[GU13a] Andrew Gelman and Antony Unwin. Infovis and statistical graphics: Different goals, different looks. Journal of Computational and Graphical Statistics, 22(1):2–28, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gelman:2013:TIG

[GU13b] Andrew Gelman and Antony Unwin. Tradeoffs in information graphics. Journal of Computational and Graphical Statistics, 22(1):45–49, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:2015:HEB

[Gu15] Chong Gu. Hazard estimation with bivariate survival data and copula density estimation. Journal of Computational and Graphical Statistics, 24 (4):1053–1073, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.964356.

Guan:2023:SPP

[Gua23] Leying Guan. Smooth and probabilistic PARAFAC model with auxiliary covariates. Journal of Computational and Graphical Statistics, 33(2):538– 550, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257783.

Guedon:2003:EHS

[Gué03]  Yann Guédon.  Es-

timating hidden semi-Markov chains from discrete sequences. Journal of Computational and Graphical Statistics, 12 (3):604–639, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-

bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s7/p604;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032030.

Guha:2008:PSG

[Guh08] Subharup Guha. Posterior simulation in the generalized linear mixed model with semiparametric random effects. Journal of Computational and Graphical Statistics, 17 (2):410–425, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Golub:1997:GCV

[Gv97] Gene H. Golub and Urs von Matt. Generalized cross-validation for largescale problems. Journal of Computational and Graphical Statistics, 6(1): 1–34, March 1997. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts/march/
golub.html;   http:/
/www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474725.

Gentleman:2001:CAC

[GV01] R. Gentleman and A. C.

Vandal. Computational algorithms for censoreddata problems using intersection graphs. Journal of Computational and Graphical Statistics, 10(3):403–421, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317114901.

Gabbanini:2004:WPM

[GVBM04] Francesco Gabbanini, Marina Vannucci, Gianni Bartoli, and Antonio Moro. Wavelet packet methods for the analysis of variance of time series with application to crack widths on the Brunelleschi dome. Journal of Computational and Graphical Statistics, 13 (3):639–658, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gelman:2008:URP

[GvDHB08] Andrew Gelman, David A. van Dyk, Zaiying Huang, and John W. Boscardin. Using redundant parameterizations to fit hierarchical models. Journal of Computational and Graphical Statistics, 17 (1):95–122, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Giakoumatos:1999:MCM

[GVDP99] S. G. Giakoumatos, I. D. Vrontos, P. Dellaportas, and D. N. Politis. A Markov chain Monte Carlo convergence diagnostic using subsampling. Journal of Computational and Graphical Statistics, 8(3):431–451, September 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474825.

Gu:1993:SSA

[GW93] Chong Gu and Grace Wahba. Smoothing spline ANOVA with componentwise Bayesian “confidence intervals”. Journal of Computational and Graphical Statistics, 2(1): 97–117, March 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474601.

Groeneboom:2001:CCD

[GW01] Piet Groeneboom and Jon A. Wellner. Computing Chernoff’s distribution. Journal of Computational and Graphical Statistics, 10(2):388– 400, June 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152627997.

Ganguli:2004:FSG

[GW04] B. Ganguli and M. P. Wand. Feature significance in geostatistics. Journal of Computational and Graphical Statistics, 13(4):954–973, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Griffin:2011:PSN

[GW11] Jim E. Griffin and Stephen G. Walker. Posterior simulation of normalized random measure mixtures. Journal of Computational and Graphical Statistics, 20(1):241–259, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Grolemund:2015:VCD

[GW15] Garrett Grolemund and Hadley Wickham. Visualizing complex data with embedded plots. Journal of Computational and Graphical Statistics, 24 (1):26–43, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Geenens:2018:LLT

[GW18] Gery Geenens and Craig Wang. Local-likelihood transformation kernel density estimation for positive random variables.

Journal of Computational and Graphical Statistics, 27(4):822–835, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:2001:CVN

[GX01] Chong Gu and Dong Xiang. Cross-validating non-Gaussian data: Generalized approximate crossvalidation revisited. Journal of Computational and Graphical Statistics, 10(3):581–591, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317114992.

Gu:2020:FNG

[GX20] Mengyang Gu and Yanxun Xu. Fast nonseparable Gaussian stochastic process with application to methylation level interpolation. Journal of Computational and Graphical Statistics, 29(2):250– 260, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1665534.

Gotway:2007:GAL

[GY07] Carol A. Gotway and Linda J. Young. A geostatistical approach to linking geographically aggregated data from different sources. Journal of Computational and Graphical Statistics, 16(1):115–135, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Gu:2022:TCL

[GY22] Jiaqi Gu and Guosheng Yin. Triangular concordance learning of networks. Journal of Computational and Graphical Statistics, 32(2):434– 447, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2099405.

Gao:2022:AAR

[GZQ^{+}22] Yuan Gao, Xuening Zhu, Haobo Qi, Guodong Li, Riquan Zhang, and Hansheng Wang. An asymptotic analysis of random partition based minibatch momentum methods for linear regression models. Journal of Computational and Graphical Statistics, 32(3):1083–1096, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2143786.

Hill:2006:BER

[HABV06] Shawndra Hill, Deepak K.

Agarwal, Robert Bell, and Chris Volinsky. Building an effective representation for dynamic networks. Journal of Computational and Graphical Statistics, 15(3):584–608, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hans:2012:CDA

[HAC^{+}12] Chris Hans, Greg M. Allenby, Peter F. Craigmile, Juhee Lee, Steven N. MacEachern, and Xinyi Xu. Covariance decompositions for accurate computation in Bayesian scale-usage models. Journal of Computational and Graphical Statistics, 21 (2):538–557, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hansen:2024:FVI

[HAPRD24] Blake Hansen, Alejandra Avalos-Pacheco, Massimiliano Russo, and Roberta De Vito. Fast variational inference for Bayesian factor analysis in single and multi-study settings. Journal of Computational and Graphical Statistics, 34(1):96–108, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2356173.

Heiberger:1992:DFR

[HB92] Richard M. Heiberger and Richard A. Becker. Design of an S function for robust regression using iteratively reweighted least squares. Journal of Computational and Graphical Statistics, 1(3):181– 196, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1992.10474580.

Huzurbazar:1998:ISV

[HB98] S. Huzurbazar and Ronald W. Butler. Importance sampling for p-value computations in multivariate tests. Journal of Computational and Graphical Statistics, 7(3):342–355, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/huzurbazar.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474780.

Hyrien:2016:FND

[HB16] Ollivier Hyrien and Andrea Baran. Fast nonparametric density-based clustering of large datasets using a stochastic approximation mean-shift algorithm. Journal of Computational and Graphical Statistics, 25(3):899– 916, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1051625.

Hyndman:1996:EVC

[HBG96] Rob J. Hyndman, David M.

Bashtannyk, and Gary K. Grunwald. Estimating and visualizing conditional densities. Journal of Computational and Graphical Statistics, 5(4): 315–336, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474715.

Hertel:2022:RHO

[HBG22] Lars Hertel, Pierre Baldi, and Daniel L. Gillen. Reproducible hyperparameter optimization. Journal of Computational and Graphical Statistics, 31 (1):84–99, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1950004.

Hobert:1998:FCM

[HC98] James P. Hobert and George Casella. Functional compatibility, Markov chains, and Gibbs sampling with improper posteriors. Journal of Computational and Graphical Statistics, 7(1):42–60, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.research.att.com/
                                        ~andreas/jcgs/abstracts/
                                        1998/march/hobert.html;
                                        http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474760.

Ho:2007:TUS

[HC07] Lai Ping Ho and Sung Nok Chiu. Testing uniformity of a spatial point pattern. Journal of Computational and Graphical Statistics, 16(2):378–398, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hart:2011:NED

[HC11] Jeffrey D. Hart and Isabel Cañette. Nonparametric estimation of distributions in random effects models. Journal of Computational and Graphical Statistics, 20(2):461–478, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hoef:2004:FSM

[HCB04] Jay M. Ver Hoef, Noel Cressie, and Ronald Paul Barry. Flexible spatial models for kriging and cokriging using moving averages and the Fast Fourier Transform (FFT). Journal of Computational and Graphical Statistics, 13(2):265–282, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hu:2023:JNR

[HCCZ23] Jie Hu, Xiao Chen, Yu Chen, and Weiping Zhang. Joint network reconstruction and community detection from rich but noisy data. Journal of Computational and Graphical Statistics, 33 (2):501–514, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2267630.

Hebert:2023:AHD

[HCE23] Florian Hébert, David Causeur, and Mathieu Emily. Adaptive handling of dependence in high-dimensional regression modeling. Journal of Computational and Graphical Statistics, 32 (1):213–225, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2076687.

Huang:2002:FRC

[HCG02] Hsin-Cheng Huang, Noel Cressie, and John Gabrosek. Fast, resolution-consistent spatial prediction of global processes from satellite data. Journal of Computational and Graphical Statistics, 11(1):63– 88, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s4/p63;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375622.

Hofmann:2011:DCT

[HCK^{+}11] Heike Hofmann, Dianne Cook, Chris Kielion, Barret Schloerke, Jon Hobbs, Adam Loy, Lawrence Mosley, David Rockoff, Yuanyuan Huang, Danielle Wrolstad, and Tengfei Yin. Delayed, canceled, on time, boarding … flying in the USA. Journal of Computational and Graphical Statistics, 20(2):287–290, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2021:FDR

[HCZMH21] Hsin-Cheng Huang, Noel Cressie, Andrew Zammit-Mangion, and Guowen Huang. False discovery rates to detect signals from incomplete spatially aggregated data. Journal of Computational and Graphical Statistics, 30 (4):1081–1094, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1873144.

Hu:2023:GVS

[HD23] Zhiyong Hu and Dipak K.

Dey. Generalized variable selection algorithms for Gaussian process models by LASSO-like penalty. Journal of Computational and Graphical Statistics, 33(2):477–486, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2256802.

Hunt:2020:ATI

[HDK^{+}20] Gregory J. Hunt, Mark A.

Dane, James E. Korkola, Laura M. Heiser, and Johann A. Gagnon-Bartsch. Automatic transformation and integration to improve visualization and discovery of latent effects in imaging data. Journal of Computational and Graphical Statistics, 29 (4):929–941, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1741379.

Holmes:2002:AMU

[HDM02] C. C. Holmes, D. G. T. Denison, and B. K. Mallick. Accounting for model uncertainty in seemingly unrelated regressions. Journal of Computational and Graphical Statistics, 11 (3):533–551, September 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL

http://lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s3/p533;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002475.

Holmes:2005:BPP

[HDRM05] C. C. Holmes, D. G. T. Denison, S. Ray, and B. K. Mallick. Bayesian prediction via partitioning. Journal of Computational and Graphical Statistics, 14(4):811–830, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Heard:2011:IRA

[Hea11] Nicholas A. Heard. Iterative reclassification in agglomerative clustering. Journal of Computational and Graphical Statistics, 20(4):920–936, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Heard:2022:SPS

[Hea22]  N. A. Heard.  Stan-

dardized partial sums and products of p-values. Journal of Computational and Graphical Statistics, 31(2):563–573, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1999822.

Heaps:2023:EST

[Hea23] Sarah E. Heaps. Enforcing stationarity through the prior in vector autoregressions. Journal of Computational and Graphical Statistics, 32 (1):74–83, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2079648.

Haynes:2017:CEC

[HEF17] Kaylea Haynes, Idris A.

Eckley, and Paul Fearnhead. Computationally efficient changepoint detection for a range of penalties. Journal of Computational and Graphical Statistics, 26 (1):134–143, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1116445.

Held:2004:SPP

[Hel04] Leonhard Held. Simultaneous posterior probability statements from Monte Carlo output. Journal of Computational and Graphical Statistics, 13(1):20–35, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Helwig:2021:SSN

[Hel21] Nathaniel E. Helwig. Spectrally sparse nonparametric regression via elastic net regularized smoothers. Journal of Computational and Graphical Statistics, 30 (1):182–191, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1806855.

Helwig:2024:VDA

[Hel24] Nathaniel E. Helwig. Versatile descent algorithms for group regularization and variable selection in generalized linear models. Journal of Computational and Graphical Statistics, 34(1):239–252, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2362232.

Hennig:2004:ALD

[Hen04] Christian Hennig. Asymmetric linear dimension reduction for classification. Journal of Computational and Graphical Statistics, 13(4):930–945, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Herrmann:1997:LBC

[Her97]  Eva Herrmann.  Lo-

cal bandwidth choice in kernel regression estimation. Journal of Computational and Graphical Statistics, 6(1):35–54, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
hermann.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474726.

Hesterberg:1995:TSL

[Hes95] Tim Hesterberg. Tailspecific linear approximations for efficient bootstrap simulations. Journal of Computational and Graphical Statistics, 4(2): 113–133, June 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1995.10474670.

Hestir:1998:CCG

[Hes98] Kevin Hestir. Conditional coding, a general method for simulating conditional distributions. Journal of Computational and Graphical Statistics, 7(1):77–91, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.research.att.com/
                                        ~andreas/jcgs/abstracts/
                                        1998/march/hestir.html;
                                        http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474762.

Hofmann:2010:ELT

[HGK10] Marc Hofmann, Cristian Gatu, and Erricos John Kontoghiorghes. An exact least trimmed squares algorithm for a range of coverage values. Journal of Computational and Graphical Statistics, 19 (1):191–204, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Harezlak:2001:CTB

[HH01] Jaroslaw Harezlak and Nancy E. Heckman. CriSP: A tool for bump hunting. Journal of Computational and Graphical Statistics, 10(4):713–729, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243412.

Heiberger:2006:MMM

[HH06a] Richard M. Heiberger and Burt Holland. Mean– mean multiple comparison displays for families of linear contrasts. Journal of Computational and Graphical Statistics, 15(4):937–955, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hunter:2006:ICE

[HH06b] David R. Hunter and Mark S. Handcock. Inference in curved exponential family models for networks. Journal of Computational and Graphical Statistics, 15(3):565–583, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hahsler:2011:DPV

[HH11] Michael Hahsler and Kurt Hornik. Dissimilarity plots: a visual exploration tool for partitional clustering. Journal of Computational and Graphical Statistics, 20(2):335–354, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Haran:2003:ACM

[HHC03] Murali Haran, James S. Hodges, and Bradley P. Carlin. Accelerating computation in Markov random field models for spatial data via structured MCMC. Journal of Computational and Graphical Statistics, 12(2):249–264, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s1/p249;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031833.

Hall:2001:NEH

[HHGG01] Peter Hall, Li-Shan Huang, James A. Gifford, and Irene Gijbels. Nonparametric estimation of hazard rate under the constraint of monotonicity. Journal of Computational and Graphical Statistics, 10(3):592–614, September 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317115135.

Hummel:2012:ISB

[HHH12] Ruth M. Hummel, David R.

Hunter, and Mark S. Handcock. Improving simulation-based algorithms for fitting ERGMs. Journal of Computational and Graphical Statistics, 21(4):920–939, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

He:2022:IEH

[HHH22] Shiyuan He, Kejun He, and Jianhua Z. Huang. Improved estimation of high-dimensional additive models using subspace learning. Journal of Computational and Graphical Statistics, 31 (3):866–876, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2034638.

Hofner:2011:FUM

[HHKS11] Benjamin Hofner, Torsten Hothorn, Thomas Kneib, and Matthias Schmid.

A framework for unbiased model selection based on boosting. Journal of Computational and Graphical Statistics, 20 (4):956–971, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hahn:2019:ESG

[HHL19] P. Richard Hahn, Jingyu He, and Hedibert F. Lopes. Efficient sampling for Gaussian linear regression with arbitrary priors. Journal of Computational and Graphical Statistics, 28(1):142–154, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hsu:2021:MAS

[HHT21] Nan-Jung Hsu, Hsin-Cheng Huang, and Ruey S. Tsay. Matrix autoregressive spatio-temporal models. Journal of Computational and Graphical Statistics, 30(4):1143– 1155, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1938587.

Huang:2010:SLA

[HHTB10] Hsin-Cheng Huang, Nan-Jung Hsu, David M. Theobald, and F. Jay Breidt. Spatial lasso with applications to GIS model selection. Journal of Computational and Graphical Statistics, 19 (4):963–983, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hothorn:2006:URP

[HHZ06] Torsten Hothorn, Kurt Hornik, and Achim Zeileis. Unbiased recursive partitioning. Journal of Computational and Graphical Statistics, 15(3):651–674, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hicks:2017:GDS

[Hic17] Stephanie Hicks. Greater data science ahead! Journal of Computational and Graphical Statistics, 26 (4):770–771, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Higdon:2001:DAD

[Hig01] Dave Higdon. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):69–74, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1198/10618600152418737.

See [vDM01a].

Hill:2011:BNM

[Hil11] Jennifer L. Hill. Bayesian nonparametric modeling for causal inference. Journal of Computational and Graphical Statistics, 20 (1):217–240, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hamura:2022:DAM

[HIS22] Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. On data augmentation for models involving reciprocal gamma functions. Journal of Computational and Graphical Statistics, 32(3):908– 916, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2119988.

Hamura:2023:GSM

[HIS23] Yasuyuki Hamura, Kaoru Irie, and Shonosuke Sugasawa. Gibbs sampler for matrix generalized inverse Gaussian distributions. Journal of Computational and Graphical Statistics, 33(2):331–340, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2258186.

Hastie:1999:SMC

[HIT99] Trevor Hastie, Debra Ikeda, and Robert Tibshirani. Statistical measures for the computeraided diagnosis of mammographic masses. Journal of Computational and Graphical Statistics, 8(3): 531–543, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474830.

Hanson:2004:BSA

[HJ04] Timothy Hanson and Wesley O. Johnson. A Bayesian semiparametric AFT model for intervalcensored data. Journal of Computational and Graphical Statistics, 13 (2):341–361, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Holmes:2009:AMG

[HJ09] Chris Holmes and Ajay Jasra. Antithetic methods for Gibbs samplers. Journal of Computational and Graphical Statistics, 18(2):401–414, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Holmes:2017:DSY

[HJ17] Susan Holmes and Julie Josse. Discussion of ‘50 Years of Data Science. Journal of Computational and Graphical Statistics, 26(4):768–769, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Don17].

Huang:2022:PQM

[HJM22] Chaofan Huang, V. Roshan Joseph, and Simon Mak. Population quasi-Monte Carlo. Journal of Computational and Graphical Statistics, 31(3):695– 708, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2034637.

Husson:2019:IMD

[HJNR19] François Husson, Julie Josse, Balasubramanian Narasimhan, and Geneviève Robin. Imputation of mixed data with multilevel singular value decomposition. Journal of Computational and Graphical Statistics, 28 (3):552–566, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hurn:2003:EMR

[HJR03] Merrilee Hurn, Ana Justel, and Christian P. Robert. Estimating mixtures of regressions. Journal of Computational and Graphical Statistics, 12(1):55–79, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s3/p55;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031329.

Huerta:2001:DAD

[HJT01] Gabriel Huerta, Wenxin Jiang, and Martin A. Tanner. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10(1):82–89, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418755.

See [vDM01a].

Hurzeler:1998:MCA

[HK98] Markus Hürzeler and Hans R. Künsch. Monte Carlo approximations for general state-space models. Journal of Computational and Graphical Statistics, 7(2):175– 193, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/hurzeler.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474769.

Hansen:2006:OFM

[HK06] Ben B. Hansen and Stephanie Olsen Klopfer. Optimal full matching and related designs via network flows. Journal of Computational and Graphical Statistics, 15 (3):609–627, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hirose:2010:ELA

[HK10] Yoshihiro Hirose and Fumiyasu Komaki. An extension of least angle regression based on the information geometry of dually flat spaces. Journal of Computational and Graphical Statistics, 19 (4):1007–1023, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hare:2017:DMS

[HK17] Eric Hare and Andee Kaplan. Designing modular software: A case study in introductory statistics. Journal of Computational and Graphical Statistics, 26(3):493–500, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Heiner:2022:ESH

[HK22] Matthew Heiner and Athanasios Kottas. Estimation and selection for high-order Markov chains with Bayesian mixture transition distribution models. Journal of Computational and Graphical Statistics, 31 (1):100–112, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1979565.

Helgason:2016:COF

[HKP16] Hannes Helgason, Stefanos Kechagias, and Vladas Pipiras. Convex optimization and feasible circulant matrix embeddings in synthesis of stationary Gaussian fields. Journal of Computational and Graphical Statistics, 25(4):1158–1175, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hunter:2012:CSM

[HKS12] David R. Hunter, Pavel N.

Krivitsky, and Michael Schweinberger. Computational statistical methods for social network models. Journal of Computational and Graphical Statistics, 21(4):856–882, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hunter:2000:QRM

[HL00a] David R. Hunter and Kenneth Lange. Quantile regression via an MM algorithm. Journal of Computational and Graphical Statistics, 9(1):60–77, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Hunter.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474866.

Hunter:2000:ROT

[HL00b] David R. Hunter and Kenneth Lange. Rejoinder: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 52–59, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474865.  See

[LHY00].

Ha:2003:EFM

[HL03] Il Do Ha and Youngjo Lee. Estimating frailty models via Poisson hierarchical generalized linear models. Journal of Computational and Graphical Statistics, 12 (3):663–681, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s9/p663;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032256.

Hannig:2006:RSE

[HL06] Jan Hannig and Thomas C. M. Lee. Robust SiZer for exploration of regression structures and outlier detection. Journal of Computational and Graphical Statistics, 15 (1):101–117, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2023:DRD

[HL23] Shih-Ting Huang and Johannes Lederer. Deep-MoM: Robust deep learning with median-of-means. Journal of Computational and Graphical Statistics, 32(1):181–195, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2090947.

Holbrook:2021:MPB

[HLB^{+}21] Andrew J. Holbrook, Philippe Lemey, Guy Baele, Simon Dellicour, Dirk Brockmann, Andrew Rambaut, and Marc A. Suchard. Massive parallelization boosts big Bayesian multidimensional scaling. Journal of Computational and Graphical Statistics, 30 (1):11–24, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1754226.

Huang:2013:MDW

[HLD^{+}13] Hanwen Huang, Yufeng Liu, Ying Du, Charles M. Perou, D. Neil Hayes, Michael J. Todd, and J. S. Marron. Multiclass distance-weighted discrimination. Journal of Computational and Graphical Statistics, 22 (4):953–969, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Holloman:2006:MGA

[HLH06] Christopher H. Holloman, Herbert K. H. Lee, and Dave M. Higdon. Multiresolution genetic algorithms and Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 15 (4):861–879, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Han:2015:CGM

[HLH15] Chao Han, Scotland Leman, and Leanna House. Covariance-guided mixture probabilistic principal component analysis (C-MPPCA). Journal of Computational and Graphical Statistics, 24 (1):66–83, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2007:ELC

[HLL07] Jianhua Z. Huang, Linxu Liu, and Naiping Liu. Estimation of large covariance matrices of longitudinal data with basis function approximations. Journal of Computational and Graphical Statistics, 16(1):189–209, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hall:2007:MPF

[HLP07] Peter Hall, Young K.

Lee, and Byeong U. Park. A method for projecting functional data onto a low-dimensional space. Journal of Computational and Graphical Statistics, 16(4):799–812, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hu:2022:GTD

[HLW22] Jiaxin Hu, Chanwoo Lee, and Miaoyan Wang. Generalized tensor decomposition with features on multiple modes. Journal of Computational and Graphical Statistics, 31 (1):204–218, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1978471.

He:2024:DLP

[HLW24] Yong He, Zichen Liu, and Yalin Wang. Distributed learning for principal eigenspaces without moment constraints. Journal of Computational and Graphical Statistics, 34(1):318–329, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2341889.

Huang:2015:SSC

[HLYM15] Hanwen Huang, Yufeng Liu, Ming Yuan, and J. S. Marron. Statistical significance of clustering using soft thresholding. Journal of Computational and Graphical Statistics, 24 (4):975–993, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.948179.

Hothorn:2005:DAB

[HLZH05] Torsten Hothorn, Friedrich Leisch, Achim Zeileis, and Kurt Hornik. The design and analysis of benchmark experiments. Journal of Computational and Graphical Statistics, 14 (3):675–699, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hu:2002:CAE

[HM02] Xiaomi Hu and Hari Mukerjee. Constructing all extremes of contingency tables with given marginals. Journal of Computational and Graphical Statistics, 11(4):910–919, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s11/p910;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002709.

Hall:2009:UGC

[HM09] Peter Hall and Hugh Miller. Using generalized correlation to effect variable selection in very high dimensional problems. Journal of Computational and Graphical Statistics, 18(3):533–550, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hall:2011:DDR

[HM11] Peter Hall and Hugh Miller. Determining and depicting relationships among components in high-dimensional variable selection. Journal of Computational and Graphical Statistics, 20 (4):988–1006, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hofert:2014:GGF

[HM14] Marius Hofert and Martin Mächler. A graphical goodness-of-fit test for dependence models in higher dimensions. Journal of Computational and Graphical Statistics, 23 (3):700–716, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Helwig:2015:FSM

[HM15] Nathaniel E. Helwig and Ping Ma. Fast and stable multiple smoothing parameter selection in smoothing spline analysis of variance models with large samples. Journal of Computational and Graphical Statistics, 24 (3):715–732, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Homrighausen:2016:NCS

[HM16] Darren Homrighausen and Daniel J. McDonald. On the Nyström and column-sampling methods for the approximate principal components analysis of large datasets. Journal of Computational and Graphical Statistics, 25(2):344–362, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Homrighausen:2020:CPL

[HM20] Darren Homrighausen and Daniel J. McDonald. Compressed and penalized linear regression. Journal of Computational and Graphical Statistics, 29(2):309–322, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1660179.

Hanson:2011:PTS

[HMJ11] Timothy E. Hanson, João V. D. Monteiro, and Alejandro Jara. The Polya tree sampler: Toward efficient and automatic independent Metropolis– Hastings proposals. Journal of Computational and Graphical Statistics, 20 (1):41–62, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hoegh:2018:MMS

[HML18] Andrew Hoegh, Dipayan Maiti, and Scotland Leman. Multiset model selection. Journal of Computational and Graphical Statistics, 27(2):436–448, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hoar:2003:WBH

[HMN^{+}03] Timothy J. Hoar, Ralph F. Milliff, Douglas Nychka, Christopher K. Wikle, and L. Mark Berliner. Winds from a Bayesian hierarchical model: Computation for atmosphereocean research. Journal of Computational and Graphical Statistics, 12(4):781–807, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/1061860032616.

Hall:1996:AAN

[HMP96] Peter Hall, David Matthews, and Eckhard Platen. Algorithms for analyzing nonstationary time series with fractal noise. Journal of Computational and Graphical Statistics, 5(4): 351–364, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474717.

Harris:2019:CVC

[HMPP19] David Harris, Gael M.

Martin, Indeewara Perera, and D. S. Poskitt. Construction and visualization of confidence sets for frequentist distributional forecasts. Journal of Computational and Graphical Statistics, 28 (1):92–104, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:1997:VRB

[HMS97] Chisheng Huang, John Alan McDonald, and Werner Stuetzle. Variable resolution bivariate plots. Journal of Computational and Graphical Statistics, 6(4):383–396, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/huang.
html;   http://www.tandfonline.

                                        com/doi/abs/10.1080/
                                        10618600.1997.10474749.

Hall:2006:RTD

[HMW06] Peter Hall, Hans-Georg Müller, and Ping-Shi Wu. Real-time density and mode estimation with application to time-dynamic mode tracking. Journal of Computational and Graphical Statistics, 15 (1):82–100, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huo:2022:UFA

[HMZ22] Shuning Huo, Jeffrey S. Morris, and Hongxiao Zhu. Ultra-fast approximate inference using variational functional mixed models. Journal of Computational and Graphical Statistics, 32(2):353– 365, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2107532.

Hsu:1998:MCG

[HN98] Jason C. Hsu and Barry Nelson. Multiple comparisons in the general linear model. Journal of Computational and Graphical Statistics, 7(1):23–41, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.research.att.com/
~andreas/jcgs/abstracts/
1998/march/hsu.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474759.

Hall:1993:ELC

[HO93] Peter Hall and Art B.

Owen. Empirical likelihood confidence bands in density estimation. Journal of Computational and Graphical Statistics, 2(3): 273–289, September 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474612.

Huang:1999:IMP

[HO99] Fuchun Huang and Yosihiko Ogata. Improvements of the maximum pseudo-likelihood estimators in various spatial statistical models. Journal of Computational and Graphical Statistics, 8(3): 510–530, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474829.

Ho:2006:BES

[Ho06] Man-Wai Ho. Bayes estimation of a symmetric unimodal density via Spaths. Journal of Computational and Graphical Statistics, 15(4):848–860, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hurley:2010:PDH

[HO10] C. B. Hurley and R. W. Oldford. Pairwise display of high-dimensional information via Eulerian tours and Hamiltonian decompositions. Journal of Computational and Graphical Statistics, 19 (4):861–886, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hobert:2001:DAD

[Hob01a] James P. Hobert. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):59–68, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/10618600152418728.

See [vDM01a].

Hobert:2001:SRA

[Hob01b] James P. Hobert. Stability relationships among the Gibbs sampler and its subchains. Journal of Computational and Graphical Statistics, 10 (2):185–205, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628130.

Hobolth:2008:MCM

[Hob08] Asger Hobolth. A Markov chain Monte Carlo expectation maximization algorithm for statistical analysis of DNA sequence evolution with neighbordependent substitution rates. Journal of Computational and Graphical Statistics, 17(1):138–162, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hurley:2022:ISV

[HOD22] Catherine B. Hurley, Mark O’Connell, and Katarina Domijan. Interactive slice visualization for exploring machine learning models. Journal of Computational and Graphical Statistics, 31 (1):1–13, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1983439.

Hoefling:2010:PAF

[Hoe10] Holger Hoefling. A path algorithm for the fused lasso signal approximator. Journal of Computational and Graphical Statistics, 19(4):984–1006, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hoff:2000:CNM

[Hof00] Peter D. Hoff. Constrained nonparametric maximum likelihood via mixtures. Journal of Computational and Graphical Statistics, 9(4):633– 641, December 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.amstat.org/publications/
                                        jcgs/abstracts00/Hoff.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474904.

Hofmann:2001:GOR

[Hof01] Heike Hofmann. Generalized odds ratios for visual modeling. Journal of Computational and Graphical Statistics, 10(4):628–640, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317243368.

Hoff:2009:SMB

[Hof09] Peter D. Hoff. Simulation of the matrix Bingham– von Mises–Fisher distribution, with applications to multivariate and relational data. Journal of Computational and Graphical Statistics, 18 (2):438–456, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hofmeyr:2019:ISC

[Hof19] David P. Hofmeyr. Improving spectral clustering using the asymptotic value of the normalized cut. Journal of Computational and Graphical Statistics, 28(4):980–992, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Holmes:1993:CSM

[Hol93] Susan Holmes. Comment on “A Model for Studying Display Methods of Statistical Graphics”. Journal of Computational and Graphical Statistics, 2(4): 349–353, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474618.  See

[Cle93a].

Holbrook:2023:QPM

[Hol23] Andrew J. Holbrook. A quantum parallel Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 32 (4):1402–1415, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2195890.

Hooker:2007:GFA

[Hoo07] Giles Hooker. Generalized functional ANOVA diagnostics for highdimensional functions of dependent variables. Journal of Computational and Graphical Statistics, 16(3):709–732, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hsu:1994:GRT

[HP94] Jason C. Hsu and Mario Peruggia. Graphical representations of Tukey’s multiple comparison method. Journal of Computational and Graphical Statistics, 3(2):143–161, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474636.

Hall:1999:DEU

[HP99] Peter Hall and Brett Presnell. Density estimation under constraints. Journal of Computational and Graphical Statistics, 8(2):259–277, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/density.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/presnell.pdf;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474813.

Helgason:2014:SWS

[HPA14] Hannes Helgason, Vladas Pipiras, and Patrice Abry. Smoothing windows for the synthesis of Gaussian stationary random fields using circulant matrix embedding. Journal of Computational and Graphical Statistics, 23 (3):616–635, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ha:2014:VSG

[HPOL14] Il Do Ha, Jianxin Pan, Seungyoung Oh, and Youngjo Lee. Variable selection in general frailty models using penalized H-likelihood. Journal of Computational and Graphical Statistics, 23 (4):1044–1060, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Heiner:2024:PAL

[HPQ24] Matthew J. Heiner, Garritt L. Page, and Fernando Andrés Quintana. A projection approach to local regression with variable-dimension covariates. Journal of Computational and Graphical Statistics, 34(1):109– 122, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2357636.

Hofert:2021:QRS

[HPZ21] Marius Hofert, Avinash Prasad, and Mu Zhu. Quasi-random sampling for multivariate distributions via generative neural networks. Journal of Computational and Graphical Statistics, 30 (3):647–670, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1868302.

Hofert:2022:DMA

[HPZ22] Marius Hofert, Avinash Prasad, and Mu Zhu. Dependence model assessment and selection with DecoupleNets. Journal of Computational and Graphical Statistics, 32 (4):1272–1286, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2157835.

Hall:1992:RFN

[HQT92] Peter Hall, Wei Qian, and D. M. Titterington. Ridge finding from noisy data. Journal of Computational and Graphical Statistics, 1(3):197–211, September 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474581.

Hardin:2005:DRD

[HR05] Johanna Hardin and David M. Rocke. The distribution of robust distances. Journal of Computational and Graphical Statistics, 14(4):928–946, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hobert:2011:CCC

[HR11] James P. Hobert and Jorge Carlos Román. Comment: To Center or Not to Center: That Is Not the Question—-An Ancillarity—Sufficiency Interweaving Strategy (ASIS) for Boosting MCMC Efficiency. Journal of Computational and Graphical Statistics, 20 (3):571–580, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [YM11b].

Hoeting:2002:BVT

[HRM02] Jennifer A. Hoeting, Adrian E. Raftery, and David Madigan. Bayesian variable and transformation selection in linear regression. Journal of Computational and Graphical Statistics, 11(3):485–507, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s1/p485;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002501.

Hubert:2012:DAR

[HRV12] Mia Hubert, Peter J. Rousseeuw, and Tim Verdonck. A deterministic algorithm for robust location and scatter. Journal of Computational and Graphical Statistics, 21 (3):618–637, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hu:2008:EED

[HS08] Jiaqiao Hu and Zheng Su. Efficient error determination in sequential clinical trial design.

Journal of Computational and Graphical Statistics, 17(4):925–948, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hyndman:2010:RPB

[HS10] Rob J. Hyndman and Han Lin Shang. Rainbow plots, bagplots, and boxplots for functional data. Journal of Computational and Graphical Statistics, 19(1):29–45, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2018:HLR

[HS18] Huang Huang and Ying Sun. Hierarchical low rank approximation of likelihoods for large spatial datasets. Journal of Computational and Graphical Statistics, 27 (1):110–118, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2018:SMS

[HSC18] Weihong Huang, Juan Shen, and Yuguo Chen. The self-multiset sampler. Journal of Computational and Graphical Statistics, 27(1):34–47, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huang:2023:TVC

[HSG23] Huang Huang, Ying Sun, and Marc G. Genton. Test and visualization of covariance properties for multivariate spatiotemporal random fields. Journal of Computational and Graphical Statistics, 32(4):1545–1555, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2191667.

Hsieh:1999:RTI

[Hsi99] Ping-Hung Hsieh. Robustness of tail index estimation. Journal of Computational and Graphical Statistics, 8(2):318– 332, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/tailindex.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/hsieh.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474816.

Hamidieh:2009:EEI

[HSM09] Kamal Hamidieh, Stilian Stoev, and George Michailidis. On the estimation of the extremal index based on scaling and resampling. Journal of Computational and Graphical Statistics, 18 (3):731–755, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hsu:1992:FAA

[Hsu92] Jason C. Hsu. The factor analytic approach to simultaneous inference in the general linear model. Journal of Computational and Graphical Statistics, 1(2):151–168, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477011.

Huser:2023:VLA

[HSZ23] Raphaël Huser, Michael L.

Stein, and Peng Zhong. Vecchia likelihood approximation for accurate and fast inference with intractable spatial maxstable models. Journal of Computational and Graphical Statistics, 33 (3):978–990, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2285332.

Heard:2017:ASM

[HT17] Nicholas A. Heard and Melissa J. M. Turcotte. Adaptive sequential Monte Carlo for multiple changepoint analysis. Journal of Computational and Graphical Statistics, 26 (2):414–423, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hecq:2022:HRM

[HTW22] Alain Hecq, Marie Ternes, and Ines Wilms. Hierarchical regularizers for mixed-frequency vector autoregressions. Journal of Computational and Graphical Statistics, 31 (4):1076–1090, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2058003.

Huber:1999:MDW

[Hub99] Peter J. Huber. Massive datasets workshop: Four years after. Journal of Computational and Graphical Statistics, 8(3): 635–652, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1999.10474838.

Huber:2000:LSD

[Hub00] Peter J. Huber. Languages for statistics and data analysis. Journal of Computational and Graphical Statistics, 9(3): 600–620, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474901.

Hughes:2015:CFR

[Hug15] John Hughes. copCAR: A flexible regression model for areal data. Journal of Computational and Graphical Statistics, 24 (3):733–755, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Huh:1995:EMD

[Huh95] Moon Yul Huh. Exploring multidimensional data with the flipped empirical distribution function. Journal of Computational and Graphical Statistics, 4(4):335–343, 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474688.

Hui:2022:GAF

[Hui22] Francis K. C. Hui. GEEassisted forward regression for spatial latent variable models. Journal of Computational and Graphical Statistics, 31 (4):1013–1024, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2058002.

Hurley:1993:PDI

[Hur93] Catherine Hurley. The plot-data interface in statistical graphics. Journal of Computational and Graphical Statistics, 2(4): 365–379, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474621.

Hurley:2000:GSD

[Hur00] Catherine B. Hurley. Graphical selection of data views. Journal of Computational and Graphical Statistics, 9(3):558–581, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Hurley.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474899.

Hurley:2004:CVM

[Hur04] Catherine B. Hurley. Clustering visualizations of multidimensional data. Journal of Computational and Graphical Statistics, 13(4):788–806, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hutson:2002:EBC

[Hut02] Alan D. Hutson. “exact” bootstrap confidence bands for the quantile function via Steck’s determinant. Journal of Computational and Graphical Statistics, 11(2):471– 482, June 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s13/p471;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180626.

Hernandez:2005:DRN

[HV05] Adolfo Hernández and Santiago Velilla. Dimension reduction in nonparametric kernel discriminant analysis. Journal of Computational and Graphical Statistics, 14 (4):847–866, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hofmann:2017:AHB

[HV17] Heike Hofmann and Susan VanderPlas. All of this has happened before. All of this will happen again: Data science. Journal of Computational and Graphical Statistics, 26(4):775–778, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Henderson:2019:DAA

[HV19] Nicholas C. Henderson and Ravi Varadhan. Damped Anderson acceleration with restarts and monotonicity control for accelerating EM and EMlike algorithms. Journal of Computational and Graphical Statistics, 28 (4):834–846, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hermes:2023:CGM

[HvHB23] Sjoerd Hermes, Joost van Heerwaarden, and Pariya Behrouzi. Copula graphical models for heterogeneous mixed data. Journal of Computational and Graphical Statistics, 33 (3):991–1005, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2289545.

Hirji:1996:ETI

[HVRA96] Karim F. Hirji, Stein E. Vollset, Isildinha M. Reis, and Abdelmonem A. Afifi. Exact tests for interaction in several 2 × 2 tables. Journal of Computational and Graphical Statistics, 5(3):209–224, September 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474706.

Hu:2021:QAP

[HW21] Jianchang Hu and Yazhen Wang. Quantum annealing via path-integral Monte Carlo with data augmentation. Journal of Computational and Graphical Statistics, 30 (2):284–296, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1814787.

Huang:2022:SCA

[HWF22] Sihan Huang, Haolei Weng, and Yang Feng. Spectral clustering via adaptive layer aggregation for multi-layer networks. Journal of Computational and Graphical Statistics, 32(3):1170– 1184, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2134874.

Hofmann:2017:LVP

[HWK17] Heike Hofmann, Hadley Wickham, and Karen Kafadar. Letter-value plots: Boxplots for large data. Journal of Computational and Graphical Statistics, 26(3):469–477, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Han:2024:MUC

[HWLB24] Yang Han, Lingjiao Wang, Wei Liu, and Frank Bretz. Multiple-use calibration for all future values and exact two-sided simultaneous tolerance intervals in linear regression. Journal of Computational and Graphical Statistics, 34 (1):155–161, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2359507.

Hui:2017:VAG

[HWO^{+}17] Francis K. C. Hui, David I. Warton, John T. Ormerod, Viivi Haapaniemi, and Sara Taskinen. Variational approximations for generalized linear latent variable models. Journal of Computational and Graphical Statistics, 26 (1):35–43, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1164708.

Hu:2015:EIU

[HWQ15] Jianhua Hu, Peng Wang, and Annie Qu. Estimating and identifying unspecified correlation structure for longitudinal data. Journal of Computational and Graphical Statistics, 24 (2):455–476, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.909733.

Haris:2016:CMI

[HWS16] Asad Haris, Daniela Witten, and Noah Simon. Convex modeling of interactions with strong heredity. Journal of Computational and Graphical Statistics, 25(4):981– 1004, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hilafu:2017:SDR

[HY17] Haileab Hilafu and Xiangrong Yin. Sufficient dimension reduction and variable selection for large-p–small-n data with highly correlated predictors. Journal of Computational and Graphical Statistics, 26 (1):26–34, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2016.1164057.

Huang:2019:CBS

[HY19] Li-Shan Huang and Chung-Hsin Yu. Classical backfitting for smoothbackfitting additive models. Journal of Computational and Graphical Statistics, 28(2):386–400, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

He:2017:MTV

[HYL^{+}17] Kevin He, Yuan Yang, Yanming Li, Ji Zhu, and Yi Li. Modeling timevarying effects with largescale survival data: An efficient quasi-Newton approach. Journal of Computational and Graphical Statistics, 26(3):635–645, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Hothorn:2021:PDM

[HZ21] Torsten Hothorn and Achim Zeileis. Predictive distribution modeling using transformation forests. Journal of Computational and Graphical Statistics, 30(4):1181– 1196, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1872581.

Heng:2023:BTF

[HZC23] Qiang Heng, Hua Zhou, and Eric C. Chi. Bayesian trend filtering via proximal Markov Chain Monte Carlo. Journal of Computational and Graphical Statistics, 32(3):938– 949, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2170089.

Ihaka:1996:RLD

[IG96] Ross Ihaka and Robert Gentleman. R: a language for data analysis and graphics. Journal of Computational and Graphical Statistics, 5(3): 299–314, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474713.

Ishwaran:2002:ADP

[IJ02] Hemant Ishwaran and Lancelot F. James. Approximate Dirichlet process computing in finite normal mixtures:

Smoothing and prior information. Journal of Computational and Graphical Statistics, 11(3):508–532, September 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s2/p508;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002411.

Imai:2008:TCF

[IKL08] Kosuke Imai, Gary King, and Olivia Lau. Toward a common framework for statistical analysis and development. Journal of Computational and Graphical Statistics, 17(4):892–913, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Izbicki:2016:NCD

[IL16] Rafael Izbicki and Ann B. Lee. Nonparametric conditional density estimation in a high-dimensional regression setting. Journal of Computational and Graphical Statistics, 25 (4):1297–1316, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Izbicki:2019:ACT

[ILP19] Rafael Izbicki, Ann B.

Lee, and Taylor Pospisil. ABC-CDE: Toward approximate Bayesian computation with complex high-dimensional data and limited simulations. Journal of Computational and Graphical Statistics, 28(3):481–492, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ionides:2008:TIS

[Ion08] Edward L. Ionides. Truncated importance sampling. Journal of Computational and Graphical Statistics, 17(2):295–311, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Iovleff:2004:ENS

[IP04] Serge Iovleff and Olivier Perrin. Estimating a nonstationary spatial structure using simulated annealing. Journal of Computational and Graphical Statistics, 13(1):90–105, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Inglis:2022:VVI

[IPH22] Alan Inglis, Andrew Parnell, and Catherine B. Hurley. Visualizing variable importance and variable interaction effects in machine learning models. Journal of Computational and Graphical Statistics, 31(3):766–778, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2007935.

Ishwaran:1999:AHM

[Ish99] Hemant Ishwaran. Applications of hybrid Monte Carlo to Bayesian generalized linear models: Quasicomplete separation and neural networks. Journal of Computational and Graphical Statistics, 8(4):779–799, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/ishwar.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474849.

Ilies:2010:PBP

[IW10] Iulian Ilies and Adalbert Wilhelm. Projectionbased partitioning for large, high-dimensional datasets. Journal of Computational and Graphical Statistics, 19(2):474–492, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Isberg:2022:ADA

[IW22] Sonja Isberg and William J.

Welch. Adaptive design and analysis via partitioning trees for emulation of a complex computer code. Journal of Computational and Graphical Statistics, 31(4):1280–1291, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2039160.

Jung:2019:POI

[JAJ19] Sungkyu Jung, Jeongyoun Ahn, and Yongho Jeon. Penalized orthogonal iteration for sparse estimation of generalized eigenvalue problem. Journal of Computational and Graphical Statistics, 28 (3):710–721, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jank:2006:IDS

[Jan06]  Wolfgang Jank.  Im-

plementing and diagnosing the stochastic approximation EM algorithm. Journal of Computational and Graphical Statistics, 15(4):803–829, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jank:2003:EMC

[JB03] Wolfgang Jank and James Booth. Efficiency of Monte Carlo EM and simulated maximum likelihood in two-stage hierarchical models. Journal of Computational and Graphical Statistics, 12(1):214–229, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s11/p214;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031338.

Jin:2022:CHV

[JB22] Honghe Jin and Lynne Billard. Copulas and histogram-valued data. Journal of Computational and Graphical Statistics, 32(2):642–657, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2107535.

Jiang:2022:ABS

[JBJ^{+}22] Wei Jiang, Małgorzata Bogdan, Julie Josse, Szymon Majewski, Błazej Miasojedow, Veronika Rocková, and TraumaBase Group. Adaptive Bayesian SLOPE:

Model selection with incomplete data. Journal of Computational and Graphical Statistics, 31 (1):113–137, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1963263.

Johns:2019:VDF

[JCZG19] Jordan T. Johns, Ciprian Crainiceanu, Vadim Zipunnikov, and Jonathan Gellar. Variable-domain functional principal component analysis. Journal of Computational and Graphical Statistics, 28 (4):993–1006, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

James:1998:ECM

[JH98] Gareth James and Trevor Hastie. The error coding method and PICTs. Journal of Computational and Graphical Statistics, 7(3):377–387, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/james.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474782.

Jauch:2021:MCS

[JHD21] Michael Jauch, Peter D. Hoff, and David B. Dunson. Monte Carlo simulation on the Stiefel manifold via polar expansion. Journal of Computational and Graphical Statistics, 30(3):622–631, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1859382.

Jara:2009:RGL

[JHL09] Alejandro Jara, Timothy E. Hanson, and Emmanuel Lesaffre. Robustifying generalized linear mixed models using a new class of mixtures of multivariate Polya trees. Journal of Computational and Graphical Statistics, 18(4):838–860, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jalaluddin:2000:ARI

[JK00] Muhammad Jalaluddin and Michael R. Kosorok. An algorithm for robust inference for the Cox model with frailties. Journal of Computational and Graphical Statistics, 9(4):642–652, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Jalaluddin.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474905.

Jurek:2021:MRF

[JK21] Marcin Jurek and Matthias Katzfuss. Multi-resolution filters for massive spatiotemporal data. Journal of Computational and Graphical Statistics, 30 (4):1095–1110, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1886938.

Jiang:1992:CCM

[JKD92] Thomas J. Jiang, Joseph B.

Kadane, and James M. Dickey. Computation of Carlson’s multiple hypergeometric function R for Bayesian applications. Journal of Computational and Graphical Statistics, 1(3):231–251, September 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474583.

Jin:2013:FSN

[JL13] Ick Hoon Jin and Faming Liang. Fitting social network models using varying truncation stochastic approximation MCMC algorithm. Journal of Computational and Graphical Statistics, 22(4):927–952, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jiang:2022:EAF

[JLAW22] Ci-Ren Jiang, Eardi Lila, John A. D. Aston, and Jane-Ling Wang. Eigenadjusted functional principal component analysis. Journal of Computational and Graphical Statistics, 31(4):1141–1152, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2067550.

Jiang:2016:CAO

[JMY^{+}16] Hui Jiang, John Chong Mu, Kun Yang, Chao Du, Luo Lu, and Wing Hung Wong. Computational aspects of optional Pólya tree. Journal of Computational and Graphical Statistics, 25(1):301– 320, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.1002927;   http:
//www.tandfonline.com/
doi/full/10.1080/10618600.
2014.1002927.

Jain:2004:SMM

[JN04] Sonia Jain and Radford M. Neal. A split-merge Markov chain Monte Carlo procedure for the Dirichlet process mixture model. Journal of Computational and Graphical Statistics, 13 (1):158–182, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jones:2002:MCM

[JO02] L. K. Jones and P. J.

O’Neil. On Markov chain Monte Carlo algorithms for computing conditional expectations based on sufficient statistics. Journal of Computational and Graphical Statistics, 11(3):660–677, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s9/p660;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002510.

Johnson:2013:DPA

[Joh13] Nicholas A. Johnson. A dynamic programming algorithm for the fused lasso and L_{0}-segmentation. Journal of Computational and Graphical Statistics, 22(2):246–260, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jongbloed:1998:ICM

[Jon98] Geurt Jongbloed. The iterative convex minorant algorithm for nonparametric estimation. Journal of Computational and Graphical Statistics, 7(3):310–321, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/jongbloed.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474778.

Jornsten:2009:SMS

[Jör09] Rebecka Jörnsten. Simultaneous model selection via rate-distortion theory, with applications to cluster and significance analysis of gene expression data. Journal of Computational and Graphical Statistics, 18(3):613–639, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Juan:1995:SMC

[JP95] Jesus Juan and Francisco J. Prieto. A subsampling method for the computation of multivariate estimators with high breakdown point. Journal of Computational and Graphical Statistics, 4(4): 319–334, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474687.

Justel:1996:GSW

[JP96] Ana Justel and Daniel Peña. Gibbs sampling will fail in outlier problems with strong masking. Journal of Computational and Graphical Statistics, 5(2):176–189, June 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474703.

Jung:2022:SIH

[JP22] Yohan Jung and Jinkyoo Park. Scalable inference for hybrid Bayesian hidden Markov model using Gaussian process emission. Journal of Computational and Graphical Statistics, 31(3):666– 683, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2023021.

Jeong:2022:BMS

[JPvD22] Seonghyun Jeong, Taeyoung Park, and David A. van Dyk. Bayesian model selection in additive partial linear models via locally adaptive splines. Journal of Computational and Graphical Statistics, 31(2):324–336, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1999827.

Jones:2022:RCN

[JRH22] Corinne Jones, Vincent Roulet, and Zaid Harchaoui. Revisiting convolutional neural networks from the viewpoint of kernel-based methods. Journal of Computational and Graphical Statistics, 32(4):1237–1247, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2163649.

Jacob:2011:UPC

[JRS11] P. Jacob, C. P. Robert, and M. H. Smith. Using parallel computation to improve independent Metropolis–Hastings based estimation. Journal of Computational and Graphical Statistics, 20 (3):616–635, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jennison:1995:TEP

[JS95] Christopher Jennison and Nuala Sheehan. Theoretical and empirical properties of the genetic algorithm as a numerical optimizer. Journal of Computational and Graphical Statistics, 4(4):296– 318, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474686.

Ji:2013:AMC

[JS13] Chunlin Ji and Scott C.

Schmidler. Adaptive Markov Chain Monte Carlo for Bayesian variable selection. Journal of Computational and Graphical Statistics, 22 (3):708–728, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Johansson:2020:RSB

[JS20] Per Johansson and Mårten Schultzberg. Rerandomization strategies for balancing covariates using pre-experimental longitudinal data. Journal of Computational and Graphical Statistics, 29 (4):798–813, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1753531.

Jin:2021:FMC

[JT21] Rui Jin and Aixin Tan. Fast Markov chain Monte Carlo for high-dimensional Bayesian regression models with shrinkage priors. Journal of Computational and Graphical Statistics, 30(3):632–646, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1864383.

Jolliffe:2003:MPC

[JTU03] Ian T. Jolliffe, Nickolay T. Trendafilov, and Mudassir Uddin. A modified principal component technique based on the LASSO. Journal of Computational and Graphical Statistics, 12 (3):531–547, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s3/p531;   http:

//www.tandfonline.com/
doi/abs/10.1198/1061860032148.

Jolliffe:2000:SCT

[JU00] Ian T. Jolliffe and Mudassir Uddin. The simplified component technique: An alternative to rotated principal components. Journal of Computational and Graphical Statistics, 9(4):689–710, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Jolliffe.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474908.

Ju:2001:HHO

[JV01] Wen-Hua Ju and Yehuda Vardi. A hybrid highorder Markov chain model for computer intrusion detection. Journal of Computational and Graphical Statistics, 10(2):277– 295, June 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628068.

Jimenez-Varon:2024:FHD

[JVSS24] Cristian F. Jiménez-Varón, Ying Sun, and Han Lin Shang. Forecasting high-dimensional functional time series: Application to sub-national age-specific mortality. Journal of Computational and Graphical Statistics, 33(4):1160–1174, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2319166.

Jaki:2008:MKL

[JW08] Thomas Jaki and R. Webster West. Maximum kernel likelihood estimation. Journal of Computational and Graphical Statistics, 17(4):976–993, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Jiang:2013:ISR

[JWB13] Liewen Jiang, Huixia Judy Wang, and Howard D. Bondell. Interquantile shrinkage in regression models. Journal of Computational and Graphical Statistics, 22(4):970–986, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Josse:2016:CAF

[JWH16] Julie Josse, Stefan Wager, and François Husson. Confidence areas for fixed-effects PCA. Journal of Computational and Graphical Statistics, 25 (1):28–48, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.950871;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.950871.

Jaeger:2023:AIO

[JWL^{+}23] Byron C. Jaeger, Sawyer Welden, Kristin Lenoir, Jaime L. Speiser, Matthew W. Segar, Ambarish Pandey, and Nicholas M. Pajewski. Accelerated and interpretable oblique random survival forests. Journal of Computational and Graphical Statistics, 33 (1):192–207, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2231048.

Jing:2018:FMA

[JYYZ18] Bingyi Jing, Guangren Yang, Xianshi Yu, and Cunhui Zhang. Fused-MCP with application to signal processing. Journal of Computational and Graphical Statistics, 27 (4):872–886, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Karagiannis:2013:AIS

[KA13] Georgios Karagiannis and Christophe Andrieu. Annealed importance sampling reversible jump MCMC algorithms. Journal of Computational and Graphical Statistics, 22 (3):623–648, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kabaila:1993:MCC

[Kab93] Paul Kabaila. A method for the computer calculation of Edgeworth expansions for smooth function models. Journal of Computational and Graphical Statistics, 2(2):199– 207, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474607.

Kang:2020:CBI

[Kan20] Yicheng Kang. Consistent blind image deblurring using jump-preserving extrapolation. Journal of Computational and Graphical Statistics, 29 (2):372–382, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1665536.

Karabatsos:2021:FSE

[Kar21] George Karabatsos. Fast search and estimation of Bayesian nonparametric mixture models using a classification annealing EM algorithm. Journal of Computational and Graphical Statistics, 30 (1):236–247, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1807995.

Kato:2010:SRP

[Kat10] Kengo Kato. Solving ℓ_{1} regularization problems with piecewise linear losses. Journal of Computational and Graphical Statistics, 19(4):1024– 1040, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kahn:1999:WSW

[KB99] Ralph Kahn and Amy Braverman. What shall we do with the data we are expecting from upcoming Earth observation satellites? Journal of Computational and Graphical Statistics, 8(3): 575–588, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474834.

Kalisch:2008:RPA

[KB08] Markus Kalisch and Peter Bühlmann. Robustification of the PCalgorithm for directed acyclic graphs. Journal of Computational and Graphical Statistics, 17 (4):773–789, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kowal:2020:BFS

[KB20] Daniel R. Kowal and Daniel C. Bourgeois. Bayesian function-on-scalars regression for high-dimensional data. Journal of Computational and Graphical Statistics, 29(3):629–638, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1710837.

Kim:2023:SVV

[KBM23] Minwoo Kim, Shrijita Bhattacharya, and Tapabrata Maiti. Statistically valid variational Bayes algorithm for Ising model parameter estimation. Journal of Computational and Graphical Statistics, 33 (1):75–84, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2023.2217869.

Kreuzer:2020:EBI

[KC20] Alexander Kreuzer and Claudia Czado. Efficient Bayesian inference for nonlinear state space models with univariate autoregressive state equation. Journal of Computational and Graphical Statistics, 29(3):523– 534, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1725523.

Kim:2019:UNM

[KCB19] Jaeoh Kim, HyungJun Cho, and Sungwan Bang. Unified noncrossing multiple quantile regressions tree. Journal of Computational and Graphical Statistics, 28(2):454–465, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kenney:2021:MBE

[KCF21] Ana Kenney, Francesca Chiaromonte, and Giovanni Felici. MIP-BOOST: Efficient and effective L_{0} feature selection for linear regression. Journal of Computational and Graphical Statistics, 30 (3):566–577, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1845184.

Krivobokova:2008:FAP

[KCK08] Tatyana Krivobokova, Ciprian M. Crainiceanu, and Göran Kauermann. Fast adaptive penalized splines. Journal of Computational and Graphical Statistics, 17(1):1–20, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kauermann:2009:BPS

[KCO09] Göran Kauermann, Gerda Claeskens, and J. D. Opsomer. Bootstrapping for penalized spline regression. Journal of Computational and Graphical Statistics, 18(1):126–146, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kim:2020:FAM

[KCSA20] Youngseok Kim, Peter Carbonetto, Matthew Stephens, and Mihai Anitescu. A fast algorithm for maximum likelihood estimation of mixture proportions using sequential quadratic programming. Journal of Computational and Graphical Statistics, 29(2):261–273, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1689985.

Kabisa:2016:OVB

[KDM16] Sylvie (Tchumtchoua)

Kabisa, David B. Dunson, and Jeffrey S. Morris. Online variational Bayes inference for highdimensional correlated data. Journal of Computational and Graphical Statistics, 25(2):426–444, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Keim:1996:POV

[Kei96] Daniel A. Keim. Pixeloriented visualization techniques for exploring very large data bases. Journal of Computational and Graphical Statistics, 5(1): 58–77, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474695.

Kelly:2011:CDA

[Kel11] Brandon C. Kelly. Comment: Data Augmentation, Internal Representation, and Unsupervised Learning. Journal of Computational and Graphical Statistics, 20 (3):584–591, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Wu11].

Ke:2024:BR

[KF24] Xiongwen Ke and Yanan Fan. Bayesian L_{1/2} regression. Journal of Computational and Graphical Statistics, 34(1):199– 210, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2374579.

Kastner:2017:EBI

[KFSL17] Gregor Kastner, Sylvia Frühwirth-Schnatter, and Hedibert Freitas Lopes. Efficient Bayesian inference for multivariate factor stochastic volatility models. Journal of Computational and Graphical Statistics, 26(4):905–917, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kent:2018:NUA

[KGM18] John T. Kent, Asaad M. Ganeiber, and Kanti V. Mardia. A new unified approach for the simulation of a wide class of directional distributions. Journal of Computational and Graphical Statistics, 27(2):291–301, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kim:2006:ACA

[KGMT06] Hae-Young Kim, Matthew J.

Gribbin, Keith E. Muller, and Douglas J. Taylor. Analytic, computational, and approximate forms for ratios of noncentral and central Gaussian quadratic forms. Journal of Computational and Graphical Statistics, 15 (2):443–459, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See corrections [Ano06a].

Kandanaarachchi:2021:DRO

[KH21] Sevvandi Kandanaarachchi and Rob J. Hyndman. Dimension reduction for outlier detection using DOBIN. Journal of Computational and Graphical Statistics, 30(1):204– 219, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1807353.

Kandanaarachchi:2022:LOK

[KH22] Sevvandi Kandanaarachchi and Rob J. Hyndman. Leave-one-out kernel density estimates for outlier detection. Journal of Computational and Graphical Statistics, 31 (2):586–599, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000425.

Kidzinski:2023:MLD

[KH23] Lukasz Kidziński and Trevor Hastie. Modeling longitudinal data using matrix completion. Journal of Computational and Graphical Statistics, 33 (2):551–566, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257257.

Kocherginsky:2005:PCI

[KHM05] Masha Kocherginsky, Xuming He, and Yunming Mu. Practical confidence intervals for regression quantiles. Journal of Computational and Graphical Statistics, 14(1):41– 55, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kai:2023:NSQ

[KHYD23] Bo Kai, Mian Huang, Weixin Yao, and Yuexiao Dong. Nonparametric and semiparametric quantile regression via a new MM algorithm. Journal of Computational and Graphical Statistics, 32 (4):1613–1623, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2184374.

Kiers:2000:DOT

[Kie00] Henk A. L. Kiers. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 21–25, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474859.  See

[LHY00].

Kitagawa:1996:MCF

[Kit96] Genshiro Kitagawa. Monte Carlo filter and smoother for non-Gaussian nonlinear state space models. Journal of Computational and Graphical Statistics, 5(1):1–25, March 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474692.

Kozintsev:2000:GSS

[KK00] Boris Kozintsev and Benjamin Kedem. Generation of “similar” images from a given discrete image. Journal of Computational and Graphical Statistics, 9(2):286–302, 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474881.

Kauermann:2009:FTP

[KK09] Göran Kauermann and Pavel Khomski. Full time or part time reemployment: a competing risk model with frailties and smooth effects using a penalty based approach. Journal of Computational and Graphical Statistics, 18(1):106–125, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Krock:2021:NMS

[KKB21] Mitchell Krock, William Kleiber, and Stephen Becker. Nonstationary modeling with sparsity for spatial data via the basis graphical lasso. Journal of Computational and Graphical Statistics, 30 (2):375–389, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1811103.

Krock:2023:MMH

[KKHB23] Mitchell L. Krock, William Kleiber, Dorit Hammerling, and Stephen Becker. Modeling massive highly multivariate nonstationary spatial data with the basis graphical lasso. Journal of Computational and Graphical Statistics, 32(4):1472–1487, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2174126.

Kim:2008:GBO

[KKK08] Jinseog Kim, Yuwon Kim, and Yongdai Kim. A gradient-based optimization algorithm for LASSO. Journal of Computational and Graphical Statistics, 17(4):994– 1009, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kaplan:2020:SMR

[KKLN20] Andee Kaplan, Mark S.

Kaiser, Soumendra N. Lahiri, and Daniel J. Nordman. Simulating Markov random fields with a conclique-based Gibbs sampler. Journal of Computational and Graphical Statistics, 29 (2):286–296, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1668800.

Kuismin:2017:PME

[KKS17] M. O. Kuismin, J. T. Kemppainen, and M. J. Sillanpää. Precision matrix estimation with ROPE. Journal of Computational and Graphical Statistics, 26(3):682–694, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kim:2003:CTB

[KL03] Hyunjoong Kim and Wei-Yin Loh. Classification trees with bivariate linear discriminant node models. Journal of Computational and Graphical Statistics, 12 (3):512–530, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s2/p512;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032049.

Klemela:2004:VMD

[Kle04] Jussi Klemelä. Visualization of multivariate density estimates with level set trees. Journal of Computational and Graphical Statistics, 13(3):599–620, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Klemela:2006:VMD

[Kle06] Jussi Klemelä. Visualization of multivariate density estimates with shape trees. Journal of Computational and Graphical Statistics, 15(2):372–397, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Klemela:2008:MTM

[Kle08] Jussi Klemelä. Mode trees for multivariate data. Journal of Computational and Graphical Statistics, 17(4):860–869, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kleppe:2019:DRH

[Kle19] Tore Selland Kleppe. Dynamically rescaled Hamiltonian Monte Carlo for Bayesian hierarchical models. Journal of Computational and Graphical Statistics, 28(3):493–507, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kleppe:2022:CDN

[Kle22] Tore Selland Kleppe. Connecting the dots: Numerical randomized Hamiltonian Monte Carlo with state-dependent event rates. Journal of Computational and Graphical Statistics, 31(4):1238– 1253, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2066679.

Kim:1999:CEO

[KLH99] Woocheol Kim, Oliver B. Linton, and Niklaus W. Hengartner. A computationally efficient oracle estimator for additive nonparametric regression with bootstrap confidence intervals. Journal of Computational and Graphical Statistics, 8(2): 278–297, June 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts99/oracle.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/linton.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474814.

Komarek:2005:AFT

[KLH05] Arnošt Komárek, Emmanuel Lesaffre, and Joan F. Hilton. Accelerated failure time model for arbitrarily censored data with smoothed error distribution. Journal of Computational and Graphical Statistics, 14 (3):726–745, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kamary:2018:WIR

[KLR18] Kaniav Kamary, Jeong Eun Lee, and Christian P. Robert. Weakly informative reparameterizations for location-scale mixtures. Journal of Computational and Graphical Statistics, 27(4):836–848, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Karr:2005:SRD

[KLSR05] Alan F. Karr, Xiaodong Lin, Ashish P. Sanil, and Jerome P. Reiter. Secure regression on distributed databases. Journal of Computational and Graphical Statistics, 14 (2):263–279, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kwon:2023:EUS

[KLZ23] Oh-Ran Kwon, Zhaosong Lu, and Hui Zou. Exactly uncorrelated sparse principal component analysis. Journal of Computational and Graphical Statistics, 33(1):231–241, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2232843.

Kuhnert:2003:RML

[KM03] Petra M. Kuhnert and Kerrie Mengersen. Reliability measures for local nodes assessment in classification trees. Journal of Computational and Graphical Statistics, 12(2):398–416, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s8/p398;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031734.

Kim:2015:GMS

[KM15] Hang J. Kim and Steven N. MacEachern. The generalized multiset sampler. Journal of Computational and Graphical Statistics, 24(4):1134–1154, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.962701.

Klumb:2023:IES

[KMGJ23] Chad Klumb, Martina Morris, Steven M. Goodreau, and Samuel M. Jenness. Improving and extending STERGM approximations based on crosssectional data and tie durations. Journal of Computational and Graphical Statistics, 33(1):166– 180, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2233593.

Kahn:1992:CPM

[KMOS92] M. H. Kahn, M. S. Mackisack, M. R. Osborne, and G. K. Smyth. On the consistency of Prony’s method and related algorithms. Journal of Computational and Graphical Statistics, 1(4):329–349, December 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474589.

Kottas:2005:NBM

[KMQ05] Athanasios Kottas, Peter Müller, and Fernando Quintana. Nonparametric Bayesian modeling for multivariate ordinal data. Journal of Computational and Graphical Statistics, 14(3):610–625, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kashlak:2023:APT

[KMS23] Adam B. Kashlak, Sergii Myroshnychenko, and Susanna Spektor. Analytic permutation testing for functional data ANOVA. Journal of Computational and Graphical Statistics, 32(1):294–303, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2069780.

Kaufman:1997:RIP

[KN97] Linda Kaufman and Arnold Neumaier. Regularization of ill-posed problems by envelope guided conjugate gradients. Journal of Computational and Graphical Statistics, 6(4):451–463, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/kaufman.
html;   http://www.tandfonline.

com/doi/abs/10.1080/
10618600.1997.10474753.

Keich:2006:FNR

[KN06] Uri Keich and Niranjan Nagarajan. A fast and numerically robust method for exact multinomial goodness-of-fit test. Journal of Computational and Graphical Statistics, 15(4):779–802, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Knight:2023:AWD

[KNH23] Marina I. Knight, Matthew A.

Nunes, and Jessica K. Hargreaves. Adaptive wavelet domain principal component analysis for nonstationary time series. Journal of Computational and Graphical Statistics, 33(3):941–954, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2301069.

Kateri:2005:BIR

[KNN05] Maria Kateri, Anna Nicolaou, and Ioannis Ntzoufras. Bayesian inference for the RC(m) association model. Journal of Computational and Graphical Statistics, 14 (1):116–138, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Klein:2021:MCD

[KNS21] Nadja Klein, David J. Nott, and Michael Stanley Smith. Marginally calibrated deep distributional regression. Journal of Computational and Graphical Statistics, 30 (2):467–483, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1807996.

Kauermann:2004:GCV

[KO04] Göran Kauermann and J. D. Opsomer. Generalized cross-validation for bandwidth selection of backfitting estimates in generalized additive models. Journal of Computational and Graphical Statistics, 13(1):66– 89, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kolbjornsen:2005:BIP

[KO05] Odd Kolbjørnsen and Henning Omre. Bayesian inversion of piecewise affine operators in a Gaussian framework. Journal of Computational and Graphical Statistics, 14 (1):56–74, March 2005.

CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Koo:1997:SED

[Koo97]  Ja-Yong Koo.  Spline

estimation of discontinuous regression functions. Journal of Computational and Graphical Statistics, 6(3):266–284, September 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/koo.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474742.

Kooperberg:1998:BDE

[Koo98] Charles Kooperberg. Bivariate density estimation with an application to survival analysis. Journal of Computational and Graphical Statistics, 7(3):322–341, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/kooperberg.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474779.

Kosara:2013:IMM

[Kos13] Robert Kosara. InfoVis is so much more: a comment on Gelman and Unwin and an invitation to consider the opportunities. Journal of Computational and Graphical Statistics, 22(1):29–32, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Koskela:2022:ZZS

[Kos22]  Jere Koskela.  Zig-

zag sampling for discrete structures and nonreversible phylogenetic MCMC. Journal of Computational and Graphical Statistics, 31(3):684– 694, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2032722.

Kou:2014:ENC

[Kou14] Jiyao Kou. Estimating the number of clusters via the GUD statistic. Journal of Computational and Graphical Statistics, 23 (2):403–417, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Knuteson:2003:SCM

[KP03] Bruce Knuteson and Paul Padley. Statistical challenges with massive datasets in particle physics. Journal of Computational and Graphical Statistics, 12(4):808–828, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032599.

Kim:2002:FAC

[KPC02] Tae Yoon Kim, Jeong Soo Park, and Dennis D. Cox. Fast algorithm for crossvalidation of the best linear unbiased predictor. Journal of Computational and Graphical Statistics, 11(4):823–835, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s5/p823;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002826.

Kan:2017:MFT

[KR17] Raymond Kan and Cesare Robotti. On moments of folded and truncated multivariate normal distributions. Journal of Computational and Graphical Statistics, 26(4):930–934, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kooperberg:1992:LDE

[KS92] Charles Kooperberg and Charles J. Stone. Logspline density estimation for censored data. Journal of Computational and Graphical Statistics, 1(4): 301–328, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474588.

Kooperberg:1999:SOM

[KS99] Charles Kooperberg and Charles J. Stone. Stochastic optimization methods for fitting polyclass and feed-forward neural network models. Journal of Computational and Graphical Statistics, 8(2):169– 189, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/stochastic.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/koop.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474809.

Kooperberg:2004:CPB

[KS04] Charles Kooperberg and Charles J. Stone. Comparison of parametric and bootstrap approaches to obtaining confidence intervals for logspline density estimation. Journal of Computational and Graphical Statistics, 13 (1):106–122, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Koekemoer:2008:TKD

[KS08] Gerhard Koekemoer and Jan W. H. Swanepoel. Transformation kernel density estimation with applications. Journal of Computational and Graphical Statistics, 17 (3):750–769, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kovac:2011:NRG

[KS11] Arne Kovac and Andrew D. A. C. Smith. Nonparametric regression on a graph. Journal of Computational and Graphical Statistics, 20(2):432–447, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kang:2022:JML

[KS22] Kai Kang and Xin Yuan Song. Joint modeling of longitudinal imaging and survival data. Journal of Computational and Graphical Statistics, 32 (2):402–412, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2022.2102027.

Kirch:2024:MAE

[KSC^{+}24] Jhessica Leticia Kirch, Acácia Mecejana Diniz Souza Spitti, Alisson Fernando Chiorato, Carlos Tadeu dos Santos Dias, and César Gonçalves de Lima. A multiattribute evaluation of genotype-environment experiments using biplots and joint plots graphics. Journal of Computational and Graphical Statistics, 33(4):1488–1496, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2325445.

Kaufman:2023:FLO

[KSH23] Jacob M. Kaufman, Alyssa J. Stenberg, and Toby D. Hocking. Functional labeled optimal partitioning. Journal of Computational and Graphical Statistics, 33(4):1131– 1138, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2293216.

Kunzel:2022:LAT

[KSLS22] Sören R. Künzel, Theo F. Saarinen, Edward W. Liu, and Jasjeet S. Sekhon. Linear aggregation in tree-based estimators. Journal of Computational and Graphical Statistics, 31(3):917–934, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2026780.

Konomi:2014:ABN

[KSM14] Bledar A. Konomi, Huiyan Sang, and Bani K. Mallick. Adaptive Bayesian nonstationary modeling for large spatial datasets using covariance approximations. Journal of Computational and Graphical Statistics, 23(3):802–829, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kim:2018:AFF

[KSM^{+}18] Janet S. Kim, Ana-Maria Staicu, Arnab Maity, Raymond J. Carroll, and David Ruppert. Additive function-on-function regression. Journal of Computational and Graphical Statistics, 27(1):234–244, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kuipers:2022:ESS

[KSM22] Jack Kuipers, Polina Suter, and Giusi Moffa.

Efficient sampling and structure learning of Bayesian networks. Journal of Computational and Graphical Statistics, 31 (3):639–650, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2020127.

Karmakar:2019:UAA

[KSR19] Bikram Karmakar, Dylan S. Small, and Paul R. Rosenbaum. Using approximation algorithms to build evidence factors and related designs for observational studies. Journal of Computational and Graphical Statistics, 28 (3):698–709, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kawaguchi:2021:SAL

[KSSL21] Eric S. Kawaguchi, Jenny I. Shen, Marc A. Suchard, and Gang Li. Scalable algorithms for large competing risks data. Journal of Computational and Graphical Statistics, 30 (3):685–693, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1841650.

Karppinen:2023:CPF

[KSV23] Santeri Karppinen, Sumeetpal S. Singh, and Matti Vihola. Conditional particle filters with bridge backward sampling. Journal of Computational and Graphical Statistics, 33 (2):364–378, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2231514.

Kauermann:2003:SNM

[KT03] Göran Kauermann and Gerhard Tutz. Semiand nonparametric modeling of ordinal data. Journal of Computational and Graphical Statistics, 12(1):176–196, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s9/p176;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031310.

Kamm:2017:ECJ

[KTS17] John A. Kamm, Jonathan Terhorst, and Yun S. Song. Efficient computation of the joint sample frequency spectra for multiple populations. Journal of Computational and Graphical Statistics, 26 (1):182–194, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1159212.

Koo:2023:PAB

[KTT23] John Koo, Minh Tang, and Michael W. Trosset. Popularity adjusted block models are generalized random dot product graphs. Journal of Computational and Graphical Statistics, 32(1):131– 144, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2081576.

Kinson:2020:LPC

[KTZQ20] Christopher Kinson, Xiwei Tang, Zhen Zuo, and Annie Qu. Longitudinal principal component analysis with an application to marketing data. Journal of Computational and Graphical Statistics, 29 (2):335–350, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1677244.

Kustra:2006:RRR

[Kus06] Rafal Kustra. Reducedrank regularized multivariate model for highdimensional data. Journal of Computational and Graphical Statistics, 15 (2):312–338, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Karcher:2001:GNM

[KW01] Peter Karcher and Yuedong Wang. Generalized nonparametric mixed effects models. Journal of Computational and Graphical Statistics, 10(4):641–655, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243377.

Kume:2010:MLE

[KW10] Alfred Kume and Max Welling. Maximum likelihood estimation for the offset-normal shape distributions using EM. Journal of Computational and Graphical Statistics, 19 (3):702–723, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Kim:2021:GSV

[KW21] Myungjin Kim and Li Wang.

Generalized spatially varying coefficient models. Journal of Computational and Graphical Statistics, 30(1):1–10, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1754225.

Kwong:1998:SSQ

[Kwo98] Koon-Shing Kwong. On sample size and quick simultaneous confidence interval estimations for multinomial proportions. Journal of Computational and Graphical Statistics, 7(2):212–222, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/kwong.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474771.

Ke:2020:DDP

[KXY20] Zheng Tracy Ke, Lingzhou Xue, and Fan Yang. Diagonally dominant principal component analysis. Journal of Computational and Graphical Statistics, 29(3):592–607, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1713798.

Kuo:1995:BCS

[KY95] Lynn Kuo and Tae Yang.

Bayesian computation of software reliability. Journal of Computational and Graphical Statistics, 4(1): 65–82, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474666.

Ko:2019:EPD

[KYW19] Seyoon Ko, Donghyeon Yu, and Joong-Ho Won. Easily parallelizable and distributable class of algorithms for structured sparsity, with optimal acceleration. Journal of Computational and Graphical Statistics, 28 (4):821–833, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lopes:2012:VMS

[LA12] Danilo Lopes and Renato Assunção. Visualizing marked spatial and origindestination point patterns with dynamically linked windows. Journal of Computational and Graphical Statistics, 21 (1):134–154, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2022:RAH

[LAB22] Yazhe Li, Niall Adams, and Tony Bellotti. A relabeling approach to handling the class imbalance problem for logistic regression. Journal of Computational and Graphical Statistics, 31 (1):241–253, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1978470.

Liang:2006:TDV

[LACRM06] Li Liang, Christine M. Anderson-Cook, Timothy J. Robinson, and Raymond H. Myers. Threedimensional variance dispersion graphs for splitplot designs. Journal of Computational and Graphical Statistics, 15 (4):757–778, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Laa:2023:NSM

[LACV23] Ursula Laa, Alex Aumann, Dianne Cook, and German Valencia. New and simplified manual controls for projection and slice tours, with application to exploring classification boundaries in high dimensions. Journal of Computational and Graphical Statistics, 32 (3):1229–1236, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2206459.

Lee:2013:HDA

[LAJ13] Myung Hee Lee, Jeongyoun Ahn, and Yongho Jeon. HDLSS discrimination with adaptive data piling. Journal of Computational and Graphical Statistics, 22(2):433–451, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Laria:2019:ISG

[LAML19] Juan C. Laria, M. Carmen Aguilera-Morillo, and Rosa E. Lillo. An iterative sparse-group lasso. Journal of Computational and Graphical Statistics, 28(3):722–731, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lang:2000:OEN

[Lan00] Duncan Temple Lang.

The Omegahat environment: New possibilities for statistical computing. Journal of Computational and Graphical Statistics, 9(3):423–451, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Lang.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474891.

Lazar:2005:AEI

[Laz05] Nicole A. Lazar. Assessing the effect of individual data points on inference from empirical likelihood. Journal of Computational and Graphical Statistics, 14(3):626–642, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lang:2004:BPS

[LB04] Stefan Lang and Andreas Brezger. Bayesian Psplines. Journal of Computational and Graphical Statistics, 13(1):183–212, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lutz:2006:CDB

[LB06] Roman Werner Lutz and Peter Bühlmann. Conjugate direction boosting. Journal of Computational and Graphical Statistics, 15(2):287–311, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lee:2015:SRN

[LB15] Sangin Lee and Patrick Breheny. Strong rules for nonconvex penalties and their implications for efficient algorithms in highdimensional regression. Journal of Computational and Graphical Statistics, 24(4):1074–1091, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.975231.

Lou:2016:SPL

[LBCG16] Yin Lou, Jacob Bien, Rich Caruana, and Johannes Gehrke. Sparse partially linear additive models. Journal of Computational and Graphical Statistics, 25(4):1026– 1040, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2014:MIO

[LBM^{+}14] Fan Li, Michela Baccini, Fabrizia Mealli, Elizabeth R. Zell, Constantine E. Frangakis, and Donald B. Rubin. Multiple imputation by ordered monotone blocks with application to the anthrax vaccine research program. Journal of Computational and Graphical Statistics, 23(3):877–892, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Luts:2014:RTS

[LBW14] J. Luts, T. Broderick, and M. P. Wand. Realtime semiparametric regression. Journal of Computational and Graphical Statistics, 23(3):589–615, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2001:IMC

[LC01] Richard A. Levine and George Casella. Implementations of the Monte Carlo EM algorithm. Journal of Computational and Graphical Statistics, 10(3):422–439, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317115045.

Leroux:2024:DPU

[LC24] Andrew Leroux and Ciprian Crainiceanu. Dynamic prediction using landmark historical functional Cox regression. Journal of Computational and Graphical Statistics, 34 (1):59–71, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2024.2374576.

Li:2019:GHE

[LCB19] Yunfan Li, Bruce A.

Craig, and Anindya Bhadra. The graphical horseshoe estimator for inverse covariance matrices. Journal of Computational and Graphical Statistics, 28(3):747–757, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Laa:2022:HGS

[LCBV22] Ursula Laa, Dianne Cook, Andreas Buja, and German Valencia. Hole or grain? a section pursuit index for finding hidden structure in multiple dimensions. Journal of Computational and Graphical Statistics, 31 (3):739–752, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2035230.

Luo:2024:HPS

[LCD^{+}24] Hengrui Luo, Younghyun Cho, James W. Demmel, Xiaoye S. Li, and Yang Liu. Hybrid parameter search and dynamic model selection for mixed-variable Bayesian optimization. Journal of Computational and Graphical Statistics, 33 (3):855–868, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2308216.

Lee:2005:PPE

[LCKL05] Eun-Kyung Lee, Dianne Cook, Sigbert Klinke, and Thomas Lumley. Projection pursuit for exploratory supervised classification. Journal of Computational and Graphical Statistics, 14 (4):831–846, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Laa:2022:BSR

[LCL22] Ursula Laa, Dianne Cook, and Stuart Lee. Burning sage: Reversing the curse of dimensionality in the visualization of highdimensional data. Journal of Computational and Graphical Statistics, 31 (1):40–49, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1963264.

Li:2023:MCL

[LCLR23] Xinmin Li, Feifei Chen, Hua Liang, and David Ruppert. Model checking for logistic models when the number of parameters tends to infinity. Journal of Computational and Graphical Statistics, 32 (1):241–251, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2084403.

Lenzi:2021:IBL

[LCRG21] Amanda Lenzi, Stefano Castruccio, Håvard Rue, and Marc G. Genton. Improving Bayesian local spatial models in large datasets. Journal of Computational and Graphical Statistics, 30(2):349– 359, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1814789.

Liang:2014:MAC

[LCSL14] Waley W. J. Liang, Jacob B. Colvin, Bruno Sansó, and Herbert K. H. Lee. Modeling and anomalous cluster detection for point processes using process convolutions. Journal of Computational and Graphical Statistics, 23(1):129–150, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2024:PWT

[LCTV24] Weihao Li, Dianne Cook, Emi Tanaka, and Susan VanderPlas. A plot is worth a thousand tests: Assessing residual diagnostics with the lineup protocol. Journal of Computational and Graphical Statistics, 33(4):1497– 1511, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2344612.

Laa:2020:STF

[LCV20] Ursula Laa, Dianne Cook, and German Valencia. A slice tour for finding hollowness in highdimensional data. Journal of Computational and Graphical Statistics, 29 (3):681–687, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1777140.

Lavine:2022:BCD

[LCW22] Isaac Lavine, Andrew Cron, and Mike West. Bayesian computation in dynamic latent factor models. Journal of Computational and Graphical Statistics, 31(3):651– 665, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2021208.

Lin:2017:LSE

[LCWW17] Zhenhua Lin, Jiguo Cao, Liangliang Wang, and Haonan Wang. Locally sparse estimator for functional linear regression models. Journal of Computational and Graphical Statistics, 26(2):306–318, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2006:NAS

[LD06] Xuefeng Liu and Michael J.

Daniels. A new algorithm for simulating a correlation matrix based on parameter expansion and reparameterization. Journal of Computational and Graphical Statistics, 15(4):897–914, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lenkoski:2011:CAR

[LD11] Alex Lenkoski and Adrian Dobra. Computational aspects related to inference in Gaussian graphical models with the G-Wishart prior. Journal of Computational and Graphical Statistics, 20 (1):140–157, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2021:MFV

[LD21] Zeda Li and Yuexiao Dong. Model-free variable selection with matrixvalued predictors. Journal of Computational and Graphical Statistics, 30 (1):171–181, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1806854.

Linero:2022:GPB

[LD22] Antonio R. Linero and Junliang Du. Gibbs priors for Bayesian nonparametric variable selection with weak learners. Journal of Computational and Graphical Statistics, 32 (3):1046–1059, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2142594.

Lu:2023:UAV

[LDHW23] Youhan Lu, Yushen Dong, Juan Hu, and Yichao Wu. A unified approach to variable selection for partially linear models. Journal of Computational and Graphical Statistics, 33(1):250– 260, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2216254.

Li:2023:MBS

[LDR23] Dongjin Li, Somak Dutta, and Vivekananda Roy. Model based screening embedded Bayesian variable selection for ultrahigh dimensional settings. Journal of Computational and Graphical Statistics, 32(1):61–73, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2074428.

Lee:2002:DEN

[Lee02] Herbert K. H. Lee. Difficulties in estimating the normalizing constant of the posterior for a neural network. Journal of Computational and Graphical Statistics, 11(1):222– 235, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s12/p222;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375703.

Lee:2007:SMN

[Lee07] Jaeyong Lee. Sampling methods of neutral to the right processes. Journal of Computational and Graphical Statistics, 16 (3):656–671, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lee:2014:E

[Lee14] Thomas C. M. Lee. Editorial. Journal of Computational and Graphical Statistics, 23(1):1–2, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lenk:2003:BSD

[Len03] Peter J. Lenk. Bayesian semiparametric density estimation and model verification using a logistic-Gaussian process. Journal of Computational and Graphical Statistics, 12 (3):548–565, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

                                        URL http://giorgio.
                                        ingentaselect.com/cgi-
                                        bin/linker?ini=asa{\&
                                        }reqidx=/cw/asa/10618600/
                                        v12n3/s4/p548;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/1061860032021.

Lenk:2009:SPB

[Len09] Peter Lenk. Simulation pseudo-bias correction to the harmonic mean estimator of integrated likelihoods. Journal of Computational and Graphical Statistics, 18(4):941–960, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2001:DAD

[Lev01] Richard A. Levine. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):51–58, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418719.

See [vDM01a].

Levine:2011:EVR

[Lev11] Richard A. Levine. Editorial: On volume 20 and on review times. Journal of Computational and Graphical Statistics, 20 (1):1–2, March 2011. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2012:EVP

[Lev12] Richard A. Levine. Editorial: Volume 21 preview and manuscript processing statistics, in brief. Journal of Computational and Graphical Statistics, 21(1):1, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2013:E

[Lev13a] Rich Levine. Editorial. Journal of Computational and Graphical Statistics, 22(3):515–517, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2013:EVP

[Lev13b] Richard A. Levine. Editorial: Volume 22 preview and manuscript processing statistics, in brief. Journal of Computational and Graphical Statistics, 22(1):1, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2016:PAB

[LF16] Han Li and Xiaodan Fan. A pivotal allocation-based algorithm for solving the label-switching problem in Bayesian mixture models. Journal of Computational and Graphical Statistics, 25(1):266– 283, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2014.983643;   http:/
                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2014.983643.

Lipsitz:2017:OSG

[LFS^{+}17] Stuart Lipsitz, Garrett Fitzmaurice, Debajyoti Sinha, Nathanael Hevelone, Jim Hu, and Louis L. Nguyen. One-step generalized estimating equations with large cluster sizes. Journal of Computational and Graphical Statistics, 26(3):734–737, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lomeli:2017:MSS

[LFT17] María Lomelí, Stefano Favaro, and Yee Whye Teh. A marginal sampler for σ-stable Poisson– Kingman mixture models. Journal of Computational and Graphical Statistics, 26(1):44–53, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1110526.

Li:1997:SPP

[LFZ97] Run-Ze Li, Kai-Tai Fang, and Li-Xing Zhu. Some Q–Q probability plots to test spherical and elliptical symmetry. Journal of Computational and Graphical Statistics, 6(4):435–450, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/li.html;
 http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474752.

Lin:1992:CMS

[LG92] D. Y. Lin and C. J. Geyer. Computational methods for semiparametric linear regression with censored data. Journal of Computational and Graphical Statistics, 1(1): 77–90, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474577.

Leon-Gonzalez:2003:PDS

[LG03] Roberto Leon-Gonzalez. A panel data simultaneous equation model with a dependent categorical variable and selectivity. Journal of Computational and Graphical Statistics, 12(1):230–242, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s12/p230;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031293.

Lau:2007:BMB

[LG07] John W. Lau and Peter J.

Green. Bayesian modelbased clustering procedures. Journal of Computational and Graphical Statistics, 16(3):526–558, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2020:BTV

[LG20] Yi Liu and Feng Guo. A Bayesian time-varying coefficient model for multitype recurrent events. Journal of Computational and Graphical Statistics, 29(2):383–395, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1686988.

Liu:2016:ESM

[LGL16] Xiaoyu Liu, Serge Guillas, and Ming-Jun Lai. Efficient spatial modeling using the SPDE approach with bivariate splines. Journal of Computational and Graphical Statistics, 25(4):1176–1194, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lamnisos:2009:TSA

[LGS09] Demetris Lamnisos, Jim E.

Griffin, and Mark F. J.

Steel. Transdimensional sampling algorithms for Bayesian variable selection in classification problems with many more variables than observations. Journal of Computational and Graphical Statistics, 18(3):592–612, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lamnisos:2013:AMC

[LGS13] Demetris Lamnisos, Jim E. Griffin, and Mark F. J. Steel. Adaptive Monte Carlo for Bayesian variable selection in regression models. Journal of Computational and Graphical Statistics, 22 (3):729–748, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liao:1995:FAC

[LH95] J. G. Liao and Charles B. Hall. Fast and accurate computation of the exact confidence limits for the common odds ratio in several 2 × 2 tables. Journal of Computational and Graphical Statistics, 4(3):173–179, 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/abs/10.1080/10618600.
1995.10474675.

Lumley:2000:GEA

[LH00] Thomas Lumley and Patrick Heagerty. Graphical exploratory analysis of survival data. Journal of Computational and Graphical Statistics, 9(4):738–749, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Quintana.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474910.

Lystig:2002:ECO

[LH02] Theodore C. Lystig and James P. Hughes. Exact computation of the observed information matrix for hidden Markov models. Journal of Computational and Graphical Statistics, 11(3):678–689, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s10/p678;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002402.

Lee:2015:LSM

[LH15a] Jason D. Lee and Trevor J.

Hastie. Learning the structure of mixed graphical models. Journal of Computational and Graphical Statistics, 24 (1):230–253, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lim:2015:LIH

[LH15b] Michael Lim and Trevor Hastie. Learning interactions via hierarchical group-lasso regularization. Journal of Computational and Graphical Statistics, 24(3):627–654, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Loy:2015:YNP

[LH15c] Adam Loy and Heike Hofmann. Are you normal? The problem of confounded residual structures in hierarchical linear models. Journal of Computational and Graphical Statistics, 24(4):1191– 1209, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2014.960084.

Loy:2017:MCD

[LHC17] Adam Loy, Heike Hofmann, and Dianne Cook. Model choice and diagnostics for linear mixedeffects models using statistics on street corners. Journal of Computational and Graphical Statistics, 26(3):478–492, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lu:2017:BRF

[LHK17] Yi Lu, Radu Herbei, and Sebastian Kurtek. Bayesian registration of functions with a Gaussian process prior. Journal of Computational and Graphical Statistics, 26 (4):894–904, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lu:2003:GCV

[LHL03] Henry Horng-Shing Lu, Su-Yun Huang, and Fang-Jiun Lin. Generalized cross-validation for wavelet shrinkage in nonparametric mixed effects models. Journal of Computational and Graphical Statistics, 12 (3):714–730, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s12/p714;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032076.

Lai:2021:MGU

[LHL21] Randy C. S. Lai, Jan Hannig, and Thomas C. M. Lee. Method G: Uncertainty quantification for distributed data problems using generalized fiducial inference. Journal of Computational and Graphical Statistics, 30 (4):934–945, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1923514.

Lange:2000:OTU

[LHY00] Kenneth Lange, David R. Hunter, and Ilsoon Yang. Optimization transfer using surrogate objective functions. Journal of Computational and Graphical Statistics, 9(1): 1–20, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.amstat.org/publications/
                                        jcgs/abstracts00/Lange.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474858.

See discussion [Kie00, dLM00, Wu00, Men00, GH00, Gel00] and rejoinder [HL00b].

Li:2005:CBM

[Li05] Jia Li. Clustering based on a multilayer mixture model. Journal of Computational and Graphical Statistics, 14(3):547–568, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liao:1998:VRG

[Lia98] J. G. Liao. Variance reduction in Gibbs sampler using quasi random numbers. Journal of Computational and Graphical Statistics, 7(3):253–266, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/liao.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474775.

Liang:2007:CCM

[Lia07] Faming Liang. Continuous contour Monte Carlo for marginal density estimation with an application to a spatial statistical model. Journal of Computational and Graphical Statistics, 16(3):608–632, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lindstrom:1999:PEF

[Lin99] Mary J. Lindstrom. Penalized estimation of freeknot splines. Journal of Computational and Graphical Statistics, 8(2): 333–352, June 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts99/splines.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/lind.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474817.

Liu:2001:DAD

[Liu01] Chuanhai Liu. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):75–81, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418746.

See [vDM01a].

Liu:2019:MST

[LJ19] Tzu-Ying Liu and Hui Jiang. Minimizing sum of truncated convex functions and its applications. Journal of Computational and Graphical Statistics, 28(1):1–10, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lindsten:2017:DCS

[LJN^{+}17] F. Lindsten, A. M. Johansen, C. A. Naesseth, B. Kirkpatrick, T. B. Schön, J. A. D. Aston, and A. Bouchard-Côté. Divide-and-conquer with sequential Monte Carlo. Journal of Computational and Graphical Statistics, 26(2):445–458, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2005:SBS

[LJZD05] W. Liu, M. Jamshidian, Y. Zhang, and J. Donnelly. Simulation-based simultaneous confidence bands in multiple linear regression with predictor variables constrained in intervals. Journal of Computational and Graphical Statistics, 14 (2):459–484, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lazar:2002:MVM

[LK02] Nicole A. Lazar and Joseph B. Kadane. Movies for the visualization of MCMC output. Journal of Computational and Graphical Statistics, 11(4):863–874, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/

                                        v11n4/s8/p863;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002376.

Londschien:2021:CPD

[LKB21] Malte Londschien, Solt Kovács, and Peter Bühlmann. Change-point detection for graphical models in the presence of missing values. Journal of Computational and Graphical Statistics, 30(3):768– 779, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1853549.

Liu:2022:DMA

[LKR22] Xiran Liu, Naama M. Kopelman, and Noah A. Rosenberg. A Dirichlet model of alignment cost in mixed-membership unsupervised clustering. Journal of Computational and Graphical Statistics, 32 (3):1145–1159, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2127739.

Lim:2009:ESO

[LKW09] Johan Lim, Seung Jean Kim, and Xinlei Wang. Estimating stochastically ordered survival functions via geometric programming. Journal of Computational and Graphical Statistics, 18(4):978–994, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liechty:2010:MNS

[LL10] Merrill W. Liechty and Jingjing Lu. Multivariate normal slice sampling. Journal of Computational and Graphical Statistics, 19(2):281–294, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lok:2011:RDB

[LL11] W. S. Lok and Stephen M. S. Lee. Robustness diagnosis for bootstrap inference. Journal of Computational and Graphical Statistics, 20(2):448–460, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lee:2012:VSC

[LL12] Hyangmin Lee and Jia Li. Variable selection for clustering by separability based on ridgelines. Journal of Computational and Graphical Statistics, 21(2):315–337, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lindberg:2015:OUC

[LL15] David V. Lindberg and Herbert K. H. Lee. Optimization under constraints by applying an asymmetric entropy measure. Journal of Computational and Graphical Statistics, 24(2):379– 393, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.901225.

Landeros:2022:ASS

[LL22] Alfonso Landeros and Kenneth Lange. Algorithms for sparse support vector machines. Journal of Computational and Graphical Statistics, 32 (3):1097–1108, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2146697.

Liechty:2009:SP

[LLM09] Merrill W. Liechty, John C. Liechty, and Peter Müller. The shadow prior. Journal of Computational and Graphical Statistics, 18 (2):368–383, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lin:2023:SDP

[LLM23] Jiahe Lin, Huitian Lei, and George Michailidis. Structural discovery with partial ordering information for time-dependent data with convergence guarantees. Journal of Computational and Graphical Statistics, 33 (3):968–977, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2301097.

Liu:2024:BHM

[LLRM24] Bolun Liu, Shane Lubold, Adrian E. Raftery, and Tyler H. McCormick. Bayesian hyperbolic multidimensional scaling. Journal of Computational and Graphical Statistics, 33 (3):869–882, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2308219.

Li:2022:RLP

[LLW^{+}22] Changcheng Li, Runze Li, Jiawei Wen, Songshan Yang, and Xiang Zhan. Regularized linear programming discriminant rule with folded concave penalty for ultrahighdimensional data. Journal of Computational and Graphical Statistics, 32 (3):1074–1082, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2143785.

Liu:2009:FEA

[LLZ09] Lei Liu, Michael Levine, and Yu Zhu. A functional EM algorithm for mixing density estimation via nonparametric penalized likelihood maximization. Journal of Computational and Graphical Statistics, 18(2):481–504, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2023:CNC

[LLZ23] Jicai Liu, Jinhong Li, and Riquan Zhang. K -CDFs: A nonparametric clustering algorithm via cumulative distribution function. Journal of Computational and Graphical Statistics, 32(1):304–318, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2091575.

Li:2019:ECM

[LM19] Zehang Richard Li and Tyler H. McCormick. An expectation conditional maximization approach for Gaussian graphical models. Journal of Computational and Graphical Statistics, 28(4):767–777, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2023:MBT

[LM23] Junge Li and Qing Mai.

Model-based tensor lowrank clustering. Journal of Computational and Graphical Statistics, 33 (1):208–218, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2205913.

Liu:2022:MFR

[LMC22] Xiaokang Liu, Shujie Ma, and Kun Chen. Multivariate functional regression via nested reducedrank regularization. Journal of Computational and Graphical Statistics, 31 (1):231–240, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1960850.

Liu:2019:FCT

[LMM19] Xiaohui Liu, Karl Mosler, and Pavlo Mozharovskyi. Fast computation of Tukey trimmed regions and median in dimension p > 2. Journal of Computational and Graphical Statistics, 28(3):682–697, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Loaiza-Maya:2019:VBE

[LMS19] Rubén Loaiza-Maya and Michael Stanley Smith. Variational Bayes estimation of discrete-margined copula models with application to time series. Journal of Computational and Graphical Statistics, 28(3):523–539, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lam:2018:FAL

[LMST18] Xin Yee Lam, J. S. Marron, Defeng Sun, and Kim-Chuan Toh. Fast algorithms for largescale generalized distance weighted discrimination. Journal of Computational and Graphical Statistics, 27(2):368–379, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2022:LRT

[LMW^{+}22] Linxi Liu, Yang Meng, Xiaoru Wu, Zhiliang Ying, and Tian Zheng. Logrank-type tests for equality of distributions in high-dimensional spaces.

Journal of Computational and Graphical Statistics, 31(4):1384–1396, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2051530.

Liang:2009:MNP

[LN09] Jiajuan Liang and Kai Wang Ng. A multivariate normal plot to detect nonnormality. Journal of Computational and Graphical Statistics, 18(1):52– 72, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lim:2016:DAP

[LO16] Yaeji Lim and Hee-Seok Oh. A data-adaptive principal component analysis: Use of composite asymmetric Huber function. Journal of Computational and Graphical Statistics, 25(4):1230– 1247, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Locascio:2011:LE

[Loc11] Joseph J. Locascio. Letter to the Editor. Journal of Computational and Graphical Statistics, 20 (3):785–786, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lock:2018:TTR

[Loc18] Eric F. Lock. Tensor-on-tensor regression. Journal of Computational and Graphical Statistics, 27 (3):638–647, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lock:2022:BDW

[Loc22] Eric F. Lock. Bayesian distance weighted discrimination. Journal of Computational and Graphical Statistics, 31 (4):1177–1188, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2069778.

Loader:2007:IRG

[LP07] Catherine Loader and Ramani S. Pilla. Iteratively reweighted generalized least squares for estimation and testing with correlated data. Journal of Computational and Graphical Statistics, 16 (4):925–945, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lin:2014:TCS

[LP14] Bingqing Lin and Zhen Pang. Tilted correlation screening learning in highdimensional data analysis. Journal of Computational and Graphical Statistics, 23(2):478–496, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liang:2023:VSS

[LP23] Jingwei Liang and Clarice Poon. Variable screening for sparse online regression. Journal of Computational and Graphical Statistics, 32(1):275– 293, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2099872.

Liao:2016:AFE

[LPHK16] Lina Liao, Cheolwoo Park, Jan Hannig, and Kee-Hoon Kang. Autocovariance function estimation via penalized regression. Journal of Computational and Graphical Statistics, 25(4):1041– 1056, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lin:2013:FRE

[LPJ13] Bingqing Lin, Zhen Pang, and Jiming Jiang. Fixed and random effects selection by REML and pathwise coordinate optimization. Journal of Computational and Graphical Statistics, 22(2):341–355, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lipsitz:1998:OML

[LPM98] Stuart R. Lipsitz, Michael Parzen, and Geert Molenberghs. Obtaining the maximum likelihood estimates in incomplete R×C contingency tables using a Poisson generalized linear model. Journal of Computational and Graphical Statistics, 7(3):356–376, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/lipsitz.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474781.

Lyzinski:2017:FEJ

[LPPT17] Vince Lyzinski, Youngser Park, Carey E. Priebe, and Michael Trosset. Fast embedding for JOFC using the raw stress criterion. Journal of Computational and Graphical Statistics, 26(4):786–802, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lin:2022:OIL

[LPQ^{+}22] Cunjie Lin, Jingfu Peng, Yichen Qin, Yang Li, and Yuhong Yang. Optimal integrating learning for split questionnaire design type data. Journal of Computational and Graphical Statistics, 32 (3):1009–1023, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2118753.

Levy:1999:SPC

[LPS99] Gad Levy, Calton Pu, and Paul D. Sampson. Statistical, physical, and computational aspects of massive data analysis and assimilation in atmospheric applications. Journal of Computational and Graphical Statistics, 8(3): 559–574, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474833.

Luo:2019:IMM

[LQ19] Ruiyan Luo and Xin Qi. Interaction model and model selection for function-on-function regression. Journal of Computational and Graphical Statistics, 28(2):309–322, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Luo:2021:FRD

[LQ21] Ruiyan Luo and Xin Qi. Functional regression for densely observed data with novel regularization. Journal of Computational and Graphical Statistics, 30(1):220–235, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1807994.

Luo:2023:GNF

[LQ23] Ruiyan Luo and Xin Qi. General nonlinear function-on-function regression via functional universal approximation. Journal of Computational and Graphical Statistics, 33(2):578–587, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2252033.

Lu:2004:OPM

[LR04] Bo Lu and Paul R. Rosenbaum. Optimal pair matching with two control groups. Journal of Computational and Graphical Statistics, 13(2):422–434, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liang:2005:GPL

[LR05] Hua Liang and Haobo Ren. Generalized partially linear measurement error models. Journal of Computational and Graphical Statistics, 14 (1):237–250, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Luss:2014:GIR

[LR14] Ronny Luss and Saharon Rosset. Generalized isotonic regression. Journal of Computational and Graphical Statistics, 23 (1):192–210, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Luss:2015:C

[LR15] Ronny Luss and Saharon Rosset. Correction. Journal of Computational and Graphical Statistics, 24 (4):1230, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2015.1121728.

Le-Rademacher:2012:SCP

[LRB12] Jennifer Le-Rademacher and Lynne Billard. Symbolic covariance principal component analysis and visualization for intervalvalued data. Journal of Computational and Graphical Statistics, 21 (2):413–432, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2003:ICS

[LRBS03] Yi Li, Louise Ryan, Scarlett Bellamy, and Glen A. Satten. Inference on clustered survival data using imputed frailties. Journal of Computational and Graphical Statistics, 12(3):640–662, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s8/p640;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032247.

Li:2021:ABS

[LRFK21] Zeda Li, Ori Rosen, Fabio Ferrarelli, and Robert T. Krafty. Adaptive Bayesian spectral analysis of high-dimensional nonstationary time series. Journal of Computational and Graphical Statistics, 30(3):794–807, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1868305.

Latouche:2018:GFL

[LRO18] Pierre Latouche, Stéphane Robin, and Sarah Ouadah. Goodness of fit of logistic regression models for random graphs. Journal of Computational and Graphical Statistics, 27 (1):98–109, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lim:2023:DLG

[LROI23] David K. Lim, Naim U.

Rashid, Junier B. Oliva, and Joseph G. Ibrahim. Deeply learned generalized linear models with missing data. Journal of Computational and Graphical Statistics, 33 (2):638–650, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2276122.

Liebl:2020:IEF

[LRR20] Dominik Liebl, Stefan Rameseder, and Christoph Rust. Improving estimation in functional linear regression with points of impact: Insights into Google AdWords. Journal of Computational and Graphical Statistics, 29 (4):814–826, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1754224.

Lange:1993:NID

[LS93] Kenneth Lange and Janet S. Sinsheimer. Normal/ independent distributions and their applications in robust regression. Journal of Computational and Graphical Statistics, 2(2): 175–198, June 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474606.

Loader:1997:RTF

[LS97] Clive Loader and Jiayang Sun. Robustness of tube formula based confidence bands. Journal of Computational and Graphical Statistics, 6(2):242– 250, June 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts/june/
                                        loader.html;   http:/
                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474739.

Li:2004:DSP

[LS04] Lei M. Li and Terence P. Speed. Deconvolution of sparse positive spikes. Journal of Computational and Graphical Statistics, 13(4):853–870, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lockwood:2005:MSC

[LS05] J. R. Lockwood and Mark J. Schervish. MCMC strategies for computing Bayesian predictive densities for censored multivariate data. Journal of Computational and Graphical Statistics, 14 (2):395–414, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2005:MLS

[LSD05] Yufeng Liu, Xiaotong Shen, and Hani Doss. Multicategory ψ-learning and support vector machine. Journal of Computational and Graphical Statistics, 14(1):219–236, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lu:2018:EIT

[LSFI18] Min Lu, Saad Sadiq, Daniel J. Feaster, and Hemant Ishwaran. Estimating individual treatment effect in observational data using random forest methods. Journal of Computational and Graphical Statistics, 27 (1):209–219, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2016:SSF

[LSH16] Gen Li, Haipeng Shen, and Jianhua Z. Huang. Supervised sparse and functional principal component analysis. Journal of Computational and Graphical Statistics, 25 (3):859–878, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1064434.

Liu:2022:PTF

[LSLX22] Ye Liu, Rui Song, Wenbin Lu, and Yanghua Xiao. A probit tensor factorization model for relational learning. Journal of Computational and Graphical Statistics, 31 (3):846–855, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2003204.

Lan:2015:MCM

[LSSG15] Shiwei Lan, Vasileios Stathopoulos, Babak Shahbaba, and Mark Girolami. Markov chain Monte Carlo from Lagrangian dynamics. Journal of Computational and Graphical Statistics, 24 (2):357–378, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.902764.

Lefebvre:2009:PSC

[LSV^{+}09] Geneviève Lefebvre, Russell Steele, Alain C. Vandal, Sridar Narayanan, and Douglas L. Arnold. Path sampling to compute integrated likelihoods: An adaptive approach. Journal of Computational and Graphical Statistics, 18(2):415–437, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lee:1994:BCI

[LT94] J. Jack Lee and Z. Nora Tu. A better confidence interval for kappa (κ) on measuring agreement between two raters with binary outcomes. Journal of Computational and Graphical Statistics, 3(3): 301–321, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474647.

LeBlanc:1998:MST

[LT98] Michael LeBlanc and Robert Tibshirani. Monotone shrinkage of trees. Journal of Computational and Graphical Statistics, 7(4):417–433, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/leblanc.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474786.

Liu:2007:SSE

[LTW07] Anna Liu, Tiejun Tong, and Yuedong Wang. Smoothing spline estimation of variance functions. Journal of Computational and Graphical Statistics, 16(2):312–329, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Levine:2010:EPA

[LTW^{+}10] Richard A. Levine, Luke Tierney, Hadley Wickham, Eric Sampson, Dianne Cook, and David A. van Dyk. Editorial: Publishing animations, 3D visualizations, and movies in JCGS. Journal of Computational and Graphical Statistics, 19(1):1–2, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lumley:2019:FGL

[Lum19] Thomas Lumley. Fast generalized linear models by database sampling and one-step polishing. Journal of Computational and Graphical Statistics, 28 (4):1007–1010, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lund:2017:PEL

[LVH17] Adam Lund, Martin Vincent, and Niels Richard Hansen. Penalized estimation in large-scale generalized linear array models. Journal of Computational and Graphical Statistics, 26(3):709–724, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Liu:2007:VSC

[LW07] Yufeng Liu and Yichao Wu. Variable selection via a combination of the L_{0} and L_{1} penalties. Journal of Computational and Graphical Statistics, 16 (4):782–798, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lange:2008:MAM

[LW08] Kenneth Lange and Tong Tong Wu. An MM algorithm for multicategory vertex discriminant analysis.

Journal of Computational and Graphical Statistics, 17(3):527–544, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Leng:2009:GAS

[LW09] Chenlei Leng and Hansheng Wang. On general adaptive sparse principal component analysis. Journal of Computational and Graphical Statistics, 18(1):201–215, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Langrene:2019:FSM

[LW19] Nicolas Langrené and Xavier Warin. Fast and stable multivariate kernel density estimation by fast sum updating. Journal of Computational and Graphical Statistics, 28 (3):596–608, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lu:2023:GAC

[LWH23] Rung-Sheng Lu, Shao-Hsuan Wang, and Su-Yun Huang. A geometric algorithm for contrastive principal component analysis in high dimension. Journal of Computational and Graphical Statistics, 33 (3):909–916, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2289542.

Lee:2017:LSS

[LWLY17] Taehoon Lee, Joong-Ho Won, Johan Lim, and Sungroh Yoon. Largescale structured sparsity via parallel fused lasso on multiple GPUs. Journal of Computational and Graphical Statistics, 26 (4):851–864, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2022:LPE

[LWLZ22] Tianxi Li, Yun-Jhong Wu, Elizaveta Levina, and Ji Zhu. Link prediction for egocentrically sampled networks. Journal of Computational and Graphical Statistics, 32 (4):1296–1319, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2163648.

Lin:2021:MNF

[LWM21] Chen Lin, Kevin Wang, and Samuel Mueller. MCVIS: A new framework for collinearity discovery, diagnostic, and visualization. Journal of Computational and Graphical Statistics, 30 (1):125–132, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1779729.

Liu:2014:SCM

[LWZ14] Han Liu, Lie Wang, and Tuo Zhao. Sparse covariance matrix estimation with eigenvalue constraints. Journal of Computational and Graphical Statistics, 23(2):439–459, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2016:SMS

[LX16] Na Li and Xingzhong Xu. Spline multiscale smoothing to control FDR for exploring features of regression curves. Journal of Computational and Graphical Statistics, 25 (2):325–343, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lyu:2022:LSM

[LXZ22] Zhongyuan Lyu, Dong Xia, and Yuan Zhang. Latent space model for higher-order networks and generalized tensor decomposition. Journal of Computational and Graphical Statistics, 32 (4):1320–1336, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2164289.

Liu:2011:RMS

[LY11] Yufeng Liu and Ming Yuan. Reinforced multicategory support vector machines. Journal of Computational and Graphical Statistics, 20 (4):901–919, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lim:2016:ESC

[LY16] Chinghway Lim and Bin Yu. Estimation stability with cross-validation (ESCV). Journal of Computational and Graphical Statistics, 25(2):464–492, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lee:2010:UGC

[LYG^{+}10] Anthony Lee, Christopher Yau, Michael B. Giles, Arnaud Doucet, and Christopher C. Holmes. On the utility of graphics cards to perform massively parallel simulation of advanced Monte Carlo methods. Journal of Computational and Graphical Statistics, 19 (4):769–789, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Lei:2015:FDM

[LYHM15] Edwin Lei, Fang Yao, Nancy Heckman, and Karin Meyer. Functional data model for genetically related individuals with application to cow growth. Journal of Computational and Graphical Statistics, 24(3):756–770, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2023:SWC

[LYLL23] Jialu Li, Guan Yu, Qizhai Li, and Yufeng Liu. Sample-wise combined missing effect model with penalization. Journal of Computational and Graphical Statistics, 32 (1):263–274, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2070172.

Li:2023:SMF

[LYM23] Tao Li, Jun Yu, and Cheng Meng. Scalable model-free feature screening via sliced-Wasserstein dependency. Journal of Computational and Graphical Statistics, 32 (4):1501–1511, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2183213.

Li:2023:EAG

[LYXM23] Mengyu Li, Jun Yu, Hongteng Xu, and Cheng Meng. Efficient approximation of Gromov-Wasserstein distance using importance sparsification. Journal of Computational and Graphical Statistics, 32(4):1512– 1523, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2165500.

Li:2008:NQR

[LZ08] Youjuan Li and Ji Zhu.

L_{1}-norm quantile regression. Journal of Computational and Graphical Statistics, 17(1):163–185, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2018:GSE

[LZ18] Jia Li and Fuqing Zhang.

Geometry-sensitive ensemble mean based on Wasserstein barycenters: Proof-of-concept on cloud simulations. Journal of Computational and Graphical Statistics, 27 (4):785–797, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Li:2023:BMT

[LZ23] Guanxun Li and Quan Zhou. Bayesian multitask variable selection with an application to differential DAG analysis. Journal of Computational and Graphical Statistics, 33(1):35–46, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2252023.

Loyal:2022:DRF

[LZCZ22] Joshua Daniel Loyal, Ruoqing Zhu, Yifan Cui, and Xin Zhang. Dimension reduction forests: Local variable importance using structured random forests. Journal of Computational and Graphical Statistics, 31(4):1104– 1113, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2069777.

Li:2024:NAM

[LZM24] Mengyu Li, Jingyi Zhang, and Cheng Meng. Nonparametric additive models for billion observations. Journal of Computational and Graphical Statistics, 33(4):1397– 1412, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2319684.

Li:2022:DIC

[LZYQ22] Yutong Li, Ruoqing Zhu, Mike Yeh, and Annie Qu. Dermoscopic image classification with neural style transfer. Journal of Computational and Graphical Statistics, 31 (4):1318–1331, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2061496.

Mallik:2019:SMT

[MA19] Abhirup Mallik and Zack W.

Almquist. Stable multiple time step simulation/prediction from lagged dynamic network regression models. Journal of Computational and Graphical Statistics, 28 (4):967–979, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Maathuis:2005:RAN

[Maa05] Marloes H. Maathuis. Reduction algorithm for the NPMLE for the distribution function of bivariate interval-censored data. Journal of Computational and Graphical Statistics, 14(2):352–362, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mahmoudian:2017:SHT

[Mah17] Behzad Mahmoudian. A skewed and heavy-tailed latent random field model for spatial extremes. Journal of Computational and Graphical Statistics, 26(3):658–670, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Maitra:1997:EPF

[Mai97] Ranjan Maitra. Estimating precision in functional images. Journal of Computational and Graphical Statistics, 6(1):132–142, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts/march/
                                        maitra.html;   http:/
                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474732.

Marron:1998:ERA

[MAJ^{+}98] J. S. Marron, S. Adak, I. M. Johnstone, M. H. Neumann, and P. Patil. Exact risk analysis of wavelet regression. Journal of Computational and Graphical Statistics, 7(3):278–309, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/marron.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474777.

Marron:1994:VUH

[Mar94] J. S. Marron. Visual understanding of higherorder kernels. Journal of Computational and Graphical Statistics, 3(4): 447–458, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474657.

Marais:1996:ESS

[Mar96] Johannes L. Marais. Extensible software systems in Oberon. Journal of Computational and Graphical Statistics, 5(3): 284–298, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1996.10474712.

Marschner:2010:SCM

[Mar10] Ian C. Marschner. Stable computation of maximum likelihood estimates in identity link Poisson regression. Journal of Computational and Graphical Statistics, 19(3):666–683, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Machler:2004:VLM

[MB04] Martin Mächler and Peter Bühlmann. Variable length Markov chains. Journal of Computational and Graphical Statistics, 13(2):435–455, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

McIntyre:2018:LBS

[MB18] Julie McIntyre and Ronald P. Barry. A lattice-based smoother for regions with irregular boundaries and holes. Journal of Computational and Graphical Statistics, 27(2):360–367, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mohler:2011:FTR

[MBGO11] George O. Mohler, Andrea L. Bertozzi, Thomas A.

Goldstein, and Stanley J. Osher. Fast TV regularization for 2D maximum penalized likelihood estimation. Journal of Computational and Graphical Statistics, 20(2):479–491, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mestres:2018:SRP

[MBM18] Adria Caballe Mestres, Natalia Bochkina, and Claus Mayer. Selection of the regularization parameter in graphical models using network characteristics. Journal of Computational and Graphical Statistics, 27(2):323–333, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mira:2004:UDS

[MBS04] José Mira, Ricardo Bolado, and Pablo Solana. The use of dimensional and similarity analyses in the propagation of uncertainties. Journal of Computational and Graphical Statistics, 13(2):467–486, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mejia:2022:TIC

[MBY^{+}22] Amanda F. Mejia, David Bolin, Yu Ryan Yue, Jiongran Wang, Brian S. Caffo, and Mary Beth Nebel. Template independent component analysis with spatial priors for accurate subject-level brain network estimation and inference. Journal of Computational and Graphical Statistics, 32 (2):413–433, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2104289.

Muller:2005:AOP

[MC05] Gernot Müller and Claudia Czado. An autoregressive ordered probit model with application to highfrequency financial data. Journal of Computational and Graphical Statistics, 14(2):320–338, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Majumder:2017:FNE

[MC17] Mahbubul Majumder and Xiaoyue Cheng. Focusing on the needs: Experiences of developing a data science program. Journal of Computational and Graphical Statistics, 26 (4):779–780, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Muller:2018:RSG

[MC18] Dominik Müller and Claudia Czado. Representing sparse Gaussian DAGs as sparse R-vines allowing for non-Gaussian dependence. Journal of Computational and Graphical Statistics, 27 (2):334–344, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Marchello:2024:DDL

[MCB24] Giulia Marchello, Marco Corneli, and Charles Bouveyron. A deep dynamic latent block model for coclustering of zero-inflated data matrices. Journal of Computational and Graphical Statistics, 33 (4):1224–1239, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2319162.

McCaffrey:1999:CSL

[McC99] Daniel F. McCaffrey.

Comment: Scatterplots for Logistic Regression. Journal of Computational and Graphical Statistics, 8(3):426–427, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474823.  See

[ET99b].

McElroy:2022:FDC

[McE22] Tucker S. McElroy. Frequency domain calculation of seasonal VARMA autocovariances. Journal of Computational and Graphical Statistics, 31 (1):301–303, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1981343.

Marrs:2020:IIN

[MCF^{+}20] Frank W. Marrs, Benjamin W. Campbell, Bailey K. Fosdick, Skyler J. Cranmer, and Tobias Böhmelt. Inferring influence networks from longitudinal bipartite relational data. Journal of Computational and Graphical Statistics, 29 (3):419–431, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1694523.

Mcintosh:1999:ATN

[Mci99] Allen A. Mcintosh. Analyzing telephone network data. Journal of Computational and Graphical Statistics, 8(3):611–619, September 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474836.

McLeod:1999:ISL

[McL99] A. Ian McLeod. Improved spread-location visualization. Journal of Computational and Graphical Statistics, 8(1):135– 141, March 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474806.

Michiels:2013:NGT

[MD13] Frederik Michiels and Ann De Schepper. A new graphical tool for copula selection. Journal of Computational and Graphical Statistics, 22 (2):471–493, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mishra:2017:SCS

[MDC17] Aditya Mishra, Dipak K.

Dey, and Kun Chen. Sequential co-sparse factor regression. Journal of Computational and Graphical Statistics, 26 (4):814–825, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Michalak:2012:DSR

[MDD^{+}12] Sarah Michalak, Andrew DuBois, David DuBois, Scott Vander Wiel, and John Hogden. Developing systems for real-time streaming analysis. Journal of Computational and Graphical Statistics, 21 (3):561–580, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Michel:2020:FEC

[MDS20] Manon Michel, Alain Durmus, and Stéphane Sénécal. Forward eventchain Monte Carlo: Fast sampling by randomness control in irreversible Markov chains. Journal of Computational and Graphical Statistics, 29 (4):689–702, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1750417.

Menictas:2023:SVI

[MDW23] Marianne Menictas, Gioia Di Credico, and Matt P. Wand. Streamlined variational inference for linear mixed models with crossed random effects. Journal of Computational and Graphical Statistics, 32(1):99–115, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2096622.

Martin:2023:BIS

[MÉG^{+}23] Alice Martin, Marie-Pierre Étienne, Pierre Gloaguen, Sylvain Le Corff, and Jimmy Olsson. Backward importance sampling for online estimation of state space models. Journal of Computational and Graphical Statistics, 32(4):1447– 1460, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2174125.

Meinshausen:2011:PM

[Mei11] Nicolai Meinshausen. Partition maps. Journal of Computational and Graphical Statistics, 20 (4):1007–1028, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Melnykov:2016:MMC

[Mel16] Volodymyr Melnykov.

Merging mixture components for clustering through pairwise overlap. Journal of Computational and Graphical Statistics, 25(1):66–90, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.978007;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.978007.

Meng:2000:DOT

[Men00] Xiao-Li Meng. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 35–43, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474862.  See

[LHY00].

Meyer:2003:EAA

[Mey03] Mary C. Meyer. An evolutionary algorithm with applications to statistics. Journal of Computational and Graphical Statistics, 12(2):265–281, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s2/p265;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031699.

Meyer:2018:CSG

[Mey18] Daniel W. Meyer. (un)conditional sample generation based on distribution element trees. Journal of Computational and Graphical Statistics, 27(4):940–946, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Maire:2019:AIM

[MFMR19] Florian Maire, Nial Friel, Antonietta Mira, and Adrian E. Raftery. Adaptive incremental mixture Markov Chain Monte Carlo. Journal of Computational and Graphical Statistics, 28(4):790–805, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mira:2001:DAD

[MG01] Antonietta Mira and Peter J. Green. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):94–97, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418773.

See [vDM01a].

Muyskens:2022:PBN

[MGF22] Amanda Muyskens, Joseph Guinness, and Montserrat Fuentes. Partitionbased nonstationary covariance estimation using the stochastic score approximation. Journal of Computational and Graphical Statistics, 31 (4):1025–1036, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2044830.

Moon:2018:PTT

[MGL18] Chul Moon, Noah Giansiracusa, and Nicole A. Lazar. Persistence terrace for topological inference of point cloud data. Journal of Computational and Graphical Statistics, 27(3):576–586, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mentch:2017:FHT

[MH17] Lucas Mentch and Giles Hooker. Formal hypothesis tests for additive structure in random forests. Journal of Computational and Graphical Statistics, 26(3):589–597, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

McNamara:2017:GDS

[MHB17] Amelia McNamara, Nicholas J. Horton, and Benjamin S.

Baumer. Greater data science at baccalaureate institutions. Journal of Computational and Graphical Statistics, 26 (4):781–783, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Moral:2020:BRP

[MHD20] Rafael A. Moral, John Hinde, and Clarice G. B. Demétrio. Bivariate residual plots with simulation polygons. Journal of Computational and Graphical Statistics, 29 (1):203–214, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1647849.

Masoudi:2019:MAC

[MHDW19] Ehsan Masoudi, Heinz Holling, Belmiro P. M. Duarte, and Weng Kee Wong. A metaheuristic adaptive cubature based algorithm to find Bayesian optimal designs for nonlinear models. Journal of Computational and Graphical Statistics, 28(4):861–876, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Martinez-Hernandez:2024:FTS

[MHG24] Israel Martínez-Hernández and Marc G. Genton. Functional time series analysis and visualization based on records. Journal of Computational and Graphical Statistics, 34 (1):72–83, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2374578.

Maadooliat:2015:IDT

[MHH15] Mehdi Maadooliat, Jianhua Z. Huang, and Jianhua Hu. Integrating data transformation in principal components analysis. Journal of Computational and Graphical Statistics, 24(1):84–103, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Moustafa:2011:MCD

[MHS11] Rida E. Moustafa, Ali S. Hadi, and Jürgen Symanzik. Multi-class data exploration using space transformed visualization plots. Journal of Computational and Graphical Statistics, 20(2):298–315, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mclean:2014:FGA

[MHS^{+}14] Mathew W. Mclean, Giles Hooker, Ana-Maria Staicu, Fabian Scheipl, and David Ruppert. Functional generalized additive models. Journal of Computational and Graphical Statistics, 23 (1):249–269, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Murrell:2000:APM

[MI00] Paul Murrell and Ross Ihaka. An approach to providing mathematical annotation in plots. Journal of Computational and Graphical Statistics, 9(3):582–599, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Murrell.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474900.

Michailidis:2012:SCB

[Mic12] George Michailidis. Statistical challenges in biological networks. Journal of Computational and Graphical Statistics, 21 (4):840–855, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Miller:2019:FAA

[Mil19] Jeffrey W. Miller. Fast and accurate approximation of the full conditional for gamma shape parameters. Journal of Computational and Graphical Statistics, 28(2):476–480, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mak:2018:MMP

[MJ18] Simon Mak and V. Roshan Joseph. Minimax and minimax projection designs using clustering. Journal of Computational and Graphical Statistics, 27(1):166–178, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Moon:2021:LMQ

[MJLK21] Sang Jun Moon, Jong-June Jeon, Jason Sang Hun Lee, and Yongdai Kim. Learning multiple quantiles with neural networks. Journal of Computational and Graphical Statistics, 30(4):1238– 1248, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1909601.

Magnusson:2018:SPC

[MJVB18] Måns Magnusson, Leif Jonsson, Mattias Villani, and David Broman. Sparse partially collapsed MCMC for parallel inference in topic models.

Journal of Computational and Graphical Statistics, 27(2):449–463, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Meyer:1992:RAC

[MK92] Michael M. Meyer and Joseph B. Kadane. Reconstructing the adjusted census for Florida: a case study in data examination. Journal of Computational and Graphical Statistics, 1(4):287– 300, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474587.

Ma:2020:FGP

[MK20] Pulong Ma and Emily L.

Kang. A fused Gaussian process model for very large spatial data. Journal of Computational and Graphical Statistics, 29 (3):479–489, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1704293.

Marchant:2021:DBD

[MKE^{+}21] Neil G. Marchant, Andee Kaplan, Daniel N. Elazar, Benjamin I. P. Rubinstein, and Rebecca C.

Steorts. d-blink: Distributed end-to-end Bayesian entity resolution. Journal of Computational and Graphical Statistics, 30 (2):406–421, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1825451.

Michaelis:2018:BMD

[MKK18] Patrick Michaelis, Nadja Klein, and Thomas Kneib. Bayesian multivariate distributional regression with skewed responses and skewed random effects. Journal of Computational and Graphical Statistics, 27(3):602–611, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mosaferi:2023:PTS

[MKN23] Sepideh Mosaferi, Mark S. Kaiser, and Daniel J. Nordman. Properties of test statistics for nonparametric cointegrating regression functions based on subsamples. Journal of Computational and Graphical Statistics, 33 (3):774–786, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2301079.

Monni:2008:VCR

[ML08] Stefano Monni and Hongzhe Li. Vertex clustering in random graphs via reversible jump Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 17 (2):388–409, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Murray:2016:PRP

[MLJ16] Lawrence M. Murray, Anthony Lee, and Pierre E. Jacob. Parallel resampling in the particle filter. Journal of Computational and Graphical Statistics, 25(3):789–805, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1062015.

Mateu:2007:DFS

[MLP07] J. Mateu, G. Lorenzo, and E. Porcu. Detecting features in spatial point processes with clutter via local indicators of spatial association. Journal of Computational and Graphical Statistics, 16(4):968–990, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Muruzabal:1997:VOS

[MM97] Jorge Muruzábal and Alberto Muñoz. On the visualization of outliers via self-organizing maps. Journal of Computational and Graphical Statistics, 6(4):355–382, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/muruzabal.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474748.

Maceachern:1998:EMD

[MM98] Steven N. Maceachern and Peter Müller. Estimating mixture of Dirichlet process models. Journal of Computational and Graphical Statistics, 7(2): 223–238, June 1998. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts98/maceachern.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474772.

Mavridis:2009:FSA

[MM09] Dimitris Mavridis and Irini Moustaki. The forward search algorithm for detecting aberrant response patterns in factor analysis for binary data. Journal of Computational and Graphical Statistics, 18(4):1016–1034, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Maitra:2010:SDS

[MM10] Ranjan Maitra and Volodymyr Melnykov. Simulating data to study performance of finite mixture modeling and clustering algorithms. Journal of Computational and Graphical Statistics, 19 (2):354–376, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mankad:2014:BTD

[MM14] Shawn Mankad and George Michailidis. Biclustering three-dimensional data arrays with plaid models. Journal of Computational and Graphical Statistics, 23(4):943–965, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Moradi:2020:FSO

[MM20] M. Mehdi Moradi and Jorge Mateu. First- and second-order characteristics of spatio-temporal point processes on linear networks. Journal of Computational and Graphical Statistics, 29 (3):432–443, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1694524.

Madigan:1997:GEB

[MMA97] David Madigan, Krzysztof Mosurski, and Russell G. Almond. Graphical explanation in belief networks. Journal of Computational and Graphical Statistics, 6(2):160– 181, June 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/june/
madigan.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474735.

Morales:2022:MFD

[MMDL22] Christian E. Galarza Morales, Larissa A. Matos, Dipak K. Dey, and Victor H. Lachos. On moments of folded and doubly truncated multivariate extended skew-normal distributions. Journal of Computational and Graphical Statistics, 31 (2):455–465, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000869.

Martin:2019:ALB

[MMF^{+}19] Gael M. Martin, Brendan P. M. McCabe, David T. Frazier, Worapree Maneesoonthorn, and Christian P. Robert. Auxiliary likelihood-based approximate Bayesian computation in state space models. Journal of Computational and Graphical Statistics, 28(3):508–522, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

McKenna:2016:CIS

[MMGG16] Sean McKenna, Miriah Meyer, Christopher Gregg, and Samuel Gerber. s-CorrPlot: An interactive scatterplot for exploring correlation. Journal of Computational and Graphical Statistics, 25 (2):445–463, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Martinez-Minaya:2022:INL

[MMLLQ^{+}22] Joaquín Martínez-Minaya, Finn Lindgren, Antonio López-Quílez, Daniel Simpson, and David Conesa. The integrated nested Laplace approximation for fitting Dirichlet regression models. Journal of Computational and Graphical Statistics, 32(3):805–823, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2144330.

McVinish:2013:RIS

[MMN^{+}13] Ross McVinish, Kerrie Mengersen, Darfiana Nur, Judith Rousseau, and Chantal Guihenneuc-Jouyaux. Recentered importance sampling with applications to Bayesian model validation. Journal of Computational and Graphical Statistics, 22 (1):215–228, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Marshall:2001:CDR

[MMO01] A. W. Marshall, J. C. Meza, and I. Olkin. Can data recognize its parent distribution? Journal of Computational and Graphical Statistics, 10(3):555–580, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317115117.

Martinez-Miranda:2008:BLB

[MMRMGMGC08] M. D. Martínez-Miranda, R. Raya-Miranda, W. González-Manteiga, and A. González-Carmona. A bootstrap local bandwidth selector for additive models. Journal of Computational and Graphical Statistics, 17 (1):38–55, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Miasojedow:2013:APT

[MMV13] Blazej Miasojedow, Eric Moulines, and Matti Vihola. An adaptive parallel tempering algorithm. Journal of Computational and Graphical Statistics, 22(3):649–664, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Minnotte:1998:BRM

[MMW98] Michael C. Minnotte, David J. Marchette, and Edward J. Wegman. The bumpy road to the mode forest. Journal of Computational and Graphical Statistics, 7(2):239– 251, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/minnotte.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474773.

Min:2022:FSE

[MMZ22] Keqian Min, Qing Mai, and Xin Zhang. Fast and separable estimation in high-dimensional tensor Gaussian graphical models. Journal of Computational and Graphical Statistics, 31(1):294– 300, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1938086.

Mau:1997:PIB

[MN97] Bob Mau and Michael A. Newton. Phylogenetic inference for binary data on dendograms using Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 6(1):122–131, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
mau.html;   http://www.
tandfonline.com/doi/
abs/10.1080/10618600.
1997.10474731.

Mockus:1998:EDS

[Moc98] Audris Mockus. Estimating dependencies from spatial averages. Journal of Computational and Graphical Statistics, 7(4):501–513, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/

jcgs/abstracts98/mockus.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474791.

Maugis:2020:TEN

[MOPW20] P.-A. G. Maugis, S. C.

Olhede, C. E. Priebe, and P. J. Wolfe. Testing for equivalence of network distribution using subgraph counts. Journal of Computational and Graphical Statistics, 29 (3):455–465, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1736085.

Moustafa:2009:QQG

[Mou09] Rida E. A. Moustafa.

QGPCP: Quantized generalized parallel coordinate plots for large multivariate data visualization. Journal of Computational and Graphical Statistics, 18(1):32–51, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Maceachern:2000:ILF

[MP00] Steven N. Maceachern and Mario Peruggia. Importance link function estimation for Markov chain Monte Carlo methods. Journal of Computational and Graphical Statistics, 9(1):99–121, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/MacEachern.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474868.

Majumdar:2016:ZEP

[MP16] Anandamayee Majumdar and Debashis Paul. Zero expectile processes and Bayesian spatial regression. Journal of Computational and Graphical Statistics, 25(3):727– 747, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1062014.

Mitra:2024:IMS

[MPC24] Anirban Mitra, Konasale Prasad, and Joshua Cape. On inference for modularity statistics in structured networks. Journal of Computational and Graphical Statistics, 33 (4):1370–1382, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2336147.

McGowan:2022:DPD

[MPH22] Lucy D’Agostino Mc-Gowan, Roger D. Peng, and Stephanie C. Hicks. Design principles for data analysis. Journal of Computational and Graphical Statistics, 32(2):754– 761, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2104290.

McGrory:2012:VBR

[MPR^{+}12] Clare A. McGrory, Anthony N. Pettitt, Robert Reeves, Mark Griffin, and Mark Dwyer. Variational Bayes and the reduced dependence approximation for the autologistic model on an irregular grid with applications. Journal of Computational and Graphical Statistics, 21 (3):781–796, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mehta:1992:ESL

[MPS92] Cyrus R. Mehta, Nitin Patel, and Pralay Senchaudhuri. Exact stratified linear rank tests for ordered categorical and binary data. Journal of Computational and Graphical Statistics, 1(1): 21–40, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1992.10474574.

Madrid-Padilla:2017:TDG

[MPS17] Oscar Hernan Madrid-Padilla and James Scott. Tensor decomposition with generalized lasso penalties. Journal of Computational and Graphical Statistics, 26(3):537– 546, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Menafoglio:2021:KRD

[MPS21] Alessandra Menafoglio, Davide Pigoli, and Piercesare Secchi. Kriging Riemannian data via random domain decompositions. Journal of Computational and Graphical Statistics, 30(3):709–727, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1853548.

Man:2023:UAP

[MPTZ23] Rebeka Man, Xiaoou Pan, Kean Ming Tan, and Wen-Xin Zhou. A unified algorithm for penalized convolution smoothed quantile regression. Journal of Computational and Graphical Statistics, 33 (2):625–637, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2275999.

Mukherjee:2015:IDL

[MQ15] Partha Sarathi Mukherjee and Peihua Qiu. Image denoising by a local clustering framework. Journal of Computational and Graphical Statistics, 24(1):254–273, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Murua:2017:SBR

[MQ17] Alejandro Murua and Fernando A. Quintana. Semiparametric Bayesian regression via Potts model. Journal of Computational and Graphical Statistics, 26(2):265–274, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Muller:2011:PPM

[MQR11] Peter Müller, Fernando Quintana, and Gary L. Rosner. A product partition model with regression on covariates. Journal of Computational and Graphical Statistics, 20 (1):260–278, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ming:2001:NOM

[MR01] Kewei Ming and Paul R. Rosenbaum. A note on optimal matching with variable controls using the assignment algorithm. Journal of Computational and Graphical Statistics, 10(3):455–463, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317114938.

Maitra:2010:MDA

[MR10] Ranjan Maitra and Ivan P. Ramler. A k-mean-directions algorithm for fast clustering of data on the sphere. Journal of Computational and Graphical Statistics, 19 (2):377–396, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mukherjee:2016:GPS

[MR16] Chiranjit Mukherjee and Abel Rodriguez. GPUpowered shotgun stochastic search for Dirichlet process mixtures of Gaussian graphical models. Journal of Computational and Graphical Statistics, 25(3):762–788, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2015.1037883.

Molstad:2019:PLM

[MR19] Aaron J. Molstad and Adam J. Rothman. A penalized likelihood method for classification with matrix-valued predictors. Journal of Computational and Graphical Statistics, 28(1):11–22, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Moradi:2018:KBI

[MRCM18] M. Mehdi Moradi, Francisco J. Rodríguez-Cortés, and Jorge Mateu. On kernel-based intensity estimation of spatial point patterns on linear networks. Journal of Computational and Graphical Statistics, 27(2):302–311, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Masak:2022:ICS

[MRP22] Tomas Masak, Tomas Rubin, and Victor M. Panaretos. Inference and computation for sparsely sampled random surfaces. Journal of Computational and Graphical Statistics, 31(4):1361–1374, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2022.2034640.

Metry:1993:CGB

[MS93a] Magdy H. Metry and Allan R. Sampson. Characterizing and generating bivariate empirical rank distributions satisfying certain positive dependence concepts. Journal of Computational and Graphical Statistics, 2(2): 209–223, June 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474608.

Minnotte:1993:MTT

[MS93b] Michael C. Minnotte and David W. Scott. The mode tree: a tool for visualization of nonparametric density features. Journal of Computational and Graphical Statistics, 2(1):51–68, 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474599.

Meng:2002:WBS

[MS02] Xiao-Li Meng and Stephen Schilling. Warp bridge sampling. Journal of Computational and Graphical Statistics, 11(3):552– 586, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s4/p552;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002457.

Miyata:2003:AFK

[MS03] Satoshi Miyata and Xiaotong Shen. Adaptive free-knot splines. Journal of Computational and Graphical Statistics, 12(1):197–213, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s10/p197;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031284.

Ma:2018:ADV

[MS18a] Li Ma and Jacopo Soriano. Analysis of distributional variation through graphical multiscale beta-binomial models. Journal of Computational and Graphical Statistics, 27(3):529–541, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Morrison:2018:RCC

[MS18b] Jean Morrison and Noah Simon. Rank conditional coverage and confidence intervals in highdimensional problems. Journal of Computational and Graphical Statistics, 27(3):648–656, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Marbac:2022:SSE

[MSBV22] Matthieu Marbac, Mohammed Sedki, Christophe Biernacki, and Vincent Vandewalle. Simultaneous semiparametric estimation of clustering and regression. Journal of Computational and Graphical Statistics, 31 (2):477–485, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000872.

Marco:2024:FMM

[MSJ^{+}24] Nicholas Marco, Damla Sentürk, Shafali Jeste, Charlotte DiStefano, Abigail Dickinson, and Donatello Telesca. Functional mixed membership models. Journal of Computational and Graphical Statistics, 33(4):1139– 1149, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2024.2304633.

Murua:2008:PMC

[MSS08] Alejandro Murua, Larissa Stanberry, and Werner Stuetzle. On Potts model clustering, kernel K-means and density estimation. Journal of Computational and Graphical Statistics, 17(3):629–658, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Morgenthaler:2000:FQD

[MT00] Stephan Morgenthaler and John W. Tukey. Fitting quantiles: Doubling, HR, HQ, and HHH distributions. Journal of Computational and Graphical Statistics, 9(1):180–195, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Morgenthaler.
htm;   https://www.math.
utah.edu/pub/bibnet/
authors/t/tukey-john-
w.bib;   http://www.
tandfonline.com/doi/
abs/10.1080/10618600.
2000.10474873.

McQuarrie:2003:ODA

[MT03] Allan D. McQuarrie and Chih-Ling Tsai. Outlier detections in autoregressive models. Journal of Computational and Graphical Statistics, 12(2):450–471, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s11/p450;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031671.

Muller:1998:NAP

[Mul98] Keith E. Muller. A new F approximation for the Pillai–Bartlett trace under H_{0}. Journal of Computational and Graphical Statistics, 7(1):131–137, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.research.att.com/
~andreas/jcgs/abstracts/
1998/march/muller.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474765.

Muller:1999:ACB

[Mül99] D. W. Müller. An algorithm for computing the best convex contrast curve of two bivariate samples. Journal of Computational and Graphical Statistics, 8(3):479–492, September 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474827.

Murrell:1999:LMA

[Mur99] Paul R. Murrell. Layouts: a mechanism for arranging plots on a page. Journal of Computational and Graphical Statistics, 8 (1):121–134, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474805.

Murrell:2013:ISG

[Mur13] Paul Murrell. InfoVis and statistical graphics: Comment. Journal of Computational and Graphical Statistics, 22(1):33–37, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Muller:1998:BIW

[MV98] Peter Müller and Brani Vidakovic. Bayesian inference with wavelets: Density estimation. Journal of Computational and Graphical Statistics, 7(4):456–468, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/muller2.
html;   http://www.tandfonline.

com/doi/abs/10.1080/
10618600.1998.10474788.

Medeiros:2002:CAP

[MVR02] Marcelo C. Medeiros, Alvaro Veiga, and Mauricio G. C. Resende. A combinatorial approach to piecewise linear time series analysis. Journal of Computational and Graphical Statistics, 11 (1):236–258, March 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s13/p236;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375712.

Manrique-Vallier:2014:BED

[MVR14] Daniel Manrique-Vallier and Jerome P. Reiter. Bayesian estimation of discrete multivariate latent structure models with structural zeros. Journal of Computational and Graphical Statistics, 23(4):1061–1079, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

McCoy:1996:WAS

[MW96] E. J. McCoy and A. T. Walden. Wavelet analysis and synthesis of stationary long-memory processes. Journal of Computational and Graphical Statistics, 5(1):26–56, March 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474693.

Marchette:1997:FMT

[MW97] David J. Marchette and Edward J. Wegman. The filtered mode tree. Journal of Computational and Graphical Statistics, 6(2):143–159, June 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/june/
marchette.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474734.

Muller:2004:CTM

[MW04] Hans-Georg Müller and Newton Wai. Change trees and mutagrams for the visualization of local changes in sequence data. Journal of Computational and Graphical Statistics, 13(3):571–585, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Murua:2014:CPC

[MW14] Alejandro Murua and Nicolas Wicker. The conditional-Potts clustering model. Journal of Computational and Graphical Statistics, 23 (3):717–739, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Mena:2015:BMM

[MW15] Ramsés H. Mena and Stephen G. Walker. On the Bayesian mixture model and identifiability. Journal of Computational and Graphical Statistics, 24(4):1155–1169, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.950376.

Molstad:2021:EMC

[MWDR21] Aaron J. Molstad, Guangwei Weng, Charles R. Doss, and Adam J. Rothman. An explicit meancovariance parameterization for multivariate response linear regression. Journal of Computational and Graphical Statistics, 30(3):612–621, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1853551.

Malsiner-Walli:2017:IMM

[MWFSG17] Gertraud Malsiner-Walli, Sylvia Frühwirth-Schnatter, and Bettina Grün. Identifying mixtures of mixtures using Bayesian estimation. Journal of Computational and Graphical Statistics, 26(2):285–295, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Melnykov:2023:SSC

[MWM^{+}23] Volodymyr Melnykov, Yang Wang, Yana Melnykov, Francesca Torti, Domenico Perrotta, and Marco Riani. On simulating skewed and clusterweighted data for studying performance of clustering algorithms. Journal of Computational and Graphical Statistics, 33 (1):303–309, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2210338.

Mao:2024:MMC

[MWWY24] Xiaojun Mao, Hengfang Wang, Zhonglei Wang, and Shu Yang. Mixed matrix completion in complex survey sampling under heterogeneous missingness. Journal of Computational and Graphical Statistics, 33(4):1320– 1328, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2319154.

Meng:2021:LDB

[MXM^{+}21] Cheng Meng, Rui Xie, Abhyuday Mandal, Xinlian Zhang, Wenxuan Zhong, and Ping Ma. LowCon: a design-based subsampling approach in a misspecified linear model. Journal of Computational and Graphical Statistics, 30(3):694– 708, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1844215.

McFarlane:1994:GSA

[MY94] Mary McFarlane and Forrest W. Young. Graphical sensitivity analysis for multidimensional scaling. Journal of Computational and Graphical Statistics, 3(1):23–33, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474628.

Meng:2022:SSA

[MYC^{+}22] Cheng Meng, Jun Yu, Yongkai Chen, Wenxuan Zhong, and Ping Ma. Smoothing splines approximation using Hilbert curve basis selection. Journal of Computational and Graphical Statistics, 31(3):802–812, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2002161.

Noguchi:2016:EAM

[NAB16] Kimihiro Noguchi, Alexander Aue, and Prabir Burman. Exploratory analysis and modeling of stock returns. Journal of Computational and Graphical Statistics, 25(2):363–381, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nychka:2015:MGP

[NBH^{+}15] Douglas Nychka, Soutir Bandyopadhyay, Dorit Hammerling, Finn Lindgren, and Stephan Sain. A multiresolution Gaussian process model for the analysis of large spatial datasets. Journal of Computational and Graphical Statistics, 24(2):579– 599, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.914946.

Nieto-Barajas:2021:PPT

[NBNA21] Luis Nieto-Barajas and Gabriel Núñez-Antonio. Projected Pólya tree. Journal of Computational and Graphical Statistics, 30(4):1197–1208, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1923515.

Nagy:2021:ID

[ND21] Stanislav Nagy and Jirí Dvorák. Illumination depth. Journal of Computational and Graphical Statistics, 30(1):78– 90, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1776717.

Neal:2000:MCS

[Nea00] Radford M. Neal. Markov chain sampling methods for Dirichlet process mixture models. Journal of Computational and Graphical Statistics, 9(2): 249–265, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474879.

Nason:2004:C

[NEL04] Martha Nason, Scott Emerson, and Michael LeBlanc. CARTscans. Journal of Computational and Graphical Statistics, 13(4):807–825, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nemeth:2016:PAS

[NFM16] Christopher Nemeth, Paul Fearnhead, and Lyudmila Mihaylova. Particle approximations of the score and observed information matrix for parameter estimation in state-space models with linear computational cost. Journal of Computational and Graphical Statistics, 25 (4):1138–1157, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nott:2014:ABC

[NFMS14] D. J. Nott, Y. Fan, L. Marshall, and S. A. Sisson. Approximate Bayesian computation and Bayes’ linear analysis: Toward high-dimensional ABC. Journal of Computational and Graphical Statistics, 23(1):65–86, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nott:2004:BVS

[NG04] David J. Nott and Peter J. Green. Bayesian variable selection and the Swendsen–Wang algorithm. Journal of Computational and Graphical Statistics, 13(1):141–157, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nagy:2017:DBR

[NGH17] Stanislav Nagy, Irène Gijbels, and Daniel Hlubinka. Depth-based recognition of shape outlying functions. Journal of Computational and Graphical Statistics, 26 (4):883–893, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ni:2020:CMC

[NJM20] Yang Ni, Yuan Ji, and Peter Müller. Consensus Monte Carlo for random subsets using shared anchors. Journal of Computational and Graphical Statistics, 29(4):703– 714, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1737085.

Nott:2004:SSB

[NL04] David J. Nott and Daniela Leonte. Sampling schemes for Bayesian variable selection in generalized linear models. Journal of Computational and Graphical Statistics, 13 (2):362–382, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nelson:2006:UPI

[NLF^{+}06] Kerrie P. Nelson, Stuart R. Lipsitz, Garrett M. Fitzmaurice, Joseph Ibrahim, Michael Parzen, and Robert Strawderman. Use of the probability integral transformation to fit nonlinear mixed-effects models with nonnormal random effects. Journal of Computational and Graphical Statistics, 15(1):39– 57, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Natarajan:1998:GSD

[NM98] Ranjini Natarajan and Charles E. McCulloch. Gibbs sampling with diffuse proper priors: a valid approach to datadriven inference? Journal of Computational and Graphical Statistics, 7(3):267–277, September 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/natarajan.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474776.

Ni:2020:SBN

[NMD^{+}20] Yang Ni, Peter Müller, Maurice Diesendruck, Sinead Williamson, Yitan Zhu, and Yuan Ji. Scalable Bayesian nonparametric clustering and classification. Journal of Computational and Graphical Statistics, 29 (1):53–65, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1624366.

Naiman:2001:CSS

[NP01] Daniel Q. Naiman and Carey E. Priebe. Computing scan statistic p values using importance sampling, with applications to genetics and medical image analysis. Journal of Computational and Graphical Statistics, 10 (2):296–328, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628194.

Nason:1994:DWT

[NS94] G. P. Nason and B. W.

Silverman. The discrete wavelet transform in S. Journal of Computational and Graphical Statistics, 3(2):163–191, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474637.

Neto:2012:PLS

[NST12] David Neto, Sylvain Sardy, and Paul Tseng. l_{1}-penalized likelihood smoothing and segmentation of volatility processes allowing for abrupt changes. Journal of Computational and Graphical Statistics, 21(1):217–233, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Naul:2017:SSC

[NT17] Brett Naul and Jonathan Taylor. Sparse Steinian covariance estimation. Journal of Computational and Graphical Statistics, 26(2):355–366, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nott:2012:RDE

[NTVK12] David J. Nott, Siew Li Tan, Mattias Villani, and Robert Kohn. Regression density estimation with variational methods and stochastic approximation. Journal of Computational and Graphical Statistics, 21(3):797–820, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Niemi:2010:AMM

[NW10] Jarad Niemi and Mike West. Adaptive mixture modeling Metropolis methods for Bayesian analysis of nonlinear state-space models. Journal of Computational and Graphical Statistics, 19 (2):260–280, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Nan:2014:VSD

[NY14] Ying Nan and Yuhong Yang. Variable selection diagnostics measures for high-dimensional regression. Journal of Computational and Graphical Statistics, 23(3):636–656, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ng:2015:LIH

[NYC15] Chi Tim Ng, Chun Yip Yau, and Ngai Hang Chan. Likelihood inferences for high-dimensional factor analysis of time series with applications in finance. Journal of Computational and Graphical Statistics, 24(3):866–884, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ofori-Boateng:2021:NAD

[OBGC21] Dorcas Ofori-Boateng, Yulia R. Gel, and Ivor Cribben. Nonparametric anomaly detection on time series of graphs. Journal of Computational and Graphical Statistics, 30(3):756–767, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1844214.

OSullivan:2001:ARP

[OC01] Finbarr O’Sullivan and Kingshuk Roy Choudhury. An analysis of the role of positivity and mixture model constraints in Poisson deconvolution problems. Journal of Computational and Graphical Statistics, 10(4):673–696, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243395.

Oehlert:1998:FAI

[Oeh98] Gary W. Oehlert. Faster adaptive importance sampling in low dimensions. Journal of Computational and Graphical Statistics, 7(2):158– 174, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts98/oehlert.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474768.

Ortner:2018:GPA

[OFR^{+}18] Thomas Ortner, Peter Filzmoser, Maia Rohm, Christian Breiteneder, and Sarka Brodinova. Guided projections for analyzing the structure of high-dimensional data. Journal of Computational and Graphical Statistics, 27(4):750–762, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Okuno:2024:INN

[OH24] Akifumi Okuno and Kazuharu Harada. An interpretable neural networkbased nonproportional odds model for ordinal regression. Journal of Computational and Graphical Statistics, 33(4):1454– 1463, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2024.2321208.

Osmundsen:2021:ISB

[OKL21] Kjartan Kloster Osmundsen, Tore Selland Kleppe, and Roman Liesenfeld. Importance samplingbased transport map Hamiltonian Monte Carlo for Bayesian hierarchical models. Journal of Computational and Graphical Statistics, 30(4):906– 919, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923519.

Oh:2011:CFN

[OLN11a] Hee-Seok Oh, Thomas C. M. Lee, and Douglas W. Nychka. Correction to “Fast Nonparametric Quantile Regression With Arbitrary Smoothing Methods” published in the Journal of Computational and Graphical Statistics, 20, 510–526. Journal of Computational and Graphical Statistics, 20(3):787–788, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [OLN11b].

Oh:2011:FNQ

[OLN11b] Hee-Seok Oh, Thomas C. M. Lee, and Douglas W. Nychka. Fast nonparametric quantile regression with arbitrary smoothing methods. Journal of Computational and Graphical Statistics, 20(2):510–526, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See correction [OLN11a].

Oman:2019:LSC

[OM19] Samuel D. Oman and Jorge Mateu. The latent scale covariogram: A tool for exploring the spatial dependence structure of nonnormal responses. Journal of Computational and Graphical Statistics, 28(1):127–141, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ong:2018:GVA

[ONS18] Victor M.-H. Ong, David J. Nott, and Michael S. Smith. Gaussian variational approximation with a factor covariance structure. Journal of Computational and Graphical Statistics, 27(3):465–478, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ooi:2002:DVM

[Ooi02] Hong Ooi. Density visualization and mode hunting using trees. Journal of Computational and Graphical Statistics, 11 (2):328–347, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s5/p328;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180545.

Osborne:2000:LD

[OPT00] Michael R. Osborne, Brett Presnell, and Berwin A. Turlach. On the LASSO and its dual. Journal of Computational and Graphical Statistics, 9(2): 319–337, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474883.

Osborne:2022:LNE

[OPV22] Nathan Osborne, Christine B. Peterson, and Marina Vannucci. Latent network estimation and variable selection for compositional data via variational EM. Journal of Computational and Graphical Statistics, 31 (1):163–175, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2021.1935971.

Opsomer:1999:RCB

[OR99] Jean D. Opsomer and David Ruppert. A rootn consistent backfitting estimator for semiparametric additive modeling. Journal of Computational and Graphical Statistics, 8(4):715–732, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts99/opsom.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1999.10474845.

Oh:2007:MBC

[OR07] Man-Suk Oh and Adrian E. Raftery. Model-based clustering with dissimilarities. Journal of Computational and Graphical Statistics, 16(3):559–585, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Otto:2022:ESW

[OS22] Philipp Otto and Rick Steinert. Estimation of the spatial weighting matrix for spatiotemporal data under the presence of structural breaks. Journal of Computational and Graphical Statistics, 32 (2):696–711, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2107530.

Olshen:2018:NAD

[OSR^{+}18] Adam B. Olshen, Robert L.

Strawderman, Gregory Ryslik, Karen Lostritto, Alice M. Arnold, and Annette M. Molinaro. Novel aggregate deletion/ substitution/addition learning algorithms for recursive partitioning. Journal of Computational and Graphical Statistics, 27 (1):146–156, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Osborne:2011:HAQ

[OT11] M. R. Osborne and B. A.

Turlach. A homotopy algorithm for the quantile regression lasso and related piecewise linear problems. Journal of Computational and Graphical Statistics, 20 (4):972–987, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Oue:1994:D

[Oue94] Shingo Oue. Discussion.

Journal of Computational and Graphical Statistics, 3(4):363–367, 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474651.

OConnell:1997:SRM

[OW97] Michael A. O’Connell and Russell D. Wolfinger. Spatial regression models, response surfaces, and process optimization. Journal of Computational and Graphical Statistics, 6(2):224– 241, June 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/june/
oconnell.html;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474738.

Ormerod:2012:GVA

[OW12] J. T. Ormerod and M. P. Wand. Gaussian variational approximate inference for generalized linear mixed models. Journal of Computational and Graphical Statistics, 21 (1):2–17, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Owen:2017:SET

[Owe17] Art B. Owen. Statistically efficient thinning of a Markov chain sampler. Journal of Computational and Graphical Statistics, 26(3):738–744, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Okuno:2023:GGE

[OY23] Akifumi Okuno and Keisuke Yano. A generalization gap estimation for overparameterized models via the Langevin functional variance. Journal of Computational and Graphical Statistics, 32(4):1287– 1295, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2197488.

Papp:2014:SCE

[PA14] Dávid Papp and Farid Alizadeh. Shape-constrained estimation using nonnegative splines. Journal of Computational and Graphical Statistics, 23 (1):211–231, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Paatero:1999:MET

[Paa99] Pentti Paatero. The multilinear engine — a tabledriven, least squares program for solving multilinear problems, including the n-way parallel factor analysis model. Journal of Computational and Graphical Statistics, 8(4):854–888, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/paatero.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474853.

Pan:1999:BLM

[Pan99a] Wei Pan. Bootstrapping likelihood for model selection with small samples. Journal of Computational and Graphical Statistics, 8(4):687–698, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/pan.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474843.

Pan:1999:EIC

[Pan99b]  Wei Pan.  Extending

the iterative convex minorant algorithm to the Cox model for intervalcensored data. Journal of Computational and Graphical Statistics, 8 (1):109–120, March 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/abs/10.1080/10618600.
1999.10474804.

Poythress:2022:LRO

[PAP22] J. C. Poythress, Jeongyoun Ahn, and Cheolwoo Park. Low-rank, orthogonally decomposable tensor regression with application to visual stimulus decoding of fMRI data. Journal of Computational and Graphical Statistics, 31(1):190–203, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1951741.

Pardoe:2001:BSA

[Par01] Iain Pardoe. A Bayesian sampling approach to regression model checking. Journal of Computational and Graphical Statistics, 10(4):617–627, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243359.

Park:2005:PLA

[Par05] Trevor Park. A penalized likelihood approach to rotation of principal components. Journal of Computational and Graphical Statistics, 14(4):867–888, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Park:2011:BAI

[Par11] Taeyoung Park. Bayesian analysis of individual choice behavior with aggregate data. Journal of Computational and Graphical Statistics, 20 (1):158–173, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Paradis:2018:MSV

[Par18] Emmanuel Paradis. Multidimensional scaling with very large datasets. Journal of Computational and Graphical Statistics, 27 (4):935–939, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Park:2022:CEL

[Par22] Gunwoong Park. Computationally efficient learning of Gaussian linear structural equation models with equal error variances. Journal of Computational and Graphical Statistics, 32(3):1060– 1073, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2154779.

Pinheiro:1995:ALL

[PB95] José C. Pinheiro and Douglas M. Bates. Approximations to the loglikelihood function in the nonlinear mixed-effects model. Journal of Computational and Graphical Statistics, 4(1):12– 35, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474663.

Padellini:2022:PFT

[PB22] Tullia Padellini and Pierpaolo Brutti. Persistence flamelets: Topological invariants for scale spaces. Journal of Computational and Graphical Statistics, 32(2):671–683, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2074427.

Pinheiro:2006:ELA

[PC06] José C. Pinheiro and Edward C. Chao. Efficient Laplacian and adaptive Gaussian quadrature algorithms for multilevel generalized linear mixed models. Journal of Computational and Graphical Statistics, 15(1):58– 81, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pircalabelu:2022:LMM

[PC22] Eugen Pircalabelu and Gerda Claeskens. Linear manifold modeling and graph estimation based on multivariate functional data with different coarseness scales. Journal of Computational and Graphical Statistics, 32 (2):378–387, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2108818.

Pratola:2014:PBA

[PCG^{+}14] Matthew T. Pratola, Hugh A. Chipman, James R. Gattiker, David M. Higdon, Robert McCulloch, and William N. Rust. Parallel Bayesian additive regression trees. Journal of Computational and Graphical Statistics, 23 (3):830–852, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pratola:2020:HBM

[PCGM20] M. T. Pratola, H. A. Chipman, E. I. George, and R. E. McCulloch. Heteroscedastic BART via multiplicative regression trees. Journal of Computational and Graphical Statistics, 29 (2):405–417, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1677243.

Pao:2011:SIR

[PCP11] Henry Pao, Glen A. Coppersmith, and Carey E. Priebe. Statistical inference on random graphs: Comparative power analyses via Monte Carlo. Journal of Computational and Graphical Statistics, 20(2):395–416, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Price:2018:BSL

[PDLN18] L. F. Price, C. C.

Drovandi, A. Lee, and D. J. Nott. Bayesian synthetic likelihood. Journal of Computational and Graphical Statistics, 27 (1):1–11, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pal:2023:FMF

[PDM23] Subrata Pal, Somak Dutta, and Ranjan Maitra. Fast matrix-free methods for model-based personalized synthetic MR imaging. Journal of Computational and Graphical Statistics, 33(3):1109– 1117, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2284208.

Peng:2017:CSY

[Pen17] Roger D. Peng. Comment on “50 Years of Data Science”. Journal of Computational and Graphical Statistics, 26 (4):767, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Don17].

Puchhammer:2023:SSR

[PF23]  Patricia  Puchhammer

and Peter Filzmoser. Spatially smoothed robust covariance estimation for local outlier detection. Journal of Computational and Graphical Statistics, 33(3):928–940, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2277875.

Piancastelli:2022:MCM

[PFBSO22] Luiza S. C. Piancastelli, Nial Friel, Wagner Barreto-Souza, and Hernando Ombao. Multivariate Conway–Maxwell– Poisson distribution: Sarmanov method and doubly intractable Bayesian inference. Journal of Computational and Graphical Statistics, 32(2):483– 500, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2116443.

Pratola:2023:IOB

[PGM23] M. T. Pratola, E. I.

George, and R. E. McCulloch. Influential observations in Bayesian regression tree models. Journal of Computational and Graphical Statistics, 33 (1):47–63, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2210180.

Price:2015:RFS

[PGR15] Bradley S. Price, Charles J.

Geyer, and Adam J. Rothman. Ridge fusion in statistical learning. Journal of Computational and Graphical Statistics, 24 (2):439–454, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.920709.

Price:2019:ARC

[PGR19] Bradley S. Price, Charles J. Geyer, and Adam J. Rothman. Automatic response category combination in multinomial logistic regression. Journal of Computational and Graphical Statistics, 28 (3):758–766, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Park:2020:FEA

[PH20] Jaewoo Park and Murali Haran. A function emulation approach for doubly intractable distributions. Journal of Computational and Graphical Statistics, 29(1):66–77, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1629941.

Park:2021:RDM

[PH21] Jaewoo Park and Murali Haran. Reduceddimensional Monte Carlo maximum likelihood for latent Gaussian random field models. Journal of Computational and Graphical Statistics, 30 (2):269–283, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2020.1811106.

Passino:2023:MEP

[PH23] Francesco Sanna Passino and Nicholas A. Heard. Mutually exciting point process graphs for modeling dynamic networks. Journal of Computational and Graphical Statistics, 32(1):116–130, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2096048.

Park:2014:NCM

[PHK14] Cheolwoo Park, Jan Hannig, and Kee-Hoon Kang. Nonparametric comparison of multiple regression curves in scale-space. Journal of Computational and Graphical Statistics, 23(3):657–677, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Papastamoulis:2010:AAB

[PI10] Panagiotis Papastamoulis and George Iliopoulos. An artificial allocations based solution to the label switching problem in Bayesian analysis of mixtures of distributions. Journal of Computational and Graphical Statistics, 19(2):313–331, June 2010.

CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Picchini:2014:ISM

[Pic14] Umberto Picchini. Inference for SDE models via approximate Bayesian computation. Journal of Computational and Graphical Statistics, 23 (4):1080–1100, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Piepho:2004:ALB

[Pie04] Hans-Peter Piepho. An algorithm for a letterbased representation of all-pairwise comparisons. Journal of Computational and Graphical Statistics, 13(2):456–466, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pittman:2002:ASG

[Pit02] Jennifer Pittman. Adaptive splines and genetic algorithms. Journal of Computational and Graphical Statistics, 11(3):615–638, September 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n3/s7/p615;   http:

//www.tandfonline.com/
doi/abs/10.1198/106186002448.

Park:2022:BSF

[PJS^{+}22] Jaewoo Park, Yeseul Jeon, Minsuk Shin, Minjeong Jeon, and Ick Hoon Jin. Bayesian shrinkage for functional network models, with applications to longitudinal item response data. Journal of Computational and Graphical Statistics, 31 (2):360–377, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1999823.

Pal:2015:IDA

[PKH15] Subhadip Pal, Kshitij Khare, and James P. Hobert. Improving the data augmentation algorithm in the two-block setup. Journal of Computational and Graphical Statistics, 24(4):1114– 1133, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.955177.

Park:2012:BAG

[PL12] Jincheol Park and Faming Liang. Bayesian analysis of geostatistical models with an auxiliary lattice. Journal of Computational and Graphical Statistics, 21(2):453–475, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pourmohamad:2022:BOB

[PL22] Tony Pourmohamad and Herbert K. H. Lee. Bayesian optimization via barrier functions. Journal of Computational and Graphical Statistics, 31 (1):74–83, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1935270.

Park:2010:MEA

[PLH10] Cheolwoo Park, Thomas C. M. Lee, and Jan Hannig. Multiscale exploratory analysis of regression quantiles using quantile sizer. Journal of Computational and Graphical Statistics, 19 (3):497–513, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Parzen:2002:EOR

[PLIK02] Michael Parzen, Stuart Lipsitz, Joseph Ibrahim, and Neil Klar. An estimate of the odds ratio that always exists. Journal of Computational and Graphical Statistics, 11 (2):420–436, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s10/p420;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180590.

Pinheiro:2001:EAR

[PLW01] José C. Pinheiro, Chuanhai Liu, and Ying Nian Wu. Efficient algorithms for robust estimation in linear mixed-effects models using the multivariate t distribution. Journal of Computational and Graphical Statistics, 10 (2):249–276, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628059.

Punzo:2016:CML

[PM16] Antonio Punzo and Antonello Maruotti. Clustering multivariate longitudinal observations: The contaminated Gaussian hidden Markov model. Journal of Computational and Graphical Statistics, 25(4):1097–1098, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Papageorgiou:2020:BSA

[PM20] Georgios Papageorgiou and Benjamin C. Marshall. Bayesian semiparametric analysis of multivariate continuous responses, with variable selection. Journal of Computational and Graphical Statistics, 29(4):896– 909, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1739534.

Paul:2012:ACM

[PMB12] Rajib Paul, Steven N. MacEachern, and L. Mark Berliner. Assessing convergence and mixing of MCMC implementations via stratification. Journal of Computational and Graphical Statistics, 21 (3):693–712, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Perrakis:2020:SBR

[PMI20] Konstantinos Perrakis, Sach Mukherjee, and The Alzheimer’s Disease Neuroimaging Initiative. Scalable Bayesian regression in high dimensions with multiple data sources. Journal of Computational and Graphical Statistics, 29(1):28–39, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1624294.

Price:2021:EMP

[PMS21] Bradley S. Price, Aaron J.

Molstad, and Ben Sherwood. Estimating multiple precision matrices with cluster fusion regularization. Journal of Computational and Graphical Statistics, 30 (4):823–834, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1874963.

Pehlivanian:2022:EOP

[PN22] Charles A. Pehlivanian and Daniel B. Neill. Efficient optimization of partition scan statistics via the consecutive partitions property. Journal of Computational and Graphical Statistics, 32 (2):712–729, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2077351.

Pham:2021:EAM

[PNLL21] Minh Pham, Anh Ninh, Hoang Le, and Yufeng Liu. An efficient algorithm for minimizing multi non-smooth component functions. Journal of Computational and Graphical Statistics, 30 (1):162–170, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1804390.

Posse:1995:TTD

[Pos95] Christian Posse. Tools for two-dimensional exploratory projection pursuit. Journal of Computational and Graphical Statistics, 4(2):83– 100, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474668.

Posse:2001:HMB

[Pos01] Christian Posse. Hierarchical model-based clustering for large datasets. Journal of Computational and Graphical Statistics, 10(3):464–486, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317115072.

Palmer:1996:RUI

[PP96] Judy L. Palmer and Lawrence I. Pettit. Risks of using improper priors with Gibbs sampling and autocorrelated errors. Journal of Computational and Graphical Statistics, 5(3):245–249, 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474709.

Perisic:2005:PPI

[PP05] Igor Perisic and Christian Posse. Projection pursuit indices based on the empirical distribution function. Journal of Computational and Graphical Statistics, 14(3):700–715, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pena:2007:CRS

[PP07] Daniel Peña and Francisco J. Prieto. Combining random and specific directions for outlier detection and robust estimation in high-dimensional multivariate data. Journal of Computational and Graphical Statistics, 16 (1):228–254, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Peng:2009:GAM

[PP09] Jie Peng and Debashis Paul. A geometric approach to maximum likelihood estimation of the functional principal components from sparse longitudinal data. Journal of Computational and Graphical Statistics, 18 (4):995–1015, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pakman:2014:EHM

[PP14] Ari Pakman and Liam Paninski. Exact Hamiltonian Monte Carlo for truncated multivariate Gaussians. Journal of Computational and Graphical Statistics, 23 (2):518–542, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Park:2022:SIM

[PPTO22] Hyung Park, Eva Petkova, Thaddeus Tarpey, and R. Todd Ogden. A single-index model with a surface-link for optimizing individualized dose rules. Journal of Computational and Graphical Statistics, 31(2):553– 562, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2021.1923521.

Page:2022:DMT

[PQD22] Garritt L. Page, Fernando A. Quintana, and David B. Dahl. Dependent modeling of temporal sequences of random partitions. Journal of Computational and Graphical Statistics, 31 (2):614–627, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1987255.

Page:2022:CPV

[PQM22] Garritt L. Page, Fernando A. Quintana, and Peter Müller. Clustering and prediction with variable dimension covariates. Journal of Computational and Graphical Statistics, 31(2):466–476, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1999824.

Paul:2024:CTM

[PRCF24] Biplab Paul, Philip T.

Reiss, Erjia Cui, and Noemi Foà. Continuoustime multivariate analysis. Journal of Computational and Graphical Statistics, 34(1):384– 394, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2374570.

Pregibon:1995:D

[Pre95] Daryl Pregibon. Discussion. Journal of Computational and Graphical Statistics, 4(4):257– 258, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1995.10474682.

Pnevmatikakis:2014:FKF

[PRHP14] Eftychios A. Pnevmatikakis, Kamiar Rahnama Rad, Jonathan Huggins, and Liam Paninski. Fast Kalman filtering and forward-backward smoothing via a low-rank perturbative approach. Journal of Computational and Graphical Statistics, 23 (2):316–339, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Philipp:2018:MSR

[PRHS18] Michel Philipp, Thomas Rusch, Kurt Hornik, and Carolin Strobl. Measuring the stability of results from supervised statistical learning. Journal of Computational and Graphical Statistics, 27 (4):685–700, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Proietti:2006:ENA

[Pro06] Tommaso Proietti. On the estimation of nonlinearly aggregated mixed models. Journal of Computational and Graphical Statistics, 15(1):18– 38, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Papaspiliopoulos:2011:WTN

[PRS11] O. Papaspiliopoulos, G. O.

Roberts, and G. Sermaidis. Whether ’tis nobler in the mind to suffer the slings and arrows of outrageous mixing problems, or to take arms against a sea of troubles, and by opposing end them? Journal of Computational and Graphical Statistics, 20(3):592–602, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Papaspiliopoulos:2013:DAD

[PRS13] Omiros Papaspiliopoulos, Gareth O. Roberts, and Osnat Stramer. Data augmentation for diffusions. Journal of Computational and Graphical Statistics, 22(3):665–688, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Price:2022:LWV

[PSDK22] Bradley S. Price, John P. Saldanha, Dariane Drake, and Katherine Kopp. Lessons from West Virginia’s pandemic response. Journal of Computational and Graphical Statistics, 32(3):763– 764, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2126481.

Priddle:2022:EBS

[PSF^{+}22] Jacob W. Priddle, Scott A. Sisson, David T. Frazier, Ian Turner, and Christopher Drovandi. Efficient Bayesian synthetic likelihood with whitening transformations. Journal of Computational and Graphical Statistics, 31 (1):50–63, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1979012.

Predescu:2023:TSS

[PSRM^{+}23] Cristian Predescu, Michael Snarski, Avi Robinson-Mosher, Duluxan Sritharan, Tamas Szalay, and David E. Shaw. Times square sampling: an adaptive algorithm for free energy estimation. Journal of Computational and Graphical Statistics, 33(3):1073–1083, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2291108.

Priebe:2015:SIE

[PSTV15] Carey E. Priebe, Daniel L.

Sussman, Minh Tang, and Joshua T. Vogelstein. Statistical inference on errorfully observed graphs. Journal of Computational and Graphical Statistics, 24(4):930–953, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.951049.

Park:2009:CMS

[PSY09] Jin-Hong Park, T. N. Sriram, and Xiangrong Yin. Central mean subspace in time series. Journal of Computational and Graphical Statistics, 18 (3):717–730, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Pilla:2003:AMS

[PTP03] Ramani S. Pilla, Peng Tao, and Carey E. Priebe. Adaptive methods for spatial scan analysis via semiparametric mixture models. Journal of Computational and Graphical Statistics, 12(2):332–353, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

                                        URL http://giorgio.
                                        ingentaselect.com/cgi-
                                        bin/linker?ini=asa{\&
                                        }reqidx=/cw/asa/10618600/
                                        v12n2/s5/p332;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/1061860031716.

Pison:2004:DPR

[PV04] Greet Pison and Stefan Van Aelst. Diagnostic plots for robust multivariate methods. Journal of Computational and Graphical Statistics, 13 (2):310–329, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Prangle:2023:DIS

[PV23] Dennis Prangle and Cecilia Viscardi. Distilling importance sampling for likelihood free inference. Journal of Computational and Graphical Statistics, 32(4):1461–1471, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2175688.

Park:2009:PCG

[PvD09] Taeyoung Park and David A.

van Dyk. Partially collapsed Gibbs samplers: Illustrations and applications. Journal of Computational and Graphical Statistics, 18(2):283–305, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Park:2005:BMW

[PVH05] Chun Gun Park, Marina Vannucci, and Jeffrey D. Hart. Bayesian methods for wavelet series in single-index models. Journal of Computational and Graphical Statistics, 14(4):770–794, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Peng:2014:CNK

[PW14] Chien-Yu Peng and C. F. Jeff Wu. On the choice of nugget in kriging modeling for deterministic computer experiments. Journal of Computational and Graphical Statistics, 23 (1):151–168, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Peng:2015:ICD

[PW15] Bo Peng and Lan Wang. An iterative coordinate descent algorithm for high-dimensional nonconvex penalized quantile regression. Journal of Computational and Graphical Statistics, 24(3):676–694, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Peng:2022:GQT

[PW22] Xiang Peng and Huixia Judy Wang. A generalized quantile tree method for subgroup identification. Journal of Computational and Graphical Statistics, 31(3):824–834, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2032723.

Poston:1997:DMR

[PWPS97] Wendy L. Poston, Edward J. Wegman, Carey E. Priebe, and Jeffrey L. Solka. A deterministic method for robust estimation of multivariate location and shape. Journal of Computational and Graphical Statistics, 6(3):300–313, September 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/poston.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474744.

Petersen:2016:FLA

[PWS16] Ashley Petersen, Daniela Witten, and Noah Simon. Fused lasso additive model. Journal of Computational and Graphical Statistics, 25(4):1005– 1025, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Qiao:2024:FRL

[QC24] Nan Qiao and Canyi Chen. Fast and robust low-rank learning over networks: a decentralized matrix quantile regression approach. Journal of Computational and Graphical Statistics, 33 (4):1214–1223, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2353640.

Qian:1996:AIR

[QE96] Shixian Qian and William F.

Eddy. An algorithm for isotonic regression on ordered rectangular grids. Journal of Computational and Graphical Statistics, 5(3):225–235, September 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1996.10474707.

Qin:2006:SSR

[QG06] Li Qin and Wensheng Guo. State space representation for smoothing spline ANOVA models. Journal of Computational and Graphical Statistics, 15(4):830–847, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Qu:2022:SFB

[QG22] Zhuo Qu and Marc G. Genton. Sparse functional boxplots for multivariate curves. Journal of Computational and Graphical Statistics, 31 (4):976–989, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2066680.

Qin:2009:TFF

[QGL09] Li Qin, Wensheng Guo, and Brian Litt. A time-frequency functional model for locally stationary time series data. Journal of Computational and Graphical Statistics, 18 (3):675–693, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Qiu:2002:NPD

[Qiu02] Peihua Qiu. A nonparametric procedure to detect jumps in regression surfaces. Journal of Computational and Graphical Statistics, 11(4):799–822, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s4/p799;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002880.

Qi:2015:SRP

[QLCZ15] Xin Qi, Ruiyan Luo, Raymond J. Carroll, and Hongyu Zhao. Sparse regression by projection and sparse discriminant analysis. Journal of Computational and Graphical Statistics, 24(2):416– 438, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.907094.

Quintana:2004:NBA

[QM04] Fernando A. Quintana and Peter Müller. Nonparametric Bayesian assessment of the order of dependence for binary sequences. Journal of Computational and Graphical Statistics, 13(1):213–231, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Quintana:2000:CAN

[QN00] Fernando A. Quintana and Michael A. Newton. Computational aspects of nonparametric Bayesian analysis with applications to the modeling of multiple binary sequences. Journal of Computational and Graphical Statistics, 9(4):711–737, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Quintana.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474909.

Qiu:2009:UCE

[QS09] Peihua Qiu and Jingran Sun. Using conventional edge detectors and postsmoothing for segmentation of spotted microarray images. Journal of Computational and Graphical Statistics, 18 (1):147–164, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Quiroz:2021:BPE

[QTV^{+}21] Matias Quiroz, Minh-Ngoc Tran, Mattias Villani, Robert Kohn, and Khue-Dung Dang. The block-Poisson estimator for optimally tuned exact subsampling MCMC. Journal of Computational and Graphical Statistics, 30(4):877–888, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1917420.

Quiroz:2018:SMD

[QTVK18] Matias Quiroz, Minh-Ngoc Tran, Mattias Villani, and Robert Kohn. Speeding up MCMC by delayed acceptance and data subsampling. Journal of Computational and Graphical Statistics, 27 (1):12–22, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Qi:2023:SAF

[QWW23] Haobo Qi, Feifei Wang, and Hansheng Wang. Statistical analysis of fixed mini-batch gradient descent estimator. Journal of Computational and Graphical Statistics, 32 (4):1348–1360, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2204130.

Qiu:1997:JDR

[QY97] Peihua Qiu and Brian Yandell. Jump detection in regression surfaces. Journal of Computational and Graphical Statistics, 6(3):332–354, September 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts/qiu.html;
                                         http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1997.10474746.

Qian:2016:TCP

[QYZ16] Wei Qian, Yi Yang, and Hui Zou. Tweedie’s compound Poisson model with grouped elastic net. Journal of Computational and Graphical Statistics, 25(2):606–625, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Riani:2001:UAO

[RA01] Marco Riani and Anthony C. Atkinson. A unified approach to outliers, inuence, and transformations in discriminant analysis. Journal of Computational and Graphical Statistics, 10(3):513–544, September 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317114965.

Riani:2023:RTM

[RAC23] Marco Riani, Anthony C.

Atkinson, and Aldo Corbellini. Robust transformations for multiple regression via additivity and variance stabilization. Journal of Computational and Graphical Statistics, 33(1):85– 100, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2205447.

Ramsay:2017:FEC

[Ram17] J. O. Ramsay. A functional estimate of covariation. Journal of Computational and Graphical Statistics, 26(1):160– 170, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1124041.

Roy:2019:LIL

[RAM19] Sandipan Roy, Yves Atchadé, and George Michailidis. Likelihood inference for large scale stochastic blockmodels with covariates based on a divide-and-conquer parallelizable algorithm with communication. Journal of Computational and Graphical Statistics, 28 (3):609–619, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raphael:2001:PES

[Rap01] Christopher Raphael. A probabilistic expert system for automatic musical accompaniment. Journal of Computational and Graphical Statistics, 10(3):487–512, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317115081.

Rocci:2002:RTA

[RB02] R. Rocci and G. Bove. Rotation techniques in asymmetric multidimensional scaling. Journal of Computational and Graphical Statistics, 11 (2):405–419, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s9/p405;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002760180581.

Rodrigues:2019:PQR

[RDBF19] T. Rodrigues, J.-L. Dortet-Bernadet, and Y. Fan. Pyramid quantile regression. Journal of Computational and Graphical Statistics, 28(3):732–746, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Royle:2007:AMM

[RDL07] J. Andrew Royle, Robert M.

Dorazio, and William A. Link. Analysis of multinomial models with unknown index using data augmentation. Journal of Computational and Graphical Statistics, 16 (1):67–85, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raykar:2010:FCK

[RDZ10] Vikas C. Raykar, Ramani Duraiswami, and Linda H. Zhao. Fast computation of kernel estimators. Journal of Computational and Graphical Statistics, 19(1):205–220, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Resin:2022:SAE

[Res22] Johannes Resin. A simple algorithm for exact multinomial tests. Journal of Computational and Graphical Statistics, 32 (2):539–550, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2102026.

Rodrigues:2017:RAN

[RF17a] T. Rodrigues and Y. Fan. Regression adjustment for noncrossing Bayesian quantile regression. Journal of Computational and Graphical Statistics, 26 (2):275–284, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rover:2017:DAM

[RF17b] Christian Röver and Tim Friede. Discrete approximation of a mixture distribution via restricted divergence. Journal of Computational and Graphical Statistics, 26 (1):217–222, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1276840.

Rumsey:2023:DAS

[RFV23] Kellin Rumsey, Devin Francom, and Scott Vander Wiel. Discovering active subspaces for high-dimensional computer models. Journal of Computational and Graphical Statistics, 33 (3):896–908, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2277297.

Robertson:2021:AVS

[RFVJ21] Nathan Robertson, James M.

Flegal, Dootika Vats, and Galin L. Jones. Assessing and visualizing simultaneous simulation error. Journal of Computational and Graphical Statistics, 30(2):324–334, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1824871.

Ruffieux:2009:AMC

[RG09] Yann Ruffieux and Peter J. Green. Alignment of multiple configurations using hierarchical models. Journal of Computational and Graphical Statistics, 18(3):756–773, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Roosen:1994:ASS

[RH94] Charles B. Roosen and Trevor J. Hastie. Automatic smoothing spline projection pursuit. Journal of Computational and Graphical Statistics, 3(3): 235–248, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474642.

Reiss:2014:MPN

[RHC^{+}14] Philip T. Reiss, Lei Huang, Yin-Hsiu Chen, Lan Huo, Thaddeus Tarpey, and Maarten Mennes. Massively parallel nonparametric regression, with an application to developmental brain mapping. Journal of Computational and Graphical Statistics, 23 (1):232–248, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ramirez-Hassan:2024:TMS

[RHF24] Andrés Ramírez-Hassan and David T. Frazier. Testing model specification in approximate Bayesian computation using asymptotic properties. Journal of Computational and Graphical Statistics, 33(3):1122–1128, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2357630.

Rusch:2018:AQC

[RHM18] Thomas Rusch, Kurt Hornik, and Patrick Mair. Assessing and quantifying clusteredness: The OP-TICS cordillera. Journal of Computational and Graphical Statistics, 27 (1):220–233, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rhodes:2003:SPE

[Rho03] Lee Rhodes. A stream processor for extracting usage intelligence from high-momentum Internet data. Journal of Computational and Graphical Statistics, 12(4):927–944, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032706.

Raynal:2022:CBF

[RHO22] Louis Raynal, Till Hoffmann, and Jukka-Pekka Onnela. Cost-based feature selection for network model choice. Journal of Computational and Graphical Statistics, 32 (3):1109–1118, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2151453.

Rossini:2004:ESS

[RHS^{+}04] A. J. Rossini, Richard M. Heiberger, Rodney A. Sparapani, Martin Mächler, and Kurt Hornik. Emacs speaks statistics. Journal of Computational and Graphical Statistics, 13 (1):247–261, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Robinson:2022:EFS

[RHV22] Emily A. Robinson, Reka Howard, and Susan VanderPlas. Eye fitting straight lines in the modern era. Journal of Computational and Graphical Statistics, 32(4):1537– 1544, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2140668.

Rao:2016:SHS

[RIC^{+}16] Shangbang Rao, Joseph G. Ibrahim, Jian Cheng, Pew-Thian Yap, and Hongtu Zhu. SR-HARDI: Spatially regularizing high angular resolution diffusion imaging. Journal of Computational and Graphical Statistics, 25(4):1195– 1211, 2016. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rendell:2021:GCM

[RJLW21] Lewis J. Rendell, Adam M.

Johansen, Anthony Lee, and Nick Whiteley. Global consensus Monte Carlo. Journal of Computational and Graphical Statistics, 30(2):249–259, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1811105.

Rodu:2022:B

[RK22] Jordan Rodu and Karen Kafadar. The q–q boxplot. Journal of Computational and Graphical Statistics, 31(1):26– 39, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1938586.

Ruczinski:2003:LR

[RKL03] Ingo Ruczinski, Charles Kooperberg, and Michael Leblanc. Logic regression. Journal of Computational and Graphical Statistics, 12(3):475–511, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/

                                        v12n3/s1/p475;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/1061860032238.

Ruggieri:2014:BCP

[RL14] Eric Ruggieri and Charles E. Lawrence. The Bayesian change point and variable selection algorithm: Application to the δ^{18}O proxy record of the Plio-Pleistocene. Journal of Computational and Graphical Statistics, 23 (1):87–110, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rothman:2010:SMR

[RLZ10] Adam J. Rothman, Elizaveta Levina, and Ji Zhu. Sparse multivariate regression with covariance estimation. Journal of Computational and Graphical Statistics, 19 (4):947–962, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rusch:2021:COP

[RMH21] Thomas Rusch, Patrick Mair, and Kurt Hornik. Cluster optimized proximity scaling. Journal of Computational and Graphical Statistics, 30 (4):1156–1167, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1869027.

Raj:2017:PBM

[RMR^{+}17] Mukund Raj, Mahsa Mirzargar, Robert Ricci, Robert M. Kirby, and Ross T. Whitaker. Path boxplots: A method for characterizing uncertainty in path ensembles on a graph. Journal of Computational and Graphical Statistics, 26 (2):243–252, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Reiss:2017:PNS

[RMWH17] Philip T. Reiss, David L.

Miller, Pei-Shien Wu, and Wen-Yu Hua. Penalized nonparametric scalar-on-function regression via principal coordinates. Journal of Computational and Graphical Statistics, 26(3):569–578, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raftery:2012:FIL

[RNHY12] Adrian E. Raftery, Xiaoyue Niu, Peter D. Hoff, and Ka Yee Yeung. Fast inference for the latent space network model using a case-control approximate likelihood. Journal of Computational and Graphical Statistics, 21 (4):901–919, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raim:2018:EGL

[RNM18] Andrew M. Raim, Nagaraj K. Neerchal, and Jorge G. Morel. An extension of generalized linear models to finite mixture outcome distributions. Journal of Computational and Graphical Statistics, 27(3):587–601, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rust:1994:CID

[RO94] Bert W. Rust and Dianne P. O’Leary. Confidence intervals for discrete approximations to ill-posed problems. Journal of Computational and Graphical Statistics, 3(1): 67–96, March 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474631.

Robinson:2019:GPS

[Rob19] Heman Robinson. Good plot symbols by default. Journal of Computational and Graphical Statistics, 28(4):1011–1016, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rosenbaum:1993:CS

[Ros93] Paul R. Rosenbaum. Confident search. Journal of Computational and Graphical Statistics, 2(4): 381–403, December 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474622.

Rosenbaum:2012:OMO

[Ros12] Paul R. Rosenbaum. Optimal matching of an optimally chosen subset in observational studies. Journal of Computational and Graphical Statistics, 21 (1):57–71, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rosenbaum:2017:IMQ

[Ros17] Paul R. Rosenbaum.

Imposing minimax and quantile constraints on optimal matching in observational studies. Journal of Computational and Graphical Statistics, 26 (1):66–78, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1152971.

Rougier:2008:EEM

[Rou08] Jonathan Rougier. Efficient emulators for multivariate deterministic functions. Journal of Computational and Graphical Statistics, 17(4):827– 843, December 2008. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Romanes:2020:DDA

[ROY20] Sarah E. Romanes, John T. Ormerod, and Jean Y. H. Yang. Diagonal discriminant analysis with feature selection for highdimensional data. Journal of Computational and Graphical Statistics, 29 (1):114–127, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1637748.

Roberts:2009:EAM

[RR09] Gareth O. Roberts and Jeffrey S. Rosenthal. Examples of adaptive MCMC. Journal of Computational and Graphical Statistics, 18(2):349–367, June 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Raymaekers:2022:SQR

[RR22] Jakob Raymaekers and Peter J. Rousseeuw. Silhouettes and quasi residual plots for neural nets and tree-based classifiers. Journal of Computational and Graphical Statistics, 31(4):1332–1343, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2050249.

Rao:2023:NFM

[RR23] Aniruddha Rajendra Rao and Matthew Reimherr. Nonlinear functional modeling using neural networks. Journal of Computational and Graphical Statistics, 32(4):1248– 1257, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2165498.

Roy:2021:SSP

[RRG^{+}21] Arkaprava Roy, Brian J.

Reich, Joseph Guinness, Russell T. Shinohara, and Ana-Maria Staicu. Spatial shrinkage via the product independent Gaussian process prior. Journal of Computational and Graphical Statistics, 30(4):1068–1080, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2021.1923512.

Rousseeuw:2018:MDO

[RRH18] Peter J. Rousseeuw, Jakob Raymaekers, and Mia Hubert. A measure of directional outlyingness with applications to image data and video. Journal of Computational and Graphical Statistics, 27 (2):345–359, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rajaratnam:2019:IDH

[RRSY19] Bala Rajaratnam, Steven Roberts, Doug Sparks, and Honglin Yu. Influence diagnostics for highdimensional lasso regression. Journal of Computational and Graphical Statistics, 28(4):877–890, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Roberts:2001:APC

[RS01] Gareth O. Roberts and Sujit K. Sahu. Approximate predetermined convergence properties of the Gibbs sampler. Journal of Computational and Graphical Statistics, 10 (2):216–229, June 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/10618600152627915.

Reich:2019:SMM

[RS19] Brian J. Reich and Benjamin A. Shaby. A spatial Markov model for climate extremes. Journal of Computational and Graphical Statistics, 28 (1):117–126, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Roy:2023:EFS

[RSDG23] Sarbojit Roy, Soham Sarkar, Subhajit Dutta, and Anil K. Ghosh. On exact feature screening in ultrahigh-dimensional binary classification. Journal of Computational and Graphical Statistics, 33 (2):448–462, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2270656.

Rajaratnam:2019:UQM

[RSKZ19] Bala Rajaratnam, Doug Sparks, Kshitij Khare, and Liyuan Zhang. Uncertainty quantification for modern high-dimensional regression via scalable Bayesian methods. Journal of Computational and Graphical Statistics, 28 (1):174–184, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ryden:1998:CBA

[RT98] Tobias Rydén and D. M. Titterington. Computational Bayesian analysis of hidden Markov models. Journal of Computational and Graphical Statistics, 7(2):194– 211, June 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/ryden.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474770.

Ramdas:2016:FFA

[RT16] Aaditya Ramdas and Ryan J. Tibshirani. Fast and flexible ADMM algorithms for trend filtering. Journal of Computational and Graphical Statistics, 25(3):839–858, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1054033.

Rohrbeck:2021:BSC

[RT21] Christian Rohrbeck and Jonathan A. Tawn. Bayesian spatial clustering of extremal behavior for hydrological variables. Journal of Computational and Graphical Statistics, 30 (1):91–105, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1777139.

Rossini:2007:SPS

[RTL07] A. J. Rossini, Luke Tierney, and Na Li. Simple parallel statistical computing in R. Journal of Computational and Graphical Statistics, 16 (2):399–420, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ruppert:1992:CER

[Rup92] David Ruppert. Computing S estimators for regression and multivariate location/dispersion. Journal of Computational and Graphical Statistics, 1(3): 253–270, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474584.

Ruppert:2002:SNK

[Rup02] David Ruppert. Selecting the number of knots for penalized splines. Journal of Computational and Graphical Statistics, 11(4):735–757, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n4/s1/p735;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002853.

Ruckstuhl:1999:RBN

[RW99] Andreas F. Ruckstuhl and Alan H. Welsh. Reference bands for nonparametrically estimated link functions. Journal of Computational and Graphical Statistics, 8(4):699–714, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts99/ruck.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1999.10474844.

Rufibach:2010:BCM

[RW10] Kaspar Rufibach and Guenther Walther. The block criterion for multiscale inference about a density, with applications to other multiscale problems. Journal of Computational and Graphical Statistics, 19(1):175–190, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Rodriguez:2014:LSB

[RW14] Carlos E. Rodríguez and Stephen G. Walker. Label switching in Bayesian mixture models: Deterministic relabeling strategies. Journal of Computational and Graphical Statistics, 23(1):25–45, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ryan:2003:EEI

[Rya03] Kenneth J. Ryan. Estimating expected information gains for experimental designs with application to the random fatigue-limit model. Journal of Computational and Graphical Statistics, 12 (3):585–603, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s6/p585;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032012.

Raudenbush:2000:MLG

[RYY00] Stephen W. Raudenbush, Meng-Li Yang, and Matheos Yosef. Maximum likelihood for generalized linear models with nested random effects via high-order, multivariate Laplace approximation. Journal of Computational and Graphical Statistics, 9(1):141–157, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Raudenbush.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474870.

Roy:2022:CPD

[RZ22] Vivekananda Roy and Lijin Zhang. Convergence of position-dependent MALA with application to conditional simulation in GLMMs. Journal of Computational and Graphical Statistics, 32 (2):501–512, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2118128.

Samarov:2015:FRL

[Sam15] Daniel V. Samarov. The fast RODEO for local polynomial regression. Journal of Computational and Graphical Statistics, 24(4):1034–1052, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2014.949724.

Sankaran:2023:BCR

[San23] Kris Sankaran. Bootstrap confidence regions for learned feature embeddings. Journal of Computational and Graphical Statistics, 32(4):1337– 1347, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2197478.

Sarkar:2022:BSC

[Sar22] Abhra Sarkar. Bayesian semiparametric covariate informed multivariate density deconvolution. Journal of Computational and Graphical Statistics, 31(4):1153–1163, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2060239.

Sardy:2004:ASW

[SAT04] Sylvain Sardy, Anestis Antoniadis, and Paul Tseng. Automatic smoothing with wavelets for a wide class of distributions. Journal of Computational and Graphical Statistics, 13(2):399–421, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sawitzki:1996:ESS

[Saw96] Günther Sawitzki. Extensible statistical software: On a voyage to Oberon. Journal of Computational and Graphical Statistics, 5(3):263–283, 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1996.10474711.

Schweinberger:2017:HDM

[SBE17] Michael Schweinberger, Sergii Babkin, and Katherine B. Ensor. Highdimensional multivariate time series with additional structure. Journal of Computational and Graphical Statistics, 26 (3):610–622, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Seifert:1994:FAN

[SBEG94] Burkhardt Seifert, Michael Brockmann, Joachim Engel, and Theo Gasser. Fast algorithms for nonparametric curve estimation. Journal of Computational and Graphical Statistics, 3(2):192– 213, June 1994. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474638.

Stoehr:2019:NHM

[SBF19] Julien Stoehr, Alan Benson, and Nial Friel. Noisy Hamiltonian Monte Carlo for doubly intractable distributions. Journal of Computational and Graphical Statistics, 28 (1):220–232, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sobczyk:2017:BDR

[SBJ17] Piotr Sobczyk, Małgorzata Bogdan, and Julie Josse. Bayesian dimensionality reduction with PCA using penalized semi-integrated likelihood. Journal of Computational and Graphical Statistics, 26 (4):826–839, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shin:2023:ANN

[SBK23] Jae-Kyung Shin, Kwan-Young Bak, and Ja-Yong Koo. An adaptive neural network regression method for structure identification. Journal of Computational and Graphical Statistics, 33 (3):749–762, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2284206.

Shen:2023:SSC

[SBL23] Hui Shen, Shankar Bhamidi, and Yufeng Liu. Statistical significance of clustering with multidimensional scaling. Journal of Computational and Graphical Statistics, 33 (1):219–230, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2219708.

Safikhani:2022:FSA

[SBM22] Abolfazl Safikhani, Yue Bai, and George Michailidis. Fast and scalable algorithm for detection of structural breaks in big VAR models. Journal of Computational and Graphical Statistics, 31 (1):176–189, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1950005.

Storlie:2010:LAP

[SBR10] Curtis B. Storlie, Howard D. Bondell, and Brian J. Reich. A locally adaptive penalty for estimation of functions with varying roughness. Journal of Computational and Graphical Statistics, 19 (3):569–589, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schalk:2022:ACG

[SBR22] Daniel Schalk, Bernd Bischl, and David Rügamer. Accelerated componentwise gradient boosting using efficient data representation and momentumbased optimization. Journal of Computational and Graphical Statistics, 32 (2):631–641, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2116446.

Sholokhov:2023:RAF

[SBS^{+}23a] Aleksei Sholokhov, James V.

Burke, Damian F. Santomauro, Peng Zheng, and Aleksandr Aravkin. A relaxation approach to feature selection for linear mixed effects models. Journal of Computational and Graphical Statistics, 33(1):261–275, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2231496.

Stringer:2023:FSA

[SBS23b] Alex Stringer, Patrick Brown, and Jamie Stafford. Fast, scalable approximations to posterior distributions in extended latent Gaussian models. Journal of Computational and Graphical Statistics, 32 (1):84–98, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2099403.

Sardy:2000:BCR

[SBT00] Sylvain Sardy, Andrew G. Bruce, and Paul Tseng. Block coordinate relaxation methods for nonparametric wavelet denoising. Journal of Computational and Graphical Statistics, 9(2):361– 379, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2000.10474885.

Salibian-Barrera:2008:FER

[SBWZ08] Matias Salibian-Barrera, Gert Willems, and Ruben Zamar. The fast-τ estimator for regression. Journal of Computational and Graphical Statistics, 17 (3):659–682, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Salibian-Barrera:2006:FAR

[SBY06] Matías Salibian-Barrera and Víctor J. Yohai. A fast algorithm for S-regression estimates.

Journal of Computational and Graphical Statistics, 15(2):414–427, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sharma:2013:CGI

[SBZ13] Dhruv B. Sharma, Howard D.

Bondell, and Hao Helen Zhang. Consistent group identification and variable selection in regression with correlated predictors. Journal of Computational and Graphical Statistics, 22(2):319–340, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schervish:1992:CSS

[SC92] Mark J. Schervish and Bradley P. Carlin. On the convergence of successive substitution sampling. Journal of Computational and Graphical Statistics, 1(2):111– 127, June 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10477008.

StAmant:1998:ISE

[SC98] Robert St. Amant and Paul R. Cohen. Intelligent support for exploratory data analysis. Journal of Computational and Graphical Statistics, 7(4):545–558, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/amant.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474794.

Scott:2008:FIS

[SC08] James G. Scott and Carlos M. Carvalho. Feature-inclusion stochastic search for Gaussian graphical models. Journal of Computational and Graphical Statistics, 17 (4):790–808, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stojkova:2019:IMI

[SC19] Biljana Jonoska Stojkova and David A. Campbell. Incremental mixture importance sampling with shotgun optimization. Journal of Computational and Graphical Statistics, 28(4):806–820, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stindl:2023:EAE

[SC23] Tom Stindl and Feng Chen. EM algorithm for the estimation of the RETAS model. Journal of Computational and Graphical Statistics, 33 (2):341–351, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2253293.

Shlomovich:2022:PEB

[SCAP22] Leigh Shlomovich, Edward A. K. Cohen, Niall Adams, and Lekha Patel. Parameter estimation of binned Hawkes processes. Journal of Computational and Graphical Statistics, 31(4):990– 1000, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2050247.

Swayne:1998:XID

[SCB98] Deborah F. Swayne, Dianne Cook, and Andreas Buja. XGobi: Interactive dynamic data visualization in the X Window System. Journal of Computational and Graphical Statistics, 7(1):113–130, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.research.att.com/
~andreas/jcgs/abstracts/
1998/march/swayne.html;
http://www.tandfonline.

                                        com/doi/abs/10.1080/
                                        10618600.1998.10474764.

Sauer:2022:VAD

[SCG22] Annie Sauer, Andrew Cooper, and Robert B. Gramacy. Vecchiaapproximated deep Gaussian processes for computer experiments. Journal of Computational and Graphical Statistics, 32 (3):824–837, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2129662.

Schonlau:2024:HPV

[Sch24] Matthias Schonlau. Hammock plots: Visualizing categorical and numerical variables. Journal of Computational and Graphical Statistics, 33 (4):1475–1487, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2322561.

Symanzik:2000:LAX

[SCLK^{+}00] Jürgen Symanzik, Dianne Cook, Nicholas Lewin-Koh, James J. Majure, and Inna Megretskaia. Linking ArcView^{T M} and XGobi: Insight behind the front end. Journal of Computational and Graphical Statistics, 9(3):470–490, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Symanzik.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474894.

Scott:1999:BAT

[Sco99] Steven L. Scott. Bayesian analysis of a two-state Markov modulated Poisson process. Journal of Computational and Graphical Statistics, 8(3): 662–670, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474840.

Scott:2003:E

[Sco03] David W. Scott. Editorial. Journal of Computational and Graphical Statistics, 12(4):757–758, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032751.

Santos:2003:CWC

[SD03] Reginaldo J. Santos and Álvaro R. De Pierro. A cheaper way to compute generalized crossvalidation as a stopping rule for linear stationary iterative methods. Journal of Computational and Graphical Statistics, 12(2):417–433, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s9/p417;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031815.

Slawski:2021:PLA

[SDBD21] Martin Slawski, Guoqing Diao, and Emanuel Ben-David. A pseudolikelihood approach to linear regression with partially shuffled data. Journal of Computational and Graphical Statistics, 30 (4):991–1003, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1870482.

Srivastava:2019:ADE

[SDL19] Sanvesh Srivastava, Glen DePalma, and Chuanhai Liu. An asynchronous distributed expectation maximization algorithm for massive data: The DEM algorithm. Journal of Computational and Graphical Statistics, 28 (2):233–243, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Steyn:2023:NNO

[SdWDL23] Matthys Lucas Steyn, Tertius de Wet, Bernard De Baets, and Stijn Luca. A nearest neighbor openset classifier based on excesses of distance ratios. Journal of Computational and Graphical Statistics, 32(1):319–328, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2096621.

Segal:1994:RCL

[Seg94] Mark R. Segal. Representative curves for longitudinal data via regression trees. Journal of Computational and Graphical Statistics, 3(2): 214–233, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474639.

Sevcikova:2004:SSP

[Šev04] Hana Ševčíková. Statistical simulations on parallel computers. Journal of Computational and Graphical Statistics, 13 (4):886–906, December 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sewell:2021:MBE

[Sew21] Daniel K. Sewell. Modelbased edge clustering. Journal of Computational and Graphical Statistics, 30(2):390–405, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1811104.

Sutton:2023:CCP

[SF23] Matthew Sutton and Paul Fearnhead. Concaveconvex PDMP-based sampling. Journal of Computational and Graphical Statistics, 32(4):1425– 1435, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2203735.

Simon:2013:SGL

[SFHT13] Noah Simon, Jerome Friedman, Trevor Hastie, and Robert Tibshirani. A sparse-group lasso. Journal of Computational and Graphical Statistics, 22 (2):231–245, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Seifert:2000:DAR

[SG00] Burkhardt Seifert and Theo Gasser. Data adaptive ridging in local polynomial regression. Journal of Computational and Graphical Statistics, 9(2): 338–360, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474884.

Scaccia:2003:BGC

[SG03] Luisa Scaccia and Peter J.

Green. Bayesian growth curves using normal mixtures with nonparametric weights. Journal of Computational and Graphical Statistics, 12(2):308–331, June 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n2/s4/p308;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031725.

Salvador:2005:JMS

[SG05] Manuel Salvador and Pilar Gargallo. Joint monitoring of several types of shocks in normal dynamic linear models. Journal of Computational and Graphical Statistics, 14 (3):590–609, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sun:2011:FB

[SG11] Ying Sun and Marc G. Genton. Functional boxplots. Journal of Computational and Graphical Statistics, 20(2):316–334, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shirota:2017:ABC

[SG17] Shinichiro Shirota and Alan E. Gelfand. Approximate Bayesian computation and model assessment for repulsive spatial point processes. Journal of Computational and Graphical Statistics, 26 (3):646–657, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Scott:2020:SDK

[SG20] James A. Scott and Axel Gandy. State-dependent kernel selection for conditional sampling of graphs. Journal of Computational and Graphical Statistics, 29(4):847–858, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1753529.

Sherlock:2023:EBI

[SG23] Chris Sherlock and Andrew Golightly. Exact Bayesian inference for discretely observed Markov jump processes using finite rate matrices. Journal of Computational and Graphical Statistics, 32 (1):36–48, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2093886.

Sherlock:2017:ADA

[SGH17] Chris Sherlock, Andrew Golightly, and Daniel A. Henderson. Adaptive, delayed-acceptance MCMC for targets with expensive likelihoods. Journal of Computational and Graphical Statistics, 26 (2):434–444, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Soong:1997:UCI

[SH97] W. C. Soong and Jason C.

Hsu. Using complex integration to compute multivariate normal probabilities. Journal of Computational and Graphical Statistics, 6(4):397–415, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/soong.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474750.

Smyth:2000:RFE

[SH00] Gordon K. Smyth and Douglas M. Hawkins. Robust frequency estimation using elemental sets. Journal of Computational and Graphical Statistics, 9(1):196–214, March 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Smyth.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474874.

Sen:2003:PWU

[SH03] Rituparna Sen and Mark H. Hansen. Predicting Web users’ next access based on log data. Journal of Computational and Graphical Statistics, 12(1):143–155, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s7/p143;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031275.

Shi:2006:ULR

[SH06] Tao Shi and Steve Horvath. Unsupervised learning with random forest predictors. Journal of Computational and Graphical Statistics, 15 (1):118–138, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shang:2017:GFT

[SH17] Han Lin Shang and Rob J.

Hyndman. Grouped functional time series forecasting: An application to age-specific mortality rates. Journal of Computational and Graphical Statistics, 26(2):330–343, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sankaran:2018:IVH

[SH18] Kris Sankaran and Susan Holmes. Interactive visualization of hierarchically structured data. Journal of Computational and Graphical Statistics, 27 (3):553–563, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shaby:2014:OFS

[Sha14] Benjamin A. Shaby. The open-faced sandwich adjustment for MCMC using estimating functions.

Journal of Computational and Graphical Statistics, 23(3):853–876, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sidrow:2024:VRS

[SHBC^{+}24] Evan Sidrow, Nancy Heckman, Alexandre Bouchard-Côté, Sarah M. E. Fortune, Andrew W. Trites, and Marie Auger-Méthé. Variance-reduced stochastic optimization for efficient inference of hidden Markov models. Journal of Computational and Graphical Statistics, 34 (1):222–238, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2350476.

Sargent:2000:SMC

[SHC00] Daniel J. Sargent, James S. Hodges, and Bradley P. Carlin. Structured Markov Chain Monte Carlo. Journal of Computational and Graphical Statistics, 9(2): 217–234, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474877.

Sorensen:2018:CSH

[SHFT18] Øystein Sørensen, Kristoffer Herland Hellton, Arnoldo Frigessi, and Magne Thoresen. Covariate selection in high-dimensional generalized linear models with measurement error. Journal of Computational and Graphical Statistics, 27(4):739–749, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shi:2024:NHD

[Shi24] Xiaoping Shi. Nonparametric high-dimensional multi-sample tests based on graph theory. Journal of Computational and Graphical Statistics, 33 (4):1446–1453, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2358156.

Solonen:2012:SBO

[SHL12] Antti Solonen, Heikki Haario, and Marko Laine. Simulation-based optimal design using a response variance criterion. Journal of Computational and Graphical Statistics, 21 (1):234–252, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Siems:2019:SCR

[SHL19] Tobias Siems, Marc Hellmuth, and Volkmar Liebscher. Simultaneous credible regions for multiple changepoint locations. Journal of Computational and Graphical Statistics, 28(2):290–298, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Scott-Hayward:2014:CRS

[SHMD^{+}14] L. A. S. Scott-Hayward, M. L. Mackenzie, C. R. Donovan, C. G. Walker, and E. Ashe. Complex Region Spatial Smoother (CReSS). Journal of Computational and Graphical Statistics, 23(2):340– 360, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Singer:2002:PEN

[Sin02] Hermann Singer. Parameter estimation of nonlinear stochastic differential equations: Simulated maximum likelihood versus extended Kalman filter and Itô-Taylor expansion. Journal of Computational and Graphical Statistics, 11(4):972–995, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n4/s14/p972;   http:

//www.tandfonline.com/
doi/abs/10.1198/106186002808.

Shmueli:2005:VOA

[SJ05] Galit Shmueli and Wolfgang Jank. Visualizing online auctions. Journal of Computational and Graphical Statistics, 14 (2):299–319, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schulz:2015:ARD

[SJH^{+}15] Jörn Schulz, Sungkyu Jung, Stephan Huckemann, Michael Pierrynowski, J. S. Marron, and Stephen M. Pizer. Analysis of rotational deformations from directional data. Journal of Computational and Graphical Statistics, 24 (2):539–560, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.914947.

Shi:2024:DSP

[SJM^{+}24] Haolun Shi, Shu Jiang, Da Ma, Mirza Faisal Beg, and Jiguo Cao. Dynamic survival prediction using sparse longitudinal images via multidimensional functional principal component analysis. Journal of Computational and Graphical Statistics, 33(4):1240– 1251, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2335182.

Sachs:2022:GMD

[SJSG22] Michael C. Sachs, Gustav Jonzon, Arvid Sjölander, and Erin E. Gabriel. A general method for deriving tight symbolic bounds on causal effects. Journal of Computational and Graphical Statistics, 32 (2):567–576, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2071905.

Scott:2010:NAC

[SK10] Clayton Scott and Eric Kolaczyk. Nonparametric assessment of contamination in multivariate data using generalized quantile sets and FDR. Journal of Computational and Graphical Statistics, 19 (2):439–456, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schuster:2021:MCI

[SK21] Ingmar Schuster and Ilja Klebanov. Markov chain importance sampling — a highly efficient estimator for MCMC. Journal of Computational and Graphical Statistics, 30 (2):260–268, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1826953.

Sischka:2024:SBS

[SK24a] Benjamin Sischka and Göran Kauermann. Stochastic block smooth graphon model. Journal of Computational and Graphical Statistics, 34(1):140– 154, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2374571.

Sun:2024:UEM

[SK24b] Thomas Y. Sun and Daniel R. Kowal. Ultraefficient MCMC for Bayesian longitudinal functional data analysis. Journal of Computational and Graphical Statistics, 34 (1):34–46, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2362227.

Skaug:2002:ADF

[Ska02] Hans J. Skaug. Automatic differentiation to facilitate maximum likelihood estimation in nonlinear random effects models. Journal of Computational and Graphical Statistics, 11(2):458–470, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n2/s12/p458;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002760180617.

Sang:2022:RKH

[SKK22] Peijun Sang, Adam B. Kashlak, and Linglong Kong. A reproducing kernel Hilbert space framework for functional classification. Journal of Computational and Graphical Statistics, 32(3):1000– 1008, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2138407.

Shen:1999:EBE

[SL99] Wei Shen and Thomas A. Louis. Empirical Bayes estimation via the smoothing by roughening approach. Journal of Computational and Graphical Statistics, 8(4):800–823, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/shen.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474850.

Song:2010:SLV

[SL10] Xin-Yuan Song and Zhao-Hua Lu. Semiparametric latent variable models with Bayesian P-splines. Journal of Computational and Graphical Statistics, 19(3):590–608, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Siden:2018:ECA

[SLBV18] Per Sidén, Finn Lindgren, David Bolin, and Mattias Villani. Efficient covariance approximations for large sparse precision matrices. Journal of Computational and Graphical Statistics, 27(4):898–909, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Su:2022:EMF

[SLC22] Zhihua Su, Bing Li, and Dennis Cook. Envelope model for function-on-function linear regression. Journal of Computational and Graphical Statistics, 32(4):1624–1635, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2163652.

Suh:2024:HDM

[SLH24] Namjoon Suh, Li-Hsiang Lin, and Xiaoming Huo. High-dimensional multivariate linear regression with weighted nuclear norm regularization. Journal of Computational and Graphical Statistics, 33(4):1264–1275, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2331020.

Seo:2022:BWV

[SLL22a] Beomseok Seo, Lin Lin, and Jia Li. Block-wise variable selection for clustering via latent states of mixture models. Journal of Computational and Graphical Statistics, 31 (1):138–150, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1982724.

Seo:2022:MLM

[SLL22b] Beomseok Seo, Lin Lin, and Jia Li. Mixture of linear models co-supervised by deep neural networks. Journal of Computational and Graphical Statistics, 31(4):1303–1317, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2107533.

Smith:2022:ICV

[SLM22] Michael Stanley Smith and Rubén Loaiza-Maya. Implicit copula variational inference. Journal of Computational and Graphical Statistics, 32 (3):769–781, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2119987.

Smith:2020:HDC

[SLMN20] Michael Stanley Smith, Rubén Loaiza-Maya, and David J. Nott. Highdimensional copula variational approximation through transformation. Journal of Computational and Graphical Statistics, 29(4):729–743, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1740097.

Srivastava:2016:RET

[SLR16] Radhendushka Srivastava, Ping Li, and David Ruppert. RAPTT: An exact two-sample test in high dimensions using random projections. Journal of Computational and Graphical Statistics, 25(3):954–970, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1062771.

Sass:2021:FFS

[SLR21] Danielle Sass, Bo Li, and Brian J. Reich. Flexible and fast spatial return level estimation via a spatially fused penalty. Journal of Computational and Graphical Statistics, 30 (4):1124–1142, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1938584.

Sire:2023:QRR

[SLR^{+}23] Charlie Sire, Rodolphe Le Riche, Didier Rullière, Jérémy Rohmer, Lucie Pheulpin, and Yann Richet. Quantizing rare random maps: Application to flooding visualization. Journal of Computational and Graphical Statistics, 32(4):1556– 1571, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2203764.

Solgi:2013:BSM

[SM13] Reza Solgi and Antonietta Mira. A Bayesian semiparametric multiplicative error model with an application to realized volatility. Journal of Computational and Graphical Statistics, 22(3):558–583, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sardy:2019:ETS

[SM19] Sylvain Sardy and Hatef Monajemi. Efficient threshold selection for multivariate total variation denoising. Journal of Computational and Graphical Statistics, 28 (1):23–35, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schelldorfer:2014:GAH

[SMB14] Jürg Schelldorfer, Lukas Meier, and Peter Bühlmann. GLMMLasso: An algorithm for high-dimensional generalized linear mixed models using l_{1}-penalization. Journal of Computational and Graphical Statistics, 23(2):460–477, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Smith:1995:DCA

[Smi95] S. P. Smith. Differentiation of the Cholesky algorithm. Journal of Computational and Graphical Statistics, 4(2):134– 147, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474671.

Smith:2001:LBA

[Smi01] S. P. Smith. Likelihoodbased analysis of linear state-space models using the Cholesky decomposition. Journal of Computational and Graphical Statistics, 10(2):350– 369, June 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152628338.

Smirnov:2005:CIM

[Smi05] Oleg A. Smirnov. Computation of the information matrix for models with spatial interaction on a lattice. Journal of Computational and Graphical Statistics, 14(4):910–927, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sinha:2015:MEP

[SMK15] Ankur Sinha, Pekka Malo, and Timo Kuosmanen. A multiobjective exploratory procedure for regression model selection. Journal of Computational and Graphical Statistics, 24(1):154–182, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Speakman:2015:SDA

[SMN15] Skyler Speakman, Edward Mcfowland III, and Daniel B. Neill. Scalable detection of anomalous patterns with connectivity constraints. Journal of Computational and Graphical Statistics, 24 (4):1014–1033, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.960926.

Sarkar:2014:BSD

[SMS^{+}14] Abhra Sarkar, Bani K.

Mallick, John Staudenmayer, Debdeep Pati, and Raymond J. Carroll. Bayesian semiparametric density deconvolution in the presence of conditionally heteroscedastic measurement errors. Journal of Computational and Graphical Statistics, 23 (4):1101–1125, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Smyth:2002:EAR

[Smy02] Gordon K. Smyth. An efficient algorithm for REML in heteroscedastic regression. Journal of Computational and Graphical Statistics, 11(4):836–847, December 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /lucia.ingentaselect.
                                        com/cgi-bin/linker?ini=
                                        asa&reqidx=/cw/asa/10618600/
                                        v11n4/s6/p836;   http:
                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002871.

Stuetzle:2010:GSL

[SN10] Werner Stuetzle and Rebecca Nugent. A generalized single linkage method for estimating the cluster tree of a density. Journal of Computational and Graphical Statistics, 19 (2):397–418, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Smith:2001:SSM

[SNB^{+}01] Andrew J. Smith, John Nelder, Andreas Buja, Zahid Malik, Lisa Tweedie, and Robert Spence. Sampling schemes for model visualization. Journal of Computational and Graphical Statistics, 10(3):545–554, September 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317115108.

Schiavon:2023:ASM

[SNC23] Lorenzo Schiavon, Bernardo Nipoti, and Antonio Canale. Accelerated structured matrix factorization. Journal of Computational and Graphical Statistics, 33(3):917– 927, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2301072.

Simpson:2017:IMC

[SNR17] Matthew Simpson, Jarad Niemi, and Vivekananda Roy. Interweaving Markov chain Monte Carlo strategies for efficient estimation of dynamic linear models. Journal of Computational and Graphical Statistics, 26(1):152– 159, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1105748.

Sodja:2020:DAT

[Sod20]  Cole Sodja.  Detect-

ing anomalous time series by GAMLSS-Akaike-weights-scoring. Journal of Computational and Graphical Statistics, 33 (1):??, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2020.1868306.

Somerville:1998:NCM

[Som98] Paul N. Somerville. Numerical computation of multivariate normal and multivariate-t probabilities over convex regions. Journal of Computational and Graphical Statistics, 7(4):529–544, December 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts98/somer.
                                        html;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1998.10474793.

Sherwood:2023:UMP

[SP23] Ben Sherwood and Bradley S. Price. On the use of minimum penalties in statistical learning. Journal of Computational and Graphical Statistics, 33 (1):138–151, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2210174.

Spall:2005:MCC

[Spa05] James C. Spall. Monte Carlo computation of the Fisher information matrix in nonstandard settings. Journal of Computational and Graphical Statistics, 14(4):889–909, December 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shen:2022:CST

[SPV22] Cencheng Shen, Sambit Panda, and Joshua T. Vogelstein. The chi-square test of distance correlation. Journal of Computational and Graphical Statistics, 31(1):254– 262, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1938585.

Solka:1995:VTS

[SPW95] Jeffrey L. Solka, Wendy L.

Poston, and Edward J. Wegman. A visualization technique for studying the iterative estimation of mixture densities. Journal of Computational and Graphical Statistics, 4(3): 180–198, September 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474676.

Sun:2007:JDR

[SQ07] Jingran Sun and Peihua Qiu. Jump detection in regression surfaces using both first-order and second-order derivatives. Journal of Computational and Graphical Statistics, 16(2):289–311, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shan:2020:JET

[SQCK20] Liang Shan, Zhilei Qiao, Lulu Cheng, and Inyoung Kim. Joint estimation of the two-level Gaussian graphical models across multiple classes. Journal of Computational and Graphical Statistics, 29 (3):562–579, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1694522.

Sebastiani:2000:BIM

[SR00] Paola Sebastiani and Marco Ramoni. Bayesian inference with missing data using bound and collapse. Journal of Computational and Graphical Statistics, 9(4):779–800, December 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Sebastiani.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474913.

Shaby:2012:TCB

[SR12] Benjamin Shaby and David Ruppert. Tapered covariance: Bayesian estimation and asymptotics. Journal of Computational and Graphical Statistics, 21(2):433–452, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shin:2023:MLE

[SR23] Yongyun Shin and Stephen W.

Raudenbush. Maximum likelihood estimation of hierarchical linear models from incomplete data: Random coefficients, statistical interactions, and measurement error. Journal of Computational and Graphical Statistics, 33 (1):112–125, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2023.2234414.

Steele:2006:CNC

[SRE06] Russell J. Steele, Adrian E. Raftery, and Mary J. Emond. Computing normalizing constants for finite mixture models via incremental mixture importance sampling (IMIS). Journal of Computational and Graphical Statistics, 15(3):712–734, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sutherland:2000:OVT

[SRL^{+}00] Peter Sutherland, Anthony Rossini, Thomas Lumley, Nicholas Lewin-Koh, Julie Dickerson, Zach Cox, and Dianne Cook. Orca: a visualization toolkit for highdimensional data. Journal of Computational and Graphical Statistics, 9(3):509–529, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Sutherland.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474896.

Shane:2001:RAC

[SS01] Karen V. Shane and Jeffrey S. Simonoff. A robust approach to categorical data analysis. Journal of Computational and Graphical Statistics, 10(1):135–157, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts01/Shane.
htm;   http://www.tandfonline.
com/doi/abs/10.1198/
10618600152418683.

Sinharay:2005:ECM

[SS05] Sandip Sinharay and Hal S. Stern. An empirical comparison of methods for computing Bayes factors in generalized linear mixed models. Journal of Computational and Graphical Statistics, 14 (2):415–435, June 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Smith:2006:EBM

[SS06] Daniel Smith and Michael Smith. Estimation of binary Markov random fields using Markov chain Monte Carlo. Journal of Computational and Graphical Statistics, 15 (1):207–227, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sahoo:2011:SDP

[SS11] Shyamsundar Sahoo and Debasis Sengupta. Some diagnostic plots and corrective adjustments for the proportional hazards regression model. Journal of Computational and Graphical Statistics, 20 (2):375–394, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Schuetter:2012:MSD

[SS12] Jared Schuetter and Tao Shi. Multiple sample data spectroscopic clustering of large datasets using Nyström extension. Journal of Computational and Graphical Statistics, 21(2):338–360, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sun:2016:SCE

[SS16] Ying Sun and Michael L. Stein. Statistically and computationally efficient estimating equations for large spatial datasets. Journal of Computational and Graphical Statistics, 25(1):187–208, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.975230;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.975230.

Scrucca:2019:PPB

[SS19] Luca Scrucca and Alessio Serafini. Projection pursuit based on Gaussian mixtures and evolutionary algorithms. Journal of Computational and Graphical Statistics, 28 (4):847–860, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Svahn:2022:SIC

[SS22] Caroline Svahn and Oleg Sysoev. Selective imputation of covariates in high dimensional censored data. Journal of Computational and Graphical Statistics, 31(4):1397– 1405, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2035233.

Shen:2014:FDA

[SSB^{+}14] Dan Shen, Haipeng Shen, Shankar Bhamidi, Yolanda Muñoz Maldonado, Yongdai Kim, and J. S. Marron. Functional data analysis of tree data objects. Journal of Computational and Graphical Statistics, 23 (2):418–438, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sienkiewicz:2017:SFD

[SSBW17] Ela Sienkiewicz, Dong Song, F. Jay Breidt, and Haonan Wang. Sparse functional dynamical models — a big data approach. Journal of Computational and Graphical Statistics, 26(2):319–329, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Scheipl:2015:FAM

[SSG15] Fabian Scheipl, Ana-Maria Staicu, and Sonja Greven. Functional additive mixed models. Journal of Computational and Graphical Statistics, 24 (2):477–501, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.901914.

Stocker:2023:FAM

[SSG23] Almond Stöcker, Lisa Steyer, and Sonja Greven. Functional additive models on manifolds of planar shapes and forms. Journal of Computational and Graphical Statistics, 32 (4):1600–1612, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2175687.

Shou:2016:SNH

[SSL^{+}16] Haochang Shou, Russell T. Shinohara, Han Liu, Daniel S. Reich, and Ciprian M. Crainiceanu. Soft null hypotheses: A case study of image enhancement detection in brain lesions. Journal of Computational and Graphical Statistics, 25 (2):570–588, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stroud:2017:BML

[SSL17] Jonathan R. Stroud, Michael L. Stein, and Shaun Lysen. Bayesian and maximum likelihood estimation for Gaussian processes on an incomplete lattice. Journal of Computational and Graphical Statistics, 26 (1):108–120, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1152970.

Speakman:2016:PFS

[SSMN16] Skyler Speakman, Sriram Somanchi, Edward Mcfowland III, and Daniel B. Neill. Penalized fast subset scanning. Journal of Computational and Graphical Statistics, 25 (2):382–404, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sidiropoulos:2018:SEC

[SSP^{+}18] Nikos Sidiropoulos, Sina Hadi Sohi, Thomas Lin Pedersen, Bo Torben Porse, Ole Winther, Nicolas Rapin, and Frederik Otzen Bagger. SinaPlot: An enhanced chart for simple and truthful representation of single observations over multiple classes. Journal of Computational and Graphical Statistics, 27(3):673–676, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Seidel:2007:TAN

[SŠS07] Wilfried Seidel, Hana Ševčíková, and Krunoslav Sever. Testing against nonparametric alternatives in mixture models. Journal of Computational and Graphical Statistics, 16(2):350–377, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Simonoff:1999:SAM

[ST99] Jeffrey S. Simonoff and Chih-Ling Tsai. Semiparametric and additive model selection using an improved Akaike information criterion. Journal of Computational and Graphical Statistics, 8(1): 22–40, March 1999. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/upcoming.html;
http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474799.

Sardy:2004:ARG

[ST04] Sylvain Sardy and Paul Tseng. AMlet, RAMlet, and GAMlet. Journal of Computational and Graphical Statistics, 13 (2):283–309, June 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

She:2019:IPS

[ST19] Yiyuan She and Shao Tang. Iterative proportional scaling revisited: A modern optimization perspective. Journal of Computational and Graphical Statistics, 28 (1):48–60, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stafford:1994:API

[Sta94] James E. Stafford. Automating the partition of indexes. Journal of Computational and Graphical Statistics, 3(3):249– 259, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474643.

Stafford:1995:ECC

[Sta95] James E. Stafford. Exact cumulant calculations for Pearson χ^{2} and Zelterman statistics for rway contingency tables. Journal of Computational and Graphical Statistics, 4(3):199–212, 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474677.

Stein:1992:PIT

[Ste92] Michael L. Stein. Prediction and inference for truncated spatial data. Journal of Computational and Graphical Statistics, 1(1):91–110, 1992. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474578.

Stewart:1995:GSG

[Ste95] G. W. Stewart. Gauss, statistics, and Gaussian elimination. Journal of Computational and Graphical Statistics, 4(1): 1–11, March 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474662.

Stein:2002:FES

[Ste02] Michael L. Stein. Fast and exact simulation of fractional Brownian surfaces. Journal of Computational and Graphical Statistics, 11(3):587–599, September 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s5/p587;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002466.

Stein:2012:SGR

[Ste12] Michael L. Stein. Simulation of Gaussian random fields with one derivative. Journal of Computational and Graphical Statistics, 21(1):155–173, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stein:2013:SPC

[Ste13] Michael L. Stein. Statistical properties of covariance tapers. Journal of Computational and Graphical Statistics, 22 (4):866–885, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Spearing:2023:MIP

[STIP23] Harriet Spearing, Jonathan Tawn, David Irons, and Tim Paulden. Modeling intransitivity in pairwise comparisons with application to baseball data. Journal of Computational and Graphical Statistics, 32(4):1383–1392, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2177299.

Salter-Townshend:2015:RAN

[STM15] Michael Salter-Townshend and Thomas Brendan Murphy. Role analysis in networks using mixtures of exponential random graph models. Journal of Computational and Graphical Statistics, 24 (2):520–538, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.923777.

Su:2006:TV

[STY06] Xiaogang Su, Chih-Ling Tsai, and Xin Yan. Treed variance. Journal of Computational and Graphical Statistics, 15(2):356–371, June 2006. CODEN ????

ISSN 1061-8600 (print), 1537-2715 (electronic).

Singh:2009:QSC

[STZM09] Kesar Singh, David E.

Tyler, Jingshan Zhang, and Somnath Mukherjee. Quantile scale curves. Journal of Computational and Graphical Statistics, 18(1):92–105, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Su:2015:VSS

[Su15] Xiaogang Su. Variable selection via subtle uprooting. Journal of Computational and Graphical Statistics, 24(4):1092– 1113, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.955176.

Suesse:2012:MER

[Sue12] Thomas Suesse. Marginalized exponential random graph models. Journal of Computational and Graphical Statistics, 21 (4):883–900, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sherlock:2023:AAP

[SUL23] Chris Sherlock, Szymon Urbas, and Matthew Ludkin. The apogee to apogee path sampler. Journal of Computational and Graphical Statistics, 32 (4):1436–1446, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2190784.

Sievert:2019:EGL

[SVC^{+}19] Carson Sievert, Susan VanderPlas, Jun Cai, Kevin Ferris, Faizan Uddin Fahad Khan, and Toby Dylan Hocking. Extending ggplot2 for linked and animated Web graphics. Journal of Computational and Graphical Statistics, 28(2):299–308, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Stoklosa:2018:GEE

[SW18] Jakub Stoklosa and David I. Warton. A generalized estimating equation approach to multivariate adaptive regression splines. Journal of Computational and Graphical Statistics, 27(1):245–253, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Storath:2023:SSD

[SW23] Martin Storath and Andreas Weinmann. Smoothing splines for discontinuous signals. Journal of Computational and Graphical Statistics, 33 (2):651–664, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2262000.

Suchard:2010:UGP

[SWC^{+}10] Marc A. Suchard, Quanli Wang, Cliburn Chan, Jacob Frelinger, Andrew Cron, and Mike West. Understanding GPU programming for statistical computation: Studies in massively parallel massive mixtures. Journal of Computational and Graphical Statistics, 19 (2):419–438, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Su:2004:MLR

[SWF04] Xiaogang Su, Morgan Wang, and Juanjuan Fan. Maximum likelihood regression trees. Journal of Computational and Graphical Statistics, 13 (3):586–598, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Shin:2023:GMP

[SWL23] Minsuk Shin, Shijie Wang, and Jun S. Liu. Generative multi-purpose sampler for weighted mestimation. Journal of Computational and Graphical Statistics, 33 (3):1084–1097, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2292668.

Stokes:2023:MSO

[SWX23] Zack Stokes, Weng Kee Wong, and Hongquan Xu. Metaheuristic solutions to order-of-addition design problems. Journal of Computational and Graphical Statistics, 33 (3):1006–1016, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2277878.

Shin:2014:TDS

[SWZ14] Seung Jun Shin, Yichao Wu, and Hao Helen Zhang. Two-dimensional solution surface for weighted support vector machines. Journal of Computational and Graphical Statistics, 23(2):383–402, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Srivastava:2021:DBI

[SX21] Sanvesh Srivastava and Yixiang Xu. Distributed Bayesian inference in linear mixed-effects models. Journal of Computational and Graphical Statistics, 30(3):594–611, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1869025.

Sit:2023:DCQ

[SX23] Tony Sit and Yue Xing.

Distributed censored quantile regression. Journal of Computational and Graphical Statistics, 32 (4):1685–1697, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2182310.

Schafer:2002:CSM

[SY02] Joseph L. Schafer and Recai M. Yucel. Computational strategies for multivariate linear mixedeffects models with missing values. Journal of Computational and Graphical Statistics, 11 (2):437–457, June 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n2/s11/p437;   http:

                                        //www.tandfonline.com/
                                        doi/abs/10.1198/106186002760180608.

Stramer:2007:SLD

[SY07] Osnat Stramer and Jun Yan. On simulated likelihood of discretely observed diffusion processes and comparison to closed-form approximation. Journal of Computational and Graphical Statistics, 16(3):672–691, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Sheng:2016:SDR

[SY16] Wenhui Sheng and Xiangrong Yin. Sufficient dimension reduction via distance covariance. Journal of Computational and Graphical Statistics, 25 (1):91–104, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        2015.1026601;   http:
                                        //www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1026601.

Saldana:2017:HMC

[SYF17] D. Franco Saldaña, Yi Yu, and Yang Feng. How many communities are there? Journal of Computational and Graphical Statistics, 26(1):171– 181, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1096790.

StThomas:2022:LSD

[SYL^{+}22] Brian St. Thomas, Kisung You, Lizhen Lin, Lek-Heng Lim, and Sayan Mukherjee. Learning subspaces of different dimensions. Journal of Computational and Graphical Statistics, 31(2):337– 350, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000420.

Salomone:2024:SVA

[SYNK24] Robert Salomone, Xuejun Yu, David J. Nott, and Robert Kohn. Structured variational approximations with skew normal decomposable graphical models and implicit copulas. Journal of Computational and Graphical Statistics, 33(4):1329– 1338, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2319159.

Shi:2022:TDF

[SYW^{+}22] Haolun Shi, Yuping Yang, Liangliang Wang, Da Ma, Mirza Faisal Beg, Jian Pei, and Jiguo Cao. Twodimensional functional principal component analysis for image feature extraction. Journal of Computational and Graphical Statistics, 31(4):1127– 1140, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2035738.

Sandri:2008:BCA

[SZ08] Marco Sandri and Paola Zuccolotto. A bias correction algorithm for the Gini variable importance measure in classification trees. Journal of Computational and Graphical Statistics, 17(3):611–628, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tjelmeland:2012:EAR

[TA12] Håkon Tjelmeland and Haakon Michael Austad. Exact and approximate recursive calculations for binary Markov random fields defined on graphs. Journal of Computational and Graphical Statistics, 21(3):758–780, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Taddy:2017:OSE

[Tad17] Matt Taddy. One-step estimator paths for concave regularization. Journal of Computational and Graphical Statistics, 26 (3):525–536, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2006:MCI

[Tan06] Zhiqiang Tan. Monte Carlo integration with acceptance-rejection. Journal of Computational and Graphical Statistics, 15 (3):735–752, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2015:RMC

[Tan15] Zhiqiang Tan. Resampling Markov chain Monte Carlo algorithms: Basic analysis and empirical comparisons. Journal of Computational and Graphical Statistics, 24 (2):328–356, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.897625.

Tan:2017:OAM

[Tan17] Zhiqiang Tan. Optimally adjusted mixture sampling and locally weighted histogram analysis. Journal of Computational and Graphical Statistics, 26 (1):54–65, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1113975.

Tarpey:1999:SCA

[Tar99] Thaddeus Tarpey. Selfconsistency algorithms. Journal of Computational and Graphical Statistics, 8(4):889–905, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/tarpe.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474854.

Tang:2017:STS

[TAS^{+}17] Minh Tang, Avanti Athreya, Daniel L. Sussman, Vince Lyzinski, Youngser Park, and Carey E. Priebe. A semiparametric twosample hypothesis testing problem for random graphs. Journal of Computational and Graphical Statistics, 26(2):344–354, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tharmaratnam:2010:EPR

[TCCSB10] Kukatharmini Tharmaratnam, Gerda Claeskens, Christophe Croux, and Matias Salibián-Barrera. S-estimation for penalized regression splines. Journal of Computational and Graphical Statistics, 19 (3):609–625, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2015:HIS

[TDH15] Aixin Tan, Hani Doss, and James P. Hobert. Honest importance sampling with multiple Markov chains. Journal of Computational and Graphical Statistics, 24(3):792–826, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Talhouk:2012:EBI

[TDM12] Aline Talhouk, Arnaud Doucet, and Kevin Murphy. Efficient Bayesian inference for multivariate probit models with sparse inverse correlation matrices. Journal of Computational and Graphical Statistics, 21(3):739–757, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tickle:2020:PCC

[TEFH20] S. O. Tickle, I. A. Eckley, P. Fearnhead, and K. Haynes. Parallelization of a common changepoint detection method. Journal of Computational and Graphical Statistics, 29(1):149–161, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1647216.

Tak:2019:RAI

[TEG19] Hyungsuk Tak, Justin A. Ellis, and Sujit K. Ghosh. Robust and accurate inference via a mixture of Gaussian and Student’s t errors. Journal of Computational and Graphical Statistics, 28(2):415–426, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2020:BVI

[TF20a] Linda S. L. Tan and Nial Friel. Bayesian variational inference for exponential random graph models. Journal of Computational and Graphical Statistics, 29(4):910–928, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1740714.

Tibshirani:2020:PL

[TF20b] Robert Tibshirani and Jerome Friedman. A pliable lasso. Journal of Computational and Graphical Statistics, 29 (1):215–225, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1648271.

Touloupou:2020:SBI

[TFS20] Panayiota Touloupou, Bärbel Finkenstädt, and Simon E. F. Spencer. Scalable Bayesian inference for coupled hidden Markov and semi-Markov models. Journal of Computational and Graphical Statistics, 29 (2):238–249, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1654880.

Thomas:2009:EJT

[TG09] Alun Thomas and Peter J. Green. Enumerating the junction trees of a decomposable graph. Journal of Computational and Graphical Statistics, 18(4):930–940, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tutz:2010:FES

[TG10] Gerhard Tutz and Jan Gertheiss. Feature extraction in signal regression: a boosting technique for functional data regression. Journal of Computational and Graphical Statistics, 19(1):154–174, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tutz:2013:LBB

[TG13] Gerhard Tutz and Andreas Groll. Likelihoodbased boosting in binary and ordinal random effects models. Journal of Computational and Graphical Statistics, 22 (2):356–378, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tibbits:2014:AFS

[TGHL14] Matthew M. Tibbits, Chris Groendyke, Murali Haran, and John C. Liechty. Automated factor slice sampling. Journal of Computational and Graphical Statistics, 23 (2):543–563, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tran:2014:CTE

[TGM^{+}14] Minh-Ngoc Tran, Paolo Giordani, Xiuyan Mun, Robert Kohn, and Michael K. Pitt. Copula-type estimators for flexible multivariate density modeling using mixtures. Journal of Computational and Graphical Statistics, 23 (4):1163–1178, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Therneau:2003:PSM

[TGP03] Terry M. Therneau, Patricia M. Grambsch, and V. Shane Pankratz. Penalized survival models and frailty. Journal of Computational and Graphical Statistics, 12(1):156–175, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s8/p156;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031365.

Tomarchio:2022:MMV

[TGPM22] Salvatore D. Tomarchio, Michael P. B. Gallaugher, Antonio Punzo, and Paul D. McNicholas. Mixtures of matrix-variate contaminated normal distributions. Journal of Computational and Graphical Statistics, 31 (2):413–421, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1999825.

Tan:2009:BGS

[TH09] Aixin Tan and James P. Hobert. Block Gibbs sampling for Bayesian random effects models with improper priors: Convergence and regeneration. Journal of Computational and Graphical Statistics, 18(4):861–878, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tuzhilina:2023:SMP

[THM^{+}23] Elena Tuzhilina, Trevor J. Hastie, Daniel J. Mc-Donald, J. Kenneth Tay, and Robert Tibshirani. Smooth multi-period forecasting with application to prediction of COVID-19 cases. Journal of Computational and Graphical Statistics, 33(3):955– 967, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2285337.

Thomas:1993:ACM

[Tho93] Neal Thomas. Asymptotic corrections for multivariate posterior moments with factored likelihood functions. Journal of Computational and Graphical Statistics, 2(3): 309–322, September 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474614.

Talagala:2020:ADS

[THSM^{+}20] Priyanga Dilini Talagala, Rob J. Hyndman, Kate Smith-Miles, Sevvandi Kandanaarachchi, and Mario A. Muñoz. Anomaly detection in streaming nonstationary temporal data. Journal of Computational and Graphical Statistics, 29 (1):13–27, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1617160.

Talagala:2021:ADH

[THSM21] Priyanga Dilini Talagala, Rob J. Hyndman, and Kate Smith-Miles. Anomaly detection in high-dimensional data. Journal of Computational and Graphical Statistics, 30(2):360–374, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2020.1807997.

Tierney:1996:RDF

[Tie96] Luke Tierney. Recent developments and future directions in LispStat. Journal of Computational and Graphical Statistics, 5(3):250–262, 1996. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1996.10474710.

Tierney:2000:CLS

[Tie00] Luke Tierney. Connecting Lisp-Stat to COM. Journal of Computational and Graphical Statistics, 9(3):459–469, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/abstracts00/Tierney.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.2000.10474893.

Tibshirani:1999:MSB

[TK99] Robert Tibshirani and Keith Knight. Model search by bootstrap “bumping”. Journal of Computational and Graphical Statistics, 8(4):671–686, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/tibshir.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474842.

Taylor:2023:FBR

[TKB23] Ian Taylor, Andee Kaplan, and Brenda Betancourt. Fast Bayesian record linkage for streaming data contexts. Journal of Computational and Graphical Statistics, 33 (3):833–844, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2283571.

Tibshirani:1992:SBD

[TL92] Robert Tibshirani and Michael LeBlanc. A strategy for binary description and classification. Journal of Computational and Graphical Statistics, 1(1): 3–20, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474573.

Theus:1999:VLM

[TL99] Martin Theus and Stephan R. W. Lauer. Visualizing loglinear models. Journal of Computational and Graphical Statistics, 8(3):396–412, September 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

                                        www.amstat.org/publications/
                                        jcgs/upcoming.html;
                                        http://www.tandfonline.
                                        com/doi/abs/10.1080/
                                        10618600.1999.10474821.

Tutz:2007:GSM

[TL07] Gerhard Tutz and Florian Leitenstorfer. Generalized smooth monotonic regression in additive modeling. Journal of Computational and Graphical Statistics, 16(1):165–188, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tran:2023:FSL

[TLLJ23] Lam Tran, Gen Li, Lan Luo, and Hui Jiang. A fast solution to the lasso problem with equality constraints. Journal of Computational and Graphical Statistics, 33 (3):804–813, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2277877.

Thind:2023:DLF

[TMC23] Barinder Thind, Kevin Multani, and Jiguo Cao. Deep learning with functional inputs. Journal of Computational and Graphical Statistics, 32 (1):171–180, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2097914.

Thompson:2020:CMV

[TMMB20] Geoffrey Z. Thompson, Ranjan Maitra, William Q. Meeker, and Ashraf F. Bastawros. Classification with the matrix-variate-t distribution. Journal of Computational and Graphical Statistics, 29 (3):668–674, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1696208.

Tak:2018:RAM

[TMvD18] Hyungsuk Tak, Xiao-Li Meng, and David A. van Dyk. A repellingattracting Metropolis algorithm for multimodality. Journal of Computational and Graphical Statistics, 27(3):479–490, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2014:VAM

[TN14a] Siew Li Tan and David J.

Nott. Variational approximation for mixtures of linear mixed models. Journal of Computational and Graphical Statistics, 23(2):564–585, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tyssedal:2014:GAA

[TN14b] John Tyssedal and Ranveig Niemi. Graphical aids for the analysis of two-level nonregular designs. Journal of Computational and Graphical Statistics, 23(3):678–699, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tran:2017:VBI

[TNK17] Minh-Ngoc Tran, David J. Nott, and Robert Kohn. Variational Bayes with intractable likelihood. Journal of Computational and Graphical Statistics, 26 (4):873–882, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tran:2016:PVB

[TNKK16] Minh-Ngoc Tran, David J. Nott, Anthony Y. C. Kuk, and Robert Kohn. Parallel variational Bayes for large datasets with an application to generalized linear mixed models. Journal of Computational and Graphical Statistics, 25(2):626–646, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tran:2020:BDN

[TNNK20] M.-N. Tran, N. Nguyen, D. Nott, and R. Kohn. Bayesian deep net GLM and GLMM. Journal of Computational and Graphical Statistics, 29 (1):97–113, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1637747.

Tokdar:2007:TFI

[Tok07] Surya T. Tokdar. Towards a faster implementation of density estimation with logistic Gaussian process priors. Journal of Computational and Graphical Statistics, 16(3):633–655, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2019:BIO

[TPA19] Ban Kheng Tan, Anastasios Panagiotelis, and George Athanasopoulos. Bayesian inference for the one-factor copula model. Journal of Computational and Graphical Statistics, 28(1):155–173, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tang:2016:MML

[TQ16] Xiwei Tang and Annie Qu. Mixture modeling for longitudinal data. Journal of Computational and Graphical Statistics, 25 (4):1117–1137, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Thouvenin:2023:DBS

[TRC23] Pierre-Antoine Thouvenin, Audrey Repetti, and Pierre Chainais. A distributed block-split Gibbs sampler with hypergraph structure for highdimensional inverse problems. Journal of Computational and Graphical Statistics, 33(3):814– 832, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2282501.

Tremmel:1993:CSM

[Tre93] Lothar Tremmel. Comment on “A Model for Studying Display Methods of Statistical Graphics”. Journal of Computational and Graphical Statistics, 2(4):345– 347, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/abs/10.1080/10618600.
1993.10474617.  See

[Cle93a].

Tremmel:1995:VSP

[Tre95] Lothar Tremmel. The visual separability of plotting symbols in scatterplots. Journal of Computational and Graphical Statistics, 4(2):101– 112, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474669.

Trendafilov:2006:DSA

[Tre06] Nickolay T. Trendafilov.

The dynamical system approach to multivariate data analysis. Journal of Computational and Graphical Statistics, 15 (3):628–650, September 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Trosset:2005:VC

[Tro05] Michael W. Trosset. Visualizing correlation. Journal of Computational and Graphical Statistics, 14 (1):1–19, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Thall:1997:VSR

[TRS97] Peter F. Thall, Kathy E.

Russell, and Richard M.

Simon. Variable selection in regression via repeated data splitting. Journal of Computational and Graphical Statistics, 6(4):416–434, December 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/thall.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474751.

Taylor-Rodriguez:2016:BVS

[TRWB16] Daniel Taylor-Rodriguez, Andrew Womack, and Nikolay Bliznyuk. Bayesian variable selection on model spaces constrained by heredity conditions. Journal of Computational and Graphical Statistics, 25(2):515–535, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tutz:2013:VCR

[TS13] Gerhard Tutz and Gunther Schauberger. Visualization of categorical response models: From data glyphs to parameter glyphs. Journal of Computational and Graphical Statistics, 22(1):156–177, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Taheri:2022:SAL

[TS22] Mohsen Taheri and Jörn Schulz. Statistical analysis of locally parameterized shapes. Journal of Computational and Graphical Statistics, 32 (2):658–670, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2116445.

Thall:1992:TBV

[TSG92] Peter F. Thall, Richard Simon, and David A. Grier. Test-based variable selection via crossvalidation. Journal of Computational and Graphical Statistics, 1(1): 41–61, 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1992.10474575.

Thorarinsdottir:2016:ACH

[TSH16] Thordis L. Thorarinsdottir, Michael Scheuerer, and Christopher Heinz. Assessing the calibration of high-dimensional ensemble forecasts using rank histograms. Journal of Computational and Graphical Statistics, 25 (1):105–122, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.977447;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.977447.

Tsionas:2002:BIN

[Tsi02] Efthymios G. Tsionas. Bayesian inference in the noncentral Student-t model. Journal of Computational and Graphical Statistics, 11(1):208– 221, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s11/p208;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375695.

Tu:1994:ADC

[Tu94] Xin M. Tu. Analysis of data with censored initiating and terminating times: a missingdata approach. Journal of Computational and Graphical Statistics, 3(1): 97–112, 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1994.10474632.

Trendafilov:2011:EFA

[TU11] Nickolay T. Trendafilov and Steffen Unkel. Exploratory factor analysis of data matrices with more variables than observations. Journal of Computational and Graphical Statistics, 20 (4):874–891, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tuchler:2008:BVS

[Tüc08] Regina Tüchler. Bayesian variable selection for logistic models using auxiliary mixture sampling. Journal of Computational and Graphical Statistics, 17(1):76–94, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tukey:1993:GCS

[Tuk93a] John W. Tukey. Graphic comparisons of several linked aspects: alternatives and suggested principles. With comments by Howard Wainer, Richard A. Becker and William S. Cleveland and a rejoinder by the author. Journal of Computational and Graphical Statistics, 2(1): 1–33, March 1993. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1080/10618600.
                                        1993.10474595. See com-

ments [Wai93, BC93] and

rejoinder [Tuk93b].

Tukey:1993:R

[Tuk93b] John W. Tukey. Rejoinder. Journal of Computational and Graphical Statistics, 2(1):49, March 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1993.10474598. Reply

to comments on Graphic comparisons of several linked aspects: Alternatives and suggested principles. See [Tuk93a].

Thompson:2024:GSS

[TV24] Ryan Thompson and Farshid Vahid. Group selection and shrinkage: Structured sparsity for semiparametric additive models. Journal of Computational and Graphical Statistics, 33(4):1286– 1297, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2330517.

Takagishi:2022:VCS

[TvdV22] Mariko Takagishi and Michel van de Velden.

Visualizing class specific heterogeneous tendencies in categorical data. Journal of Computational and Graphical Statistics, 31 (3):790–801, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2035737.

Tansey:2022:HRT

[TVZ^{+}22] Wesley Tansey, Victor Veitch, Haoran Zhang, Raul Rabadan, and David M. Blei. The holdout randomization test for feature selection in black box models. Journal of Computational and Graphical Statistics, 31(1):151– 162, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923520.

Turlach:1996:FCA

[TW96] B. A. Turlach and M. P.

Wand. Fast computation of auxiliary quantities in local polynomial regression. Journal of Computational and Graphical Statistics, 5(4):337–350, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/abs/10.1080/10618600.
                                        1996.10474716.

Tibshirani:2005:CVP

[TW05] Robert Tibshirani and Guenther Walther. Cluster validation by prediction strength. Journal of Computational and Graphical Statistics, 14 (3):511–528, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tan:2014:SBT

[TW14] Kean Ming Tan and Daniela M. Witten. Sparse biclustering of transposable data. Journal of Computational and Graphical Statistics, 23 (4):985–1008, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tak:2020:DTA

[TYG^{+}20] Hyungsuk Tak, Kisung You, Sujit K. Ghosh, Bingyue Su, and Joseph Kelly. Data transforming augmentation for heteroscedastic models. Journal of Computational and Graphical Statistics, 29(3):659–667, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1704295.

Tang:2006:MPS

[TZ06] Yongqiang Tang and Hao Helen Zhang. Multiclass proximal support vector machines. Journal of Computational and Graphical Statistics, 15 (2):339–355, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Tang:2012:SSC

[TZ12] Liansheng Larry Tang and Xiao-Hua Zhou. A semiparametric separation curve approach for comparing correlated ROC data from multiple markers. Journal of Computational and Graphical Statistics, 21(3):662–676, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Unwin:1996:IGD

[UHHS96] Antony Unwin, George Hawkins, Heike Hofmann, and Bernd Siegl. Interactive graphics for data sets with missing values — MANET. Journal of Computational and Graphical Statistics, 5(2): 113–122, 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474700.

Umlauf:2018:BBA

[UKZ18] Nikolaus Umlauf, Nadja Klein, and Achim Zeileis. BAMLSS: Bayesian additive models for location, scale, and shape (and beyond). Journal of Computational and Graphical Statistics, 27(3):612–627, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Unwin:2019:MOO

[Unw19] Antony Unwin. Multivariate outliers and the O3 plot. Journal of Computational and Graphical Statistics, 28(3):635–643, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Ushakova:2022:MMC

[UTK22] Anastasia Ushakova, Simon A. Taylor, and Rebecca Killick. Micromacro changepoint inference for periodic data sequences. Journal of Computational and Graphical Statistics, 32(2):684– 695, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2104288.

Unwin:2018:EG

[UVM18] Antony Unwin and Pedro Valero-Mora. Ensemble graphics. Journal of Computational and Graphical Statistics, 27(1):157–165, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Valk:2021:SIH

[VC21] Marcio Valk and Gabriela Bettella Cybis. U-statistical inference for hierarchical clustering. Journal of Computational and Graphical Statistics, 30 (1):133–143, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1796398.

vanDyk:2000:FME

[vD00] David A. van Dyk. Fitting mixed-effects models using efficient EMtype algorithms. Journal of Computational and Graphical Statistics, 9(1): 78–98, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.amstat.org/publications/
jcgs/abstracts00/VanDyk.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474867.

vanDyk:2008:E

[vD08] David van Dyk. Editorial. Journal of Computational and Graphical Statistics, 17(2):269, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

vandenBoom:2022:WWP

[vdBBD22] Willem van den Boom, Alexandros Beskos, and Maria De Iorio. The G-Wishart weighted proposal algorithm: Efficient posterior computation for Gaussian graphical models. Journal of Computational and Graphical Statistics, 31(4):1215– 1224, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2050250.

Vono:2021:AED

[VDC21] Maxime Vono, Nicolas Dobigeon, and Pierre Chainais. Asymptotically exact data augmentation: Models, properties, and algorithms. Journal of Computational and Graphical Statistics, 30(2):335–348, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1826954.

vanDelft:2019:DAE

[vDE19] Anne van Delft and Michael Eichler. Dataadaptive estimation of time-varying spectral densities. Journal of Computational and Graphical Statistics, 28(2):244–255, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

vanDyk:2001:ADA

[vDM01a] David A. van Dyk and Xiao-Li Meng. The art of data augmentation. Journal of Computational and Graphical Statistics, 10 (1):1–50, 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418584.

See discussion [Lev01, Hob01a, Hig01, Liu01, HJT01, WZ01, MG01] and rejoinder [vDM01b].

vanDyk:2001:RAD

[vDM01b] David A. van Dyk and Xiao-Li Meng. Rejoinder: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):98–111, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418782.

vandeWiel:2021:FCV

[vdWvNR21] Mark A. van de Wiel, Mirrelijn M. van Nee, and Armin Rauschenberger. Fast crossvalidation for multipenalty high-dimensional ridge regression. Journal of Computational and Graphical Statistics, 30 (4):835–847, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1904962.

Velilla:2008:MDR

[Vel08]  Santiago Velilla.  A

method for dimension reduction in quadratic classification problems. Journal of Computational and Graphical Statistics, 17 (3):572–589, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Vivar:2009:SMG

[VF09] Juan C. Vivar and Marco A. R. Ferreira. Spatiotemporal models for Gaussian areal data. Journal of Computational and Graphical Statistics, 18 (3):658–674, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Vansteelandt:2001:ASG

[VG01] Stijn Vansteelandt and Els Goetghebeur. Analyzing the sensitivity of generalized linear models to incomplete outcomes via the IDE algorithm. Journal of Computational and Graphical Statistics, 10(4):656–672, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/106186001317243386.

VanderPlas:2023:PGP

[VGUH23] Susan VanderPlas, Yawei Ge, Antony Unwin, and Heike Hofmann. Penguins go parallel: a grammar of graphics framework for generalized parallel coordinate plots. Journal of Computational and Graphical Statistics, 32 (4):1572–1587, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2195462.

VanderPlas:2015:SSI

[VH15] Susan VanderPlas and Heike Hofmann. Signs of the sine illusion — why we need to care. Journal of Computational and Graphical Statistics, 24 (4):1170–1190, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.951547.

VanderPlas:2017:CBT

[VH17] Susan VanderPlas and Heike Hofmann. Clusters beat trend!? testing feature hierarchy in statistical graphics. Journal of Computational and Graphical Statistics, 26 (2):231–242, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Vargovich:2022:LTD

[VH22] Joseph Vargovich and Toby Dylan Hocking. Linear time dynamic programming for computing breakpoints in the regularization path of models selected from a finite set. Journal of Computational and Graphical Statistics, 31(2):313–323, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2000422.

Vettori:2020:BMA

[VHSG20] Sabrina Vettori, Raphaël Huser, Johan Segers, and Marc G. Genton. Bayesian model averaging over tree-based dependence structures for multivariate extremes. Journal of Computational and Graphical Statistics, 29 (1):174–190, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1647847.

Viele:2000:PSP

[Vie00] Kert Viele. Posterior simulation with priors specified on functionals. Journal of Computational and Graphical Statistics, 9(2): 235–248, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474878.

VanDyk:2015:MHW

[VJ15] David A. Van Dyk and Xiyun Jiao. Metropolis– Hastings within partially collapsed Gibbs samplers. Journal of Computational and Graphical Statistics, 24(2):301–327, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.930041.

Vaida:2009:FIN

[VL09] Florin Vaida and Lin Liu. Fast implementation for normal mixed effects models with censored response. Journal of Computational and Graphical Statistics, 18(4):797–817, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

vanDyk:1997:OGC

[vlM97] David A. van Dyk and Xiao li Meng. On the orderings and groupings of conditional maximizations within ECM-Type algorithms. Journal of Computational and Graphical Statistics, 6(2): 202–223, June 1997. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.amstat.org/publications/
                                        jcgs/abstracts/june/
                                        vanDyk.html;   http:/
                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1997.10474737.

Virta:2018:JTV

[VLNO18] Joni Virta, Bing Li, Klaus Nordhausen, and Hannu Oja. JADE for tensor-valued observations. Journal of Computational and Graphical Statistics, 27(3):628–637, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Vallejos:2004:RAR

[VM04] Ronny O. Vallejos and Tomislav J. Mardesic. A recursive algorithm to restore images based on robust estimation of NSHP autoregressive models.

Journal of Computational and Graphical Statistics, 13(3):674–682, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Vatter:2018:GAM

[VN18] Thibault Vatter and Thomas Nagler. Generalized additive models for pair-copula constructions. Journal of Computational and Graphical Statistics, 27(4):715–727, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

vanNee:2022:FML

[vNvdBvdW22] Mirrelijn M. van Nee, Tim van de Brug, and Mark A. van de Wiel. Fast marginal likelihood estimation of penalties for group-adaptive elastic net. Journal of Computational and Graphical Statistics, 32(3):950– 960, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2128809.

VanderPlas:2019:FRR

[VRH19] Susan VanderPlas, Goluch C.

Ryan, and Heike Hofmann. Framed! Reproducing and revisiting 150year-old charts. Journal of Computational and Graphical Statistics, 28 (3):620–634, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

VanDerwerken:2017:MJC

[VS17] Douglas VanDerwerken and Scott C. Schmidler. Monitoring joint convergence of MCMC samplers. Journal of Computational and Graphical Statistics, 26(3):558–568, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Valz:1994:EIK

[VT94] Paul D. Valz and Mary E. Thompson. Exact inference for Kendall’s S and Spearman’s ρ with extension to Fisher’s exact test in r × c contingency tables. Journal of Computational and Graphical Statistics, 3(4):459–472, December 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474658.

vanWieringen:2019:GRE

[vW19] Wessel N. van Wieringen. The generalized ridge estimator of the inverse covariance matrix. Journal of Computational and Graphical Statistics, 28 (4):932–942, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

vanWieringen:2022:SLR

[vWB22] Wessel N. van Wieringen and Harald Binder. Sequential learning of regression models by penalized estimation. Journal of Computational and Graphical Statistics, 31 (3):877–886, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2035231.

Wainer:1993:GMC

[Wai93] Howard Wainer. Graphing multiple comparisons: Some comments on Tukey. Journal of Computational and Graphical Statistics, 2(1):35– 40, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474596.  See

[Tuk93a].

Wainer:2011:CWT

[Wai11] Howard Wainer. Comment: Why Tables Are Really Much Better Than Graphs. Journal of Computational and Graphical Statistics, 20(1):8–15, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Gel11b].

Wand:1994:FCM

[Wan94] M. P. Wand. Fast computation of multivariate kernel estimators. Journal of Computational and Graphical Statistics, 3(4): 433–445, December 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474656.

Wang:2008:BFK

[Wan08] Xiao Wang. Bayesian free-knot monotone cubic spline regression. Journal of Computational and Graphical Statistics, 17 (2):373–387, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2013:BGM

[Wan13] Junhui Wang. Boosting the generalized margin in cost-sensitive multiclass classification. Journal of Computational and Graphical Statistics, 22 (1):178–192, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2018:QMN

[Wan18] Zhu Wang. Quadratic majorization for nonconvex loss with applications to the boosting algorithm. Journal of Computational and Graphical Statistics, 27(3):491–502, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2020:GAI

[Wan20]  Tao Wang.  Graph-

assisted inverse regression for count data and its application to sequencing data. Journal of Computational and Graphical Statistics, 29(3):444– 454, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1705309.

Wang:2023:MCN

[Wan23] Wan-Lun Wang. Multivariate contaminated normal censored regression model: Properties and maximum likelihood inference. Journal of Computational and Graphical Statistics, 32(4):1671– 1684, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2023.2184375.

Walther:2022:CBL

[WASB22] Guenther Walther, Alnur Ali, Xinyue Shen, and Stephen Boyd. Confidence bands for a logconcave density. Journal of Computational and Graphical Statistics, 31 (4):1426–1438, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2076688.

Watnik:2011:ECS

[Wat11] Mitchell Watnik. Early computational statistics. Journal of Computational and Graphical Statistics, 20(4):811–817, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wallin:2018:EAM

[WB18a] Jonas Wallin and David Bolin. Efficient adaptive MCMC through precision estimation. Journal of Computational and Graphical Statistics, 27 (4):887–897, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Willis:2018:UPT

[WB18b] Amy Willis and Rayna Bell. Uncertainty in phylogenetic tree estimates. Journal of Computational and Graphical Statistics, 27(3):542–552, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2020:LFM

[WBQ20] Yu Wang, Xuan Bi, and Annie Qu. A logistic factorization model for recommender systems with multinomial responses. Journal of Computational and Graphical Statistics, 29(2):396–404, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1665535.

Whitaker:2021:LRM

[WBS21] T. Whitaker, B. Beranger, and S. A. Sisson. Logistic regression models for aggregated data. Journal of Computational and Graphical Statistics, 30 (4):1049–1067, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1895816.

Warne:2022:RBI

[WBS22] David J. Warne, Ruth E.

Baker, and Matthew J. Simpson. Rapid Bayesian inference for expensive stochastic models. Journal of Computational and Graphical Statistics, 31 (2):512–528, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2000419.

Wycoff:2021:SLA

[WBW21] Nathan Wycoff, Mickaël Binois, and Stefan M. Wild. Sequential learning of active subspaces. Journal of Computational and Graphical Statistics, 30 (4):1224–1237, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1874962.

Wood:1994:SSG

[WC94] Andrew T. A. Wood and Grace Chan. Simulation of stationary Gaussian processes in [0, 1]^{d}. Journal of Computational and Graphical Statistics, 3(4): 409–432, December 1994. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1994.10474655.

Wathen:2006:IBI

[WC06] J. Kyle Wathen and J. Andrés Christen. Implementation of backward induction for sequentially adaptive clinical trials. Journal of Computational and Graphical Statistics, 15(2):398–413, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2023:FNL

[WC23] Haixu Wang and Jiguo Cao. Functional nonlinear learning. Journal of Computational and Graphical Statistics, 33 (1):181–191, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2233581.

Whitney:2003:TRA

[WCDS03] Paul Whitney, Dennis Cox, Don Daly, and J. McLean Sloughter. Toward the routine analysis of diverse data types. Journal of Computational and Graphical Statistics, 12(4):915–926, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032535_
a.

Wang:2020:CBG

[WCH20a] Earo Wang, Dianne Cook, and Rob J. Hyndman.

Calendar-based graphics for visualizing people’s daily schedules. Journal of Computational and Graphical Statistics, 29 (3):490–502, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1715226.

Wang:2020:NTD

[WCH20b] Earo Wang, Dianne Cook, and Rob J. Hyndman. A new tidy data structure to support exploration and modeling of temporal data. Journal of Computational and Graphical Statistics, 29 (3):466–478, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1695624.

Wang:2019:PWA

[WCKL19] Yu-Bo Wang, Ming-Hui Chen, Lynn Kuo, and Paul O. Lewis. Partition weighted approach for estimating the marginal posterior density with applications. Journal of Computational and Graphical Statistics, 28 (2):334–349, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Woody:2021:MIT

[WCM21] Spencer Woody, Carlos M. Carvalho, and Jared S. Murray. Model interpretation through lower-dimensional posterior summarization. Journal of Computational and Graphical Statistics, 30 (1):144–161, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1796684.

Woodard:2013:HAR

[WCR13] Dawn B. Woodard, Ciprian Crainiceanu, and David Ruppert. Hierarchical adaptive regression kernels for regression with functional predictors. Journal of Computational and Graphical Statistics, 22(4):777–800, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2021:OUS

[WCSY21] Jing Wu, Ming-Hui Chen, Elizabeth D. Schifano, and Jun Yan. Online updating of survival analysis. Journal of Computational and Graphical Statistics, 30(4):1209– 1223, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1870481.

Wang:2011:FBI

[WD11] Lianming Wang and David B. Dunson. Fast Bayesian inference in Dirichlet process mixture models. Journal of Computational and Graphical Statistics, 20(1):196–216, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wood:1996:SLE

[WDB96] A. T. A. Wood, K.-A. Do, and B. M. Broom. Sequential linearization of empirical likelihood constraints with application to U-statistics. Journal of Computational and Graphical Statistics, 5(4): 365–385, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1080/10618600.
                                        1996.10474718.

Wang:2022:EMS

[WDMW22] Zhanfeng Wang, Hao Dong, Ping Ma, and Yuedong Wang. Estimation and model selection for nonparametric functionon-function regression. Journal of Computational and Graphical Statistics, 31(3):835–845, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2037434.

Wegman:1995:HDS

[Weg95] Edward J. Wegman. Huge data sets and the frontiers of computational feasibility. Journal of Computational and Graphical Statistics, 4(4):281– 295, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474685.

Wang:2023:AUE

[WFCT23] Wenjia Wang, Yusi Fang, Chung Chang, and George C. Tseng. Accurate and ultra-efficient p-value calculation for higher criticism tests. Journal of Computational and Graphical Statistics, 33 (2):463–476, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2270720.

Witten:2011:NIF

[WFS11] Daniela M. Witten, Jerome H.

Friedman, and Noah Simon. New insights and faster computations for the graphical lasso. Journal of Computational and Graphical Statistics, 20 (4):892–900, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2022:ASB

[WGDW22] Shijia Wang, Shufei Ge, Renny Doig, and Liangliang Wang. Adaptive semiparametric Bayesian differential equations via sequential Monte Carlo. Journal of Computational and Graphical Statistics, 31(2):600–613, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1987252.

Wolock:2024:FLM

[WGSC24] Charles J. Wolock, Peter B. Gilbert, Noah Simon, and Marco Carone. A framework for leveraging machine learning tools to estimate personalized survival curves. Journal of Computational and Graphical Statistics, 33 (3):1098–1108, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2304070.

Werthenbach:1998:FSU

[WH98] Claudia Werthenbach and Eva Herrmann. A fast and stable updating algorithm for bivariate nonparametric curve estimation. Journal of Computational and Graphical Statistics, 7(1):61–76, March 1998. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.research.att.com/
~andreas/jcgs/abstracts/
1998/march/werthenbach.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1998.10474761.

Wang:2014:BVC

[WH14] Jianqiang C. Wang and Trevor Hastie. Boosted varying-coefficient regression models for product demand prediction. Journal of Computational and Graphical Statistics, 23 (2):361–382, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2017:RPC

[WH17] Wen-Ting Wang and Hsin-Cheng Huang. Regularized principal component analysis for spatial data. Journal of Computational and Graphical Statistics, 26(1):14– 25, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2016.1157483.

Wang:2023:LGP

[WHH23] Meijia Wang, Jingyu He, and P. Richard Hahn. Local Gaussian process extrapolation for BART models with applications to causal inference. Journal of Computational and Graphical Statistics, 33 (2):724–735, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2240384.

Whitcher:2001:SGS

[Whi01] Brandon Whitcher. Simulating Gaussian stationary processes with unbounded spectra. Journal of Computational and Graphical Statistics, 10 (1):112–134, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.amstat.org/publications/
                                        jcgs/abstracts01/Whitcher.
                                        htm;   http://www.tandfonline.
                                        com/doi/abs/10.1198/
                                        10618600152418674.

Wickham:2010:LGG

[Wic10] Hadley Wickham. A layered grammar of graphics. Journal of Computational and Graphical Statistics, 19(1):3–28, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wickham:2011:TTA

[Wic11a] Charlotte Wickham. A tale of two airports: Exploring flight traffic at SFO and OAK. Journal of Computational and Graphical Statistics, 20 (2):291–293, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wickham:2011:ADE

[Wic11b] Hadley Wickham. ASA 2009 Data Expo. Journal of Computational and Graphical Statistics, 20 (2):281–283, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wicklin:2011:VAD

[Wic11c] Rick Wicklin. Visualizing airline delays and cancelations. Journal of Computational and Graphical Statistics, 20(2):284–286, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wickham:2013:GCS

[Wic13] Hadley Wickham. Graphical criticism: Some historical notes. Journal of Computational and Graphical Statistics, 22 (1):38–44, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wilkinson:1993:CSM

[Wil93] Leland Wilkinson. Comment on “A Model for Studying Display Methods of Statistical Graphics”. Journal of Computational and Graphical Statistics, 2(4):355– 360, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474619.  See

[Cle93a].

Wilks:1996:CBR

[Wil96] Allan R. Wilks. Constraintbased representations of statistical graphs. Journal of Computational and Graphical Statistics, 5(4): 400–414, December 1996. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1996.10474720.

Wills:1999:NIV

[Wil99] Graham J. Wills. NicheWorks — interactive visualization of very large graphs. Journal of Computational and Graphical Statistics, 8(2):190– 212, June 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/nicheworks.
html;   http://www.amstat.
org/publications/jcgs/
pdf99/wills.pdf;   http:
//www.tandfonline.com/
doi/abs/10.1080/10618600.
1999.10474810.

Wills:2000:NSI

[Wil00] Graham J. Wills. Natural selection: Interactive subset creation. Journal of Computational and Graphical Statistics, 9(3):544–557, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Wills.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474898.

Wills:2011:CWT

[Wil11] Graham Wills. Comment: Why Tables Are Really Much Better Than Graphs. Journal of Computational and Graphical Statistics, 20(1):28– 35, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Gel11b].

Wiper:2001:MGD

[WIR01] Michael Wiper, David Rios Insua, and Fabrizio Ruggeri. Mixtures of gamma distributions with applications. Journal of Computational and Graphical Statistics, 10(3):440–454, September 2001. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317115054.

Whiteley:2011:MCF

[WJG11] Nick Whiteley, Adam M. Johansen, and Simon Godsill. Monte Carlo filtering of piecewise deterministic processes. Journal of Computational and Graphical Statistics, 20 (1):119–139, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2023:VCM

[WJL23] Xufei Wang, Bo Jiang, and Jun S. Liu. Varying coefficient model via adaptive spline fitting. Journal of Computational and Graphical Statistics, 33(2):614–624, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2267616.

Willems:2009:DMO

[WJZ09] Gert Willems, Harry Joe, and Ruben Zamar. Diagnosing multivariate outliers detected by robust estimators. Journal of Computational and Graphical Statistics, 18 (1):73–91, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wright:2000:IAA

[WK00] Kevin Wright and William J.

Kennedy. An interval analysis approach to the EM algorithm. Journal of Computational and Graphical Statistics, 9(2): 303–318, 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474882.

Wang:2009:SSS

[WK09] Yuedong Wang and Chunlei Ke. Smoothing spline semiparametric nonlinear regression models. Journal of Computational and Graphical Statistics, 18 (1):165–183, March 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wilson:2017:AST

[WK17] Huon Wilson and Uri Keich. Accurate small tail probabilities of sums of iid lattice-valued random variables via FFT. Journal of Computational and Graphical Statistics, 26 (1):223–229, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1276841.

Wood:2008:LAN

[WKC^{+}08] Sally A. Wood, Robert Kohn, Remy Cottet, Wenxin Jiang, and Martin Tanner. Locally adaptive nonparametric binary regression. Journal of Computational and Graphical Statistics, 17 (2):352–372, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wilms:2022:GID

[WKM22] Ines Wilms, Rebecca Killick, and David S. Matteson. Graphical influence diagnostics for changepoint models. Journal of Computational and Graphical Statistics, 31 (3):753–765, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2000873.

Wu:2024:CRM

[WKT^{+}24] Wenbo Wu, John D.

Kalbfleisch, Jeremy M. G. Taylor, Jian Kang, and Kevin He. Competing risk modeling with bivariate varying coefficients to understand the dynamic impact of COVID-19. Journal of Computational and Graphical Statistics, 33 (4):1252–1263, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2304089.

Wegman:2002:MCG

[WL02] Edward J. Wegman and Qiang Luo. On methods of computer graphics for visualizing densities. Journal of Computational and Graphical Statistics, 11(1):137– 162, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s7/p137;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375659.

Wang:2008:ISS

[WL08] Xiao Wang and Feng Li. Isotonic smoothing spline regression. Journal of Computational and Graphical Statistics, 17 (1):21–37, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wagaman:2009:DSC

[WL09] A. S. Wagaman and E. Levina. Discovering sparse covariance structures with the Isomap. Journal of Computational and Graphical Statistics, 18(3):551–572, September 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2013:AWL

[WL13a] Yichao Wu and Yufeng Liu. Adaptively weighted large margin classifiers. Journal of Computational and Graphical Statistics, 22(2):416–432, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2013:FRS

[WL13b] Yichao Wu and Yufeng Liu. Functional robust support vector machines for sparse and irregular longitudinal data. Journal of Computational and Graphical Statistics, 22 (2):379–395, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wilkinson:2022:DPM

[WL22] Leland Wilkinson and Hengrui Luo. A distancepreserving matrix sketch. Journal of Computational and Graphical Statistics, 31(4):945–959, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2050246.

Wei:2023:HDM

[WL23] Zhenyu Wei and Thomas C. M. Lee. Highdimensional multi-task learning using multivariate regression and generalized fiducial inference. Journal of Computational and Graphical Statistics, 32(1):226–240, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2090946.

Wei:2024:VBN

[WL24] Susan Wei and Edmund Lau. Variational Bayesian neural networks via resolution of singularities. Journal of Computational and Graphical Statistics, 33(4):1464–1474, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2024.2325455.

Wu:2010:LSI

[WLM10] Qiang Wu, Feng Liang, and Sayan Mukherjee. Localized sliced inverse regression. Journal of Computational and Graphical Statistics, 19(4):843–860, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wei:2016:DPP

[WLWM16] Susan Wei, Chihoon Lee, Lindsay Wichers, and J. S. Marron. Directionprojection-permutation for high-dimensional hypothesis tests. Journal of Computational and Graphical Statistics, 25 (2):549–569, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2024:CEN

[WLZ^{+}24] Caixing Wang, Tao Li, Xinyi Zhang, Xingdong Feng, and Xin He. Communication-efficient nonparametric quantile regression via random features. Journal of Computational and Graphical Statistics, 33(4):1175– 1184, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2308798.

Wang:2007:REU

[WLZB07] You-Gan Wang, Xu Lin, Min Zhu, and Zhidong Bai. Robust estimation using the Huber function with a data-dependent tuning constant. Journal of Computational and Graphical Statistics, 16 (2):468–481, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Walker:2011:DS

[WLZD11] Stephen G. Walker, Purushottam W. Laud, Daniel Zantedeschi, and Paul Damien. Direct sampling. Journal of Computational and Graphical Statistics, 20(3):692–713, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wegman:2003:STS

[WM03] Edward J. Wegman and David J. Marchette. On some techniques for streaming data: A case study of Internet packet headers. Journal of Computational and Graphical Statistics, 12(4):893–914, December 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/1061860032625.

Westling:2019:BPF

[WM19] T. Westling and T. H. McCormick. Beyond prediction: A framework for inference with variational approximations in mixture models. Journal of Computational and Graphical Statistics, 28 (4):778–789, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Weylandt:2020:DVF

[WNA20] Michael Weylandt, John Nagorski, and Genevera I. Allen. Dynamic visualization and fast computation for convex clustering via algorithmic regularization. Journal of Computational and Graphical Statistics, 29 (1):87–96, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1629943.

Wang:2013:CSP

[WP13] Hao Wang and Natesh S. Pillai. On a class of shrinkage priors for covariance matrix estimation. Journal of Computational and Graphical Statistics, 22(3):689–707, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Woodruff:1993:HSA

[WR93] David L. Woodruff and David M. Rocke. Heuristic search algorithms for the minimum volume ellipsoid. Journal of Computational and Graphical Statistics, 2(1):69– 95, 1993. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1993.10474600.

Walsh:2005:CMS

[WR05] Daniel C. I. Walsh and Adrian E. Raftery. Classification of mixtures of spatial point processes via partial Bayes factors. Journal of Computational and Graphical Statistics, 14(1):139–154, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wilkinson:2000:LG

[WRCR00] Leland Wilkinson, Daniel J.

Rope, Daniel B. Carr, and Matthew A. Rubin. The language of graphics. Journal of Computational and Graphical Statistics, 9(3):530–543, September 2000. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts00/Wilkinson.
htm;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.2000.10474897.

Wood:2011:BMA

[WRK11] Sally Wood, Ori Rosen, and Robert Kohn. Bayesian mixtures of autoregressive models. Journal of Computational and Graphical Statistics, 20(1):174–195, March 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2021:EAV

[WRT21] Qi Wang, Vinayak Rao, and Yee Whye Teh. An exact auxiliary variable Gibbs sampler for a class of diffusions. Journal of Computational and Graphical Statistics, 30 (2):297–311, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1816177.

Wegman:2002:IPT

[WS02a] Edward J. Wegman and Jürgen Symanzik. Immersive projection technology for visual data mining. Journal of Computational and Graphical Statistics, 11(1):163– 188, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s8/p163;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375668.

White:2002:RTC

[WS02b] Denis White and Jean C.

Sifneos. Regression tree cartography. Journal of Computational and Graphical Statistics, 11(3):600–614, September 2002. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n3/s6/p600;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002484.

Williams:2024:OCS

[WS24] Matthew R. Williams and Terrance D. Savitsky. Optimization for calibration of survey weights under a large number of conflicting constraints. Journal of Computational and Graphical Statistics, 33 (3):1047–1060, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2024.2304071.

Wang:2024:GQR

[WSB24] Shijie Wang, Minsuk Shin, and Ray Bai. Generative quantile regression with variability penalty. Journal of Computational and Graphical Statistics, 33(4):1202–1213, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2327824.

Weinhold:2020:RFA

[WSM^{+}20] Leonie Weinhold, Matthias Schmid, Richard Mitchell, Kelly O. Maloney, Marvin N. Wright, and Moritz Berger. A random forest approach for bounded outcome variables. Journal of Computational and Graphical Statistics, 29 (3):639–658, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2019.1705310.

Wu:2007:BC

[WTW07] Yuhong Wu, Håkon Tjelmeland, and Mike West. Bayesian CART. Journal of Computational and Graphical Statistics, 16 (1):44–66, March 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2000:DOT

[Wu00] Ying Nian Wu. Discussion: Optimization Transfer Using Surrogate Objective Functions. Journal of Computational and Graphical Statistics, 9(1): 32–34, March 2000. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2000.10474861.  See

[LHY00].

Wu:2008:KSI

[Wu08]  Han-Ming Wu.  Ker-

nel sliced inverse regression with applications to classification. Journal of Computational and Graphical Statistics, 17 (3):590–610, September 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2011:DAI

[Wu11] Ying Nian Wu. Data augmentation, internal representation, and unsupervised learning. Journal of Computational and Graphical Statistics, 20 (3):581–583, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See [Kel11].

Wilkinson:2008:SD

[WW08] Leland Wilkinson and Graham Wills. Scagnostics distributions. Journal of Computational and Graphical Statistics, 17 (2):473–491, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Walsh:2012:GAD

[WW12] Stephen Walsh and Paul Whitney. A graphical approach to diagnosing the validity of the conditional independence assumptions of a Bayesian network given data. Journal of Computational and Graphical Statistics, 21 (4):961–978, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2019:FIM

[WW19] Jiacheng Wu and Daniela Witten. Flexible and interpretable models for survival data. Journal of Computational and Graphical Statistics, 28 (4):954–966, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:1997:NTO

[WWG^{+}97] Suojin Wang, Wayne A. Woodward, H. L. Gray, Stephen Wiechecki, and Stephan R. Sain. A new test for outlier detection from a multivariate mixture distribution. Journal of Computational and Graphical Statistics, 6(3):285–299, September 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/wang.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1997.10474743.

Wen:2022:SAR

[WWJ22] Canhong Wen, Qin Wang, and Yuan Jiang. Stability approach to regularization selection for reduced-rank regression. Journal of Computational and Graphical Statistics, 32(3):974–984, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2119986.

Wang:2023:JIC

[WWL^{+}23] Peiyao Wang, Haodong Wang, Quefeng Li, Dinggang Shen, and Yufeng Liu. Joint and individual component regression. Journal of Computational and Graphical Statistics, 33(3):763–773, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2284227.

Wu:2015:SED

[WY15] Wenbo Wu and Xiangrong Yin. Stable estimation in dimension reduction. Journal of Computational and Graphical Statistics, 24(1):104–120, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2022:TLI

[WY22] Lili Wu and Shu Yang. Transfer learning of individualized treatment rules from experimental to real-world data. Journal of Computational and Graphical Statistics, 32 (3):1036–1045, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2141752.

Wong:2014:REG

[WYL14] Raymond K. W. Wong, Fang Yao, and Thomas C. M. Lee. Robust estimation for generalized additive models. Journal of Computational and Graphical Statistics, 23 (1):270–289, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wong:2019:ODM

[WYZ19] Weng Kee Wong, Yue Yin, and Julie Zhou. Optimal designs for multiresponse nonlinear regression models with several factors via semidefinite programming. Journal of Computational and Graphical Statistics, 28 (1):61–73, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2001:DAD

[WZ01] Ying Nian Wu and Song Chun Zhu. Discussion: The Art of Data Argumentation. Journal of Computational and Graphical Statistics, 10 (1):90–93, March 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1198/10618600152418764.

See [vDM01a].

Wang:2016:SDW

[WZ16] Boxiang Wang and Hui Zou. Sparse distance weighted discrimination. Journal of Computational and Graphical Statistics, 25(3):826–838, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1049700.

Wong:2022:UCC

[WZ22] Weng Kee Wong and Julie Zhou. Using CVX to construct optimal designs for biomedical studies with multiple objectives. Journal of Computational and Graphical Statistics, 32 (2):744–753, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2104858.

Wang:2021:ILU

[WZLR21] Haiying Wang, Dixin Zhang, Hua Liang, and David Ruppert. Iterative likelihood: a unified inference tool. Journal of Computational and Graphical Statistics, 30 (4):920–933, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.1904961.

Wang:2018:SCC

[WZSF18] Binhuan Wang, Yilong Zhang, Will Wei Sun, and Yixin Fang. Sparse convex clustering. Journal of Computational and Graphical Statistics, 27 (2):393–403, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wang:2019:MPE

[WZW19] Xin Wang, Hao Helen Zhang, and Yichao Wu. Multiclass probability estimation with support vector machines. Journal of Computational and Graphical Statistics, 28 (3):586–595, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Wu:2024:CDL

[WZXW24] Shuyuan Wu, Jing Zhou, Ke Xu, and Hansheng Wang. Class-distributed learning for multinomial logistic regression with high dimensional features and a large number of classes. Journal of Computational and Graphical Statistics, 34(1):175– 186, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2362230.

Wang:2024:MSB

[WZZ^{+}24] Zhiwei Wang, Fa Zhang, Cong Zheng, Xianghong Hu, Mingxuan Cai, and Can Yang. MFAI: a scalable Bayesian matrix factorization approach to leveraging auxiliary information. Journal of Computational and Graphical Statistics, 33(4):1339– 1349, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2319160.

Xu:2015:IEU

[XBW15] Cong Xu, Paul Baines, and Jane-Ling Wang. Improved estimation and uncertainty quantification using Monte Carlo-based optimization algorithms. Journal of Computational and Graphical Statistics, 24(3):771–791, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xiong:2022:BLD

[XC22] Xin Xiong and Ivor Cribben. Beyond linear dynamic functional connectivity: a vine copula change point model. Journal of Computational and Graphical Statistics, 32 (3):853–872, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2127738.

Xu:2016:PEA

[XCSP16] Lizhen Xu, Radu V. Craiu, Lei Sun, and Andrew D. Paterson. Parameter expanded algorithms for Bayesian latent variable modeling of genetic pleiotropy data. Journal of Computational and Graphical Statistics, 25(2):405–425, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xia:2020:SHS

[XD20] Wei Xia and Hani Doss.

Scalable hyperparameter selection for latent Dirichlet allocation. Journal of Computational and Graphical Statistics, 29 (4):875–895, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1741378.

Xu:2011:IIP

[XGH11] Ping-Feng Xu, Jianhua Guo, and Xuming He. An improved iterative proportional scaling procedure for Gaussian graphical models. Journal of Computational and Graphical Statistics, 20 (2):417–431, June 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2015:LII

[XGT15] Ping-Feng Xu, Jianhua Guo, and Man-Lai Tang. A localized implementation of the iterative proportional scaling procedure for Gaussian graphical models. Journal of Computational and Graphical Statistics, 24 (1):205–229, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2018:MRS

[XGTB18] Hui Xu, Xiangdong Gu, Mahlet G. Tadesse, and Raji Balasubramanian. A modified random survival forests algorithm for high dimensional predictors and self-reported outcomes. Journal of Computational and Graphical Statistics, 27(4):763–772, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xue:2017:RMF

[XL17] Jingnan Xue and Faming Liang. A robust model-free feature screening method for ultrahighdimensional data. Journal of Computational and Graphical Statistics, 26 (4):803–813, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2013:TGR

[XMY13] Xiaojin Xu, Xiao-Li Meng, and Yaming Yu. Thank God that regressing Y on X is not the same as regressing X on Y: Direct and indirect residual augmentations. Journal of Computational and Graphical Statistics, 22(3):598–622, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2016:CSR

[XNN16] Ruo Xu, Dan Nettleton, and Daniel J. Nordman. Case-specific random forests. Journal of Computational and Graphical Statistics, 25 (1):49–65, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.983641;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.983641.

Xiao:2007:CCG

[XP07] Guanghua Xiao and Wei Pan. Consensus clustering of gene expression data and its application to gene function prediction. Journal of Computational and Graphical Statistics, 16(3):733–751, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2019:DGC

[XSC19] Ganggang Xu, Zuofeng Shang, and Guang Cheng.

Distributed generalized cross-validation for divideand-conquer kernel ridge regression and its asymptotic optimality. Journal of Computational and Graphical Statistics, 28 (4):891–908, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xie:2022:MDM

[XTPO22] Shanghong Xie, Thaddeus Tarpey, Eva Petkova, and R. Todd Ogden. Multiple domain and multiple kernel outcome-weighted learning for estimating individualized treatment regimes. Journal of Computational and Graphical Statistics, 31(4):1375– 1383, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2067552.

Xu:2018:DRA

[XW18] Danqing Xu and Yuedong Wang. Divide and recombine approaches for fitting smoothing spline models with large datasets. Journal of Computational and Graphical Statistics, 27 (3):677–683, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xu:2020:APV

[XW20] Danli Xu and Yong Wang.

Area-proportional visualization for circular data. Journal of Computational and Graphical Statistics, 29(2):351–357, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1654881.

Xu:2024:BCI

[XWS24] Mingwei Xu, Samuel W. K. Wong, and Peijun Sang. A Bayesian collocation integral method for parameter estimation in ordinary differential equations. Journal of Computational and Graphical Statistics, 33 (3):845–854, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2302528.

Xiao:2015:CEL

[XWZ15] Wei Xiao, Yichao Wu, and Hua Zhou. ConvexLAR: An extension of least angle regression. Journal of Computational and Graphical Statistics, 24(3):603–626, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xue:2014:SDF

[XY14] Yuan Xue and Xiangrong Yin. Sufficient dimension folding for regression mean function. Journal of Computational and Graphical Statistics, 23 (4):1028–1043, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xin:2012:SSE

[XZ12] Lu Xin and Mu Zhu. Stochastic stepwise ensembles for variable selection. Journal of Computational and Graphical Statistics, 21(2):275–294, 2012. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Xing:2024:PTD

[XZYZ24] Xin Xing, Peng Zeng, Youhui Ye, and Wenxuan Zhong. Parsimonious tensor dimension reduction. Journal of Computational and Graphical Statistics, 34(1):343–359, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2024.2362220.

Yang:2021:ARM

[Yan21] Lu Yang. Assessment of regression models with discrete outcomes using quasi-empirical residual distribution functions. Journal of Computational and Graphical Statistics, 30(4):1019–1035, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1910042.

Yang:2024:DPI

[Yan24] Lu Yang. Double probability integral transform residuals for regression models with discrete outcomes. Journal of Computational and Graphical Statistics, 33(3):787– 803, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2303336.

Yang:2020:DDA

[YASR20] Ran Yang, Daniel W. Apley, Jeremy Staum, and David Ruppert. Density deconvolution with additive measurement errors using quadratic programming. Journal of Computational and Graphical Statistics, 29(3):580– 591, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1704294.

Yatracos:2023:EGT

[Yat23] Yannis G. Yatracos. EDI-Graphic: a tool to study parameter discrimination and confirm identifiability in blackbox models, and to select data-generating machines. Journal of Computational and Graphical Statistics, 33(1):126– 137, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2205483.

Yonekura:2022:OSD

[YB22] Shouto Yonekura and Alexandros Beskos. Online smoothing for diffusion processes observed with noise. Journal of Computational and Graphical Statistics, 31 (4):1344–1360, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2027243.

Yang:2002:CMI

[YBDS02] Yee Hwa Yang, Michael J. Buckley, Sandrine Dudoit, and Terence P. Speed. Comparison of methods for image analysis on cDNA microarray data. Journal of Computational and Graphical Statistics, 11(1):108– 136, March 2002. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/lucia.ingentaselect.
com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/10618600/
v11n1/s6/p108;   http:
//www.tandfonline.com/
doi/abs/10.1198/106186002317375640.

Yin:2004:DRM

[YC04] Xiangrong Yin and R. Dennis Cook. Dimension reduction via marginal fourth moments in regression. Journal of Computational and Graphical Statistics, 13(3):554–570, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yen:2020:SFP

[Yen20] Tso-Jung Yen. Solving fused penalty estimation problems via block splitting algorithms. Journal of Computational and Graphical Statistics, 29 (2):297–308, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1660178.

Yu:2014:MCV

[YF14] Yi Yu and Yang Feng.

Modified cross-validation for penalized high-dimensional linear regression models. Journal of Computational and Graphical Statistics, 23(4):1009–1027, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yuen:2022:EDB

[YF22] Christine Yuen and Piotr Fryzlewicz. Exploiting disagreement between high-dimensional variable selectors for uncertainty visualization. Journal of Computational and Graphical Statistics, 31 (2):351–359, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2000421.

Yuan:2022:DMM

[YG22] Dongbang Yuan and Irina Gaynanova. Doublematched matrix decomposition for multi-view data. Journal of Computational and Graphical Statistics, 31(4):1114– 1126, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2067860.

Ye:2022:MCC

[YGD22] Chenglong Ye, Reza Ghanadan, and Jie Ding. Meta clustering for collaborative learning. Journal of Computational and Graphical Statistics, 32 (3):1160–1169, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2126484.

Yi:2017:SNC

[YH17] Congrui Yi and Jian Huang. Semismooth Newton coordinate descent algorithm for elastic-net penalized Huber loss regression and quantile regression. Journal of Computational and Graphical Statistics, 26(3):547–557, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yang:2023:MSE

[YHH^{+}23] Xi Yang, Jan Hannig, Katherine A. Hoadley, Iain Carmichael, and J. S. Marron. Measure of strength of evidence for visually observed differences between subpopulations. Journal of Computational and Graphical Statistics, 33(2):736– 748, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2023.2276113.

Yin:2023:APB

[YHS23] Fan Yin, Guanyu Hu, and Weining Shen. Analysis of professional basketball field goal attempts via a Bayesian matrix clustering approach. Journal of Computational and Graphical Statistics, 32 (1):49–60, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2085727.

Yang:2001:BBS

[YK01] Tae Young Yang and Lynn Kuo. Bayesian binary segmentation procedure for a Poisson process with multiple changepoints. Journal of Computational and Graphical Statistics, 10(4):772–785, December 2001. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/abs/10.1198/106186001317243449.

Yanev:2009:GBS

[YK09] Petko I. Yanev and Erricos J. Kontoghiorghes. Graph-based strategies for performing the exhaustive and random k-fold cross-validations. Journal of Computational and Graphical Statistics, 18(4):894–914, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yau:2003:BVS

[YKW03] Paul Yau, Robert Kohn, and Sally Wood. Bayesian variable selection and model averaging in highdimensional multinomial nonparametric regression. Journal of Computational and Graphical Statistics, 12(1):23–54, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s2/p23;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031301.

Young:2018:MVT

[YKZ18] Derek S. Young, Chenlu Ke, and Xiaoxue Zeng. The mixturegram: A visualization tool for assessing the number of components in finite mixture models. Journal of Computational and Graphical Statistics, 27(3):564–575, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Young:1995:GDA

[YL95] Forrest W. Young and David J. Lubinsky. Guiding data analysts with visual statistical strategies. Journal of Computational and Graphical Statistics, 4(4):229–250, 1995. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474680.

Yu:1999:FTF

[YL99] Yan Yu and Diane Lambert. Fitting trees to functional data, with an application to time-of-day patterns. Journal of Computational and Graphical Statistics, 8(4):749–762, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/yu2.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474847.

Yan:2022:BTS

[YL22] Yinqiao Yan and Xiangyu Luo. Bayesian tree-structured two-level clustering for nested data analysis. Journal of Computational and Graphical Statistics, 32(3):1185– 1194, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

doi/full/10.1080/10618600.
2022.2130927.

Yang:2019:ACW

[YLCR19] Jinyoung Yang, Evgeny Levi, Radu V. Craiu, and Jeffrey S. Rosenthal. Adaptive component-wise multiple-try Metropolis sampling. Journal of Computational and Graphical Statistics, 28 (2):276–289, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yi:2023:GLS

[YLLZ23] Si-Yu Yi, Ze Liu, Min-Qian Liu, and Yong-Dao Zhou. Global likelihood sampler for multimodal distributions. Journal of Computational and Graphical Statistics, 32 (3):927–937, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2165499.

Yu:2017:PNS

[YLM17] Qunqun Yu, Xiaosun Lu, and J. S. Marron. Principal nested spheres for time-warped functional data analysis. Journal of Computational and Graphical Statistics, 26 (1):144–151, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1115359.

Yu:2017:PAL

[YLW17] Liqun Yu, Nan Lin, and Lan Wang. A parallel algorithm for largescale nonconvex penalized quantile regression. Journal of Computational and Graphical Statistics, 26 (4):935–939, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yu:2011:RAO

[YM11a] Yaming Yu and Xiao-Li Meng. Rejoinder: Be all our insomnia remembered. . .. Journal of Computational and Graphical Statistics, 20 (3):603–615, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yu:2011:CCQ

[YM11b] Yaming Yu and Xiao-Li Meng. To center or not to center: That is not the question—an ancillarity– sufficiency interweaving strategy (ASIS) for boosting MCMC efficiency. Journal of Computational and Graphical Statistics, 20(3):531–570, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). See comment [HR11].

Yang:2014:SSV

[YMB14] Dan Yang, Zongming Ma, and Andreas Buja. A sparse singular value decomposition method for high-dimensional data. Journal of Computational and Graphical Statistics, 23(4):923–942, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yoon:2021:FCL

[YMG21] Grace Yoon, Christian L.

Müller, and Irina Gaynanova. Fast computation of latent correlations. Journal of Computational and Graphical Statistics, 30(4):1249–1256, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1882468.

Yin:2011:SGM

[YMLY11] Guosheng Yin, Yanyuan Ma, Faming Liang, and Ying Yuan. Stochastic generalized method of moments. Journal of Computational and Graphical Statistics, 20 (3):714–727, September 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yang:2018:NAC

[YNH18] Xiaorong Yang, Naveen Naidu Narisetty, and Xuming He. A new approach to censored quantile regression estimation. Journal of Computational and Graphical Statistics, 27 (2):417–425, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yu:2021:AAA

[YNTK21] Xuejun Yu, David J. Nott, Minh-Ngoc Tran, and Nadja Klein. Assessment and adjustment of approximate inference algorithms using the law of total variance. Journal of Computational and Graphical Statistics, 30 (4):977–990, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1880921.

You:2022:ACV

[YOL^{+}22] Chong You, John T. Ormerod, Xiangyang Li, Cheng Heng Pang, and Xiao-Hua Zhou. An approximated collapsed variational Bayes approach to variable selection in linear regression. Journal of Computational and Graphical Statistics, 32(3):782–792, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2149539.

Young:1995:R

[You95] Forrest W. Young. Rejoinder. Journal of Computational and Graphical Statistics, 4(4):259– 260, 1995. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
1995.10474683.

Yang:2020:ETV

[YP20] Jilei Yang and Jie Peng.

Estimating time-varying graphical models. Journal of Computational and Graphical Statistics, 29 (1):191–202, 2020. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1647848.

Yang:2021:SDC

[YPPM21] Congyuan Yang, Carey E.

Priebe, Youngser Park, and David J. Marchette. Simultaneous dimensionality and complexity model selection for spectral graph clustering. Journal of Computational and Graphical Statistics, 30(2):422–441, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1824870.

Yu:2022:GML

[YR22] Ruoqi Yu and Paul R. Rosenbaum. Graded matching for large observational studies. Journal of Computational and Graphical Statistics, 31 (4):1406–1415, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2058001.

Yue:2010:NSG

[YS10] Yu Yue and Paul L. Speckman. Nonstationary spatial Gaussian Markov random fields. Journal of Computational and Graphical Statistics, 19 (1):96–116, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yildirim:2013:OEM

[YSD13] Sinan Yildirim, Sumeetpal S. Singh, and Arnaud Doucet. An online expectation-maximization algorithm for changepoint models. Journal of Computational and Graphical Statistics, 22(4):906–926, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yildirim:2015:PEH

[YSDJ15] Sinan Yildirim, Sumeetpal S. Singh, Thomas Dean, and Ajay Jasra. Parameter estimation in hidden Markov models with intractable likelihoods using sequential Monte Carlo. Journal of Computational and Graphical Statistics, 24 (3):846–865, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

You:2024:WMP

[YSG24] Kisung You, Dennis Shung, and Mauro Giuffrè. On the Wasserstein median of probability measures. Journal of Computational and Graphical Statistics, 34(1):253– 266, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2374580.

Yang:2023:MPM

[YSJS23] Jianxiao Yang, Martijn J.

Schuemie, Xiang Ji, and Marc A. Suchard. Massive parallelization of massive sample-size survival analysis. Journal of Computational and Graphical Statistics, 33(1):289– 302, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2213279.

Yeon:2010:MAP

[YSK^{+}10] Kyupil Yeon, Moon Sup Song, Yongdai Kim, Hosik Choi, and Cheolwoo Park. Model averaging via penalized regression for tracking concept drift. Journal of Computational and Graphical Statistics, 19(2):457–473, June 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yu:2024:DBC

[YSWS24] Ping Yu, Gongming Shi, Chunjie Wang, and Xinyuan Song. Distancebased clustering of functional data with derivative principal component analysis. Journal of Computational and Graphical Statistics, 34 (1):47–58, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2366499.

Yu:1999:ASS

[YT99] Joan Z. Yu and Martin A. Tanner. An analytical study of several Markov chain Monte Carlo estimators of the marginal likelihood. Journal of Computational and Graphical Statistics, 8(4):839–853, December 1999. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts99/yu.
html;   http://www.tandfonline.
com/doi/abs/10.1080/
10618600.1999.10474852.

Yang:2021:HTV

[YT21a] Ting Yang and Zhiqiang Tan. Hierarchical total variations and doubly penalized ANOVA modeling for multivariate nonparametric regression. Journal of Computational and Graphical Statistics, 30 (4):848–862, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923513.

Yi:2021:SCM

[YT21b] Mengxi Yi and David E.

Tyler. Shrinking the covariance matrix using convex penalties on the matrix-log transformation. Journal of Computational and Graphical Statistics, 30(2):442– 451, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1814788.

Yu:2012:MOE

[Yu12] Yaming Yu. Monotonically overrelaxed EM algorithms. Journal of Computational and Graphical Statistics, 21 (2):518–537, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yuan:2008:ECR

[Yua08] Ming Yuan. Efficient computation of ℓ_{1} regularized estimates in Gaussian graphical models. Journal of Computational and Graphical Statistics, 17(4):809–826, December 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Young:2003:GAS

[YVMFB03] Forrest W. Young, Pedro Valero-Mora, Richard A. Faldowski, and Carla M. Bann. Gossip: The architecture of SpreadPlots. Journal of Computational and Graphical Statistics, 12(1):80–100, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s4/p80;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031356.

Yu:2022:NGP

[YWBA22] Weichang Yu, Sara Wade, Howard D. Bondell, and Lamiae Azizi. Nonstationary Gaussian process discriminant analysis with variable selection for high-dimensional functional data. Journal of Computational and Graphical Statistics, 32 (2):588–600, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2098136.

Yu:2015:HDF

[YWL^{+}15] Donghyeon Yu, Joong-Ho Won, Taehoon Lee, Johan Lim, and Sungroh Yoon. High-dimensional fused lasso regression using majorization-minimization and parallel processing. Journal of Computational and Graphical Statistics, 24(1):121–153, 2015. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yin:2006:EAC

[YZ06] Guosheng Yin and Donglin Zeng. Efficient algorithm for computing maximum likelihood estimates in linear transformation models. Journal of Computational and Graphical Statistics, 15(1):228–245, March 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yin:2007:EDE

[YZ07] Xiangrong Yin and Lixing Zhu. Estimating directions in extending generalized partially linear single-index models. Journal of Computational and Graphical Statistics, 16(2):330–349, June 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Yang:2013:EAC

[YZ13] Yi Yang and Hui Zou. An efficient algorithm for computing the HHSVM and its generalizations. Journal of Computational and Graphical Statistics, 22(2):396–415, 2013. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:IAN

[ZA23] Shengtong Zhang and Daniel W. Apley. Interpretable architecture neural networks for function visualization. Journal of Computational and Graphical Statistics, 32 (4):1258–1271, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2195461.

Zaslavsky:2010:PPO

[ZB10] Alan M. Zaslavsky and Eric T. Bradlow. Posterior predictive outlier detection using sample reweighting. Journal of Computational and Graphical Statistics, 19 (4):790–807, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2006:SCM

[ZBB06] Xiao Zhang, W. John Boscardin, and Thomas R. Belin. Sampling correlation matrices in Bayesian models with correlated latent variables. Journal of Computational and Graphical Statistics, 15 (4):880–896, December 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zheng:2021:TCM

[ZBS^{+}21] Peng Zheng, Ryan Barber, Reed J. D. Sorensen, Christopher J. L. Murray, and Aleksandr Y. Aravkin. Trimmed constrained mixed effects models: Formulations and algorithms. Journal of Computational and Graphical Statistics, 30 (3):544–556, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2020.1868303.

Zipunnikov:2009:CTU

[ZBY09] Vadim Zipunnikov, James G. Booth, and Ruriko Yoshida. Counting tables using the double-saddlepoint approximation. Journal of Computational and Graphical Statistics, 18 (4):915–929, December 2009. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2015:MCA

[ZC15] Jingfei Zhang and Yuguo Chen. Monte Carlo algorithms for identifying densely connected subgraphs. Journal of Computational and Graphical Statistics, 24(3):827–845, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2017:BMA

[ZCI^{+}17] Danjie Zhang, Ming-Hui Chen, Joseph G. Ibrahim, Mark E. Boye, and Wei Shen. Bayesian model assessment in joint modeling of longitudinal and survival data with applications to cancer clinical trials. Journal of Computational and Graphical Statistics, 26(1):121– 133, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1117472.

Zhao:2015:WBW

[ZCO15] Yihong Zhao, Huaihou Chen, and R. Todd Ogden. Wavelet-based weighted LASSO and screening approaches in functional linear regression. Journal of Computational and Graphical Statistics, 24(3):655–675, 2015. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2019:NFA

[ZCS^{+}19] Yun Zhang, Kehui Chen, Allan Sampson, Kai Hwang, and Beatriz Luna. Node features adjusted stochastic block model. Journal of Computational and Graphical Statistics, 28(2):362–373, 2019. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zipunnikov:2011:MFP

[ZCY^{+}11] Vadim Zipunnikov, Brian Caffo, David M. Yousem, Christos Davatzikos, Brian S. Schwartz, and Ciprian Crainiceanu. Multilevel functional principal component analysis for highdimensional data. Journal of Computational and Graphical Statistics, 20 (4):852–873, December 2011. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2020:SBA

[ZDM20] Tianjian Zhou, Michael J. Daniels, and Peter Müller. A semiparametric Bayesian approach to dropout in longitudinal studies with auxiliary covariates. Journal of Computational and Graphical Statistics, 29 (1):1–12, 2020. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2019.1617159.

Zhu:2022:FTD

[ZDM22] Yifan Zhu, Fan Dai, and Ranjan Maitra. Fully three-dimensional radial visualization. Journal of Computational and Graphical Statistics, 31 (3):935–944, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.2020129.

Zhu:2018:NAM

[ZF18] Wanchuang Zhu and Yanan Fan. A novel approach for Markov random field with intractable normalizing constant on large lattices. Journal of Computational and Graphical Statistics, 27 (1):59–70, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhai:2022:BSV

[ZG22] Ruoshui Zhai and Roee Gutman. A Bayesian singular value decomposition procedure for missing data imputation. Journal of Computational and Graphical Statistics, 32 (2):470–482, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2107534.

Zhang:2022:RSC

[ZGC22] Hai Zhang, Xiao Guo, and Xiangyu Chang. Randomized spectral clustering in large-scale stochastic block models. Journal of Computational and Graphical Statistics, 31 (3):887–906, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2022.2034636.

Zemlianskaia:2022:SHL

[ZGL22]  Natalia  Zemlianskaia,

W. James Gauderman, and Juan Pablo Lewinger. A scalable hierarchical lasso for gene-environment interactions. Journal of Computational and Graphical Statistics, 31 (4):1091–1103, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2039161.

Zoubouloglou:2022:STP

[ZGPM22] Pavlos Zoubouloglou, Eduardo García-Portugués, and J. S. Marron. Scaled torus principal component analysis. Journal of Computational and Graphical Statistics, 32 (3):1024–1035, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2119985.

Zhou:2021:MOA

[ZGTG21] Bingxin Zhou, Junbin Gao, Minh-Ngoc Tran, and Richard Gerlach. Manifold optimizationassisted Gaussian variational approximation. Journal of Computational and Graphical Statistics, 30(4):946–957, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923516.

Zhu:2003:FEN

[ZH03] Mu Zhu and Trevor J.

Hastie. Feature extraction for nonparametric discriminant analysis. Journal of Computational and Graphical Statistics, 12(1):101–120, March 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n1/s5/p101;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860031220.

Zhu:2005:KLR

[ZH05] Ji Zhu and Trevor Hastie.

Kernel logistic regression and the import vector machine. Journal of Computational and Graphical Statistics, 14(1):185–205, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zimmerman:2017:TFV

[ZH17] Dale L. Zimmerman and Jay M. Ver Hoef. The torgegram for fluvial variography: Characterizing spatial dependence on stream networks. Journal of Computational and Graphical Statistics, 26 (2):253–264, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhai:2022:DIO

[ZH22] Yuqi Zhai and Peisong Han. Data integration with oracle use of external information from heterogeneous populations. Journal of Computational and Graphical Statistics, 31(4):1001–1012, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2050248.

Zhang:1997:MAS

[Zha97] Heping Zhang. Multivariate adaptive splines for analysis of longitudinal data. Journal of Computational and Graphical Statistics, 6(1):74–91, March 1997. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http://

www.amstat.org/publications/
jcgs/abstracts/march/
zhang.html;   http:/
/www.tandfonline.com/
doi/abs/10.1080/10618600.
1997.10474728.

Zhang:2003:OIA

[Zha03] Hao Zhang. Optimal interpolation and the appropriateness of crossvalidating variogram in spatial generalized linear mixed models. Journal of Computational and Graphical Statistics, 12 (3):698–713, September 2003. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

URL http://giorgio.
ingentaselect.com/cgi-
bin/linker?ini=asa{\&
}reqidx=/cw/asa/10618600/
v12n3/s11/p698;   http:
//www.tandfonline.com/
doi/abs/10.1198/1061860032265.

Zhang:2007:CGG

[Zha07] Cun-Hui Zhang. Continuous generalized gradient descent. Journal of Computational and Graphical Statistics, 16(4):761–781, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2022:MLA

[Zha22] Tonglin Zhang. Maximum likelihood algorithm for spatial generalized linear mixed models without numerical evaluations of intractable integrals. Journal of Computational and Graphical Statistics, 32 (4):1636–1648, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2163651.

Zheng:2017:TDS

[Zhe17] Tian Zheng. Teaching data science in a statistical curriculum: Can we teach more by teaching less? Journal of Computational and Graphical Statistics, 26(4):772–774, 2017. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zeileis:2008:MBR

[ZHH08] Achim Zeileis, Torsten Hothorn, and Kurt Hornik. Model-based recursive partitioning. Journal of Computational and Graphical Statistics, 17 (2):492–514, June 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2016:ATS

[ZHJ16] Chunming Zhang, Yu Han, and Shengji Jia. Accounting for time series errors in partially linear model with singleor multiple-runs. Journal of Computational and Graphical Statistics, 25 (1):123–143, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/abs/10.1080/10618600.
2014.966107;   http:/
/www.tandfonline.com/
doi/full/10.1080/10618600.
2014.966107.

Zhou:2005:ELR

[Zho05] Mai Zhou. Empirical likelihood ratio with arbitrarily censored/truncated data by EM algorithm. Journal of Computational and Graphical Statistics, 14(3):643–656, September 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zou:2006:SPC

[ZHT06] Hui Zou, Trevor Hastie, and Robert Tibshirani. Sparse principal component analysis. Journal of Computational and Graphical Statistics, 15 (2):265–286, June 2006. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhu:2017:AAA

[Zhu17] Yunzhang Zhu. An augmented ADMM algorithm with application to the generalized lasso problem. Journal of Computational and Graphical Statistics, 26(1):195–204, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/

                                        doi/full/10.1080/10618600.
                                        2015.1114491.

Zeng:2012:LTA

[ZHZ12] Peng Zeng, Tianhong He, and Yu Zhu. A lassotype approach for estimation and variable selection in single index models. Journal of Computational and Graphical Statistics, 21(1):92–109, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2019:MAV

[ZHZL19] Hua Zhou, Liuyi Hu, Jin Zhou, and Kenneth Lange. MM algorithms for variance components models. Journal of Computational and Graphical Statistics, 28(2):350–361, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2022:ELQ

[ZHZW22] Wei Zhou, Xin He, Wei Zhong, and Junhui Wang. Efficient learning of quadratic variance function directed acyclic graphs via topological layers. Journal of Computational and Graphical Statistics, 31(4):1269– 1279, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/

doi/full/10.1080/10618600.
2022.2069776.

Zhu:2012:BCI

[ZICT12] Hongtu Zhu, Joseph G.

Ibrahim, Hyunsoon Cho, and Niansheng Tang. Bayesian case influence measures for statistical models with missing data. Journal of Computational and Graphical Statistics, 21(1):253–271, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2004:ANM

[ZJ04] Ying Zhang and Mortaza Jamshidian. On algorithms for the nonparametric maximum likelihood estimator of the failure function with censored data. Journal of Computational and Graphical Statistics, 13 (1):123–140, March 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2016:TAM

[ZJA16] Yan Zhou, Adam M. Johansen, and John A. D. Aston. Toward automatic model comparison: An adaptive sequential Monte Carlo approach. Journal of Computational and Graphical Statistics, 25(3):701–726, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2015.1060885.

Zhang:2010:MDH

[ZJRI10] Qiang Zhang, Alison Snow Jones, Frank Rijmen, and Edward H. Ip. Multivariate discrete hidden Markov models for domain-based measurements and assessment of risk factors in child development. Journal of Computational and Graphical Statistics, 19(3):746–765, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhan:2022:FST

[ZK22] Tianyu Zhan and Jian Kang. Finite-sample twogroup composite hypothesis testing via machine learning. Journal of Computational and Graphical Statistics, 31(3):856– 865, 2022. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2020128.

Zhang:2018:PAS

[ZKS18] Zhengwu Zhang, Eric Klassen, and Anuj Srivastava. Phase-amplitude separation and modeling of spherical trajectories.

Journal of Computational and Graphical Statistics, 27(1):85–97, 2018. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zheng:2022:CES

[ZKS22a] Xiaotian Zheng, Athanasios Kottas, and Bruno Sansó. On construction and estimation of stationary mixture transition distribution models. Journal of Computational and Graphical Statistics, 31(1):283–293, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1981342.

Zhou:2022:ADD

[ZKS22b] Jiayuan Zhou, Kshitij Khare, and Sanvesh Srivastava. Asynchronous and distributed data augmentation for massive data settings. Journal of Computational and Graphical Statistics, 32 (3):895–907, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2130928.

Zhou:2010:MAS

[ZL10] Hua Zhou and Kenneth Lange. MM algorithms for some discrete multivariate distributions. Journal of Computational and Graphical Statistics, 19(3):645–665, September 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2013:PAC

[ZL13] Hua Zhou and Kenneth Lange. A path algorithm for constrained estimation. Journal of Computational and Graphical Statistics, 22(2):261–283, 2013. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhao:2016:APF

[ZL16] Tuo Zhao and Han Liu. Accelerated pathfollowing iterative shrinkage thresholding algorithm with application to semiparametric graph estimation. Journal of Computational and Graphical Statistics, 25 (4):1272–1296, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2022:VID

[ZL22] Muzi Zhang and Dennis K. J. Lin. Visualization for interval data. Journal of Computational and Graphical Statistics, 31 (4):960–975, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2066678.

Zhang:2023:IRL

[ZLP23] Peng Zhang, Ben Liu, and Jingjing Pan. Iteratively reweighted least squares method for estimating polyserial and polychoric correlation coefficients. Journal of Computational and Graphical Statistics, 33(1):316–328, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2257251.

Zheng:2022:VNB

[ZLPT22] Runbing Zheng, Vince Lyzinski, Carey E. Priebe, and Minh Tang. Vertex nomination between graphs via spectral embedding and quadratic programming. Journal of Computational and Graphical Statistics, 31 (4):1254–1268, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2060238.

Zhang:2018:LGP

[ZLR18] Ru Zhang, C. Devon Lin, and Pritam Ranjan. Local Gaussian process model for large-scale dynamic computer experiments. Journal of Computational and Graphical Statistics, 27(4):798–807, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zou:2023:BHH

[ZLS23] Yudan Zou, Yiqi Lin, and Xinyuan Song. Bayesian heterogeneous hidden Markov models with an unknown number of states. Journal of Computational and Graphical Statistics, 33 (1):15–24, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2231055.

Zhu:2021:LSA

[ZLW21] Xuening Zhu, Feng Li, and Hansheng Wang. Least-square approximation for a distributed system. Journal of Computational and Graphical Statistics, 30(4):1004– 1018, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1923517.

Zhang:2016:RAB

[ZLWZ16] Chong Zhang, Yufeng Liu, Junhui Wang, and Hongtu Zhu. Reinforced angle-based multicategory support vector machines. Journal of Computational and Graphical Statistics, 25 (3):806–825, 2016. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2015.1043010.

Zeileis:2007:RBS

[ZMH07] Achim Zeileis, David Meyer, and Kurt Hornik. Residual-based shadings for visualizing (conditional) independence. Journal of Computational and Graphical Statistics, 16 (3):507–525, September 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2007:SVD

[ZMSZ07] Lingsong Zhang, J. S.

Marron, Haipeng Shen, and Zhengyuan Zhu. Singular value decomposition and its visualization. Journal of Computational and Graphical Statistics, 16(4):833–854, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:OTA

[ZMY^{+}23] Jingyi Zhang, Cheng Meng, Jun Yu, Mengrui Zhang, Wenxuan Zhong, and Ping Ma. An optimal transport approach for selecting a representative subsample with application in efficient kernel density estimation. Journal of Computational and Graphical Statistics, 32 (1):329–339, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2084404.

Zhang:2021:OSG

[ZNR21] Tao Zhang, Yang Ning, and David Ruppert. Optimal sampling for generalized linear models under measurement constraints. Journal of Computational and Graphical Statistics, 30(1):106–114, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1778483.

Zhang:2014:SAF

[ZNYJ14] Xiaole Zhang, David J. Nott, Christopher Yau, and Ajay Jasra. A sequential algorithm for fast fitting of Dirichlet process mixture models. Journal of Computational and Graphical Statistics, 23 (4):1143–1162, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhao:2012:WBL

[ZOR12] Yihong Zhao, R. Todd Ogden, and Philip T. Reiss. Wavelet-based LASSO in functional linear regression. Journal of Computational and Graphical Statistics, 21 (3):600–617, 2012. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhou:2014:PCA

[ZP14] Lan Zhou and Huijun Pan. Principal component analysis of twodimensional functional data. Journal of Computational and Graphical Statistics, 23(3):779–801, 2014. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2021:EPS

[ZR21] Boqian Zhang and Vinayak Rao. Efficient parameter sampling for Markov jump processes. Journal of Computational and Graphical Statistics, 30 (1):25–42, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1763807.

Zhao:2014:PSR

[ZRL14] Tuo Zhao, Kathryn Roeder, and Han Liu. Positive semidefinite rankbased correlation matrix estimation with application to semiparametric graph estimation. Journal of Computational and Graphical Statistics, 23 (4):895–922, 2014. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:MBS

[ZSBS23] Ziang Zhang, Alex Stringer, Patrick Brown, and Jamie Stafford. Model-based smoothing with integrated Wiener processes and overlapping splines. Journal of Computational and Graphical Statistics, 33(3):883–895, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2289532.

Zeng:2019:SRC

[ZSK19] Pengcheng Zeng, Jian Qing Shi, and Won-Seok Kim. Simultaneous registration and clustering for multidimensional functional data. Journal of Computational and Graphical Statistics, 28(4):943–953, 2019. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:SPC

[ZSK23] Xinyi Zhang, Qiang Sun, and Dehan Kong. Supervised principal component regression for functional responses with high dimensional predictors. Journal of Computational and Graphical Statistics, 33(1):242–249, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2250411.

Zhang:2022:GFI

[ZSL^{+}22] Xiao Zhang, Xingjie Shi, Yiming Liu, Xu Liu, and Shuangge Ma. A general framework for identifying hierarchical interactions and its application to genomics data. Journal of Computational and Graphical Statistics, 32 (3):873–883, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2152034.

Zhang:2023:GCM

[ZSL23] Jingfei Zhang, Will Wei Sun, and Lexin Li. Generalized connectivity matrix response regression with applications in brain connectivity studies. Journal of Computational and Graphical Statistics, 32 (1):252–262, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2074434.

Zhang:2022:FAA

[ZSW22] Hong Zhang, Judong Shen, and Zheyang Wu. A fast and accurate approximation to the distributions of quadratic forms of Gaussian variables. Journal of Computational and Graphical Statistics, 31 (1):304–311, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2021.2000423.

Zhao:2004:IUV

[ZW04] Yanli Zhao and Melanie M. Wall. Investigating the use of the variogram for lattice data. Journal of Computational and Graphical Statistics, 13 (3):719–738, September 2004. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhu:2007:EST

[ZW07] Haihong Zhu and Hulin Wu. Estimation of smooth time-varying parameters in state space models. Journal of Computational and Graphical Statistics, 16(4):813–832, December 2007. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhu:2010:EPC

[ZW10] Zhengyuan Zhu and Yichao Wu. Estimation and prediction of a class of convolution-based spatial nonstationary models for large spatial data. Journal of Computational and Graphical Statistics, 19(1):74–95, March 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2022:BKT

[ZWF^{+}22] Qinyi Zhang, Veit Wild, Sarah Filippi, Seth Flaxman, and Dino Sejdinovic. Bayesian kernel two-sample testing. Journal of Computational and Graphical Statistics, 31 (4):1164–1176, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2022.2067547.

Zhao:2017:LPP

[ZWLZ17] Yunpeng Zhao, Yun-Jhong Wu, Elizaveta Levina, and Ji Zhu. Link prediction for partially observed networks. Journal of Computational and Graphical Statistics, 26 (3):725–733, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhu:2023:GOS

[ZWS23] Jiaqing Zhu, Lin Wang, and Fasheng Sun. Grouporthogonal subsampling for hierarchical data based on linear mixed models. Journal of Computational and Graphical Statistics, 33(3):1037– 1046, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2301093.

Zhang:2024:IES

[ZWZW24] Yi Zhang, Lin Wang, Xiaoke Zhang, and HaiYing Wang. Independenceencouraging subsampling for nonparametric additive models. Journal of Computational and Graphical Statistics, 33 (4):1424–1433, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2326136.

Zhu:2005:MTS

[ZY05] Jun Zhu and Wei Yue.

A multiresolution treestructured spatial linear model. Journal of Computational and Graphical Statistics, 14(1):168–184, March 2005. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhao:2021:APD

[ZY21] Zifeng Zhao and Chun Yip Yau. Alternating pruned dynamic programming for multiple epidemic changepoint estimation. Journal of Computational and Graphical Statistics, 30 (3):808–821, 2021. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2020.1868304.

Zhang:2024:KAD

[ZY24] Yilin Zhang and Songshan Yang. Kernel angle dependence measures in metric spaces. Journal of Computational and Graphical Statistics, 34 (1):280–289, 2024. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2357620.

Zhang:2008:SCA

[ZYKC08] Zhihua Zhang, Dit-Yan Yeung, James T. Kwok, and Edward Y. Chang. Sliced coordinate analysis for effective dimension reduction and nonlinear extensions. Journal of Computational and Graphical Statistics, 17(1):225–242, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhu:2010:DRA

[ZYZ10] Li-Ping Zhu, Xiangrong Yin, and Li-Xing Zhu. Dimension reduction via an alternating inverse regression. Journal of Computational and Graphical Statistics, 19(4):887–899, December 2010. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zheng:2008:MCM

[ZZ08] Yanbing Zheng and Jun Zhu. Markov chain Monte Carlo for a spatialtemporal autologistic regression model. Journal of Computational and Graphical Statistics, 17 (1):123–137, March 2008. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:FCD

[ZZD23] Maoyu Zhang, Jingfei Zhang, and Wenlin Dai. Fast community detection in dynamic and heterogeneous networks. Journal of Computational and Graphical Statistics, 33 (2):487–500, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2232852.

Zhang:2024:FLM

[ZZF24] Xiaochen Zhang, Qingzhao Zhang, and Kuangnan Fang. Functional linear model with prior information of subjects’ network. Journal of Computational and Graphical Statistics, 33(4):1150– 1159, 2024. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2024.2319163.

Zhou:2018:CEE

[ZZL18] Jie Zhou, Jiajia Zhang, and Wenbin Lu. Computationally efficient estimation for the generalized odds rate mixture cure model with intervalcensored data. Journal of Computational and Graphical Statistics, 27 (1):48–58, 2018. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic).

Zhang:2023:CPD

[ZZSW23] Yuzhao Zhang, Jingnan Zhang, Yifan Sun, and Junhui Wang. Change point detection in dynamic networks via regularized tensor decomposition. Journal of Computational and Graphical Statistics, 33(2):515– 524, 2023. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2023.2240864.

Zhou:2022:CEH

[ZZT22] Junyi Zhou, Ying Zhang, and Wanzhu Tu. clusterMLD: an efficient hierarchical clustering method for multivariate longitudinal data. Journal of Computational and Graphical Statistics, 32 (3):1131–1144, 2022. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

                                        /www.tandfonline.com/
                                        doi/full/10.1080/10618600.
                                        2022.2149540.

Zeng:2021:MBM

[ZZW21] Yanyan Zeng, Hongyu Zhao, and Tao Wang. Model-based microbiome data ordination: a variational approximation approach. Journal of Computational and Graphical Statistics, 30(4):1036– 1048, 2021. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2021.1882467.

Zhang:2023:APL

[ZZWS23] Haixiang Zhang, Lulu Zuo, HaiYing Wang, and Liuquan Sun. Approximating partial likelihood estimators via optimal subsampling. Journal of Computational and Graphical Statistics, 33 (1):276–288, 2023. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2023.2216261.

Zhang:2017:RMM

[ZZZS17] Yiwen Zhang, Hua Zhou, Jin Zhou, and Wei Sun. Regression models for multivariate count data. Journal of Computational and Graphical Statistics, 26(1):1–13, 2017. CO-DEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:/

/www.tandfonline.com/
doi/full/10.1080/10618600.
2016.1154063.